A Selected Bibliography of Publications by, and about, Enrico Fermi

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

16 October 2025 Version 1.257

Title word cross-reference

+ [AFRW42, FF48]. −1/2 [CT67]. 1/2 [CT67]. $13.95 [Ano96b, Ano95b]. $14.95 [Ano95c]. $17.50 [Bro68]. 1s [FA34]. $21.00 [Gol99a]. $29.99 [Seg15]. $3.50 [Gra61]. 3/2 [CT67]. $30.00 [Wes16]. 4/3 [BGJR24, Boc15]. $49.95 [Stu06]. 4d [Fer29g, Fer30k]. $50 [Ano62, G.64]. $6.95 [Bro73c]. 93^{237} [WS48]. ^{237} [ML48]. ^{239} [SP48]. ^{o} [Mon66]. _{93} [ML48]. α [OVPL15]. β [Fer33b, FBE^{+}34, Fer34n, Fer34o, HS19, Yan13b]. F_{1/2}(x) [WR63]. f_{t} [Fer42d]. ≫ [Bas01, For01]. h/k [FR27b, FR27a]. k [Fer42c, Fer42s]. ≪ [Bas01, For01]. ν [FMM45]. p [FS34a, FS34b]. s [Fer28e, Fer28g].

-Decay [Yan13b, FBE^{+}34]. -Fermi [OVPL15]. -rays [FS34a, FS34b, Fer34o]. -Strahlen [Fer34o]. -Terme [Fer28e, Fer28e]. -Terms [Fer28g]. -Value [FMM45].

/2nd [KGSO55]. /FA [KGSO55]. /The [KGSO55].

0 [Bat06]. 0-226-12111-9 [Bat06]. 0-226-81664-8 [Lau13a]. 0511222 [Esp05].

1 [Ano22, Gan02]. 11 [Fer42j]. 11.25 [Ano96b]. 121 [FSA41]. 1600-1980 [GHMP87]. 19 [Bro68]. 1911 [Meh75]. 1922/23 [CS99]. 1930s [BR94, De 05c, Kra92, Orl98, Stu84, Tur06b, Stu79]. 1933 [CCJ^{+}34, Fer33c]. 1938 [Fer38e, Fer39b, Fer62a, Fer70a, All62a, Rab63]. 1939 [Sei90]. 1939-1945 [Sei90]. 1939/1946 [HA62, HA69, HA90]. 1941 [Ano41a]. 1942 [Ano11, Lau46a, Lau46b, Wat82]. 1942-1952 [Fit13]. 1943 [F^{+}43g]. 1945 [Hah75a, Pau45, Sei90]. 1946 [CSW97b]. 1947 [HA69, Sei90]. 1947-1977 [Sei90]. 1947/1952 [HAD90]. 1949 [FR76]. 1950s [BDH89, Sei90]. 1954 [ASA55b, All57, Ano54b, Bai55, BC55, Cam54, Fer65b, Gla02, Wig55, Pal69, Sha67]. 1955 [Bet55]. 1960 [Zuc70]. 1960s [Mla98]. 1968 [Ano68b]. 1969 [WH72]. 1972 [Meh73]. 1974 [Ano09d]. 1976 [MP77]. 1977 [Gro78, Sei90]. 1982 [MR86]. 1983 [GHMP87]. 1988 [DGS89, Rig00]. 1992 [GER^{+}92]. 1994 [Ano94b]. 1996 [Ano96a]. 1st [Fol86, GHMP87].

2 [Ano98, Ish19, Rab66]. 2001 [Ber02a]. 2002 [Bar05]. 20th [APB96, ABP98, Meh73]. 22 [CCJ^{+}34]. 25th [Ano05]. 29 [Ber02a]. 29.00 [Cal13]. 2nd [Gri96, KGSO55].

3 [FZ44]. 3.15 [Buc73b]. 368pp [Wes16]. 3D [KHB12].

40th [ATFF82, MKR87]. 41 [Zuc70, Bra69, Cer69, Rei69, Smi68].

5 [Wes16]. 50th [Ano95a]. 54 [Laf71]. 5th [KGSO55].

60-year-old [Ano15b].

7 [Tur08].

80th [FPLR59].

9 [Bat06, Mal08]. 93 [Nod34]. 93d [Ano34c]. 93rd [Ano34f]. 978 [Ish19, Mal08, Tur08, Wes16]. 978-0-226-81664-7 [Cal13]. 978-1-107-16365-2 [Ish19]. 978-1-62779-005-5 [Wes16]. 978-88-339-1746-7 [Tur08]. 978-88-518-0081-9 [Mal08].

A. [Ano55b, Fer35a]. AAPT [Fer52g]. Abschlusses [Pau25]. absence [Har75]. Absolute [Fer23j]. Absorbed [Fer42p]. Absorption [AF35a, AF35b, AF36a, AF36b, AF36d, AFS39, AF40, AFW42, AF42a, AF44a, AFW^{+}47, AF4x, BJ41, Fer39a, FW41, Fer42a, Fer42f, Fer44a, AFW^{+}57, Fer29d]. abstract [LFA^{+}52, NAF^{+}52]. Academician [Ano34c]. Academy [Ano41a, Ano41b, Bar05, WH72]. Accelerator [Cre10, Wal74b, Wal74a, Ano51, Ano69]. Accelerators [Pes03]. Accidental [Dav82]. According [BM35, Fer27d, Geo32, Som27]. Account [Jon85b, Sei90]. Accurate [ABF14]. achieve [Ste59]. Achieved [Kae35]. Action [Ano34c, FR38, FR33b, FR33c, FAP^{+}34]. Activation [AF44a]. Active [HS89, PAF^{+}42, Fer42t, HS39a]. Activities [F^{+}43a, F^{+}43b, F^{+}43g, F^{+}43c, F^{+}43d, F^{+}43e, F^{+}43f, Fer44d, F^{+}44c, F^{+}44d, F^{+}44e, F^{+}44f, F^{+}44g, F^{+}44a, F^{+}44b, Tur06b]. actuel [Fer32a]. ad [Fer52b, Seg86a]. Adams [Hol90d]. adaptation [JB83]. adaptive [ZZWK12]. adiabatic [Fer23h, FP26, LC02]. adiabatiche [Fer23h, FP26]. adiabatici [LC02]. Advances [Kae35, Bar87]. adventurous [ABES12]. Advisory [Ano41a, OBC^{+}49, Sei90]. AEC [Ano49, Ano63b, Ano68c, Ano70, Bet63, OBC^{+}49, Wig62]. Aestivation [BHR19, SAĆ16]. Affair [Cal13, Lau13a, Tur12]. Affairs [GR63]. Affinità [Fer29a]. affinity [Fer29a]. after [Cal01, Cra00, GHH02, Kae48, Mal03a]. Again [Esp05]. against [Bax46, Bax68, Fer26a, Fer27b]. Age [Ano67a, GR63, Lau46a, Sch17, Stu18, Wat93, Wat04, Her79, Mag09, Mar96, Rif06, RK95, SH16, Ano67c, Ban04, Lau46b, Ste59, Gar18, Lan17, Sei90, Wes16, Lan18]. Agency [Fis97, IAE97]. Ages [VD15, Oak18]. Agnew [Ano12b]. ago [PZHC09]. ai [Ama02b, Ano02, FV02, Gan02, Seg83]. Air [Fer39a, FS58a, FS61]. Air-cooled [FS61]. aktiver [HS39a]. al [Bat03b, Fer49c, Ver01a, FM47f]. Alamogordo [LW71]. Alamos [Fer51d, Fer54c, Fit99, Fit13, MKR87, Ree17, Agn04, Ano75, BHB80, Esp08b, Fer80, Fey80, Gar04, HTS83, SR92, Sza92]. Alan [Sol68]. Alarm [Rid84]. Albedo [FAW38, Fer42e]. Albert [Ein39, Her79]. alcalini [Fer30m]. alcune [Fer22h, Fer27e, Fer28f, Fer28g]. Alcuni [Fer23a, Pon92, Pon13, Sal01]. Alfred [Mon66]. algorithm [ZZWK12]. Alien [Reg87, Sho09]. Aliens [AK16, AK17, SM09, Tar03, Web15, Gra16a]. alienswhere [Web02]. alkali [Fer30o, Fer30m]. Alkalien [Fer30o]. alkaline [HS39b]. alla [CS00c, Fer28c, Fer28d, Fer28f, Sal02b]. alle [GER^{+}92]. Allen [Fri55b]. allgemeinen [Fer23b]. Allgemeines [Som28a]. allievi [Pon93]. Allison [Fer42l]. allow [Fer23h]. allowed [HS19]. Alloy [Fer42x]. Alone [Gri11, Gri18, Pap78, Cra00]. Alpha [Ano34c, FS41]. Alpha-Particles [FS41]. Alpher [AM23]. Also [Kat75]. Alternating [FR25a, FR25b, FR25c, FR25d]. alternato [FR25b, FR25c]. Aluminum [FM47d]. Always [KS99, Seg93]. Alwyn [Sei90]. Amadesi [Dra72a]. Amaldi [Bad67, Bat03b, Seg79]. Amelio [AS96]. America [Sei90, Zuc70, Kev77, Kev87, Kev95, Mal03b, Sop76, Sop80, Sop88, Stu84, Stu85, Tel02, Zuc70]. American [Bet55, Sei90, Sol68, WH72, Bad05, CHW91, GC99, Gon96a, Val06, WC84, Rom91]. americana [Val06]. Amherst [Bar05]. amici [Pon93]. ammettono [Fer23h]. ammonia [Fer32c, Fer32e, Fer32f]. among [FM03, Pin03]. Amount [Ano39a]. Analisi [Der64]. analitica [Fer23a]. Analogue [Mar59]. Analyses [Mon66]. Analysis [Fer44c, FMA54, D’A03, Der64]. Analytic [TSH^{+}19]. analytical [Fer23a]. Anders [Sei90]. angolari [Fer23h]. Anguished [Sal03b]. Angular [AFNY52a, AFMN53, Fer51a, Fer23h]. anni [Ama02b, CS00b, CDS01, Dra76, PS07]. Annihilation [BM35]. Anniversary [ATFF82, Ano95a, FB64, MKR87, BRV84, Pon55]. Announced [Ano39d]. Annual [Ano63b, Ano05, Bet63, G.64]. anomala [FR27b]. anomale [FR27a]. Anomalous [FR27b, Fer28a, FR27a]. answers [KS99]. antimateria [Ama61]. antimatter [Ama61]. Anton [Fis10a, Fis12a, Fis10a, Fis12a]. Anwendung [Fer28i, Fer28b, KGSO55]. Any [GB89]. Anyone [AK16, DS92]. aos [Seg87a]. Apfel [Büh98b]. Appendix [FSA41]. apple [Büh98b]. Application [CIR04, Fer28i, FMM47, GM85a, GM85b, GP31, Fer28b, KGSO55, Roq92]. Applications [Bro68, Sei55, Fer22h, Fer38f, SS11]. Applicazione [Fer28f]. applicazioni [Fer22h, Fer38f]. Applied [De 03]. Applying [Fer28f]. appreciation [Gol99b]. Approach [BPP18, De 05b, SN83, BP68, De 06b, EFGP99]. approximant [EFGP99]. Approximation [RNKS04, SZ62, WR63, SS11]. Approximations [CT67, MP56]. April [Bet55]. Arcangelo [Min02]. Archibald [Ano68b]. Archives [Bar05, Bat03b, Cra02, LRS00]. archivi [Bat03b]. Argomenti [Fer26a, Fer27b]. Argonne [AFMW42, Fer42s, F^{+}42b]. Arguments [Fer26a, Fer27b]. Arms [CF53a, Sei90, Sol68]. Army [Sei90]. Arnold [Nis18]. Article [Ano95a, FV02]. Articles [GR63, Sei90]. articolo [FV02]. Artificial [AGR04, ADF^{+}35d, Ano34a, Dys60, Fer34a, FAD^{+}34a, Fer35d, Fer35e, Fer39b, Fer65a, Fer70a, GLR06, GR09, GLR20, AFR37, Fer38f, ADF^{+}35a, FAD^{+}34b, Fer34d, Fer35a]. artificiale [AFR37, Fer34d, Fer35a, Fer35d, Fer35e, Fer38f]. artist [KS99]. Arts [WH72]. Ascent [Bro73a]. asintotica [Som32b]. Asks [AK16]. Aspects [SW72, Meh75]. Associate [Ama02a]. associated [Fer28d]. Association [Ano20]. assoluta [Fer23j]. assorbimento [AF36b]. Assume [G.64]. Astbury [Bad67]. Asteroid [Pap78]. Astronautical [KGSO55]. Astronautische [KGSO55]. Astronomy [Bro68, Kae35, Dea13, Mor58]. Astrophysics [Tau63]. Asymptotic [Din58, Som32a, Som32b]. Asymptotische [Som32a]. atmosfera [Fer38b]. atmosphere [Fer38b]. Atmospheric [Fer42g]. Atom [Ano34b, Ano39b, Ano39a, Ano39d, Ano54a, Ano55c, Bor57, Bor63, Bos50, Dro20, Kae45, Lau51, Mon66, Pau25, Rod19, Sei90, Som32a, Ano51, Bru11, CHW91, Fer29a, SW87, WC84, dL63]. Atombaues [Fer28i]. Atome [Som33, dL63]. Atomen [Fer24g]. Atomi [Fer30a, Fer25f, Fer25d]. Atomic [Ano34b, Ano34d, Ano39c, Ano39a, Ano41a, Ano41b, Ano50b, Ano53a, Ano53b, Ano67a, Ano16, Bro68, CBH09, Coc62, Com53, Com56, DJ99, Fer26b, Fer27e, Fer28i, Fer28g, Fer30f, Fer30h, Fer34g, FRD34, Fer46a, Fer46f, Fer57b, FR13, Fis97, Fri54, Fri55b, Gol80, Gow64b, GR63, HA69, HDHA89, HAD90, IAE97, Kow53, Lan94, Lan17, Lau46a, Lau51, Lef95, Man57, Mon66, Mor57, PG86, Rho86, Rho99, Sei90, Smy45b, Smy45a, Swi45, Tur06a, Wes16, Zin55, Bak76, Bar87, Bru11, CCJ^{+}34, Cra69, EE70, Fer28c, Fer28f, Fer28b, Fer30p, Fer32a, Fer61d, Flü39, FPLR59, Gol82b, GS22, Gow64a, GA74a, GA74b, Gro67a, Gro67b, Her79, Kra15, Lan01, LW71, Mac86, Mac87, Pur63, SH16, SR92, Tho27, TSH^{+}19, Van03, Ano67c, Ban04, Buc83, Kie13, Lau46b]. Atomic [Ste59, Sei90]. Atomica [Fer28c, Fer50a, Bru11, Fer30b]. atomici [Fer30h]. atomico [FRD34]. atomique [Fer32a, Gol80, Mon66]. atomiques [CCJ^{+}34]. Atomkerne [Flü39, Fer30p]. Atomreaktor [Lan01]. Atoms [DE04, Fer25d, Fer26c, Fer28b, Fer54k, Fer57b, Fer82c, Fer87, LS77, Rut20, Sol03, Fer24g, Fer25f, Fer01, Lie81, Lie82, Mar75, Som33, Spr91, Fer30a, Man57, Som32a, Ano95b, Cam54, Fri55b]. Attempt [Fer33b, Fer34o]. Attention [Ken54]. attuale [Fer29d, Fer32c]. attuali [Fer29e]. audience [Ano12b]. auspices [CCJ^{+}34, Smy45b]. Austrittsvorgänge [Som28a]. autobiography [Seg93]. autoionization [ACM10]. automata [Str98]. aux [Seg84]. available [Flü39]. Avvengono [Fer22c]. avventurosa [Tur08]. Award [Ano61, Ano67b, Ano68b, Ano90, Ano09b, G.64, Wal64a, Ano58, Ano59, Ano62, Ano63a, Ano63b, Ano65a, Ano66, Ano68c, Ano73, Ano79, Ano81, Ano93b, Ano94b, Ano96a, Bet63, Hol90a, WB64, Wig62]. Awarded [Ano50b]. Awarding [KGSO55]. Awards [Bar87]. Azione [FR33b, FR33c, FAP^{+}34, FPR34a, FR38].

B [Bro68, Mag61, Wil71]. B. [Dra72a]. back [Mei64a, Mei64b]. Backyard [Sei90]. Baeyer [Ano94a, Ste93]. Bagels [Win86]. Bailyn [Zuc70]. Bakerian [Rut20]. Ball [Ano39a, Sei90]. bande [Fer25g, Fer32e, Fer32f]. Bands [Fer25g, Fer32e, Fer32f]. Bare [Ano39a]. Barium [HS89, HS39a]. Bariumisotope [HS39a]. Barn [Gou59]. Based [FMT49b, BDH89, FMT49a]. Basic [Gra61, Seg15]. Basin [Car15]. basis [Som28a, Som28b]. Battle [Kae45]. Bayes [D’A05]. Bayesian [D’A03]. be [Döb22, Flü39, Gro62, Gro83, KS99, Pap78, Rig84, Sch15, AFW42, AFRW42, FF48, FM47f]. Beam [FZ44]. Beams [Ano74]. Beck [Fer34p]. Becquerel [Mon66]. Began [Ano67c, FW67, Kae48, Fri54]. Beginnings [Gon96b, Led03]. Behind [Bar05, De 06a, Fri01, LS92]. being [De 05a]. Beings [Sch81, Tip80, Tip81b, Tip81c]. Belloni [Goo91]. Below [Fey80]. Belt [Pap78]. Bemerkung [Fer34p]. Bendiner [Sol68]. Benedict [Ano73]. Benson [Sol68]. Berechnung [Fer28e]. Berekingen [Fer24a]. Berlin [Stu06, Sim03]. Bernard [Zuc70, Ano93a]. Bernardini [Gra61, Mal05, Stu06]. berücksichtigt [CSW97a]. Berylium [Ano34c]. Beryllium [AF42b, FF42]. best [Pur63]. best-kept [Pur63]. Bestandsaufnahme [GHH02]. Bestimmung [Fer28b]. Bestrahlung [HS39b]. Beta [GR09, HS19, Jen00, Kon55, Rob83, Wil68, BRV84, Fer33b]. Beta-Decay [Kon55, BRV84]. Bethe [Ano61, Ber80]. Better [Sei90]. Bettina [Lan17, Wes16]. Between [Arb03, FM47c, Gol89, Stu18, Arb00, Fer22a, Fer22f, Fer23c, Fer24g, Fer25g, Fer25f, Fer25d, Fer42f, FU43b, Fer01, Mai91, MP56, Ste77, Tam34, Val06]. Beweis [Fer23b]. Beyond [Ano05, JAH09, Mer01, Rod19, RNKS04, Dro20]. Bi [FM47f]. Bias [Hoo03, Ore03b]. Bibliografia [Ano02, Dra72b]. Bibliography [Ano02, Dra72b]. Bibliothèque [Mon66]. Big [Ano39d, Ano17]. billion [Bai17a]. billion-year-old [Bai17a]. Binding [Fer42i]. biografia [Mal03b]. biographia [Seg71, Seg87b, Seg09]. Biographical [Bar05, Seg62, Sei90, Kur02]. Biographie [Kra88]. biographies [Ano65b]. Biography [Sei90, GC99, GH90, Hol90d, LS92, Mal03b, Seg71, Seg87b, Seg09]. Biology [Kae35]. Bird [Sol68]. Birth [Ama03, BC03, BH83a, BH83b, Com53, FB64, Hol03, Kra78, Lan17, Lau46a, Lau46b, Mal03a, PZHC09, Sal02a, Wat93, Wes16, ACM10, Bas01, Ber02a, BH82, SH16]. birthday [FPLR59]. bis [Büh98b, Ern92, Fis10a, Fis12a]. Bismuth [Fer42h, FR33a]. bismuto [FR33a]. Bodies [Fer23k, Fer25b]. body [Fer29h]. Bohr [Ano39a, Lau13b, Sch00]. Boiler [Bun83]. boilers [Fer47a]. Bollati [Tur08]. Bologna [Stu06]. bolschich [Fer52h]. Boltzmann [Faz15]. Bomb [Ano54a, Ber88, FS95, Gol92, Gon96b, Kae45, Ken54, Lef95, MD67, Moo95, Mor57, Rho86, Rho95, Sei90, Swi45, Wol09, Bak76, Bru11, FR76, Gro67a, Gro67b, LS92, Mac86, Mac87, Mer01, Pur63, Sch15, SR92, Smy45b, Sti47, Sti76, Van03, Sei90]. bomba [Bru11]. Bombardamento [AFRS34, Fer34k, ADF^{+}34b, ADF^{+}34c, ADF^{+}34d, ADF^{+}34a, ADF^{+}35e, ADF^{+}35b, ADF^{+}35c, Fer34j, Fer34l]. Bombarded [AFH39]. bombarding [Ano34c]. bombardirovkoi [ADF^{+}35a, FAD^{+}34b]. Bombardment [Fer34a, FAD^{+}34a, Fer34h, Fer34k, Fer34i, Fer39b, Fer70a, Fri39, AFRS34, ADF^{+}34b, ADF^{+}34c, ADF^{+}34d, ADF^{+}34a, ADF^{+}35d, ADF^{+}35a, ADF^{+}35e, ADF^{+}35b, ADF^{+}35c, FAD^{+}34b, Fer34j, Fer34l, Fer65a, MF39b]. Bombs [Fer70d, Fer71b, Sei90]. Bonolis [Stu06, Mal05]. Book [All62a, Ama73, Ano94a, Ano95b, Ano95c, Ano96b, Bad67, Bad71, Bal63, Bat06, Bra69, Bro68, Bro73c, Buc73a, Buc73b, Cal13, Cam54, Cer69, Cre10, Dro20, Fal73, FBL64, Fra95, Fri55b, Gar18, Gol70, Gol99a, Goo91, Gra61, Gri96, Hay16, Hub13, Ish19, Kai04, Lan18, Lau13a, Lut66, Mag61, Mal05, Mal08, Man57, Min02, Moo95, Pal69, Rab63, Rab66, Rei69, Seg15, Sei90, Sha67, Smi68, Sol68, Ste93, Str98, Stu06, Tal17, Tar03, Tur08, Wei71, Wes16, Wil71, Win86, Zuc70, Dea13, Dra72a, Dra72b, Win86]. Books [Bar05, Gra61, Seg15, Sei90]. Boorse [Sei90]. Boringhieri [Tur08]. Born [Gre05]. boro [FR38]. Boron [AF42a, AF44a, AF4x, BFMM44, FR38, FMM47]. Bose [And99, Bul81, Din58, Faz15, GRJ03, Kap40, Mos50, Sch89]. both [Cra88]. Bothe [Six88]. bound [Pai86]. Boundaries [Cre10]. Boundary [FvN51]. Boys [AFM88, BCR09, Cod88, AS96, Ano01b, Bia90, FM03, Lom93]. Branching [AFG41]. brauche [Büh98a]. Brecht [Sch99b]. Breeder [FWN^{+}45]. Breeding [Fer44b, Fer45]. Brief [Ano66, Tip81a]. Briefe [Ern92]. brilliant [Mag09]. Brillouin [MP56]. Bringing [Stu85]. Britain [Ano50a, Esp08b, Gow64a, Gow64b, GA74a, GA74b, Rid84]. British [Ano20, Coc62, Fak83, Sza92]. broadcast [Swi45]. Broadway [Sch00]. Brookhaven [Gol89]. Bruchstücke [HS39a]. Bruno [Hay16, Seg15, B^{+}13, Bon05, Clo15, LMR06, Maf92]. Brussels [CCJ^{+}34]. Bruxelles [CCJ^{+}34, Far01]. Bruzzaniti [Tal17, Bru11]. Build [GB89, Gol92, SR92]. Builders [MD67]. Building [Wat74]. bulk [Spr91, vN96]. Bull [ED83]. Bulletin [Ano93a, Ano93b, GR63]. Buried [Ano09a]. Burns [Sei90]. Bursts [Cow09]. bury [vHEGM12]. Bush [Gol92, Hol90b]. businessman [Tur06b]. Byte [Lan12].

C [Fal73, FU43b, Sei90, Fer47a]. C. [Ano55b]. calcolo [Fer22h, Fer26e, Fer30i, Fer30j, Fer31d, Fer28g]. Calculating [Fer42x, Fer26e]. Calculation [Fer30i, Fer30j, Fer42b, ABF14, Fer28e, Fer31d, Tho27, Fer28g]. Calculations [FR48, Fer24a]. Called [Gra16b]. Cambridge [Bar05, Ish19, Ano34b]. came [Sch15]. cammino [AF36c]. campi [Fer26d]. Campo [Fer21b, FR25b, FR25c, FR33b, FR33c]. can [Gro62, Gro83, Kow15, Flü39]. Canadian [Coc62]. Cancer [Kae39]. Candid [HH04a]. Cannon [Ano39a]. Cannonball [Ano39d]. canonical [Fer23g]. canoniche [Fer23g]. Capital [Kai04]. Capture [AF39, AF52, De 05b, FSA41, Fer42e, FT47, De 06b]. caratteristici [FR38]. Carbon [AF40, Fer42f, Fer31e]. Cariche [Fer21a]. Carlo [Mal05, Fer46g, FR48, Stu06]. Caroline [Sol68]. Carried [F^{+}42g, F^{+}43h, F^{+}43i]. Case [GLR20, Ken54, Hol78, Seb97, SW87]. cases [DE06b, DE07]. caso [CM02, Mal08, Seb97]. casuality [Geo32]. Catalonia [GHMP87]. Catastrophe [Sag83, Kow15]. causalità [Fer30d, Fer30e, Fer30l]. causalité [Geo32]. Causality [Fer30l, Fer30d, Fer30e]. Cause [Sol68]. caused [ADF^{+}34b, Fer34l]. Cavendish [Hug03]. Cd [Fer42d]. Celebrations [Bas01, BC03]. Celebrazioni [Bas01]. cellular [Str98]. Census [Bar05, FR48, CHU87]. Census-Taking [FR48]. centenario [Bas01, Ber02a]. Centenary [BC03, Kra78, Bas01, Ber02a]. Centennial [Ber03, Led04]. Center [Ano05, Pie01, Pie01]. centrali [Cal75]. Centre [Meh73, Fer76]. Centro [CL03, Pie01]. Century [ABP98, Bar05, Buc73b, Sal03b, APB96, Bra09, Dar04, Kra99, Meh73, Sal01, Ste77, Wei77]. Ceres [Kow15]. certain [Gre05]. Cesarina [CM02]. CETI [Sag73]. Chadwell [Sei90]. Chadwick [Bro73b]. Chain [All62b, AFW^{+}47, And73b, Ano68d, Ano95a, Ano07, Fer41, F^{+}42a, Fer42k, Fer42m, Fer42l, FU43a, Fer46d, Fer47b, Fer52a, FL57, FS58b, FL58, Fer70b, Gol62, Wat74, AFW^{+}57, And73a, Ano11, F^{+}42g, F^{+}43i, Fer47a, Ste59, SF44]. Chain-Reacting [AFW^{+}47, Fer47b, AFW^{+}57]. Chains [CBH09]. Challenge [Ćir09]. Chance [SM09]. Change [Fer42y, KS99]. Changes [CFW42, Fer42g, Wal64a]. Channels [FW42]. Characteristic [FR38]. charged [Fer24g, Fer25f, Fer01]. charges [Fer21a]. Chatting [Hof90]. Che [FB69, CS00c, CS00b, Fer22c, Fer23d, Fer23e, Fer23h, Fer24b, Maf92]. Chebyshev [CT67]. Chemical [Fer42i, SW48]. Chemie [Fal28]. Chemistry [Fer26b, Kae35, Sch15, AS11, CHU87, Fal28, Bar05]. Chicago [Ano95b, Ano95c, Ano96b, Ano22, Bat06, Bro73c, Buc73b, Cal13, Cam54, Lau13a, Wei63, Wil71, Agn04, AAF^{+}52, Ano95a, Ano07, Fer33c, Fer49e, Fer50d, Fer50e, Fer65c, Gan02, Gar04, Gol82a, Pri95, Shi91, Tel04]. chimica [Fer26b]. Chomet [Bro68]. chose [Maf92]. Christian [Ano94a, Ste93, Kra92]. Christophe [dL63]. Christopher [dL63]. Churchill [Ano17]. Chwolson [Ber01]. ci [Mon66]. ci-dessus [Mon66]. Cinema [EL07]. cinematography [KGSO55]. Circulating [Fer44c]. Citizen [Sei90]. Città [Lom93]. city [Lom93, Kie13]. civilization [KB89]. Civilizations [SW66, Kre71, Oli75]. Civilized [Bau85]. classica [Gal02]. ’Classical [Mar59, BGJR24, Jac62, Mar59, SZ62, Wri64, Wri65, Gal02]. Classics [Mon66]. Climate [Gol92]. Climatic [Sag83]. Close [Seg15, Hay16]. Closed [Goe48, Goe49]. Closer [Ano12c]. cloth [Gol99a, Seg15, Stu06]. Cloud [Sei90]. club [Ber96]. clues [Dug61]. cmerti [Pon55]. Co [Wes16]. codiscoverer [CSW97b]. Coefficient [CK08, F^{+}43h]. coi [Fer23f]. Cold [Cal13, Lau13a, Maf92, Tur12]. collaboration [AM23]. Collapse [MRR^{+}02]. Collected [All62a, Pal69, Rab63, Rab66, Sha67, Tau63, Wig96b, Bal63, Dir03, Fer62a, Fer65b, Mor63a]. Collection [Ano09d]. Collimation [FZ44]. collision [Fer26h]. Collisions [Fer25d, Fer51a, Fer53a, Fer54f, Fer54e, Fer24g, Fer25f, Fer01, Hei55, Hei83, Kra92]. Colloquium [ACM10]. Colomb [dL63]. Colonization [Jon76, O’N74]. colored [Dys10]. Columbia [Ano64, Agn04, Ano39a, Ano39d, Fer42q, Fer55, Fer70c, Fer85, Lau51]. Columbus [dL63]. Column [FZ44]. come [CS99, DS07]. Coming [Ano67a]. Commemoration [Per02, Ama02a]. Commemorazione [Ama02a, Per02]. Comment [BHR19, CFU47, Gli04, Wol04, Esp05, Sal02b]. commentaries [Yan13c]. Commentary [Ern92, DP97]. commento [Sal02b]. Comments [Led04, Ros04]. Commercial [Fer38a]. Commerciali [Fer38a]. Commission [HA69, Sei90, Buc83, HDHA89]. Committee [ACF^{+}42, Ano41a, Ano41b, Sei90, FSA41, G.64, OBC^{+}49]. Commodore [Win86]. Common [Ano99]. Communication [Cre10, Sel53, Sag73]. Communications [CM59, Szi71]. communicative [Kre71]. Community [Hab69, Kev77, Kev87, Kev95]. Company [Lan17]. Comparison [FHN44]. Compensated [Ano53a]. Competency [Nis18]. Competition [Pin03]. Compiled [Ano95c]. complesso [Fer29g, Fer30k]. Complete [Gol99b]. completion [Pau25]. Complex [Gol80, Fer29g, Fer30k, Gol82b, Pau25, Str98]. complexe [Gol80]. Composition [Fer36a]. Compounds [ML48]. Computation [Fuk14, MS38]. computers [Faz15]. Computing [Lew21, Sei03, BDM81]. Con [AFRS34, Car05]. Conant [Wil71]. Concept [BR96, Fal73, Tip81a, CW73, FP26, GS22, Hoc83]. conception [Meh73]. Concerning [Bri83, HS39b]. Concluding [Sal03a]. Condensate [Sei04, GRJ03]. Condensed [Fer39a, Fer40a]. conducted [PS07]. Conduction [DuM28]. Conference [Ano03, Ano05, DGS89, FU43b, Man57, Pau30, SA79, Ber03, Far01, Fer57b]. conferences [Meh75, WH72]. Conferencias [Fer34b]. Conferenze [Fer50a]. Confessions [Sho09]. Confirmation [CRH^{+}56, DuM28]. Confirms [Sei04]. Congress [KGSO55, Wal64b]. Congressional [G.64]. conjectures [San81]. connessi [Fer28d]. Conoscere [Ber02a, BB04a]. conscience [Wil70]. conseil [CCJ^{+}34]. Consensus [Jen00]. Consequences [Moo92a]. conservati [Der64]. Conservation [GLR14]. consideration [CSW97a]. Considerations [Fer24b]. Considerazioni [Fer24b]. Consiglio [Fer49c]. consistent [Mar75]. Constancy [Fer23j]. constant [Sch99b]. Constants [Fer25g, Fer44e]. constituents [Fer33a, Fer34e]. Constitution [Rut20]. constrain [SHM19]. contact [Nea14]. contain [Fer24b]. contemporanea [Seg76, Seg83]. Contemporary [Fer29e, Seg76]. contengono [Fer24b]. content [Flü39]. contenuti [Ano02]. continues [PZHC09]. continuitá [DE06a, DE15]. continuity [DE06a, DE15]. Contract [FSA41]. contraddizione [Fer23c]. contradiction [Fer23c]. Contribution [Gol62, Rub03, FV02]. Contributions [Ano56a, Wig62, Orl98]. contributo [FV02]. contro [Fer26a, Fer27b]. Control [Fer42u, G.64, Ano51, F^{+}43h, KGSO55]. Controversy [Bri83, Jen00]. Converging [Din58]. conversation [Goo01]. conversations [HH04a]. Cooled [FS58a, ACF^{+}42, FS61, LCM^{+}42]. Cooling [Fer42m, Fer42l, Wig43]. Cooper [Gol99a]. Cooperation [Pin03]. coordinate [Ber02c, Fer23h]. coordinates [Ber02c, Fer23h]. Copenhagen [Sch00]. Copernican [Bai18]. Copernicanism [Ćir09]. Corbino [Fer37b]. Cornell [Cor04]. Corner [Sol68]. corpi [Fer23k, Fer25b]. corpo [Fer29h]. corpuscoli [Fer25f]. corrected [SHL^{+}18]. correction [Fer28e, SW87, TSH^{+}19, Fer22a, Fer23c]. Corrections [Fer28g]. Correlate [Kae35]. Correspondence [SF44]. Correzione [Fer22a, Fer23c, Fer28g]. Corso [Fer25a]. cosa [FB69]. Cosmic [Fer49f, Fer49g, Fer54b, Tar03, Fer49b]. cosmological [AM23]. Cosmology [Rei72, Wes90, KS99, MR86]. cosmos [LBS17]. Cosmotron [Fer53a, Fer54e]. Cost [Sch94]. Costandina [Sei90]. costante [Fer23j]. costanti [Fer25g]. costitutive [Fer33a, Fer34e]. could [Pap78]. Coulomb [LT05]. Council [Fer49c]. Count [Ano34d]. Countdown [Sei90]. counter [LMR06]. coupling [Tre76]. Course [Fer25a, Fer46c, Fer46b, MR86, Ano95c, Fer46i, FH46, Fer49e, Fer50d, Fer50e, Fer65c]. cousins [And99]. covering [Dea13]. CP [FZ44]. CP-3 [FZ44]. Cracknell [Fal73]. Craters [Szi71]. Crawford [Bar05]. created [Ano34e]. Creation [Kra18]. Credited [Ano39b]. cristalli [Fer22e, Fer32b, Fer34f, Fer82a, Lut66]. cristello [FR33a]. Critical [Fer42c, D’A03]. Cronin [Bat06, Kai04, Mal05]. Cronologia [Bon02]. Crookes [Mon66]. Cross [AFW42, AF42a, AF42b, AFW^{+}47, AF4x, AFNY52b, AFL^{+}52, AFLN52, BFMM44, Fer42f, FMM45, FMM47, AFW^{+}57]. Cross-section [AF42a, AF42b, BFMM44, FMM45]. Cross-sections [AFW42, Fer42f]. crowned [De 05a]. Cruz [Sei90]. Crystals [Fer32b, Fer66a, Fer22e, FR33a, Fer34f, Fer38c, Fer82a]. cultura [Mai91]. cultural [JB83]. culture [Mai91]. cultures [Ste77]. cura [Fer35a, Min02]. Curie [Mon66, Six88]. Curies [Kae48]. current [Fer29d, Fer32a, Fer32c]. Curve [Sei04]. Cylindrical [FW42]. Cylindrically [Fer51h].

D [Ber01, Ish19, Lan18, Sei90]. dagegen [CSW97a]. dagli [Ama02b, Bat03b]. D’Agostino [Min02, FM03]. dai [Car01, Seg83]. dal [Bat03b]. Dalla [Gan02]. Dan [Sei90, Gol99a]. dans [Dem15]. Dao [Ano64]. d’après [Geo32]. Dardo [Bar05]. Dark [Rho95]. darkness [Mag09]. daß [Fer23b]. data [D’A03]. Database [KHB12]. Dates [Ano97]. dattiloscritto [Fer25a]. Daunting [Ćir16]. David [Mot81, Sol68, Gar18]. dawn [Cah95, Rif06]. Day [Lam65, Lam85, Lau51, Ama02b, Bas01, HHW99, Sei90]. Days [And73b, Coc62, Hol03]. Dead [Fri50, SAĆ16]. Dean [Sei90, Sei90]. Death [Ano54c, Pon55]. Debate [Cro99, Tar03, Cro86]. deboli [Cab02]. debut [CS99, DS07]. debutto [CS99, DS07]. Debye [Fer22g]. Dec. [Ano54b, Lau46b]. Decade [Ano52]. Decades [Sei90]. Decay [FTW47, FTW65, GR09, Jen00, Kon55, Rob50, Rob83, Wil68, Yan13b, BRV84, FBE^{+}34]. December [Fer39b, Fer70a, MP77, Ano98, Ano11, Lau46a, Wat82]. Decision [GB89, Gol92, Lef95, Mor57, Bak76, Sti47, Sti76]. decisions [Str63b, Str63c]. Decomposition [Ano34c]. découvert [Six88]. découverte [Mon66]. découvertes [Mon66, Seg84]. dedicata [Bas01]. dedicated [Bas01, Cif08, FPLR59, SA79]. Deduction [Fer28g]. deduzione [Fer28f, Fer28g]. deep [Ree00]. defeat [Ols63]. Defection [Cal13, Lau13a, Tur12]. Defects [Ano09a]. Defended [WB64]. Defense [FSA41]. Defining [Mla98]. Degeneracy [DJ99]. Degenerate [DuM28, MRR^{+}02]. degli [Dra76, Fer23m, Fer24b, Fer28f, Fer30i, Fer30j, Fer31d, FA34]. Del [Tur08, Ama02a, Bas01, Fer23g, Fer23l, Fer25a, Fer26f, FR27b, Fer28f, Fer28h, Fer29d, Fer30d, Fer30e, FR33b, FR33c, Fer35a, FR38, GER^{+}92, Lee01, Maj37, MS01, Sal01, Wei77]. Delano [Ein39]. Delay [Ano54a]. Delayed [Fer42r, FT43]. delights [Ano12b]. delivered [Fer39b, Fer70a]. dell [Maj37]. della [Bas01, Bel87, Bel97, Ber02a, Car01, CL03, CM02, DE06a, DE15, De 01, Dra76, Fer22a, Fer22g, Fer22e, Fer23i, FP23, Fer23j, FR25b, FR25c, Fer25a, Fer25e, FR27b, Fer28g, Fer29e, Fer29g, Fer30b, Fer30k, Fer33a, Fer34e, Fer35d, Fer35e, Fer38f, Gan02, Goo91, Kop23, Mai91, Min02, Mir34, PS07, Seb97, Wei77]. dell’agitazione [Fer22g]. dell’ammoniaca [Fer32c, Fer32e, Fer32f]. dell’assorbimento [Fer29d]. dell’atomo [Fer26b, Fer27e, Fer28f, Fer28g, Fer29a]. dell’effetto [Fer23n]. dell’elettricità [Fer22a]. dell’emissione [Fer27d, Fer33b]. dell’energia [Sal02a, For01]. dell’entropia [Fer23j]. dell’equazione [Som32b]. dell’influenza [Fer22g, Fer22e]. dell’Intervento [For01]. dell’irradiazione [Fer29d]. dello [Bas02, FR38]. dell’omonimo [AS96]. dell’Opera [Bon02]. dell’urto [Fer25f]. Demonstration [War61, Fer23d, Fer23e]. Density [Fer42e, Rij66, SHL^{+}18, SS11]. Department [FSA41]. Dependence [Fer42r, FM47f, FM49, Dug61]. Deposits [Ano34d]. Depression [Sol68, Wei70]. Derivation [Bro61, Sch89]. derive [Cra88]. Derived [Ano53a]. descobertas [Seg87a]. desert [Cah95]. deserved [Fer54c]. Design [ACF^{+}42, Fer42d, Wal74b]. Designer [Fri55b]. Designs [Fer42l]. Desperately [BB15a, BB15b]. Despite [Gra16a]. dessus [Mon66]. Detectability [SW66]. Detection [CRH^{+}56, RC53]. Detects [Cow09].

Determination [Fal73, Fer27e, Fer42e, Fer42d, Fer42f, ML48, CW73, Fer28b]. determinazione [Fer27e]. deterrence [GA74a, GA74b]. detto [FB69]. Deuterium [AFNY52b, FS50]. deuteron [Stu86]. Deuterons [FM49, SMKW46, SWK46]. Deutsches [Sim10]. Developing [Win86]. Development [Fer46d, Fer70b, Fra86, Meh73, Cum02, FR49, Meh75, Smy45b]. Developments [Ano97, CW32a]. Device [Ano50b]. devices [MZF^{+}12]. Devising [AG14]. Diamond [Sol68]. Diary [Sei90]. Dictionary [GH90, Hol90d]. did [CSW97b, FB69, Ols63]. Didatta [Bas01]. didattica [Seb97]. Didn’t [Wei94]. Dies [Fol86]. Different [Fer04b]. Differential [Som32a, Mir34, Som32b]. Differentialgleichung [Som32a]. differenziale [Mir34, Som32b]. Difficulties [Gou32]. diffraction [Fer22g, Fer22e]. diffrazione [Fer22g, Fer22e]. Diffusion [AF36a, AF36b, Fer24d, FW42, Fer42v, AF36e, Fer24c]. diffusione [AF36b, AF36e, Fer24d]. Dimostrazione [Fer23d, Fer23e]. Dinamica [Fer21a]. dioxide [Fer31e]. Dip [Ano06]. Dirac [BY90, Ama87, BDM81, Bel82, Bel94, CT67, Cra88, Din58, DuM28, Faz15, Fer49a, Fuk14, GG51, Kap40, KS42, KaKKZL05, MS38, Rho50, Rij66, SS11]. Diraka [Bel82]. Directions [MKR87]. Director [Sei90]. disappearance [GHH02, Mag09]. disappearing [Kea10]. discharge [Som28a]. Discover [Ser12]. discovered [Six88]. discoverer [Bar00]. Discoverers [Arb00, Arb03]. Discoveries [Kae35, Bra09, Seg76, Seg80, Seg84, Seg87a, Seg07]. Discovery [AGR04, Ama03, Ano39a, Ano82, Arb00, BHT86, FW67, Gol62, Gra64, GLR06, GR09, GLR20, Hah58, Mei62, Mon66, Rog13, Sim00, Ama84, Ano01a, CHW91, Car01, De 05a, De 06a, Dra76, FW85, Gou71, GA71, Kae48, Pen13, PZHC09, Rob12, Sim89, WC84]. discrepancy [Fer22a, Fer22f]. discrepanza [Fer22a]. Discussion [FBE^{+}34, Fer44b, FWN^{+}45, LCM^{+}42, Pau34]. discussions [CCJ^{+}34]. Disembarked [Ano82]. Disintegration [Boh39, MF39a, CW32b]. Disintegrations [Fer40b]. diskette [Win86]. Dispersion [FR27b, FR27a]. dispersione [FR27b]. Display [Sim10]. Dispute [Kra18]. Disputes [Ken54]. dissemination [PS07]. Dissent [Pri95]. Dissociation [AF44b]. Dissolving [SDO18]. Distinctions [Ano68a, Phi93]. Distribution [AFNY52a, AFMN53, Fer35c, Fer42u, Fer51a, Fuk14, Sch89]. distributions [Cra88]. District [Fer54l, Fer54m, HTS83]. d’istruzione [Ano01b]. Divergent [Fer52a]. Divided [Hay16, Seg15, Clo15]. Divine [GLR20]. Division [F^{+}43b, F^{+}43c, F^{+}43d, F^{+}43e, F^{+}43f, F^{+}44c, F^{+}44d, F^{+}44e, F^{+}44f, F^{+}44g, F^{+}44a, F^{+}44b, Fri39, F^{+}42f, F^{+}42e, F^{+}42g, F^{+}43a, F^{+}43g, F^{+}43h, F^{+}43i]. divulgatore [Sci01]. divulgazione [PS07]. Do [Tip80, Fer23h, Tip81b, Tip81c]. documentary [CHW91, GA71, WC84]. Documenti [Ver02, Ver01b]. Documents [Agn03, EP08, Ver02, CSW97b, Ore03a, Ver01b]. Does [Kap40]. Domestic [Fer54l, Fer54m]. Domus [Ama59, Der64, LRS00]. Donald [Zuc70]. Donegani [Fer49d, Fer50a]. Don’t [Rob08, Sch81, Gra16a]. doppietti [Fer30m]. dott [Fer35a]. doublet [Fer30o]. doublets [Fer30m]. Doubtful [Kae39]. Douglas [Oak18]. d’ouvrages [Mon66]. Dowd [Oak18]. Down [Fer42v, Fer42w, FMM43]. Downwind [Sei90]. Dozen [Ano34a]. Dr. [Ano62, Fer35a, Ken54]. Draft [Fer34n]. Drake [Oak18, VD15]. Drama [Ken54, Sol68]. Drop [Lef95]. druzej [PP72]. Dublettkomponenten [Fer30o]. Due [AF44b, Fer42x, CS00c, Mai91]. durch [FR27a, HS39a]. during [De 03, Fer51d, Pin03]. Dutch [Fer24a]. duties [SA79]. Dynamics [BPP18, Fer21a, Mar59, Wei97, Str98]. Dyson [Ano94b, SW66].

E. [Bel82, Dra72b]. Each [Ano39d]. Earlier [Blu01]. Early [Agn82, And73b, Ber88, Boc15, Bun83, Coc62, Hol03, McM51, Pes03, CS99, De 05c, Esp08b, Gal02, Kra92, Roq92, Stu93]. earth [HS39b, FR33b, FR33c, Gri11, Har75]. Economy [Bat03a]. ed [Bat06, Fer22h, Fer23h, GER^{+}92, Sal02b, Stu06]. edited [CW71, Fer35a, Jam11, Mal05, Mot81]. Edition [Ern92, Ano95c, Gri96]. Editor [Bar05, Hub13, Moo92a, BR57, Fer57c]. Editorial [Rig84, Rom91]. Edoardo [Bat03b]. eds [Zuc70]. Education [Ano01b, EL07, Seb97]. Edward [Sei90]. Effect [AFW^{+}47, CFW42, DRWZ34, FR25a, FR25b, FR25c, Fer32b, Fer42g, Fer42h, Fer42i, Fer42r, Fer42y, FK08, AFW^{+}57, Fer23n, Fer31e, FR31, Som27, Fer42y]. effective [F^{+}42g, F^{+}43h]. Effects [AMS19, Fer27c]. Effektes [Som27]. effetti [Fer27c]. Effetto [FR25b, FR25c, FPR34b]. Effort [Gon96a, Lew21]. Eigenschaften [Fer28b]. einatomigen [Fer26g]. Einfluß [FR25d]. einiger [Fer28b]. Einstein [Cra88, Din58, Ein39, Faz15, GM85a, GM85b, GRJ03, Her79, Kap40, Lau13b, San81]. Einsteinian [Kop23]. Einsteiniana [Kop23]. Elastic [Fer23k, Fer54g]. elastici [Fer23k]. electric [Fer21a, Som28b]. Electrical [Dug61]. electrically [Fer24g, Fer25f, Fer01]. electrically-charged [Fer24g, Fer25f]. electricity [Fer22a]. electro [Som28b, Fer27c]. Electro- [Fer27c]. electro-magnetic [Som28b]. electrodynamic [Fer22f]. Electrodynamics [Fer25a, Fer29f, Fer30g, Jac62, Sch02, Sch03a, Sch58, Sch03b, Cin02, Fer23c, Fer31b, Ferxx, Roq92, Fer25a]. electrographite [BJ41]. Electromagnetic [Fer21b, Fer22a, Fer22f, Fer23c, Fer31b, Fer38b]. Electron [Ano39d, BGJR24, LT05, LM83, RF26, RNKS04, WWL^{+}09, And33, Gou71, Hoc83, Maj37, Som28a, Som28b, Fer29a, Som27]. Electrons [BM35, DuM28, Fal73, FU33a, FM47c, Zim62a, Zim62b, Zim62c, Zim62d, Zim62e, BF32, FU33b, Pau25, Mot81]. Electrostatics [Fer21b]. elektrisch [Fer24g]. elektrische [Som28b]. elektrodynamischen [Fer22f]. Elektrographit [BJ41]. elektromagnetischen [Fer22f]. Elektronen [BF32]. Elektronengruppen [Pau25]. Elektronentheorie [Som28a, Som28b, Som27]. Element [Ano34d, Ano34f, Fer34c, GA34, ML48, MA40, Nod34, SMKW46, SWK46, Ano34c, Ano34e]. elementari [Fer52f].

Elementarnaya [Fer47a]. Elementary [BH83a, Fer46e, Fer47b, FY49, Fer51b, Fer51c, Lar19, Nam85, BH82, Fer47a, Fer52f, Hub13]. elementary-particle [BH82]. Elemente [Fer28b]. elementi [Fer24b, Fer28f, FRD34, FA34]. Elements [Ano34a, Ano39a, Fer28a, Fer34g, FRD34, Fer40c, Fer40d, FMT49b, Fit99, Kra18, McM51, SW48, SP48, Sim00, CW32b, Fer24b, Fer28f, FA34, FMT49a, Fit13, Kea10, Pen79, Fer28b]. Eletrodinamica [Fer25a]. Elettrica [MF37]. Elettriche [Fer21a]. elettrici [Fer25f]. elettro [Fer27c]. Elettrodinamica [Fer25a, Cin02, Fer23c, Fer29f, Fer30g, Fer31b, Fer31c, Ferxx]. Elettromagnetiche [Fer21b, Fer22a, Fer23c, Fer31b, Fer38b]. elettrone [Maj37, RF26]. elettroni [FU33b]. elettronica [Fer29a]. Elevate [Fer34m]. elio [Fer29g, Fer30k]. Elisabeth [Bar05, Bar05]. elusive [KS99]. Emanation [GLR20]. Emergence [GRJ03, Gri11]. Emergent [Vol08]. Emigration [Rid84]. Émigrés [Stu84]. Emilio [Ama73, Ano96b, Bad71, Bro73c, Buc73a, Buc73b, Gol70, Kle71, Pal69, Wei71, Wil71, Seg93]. Emilo [All62a, Rab63]. Emission [CBH09, Fer42r, SZ39, Fer27d, Fer33b]. Emitted [AFRW42, FF42, FF48, Fer42p]. Emphasized [Kae35]. Empty [FP23]. Enciclopedia [Min02]. encyclopedia [Kur02, Bar05, DR01]. End [Lan12, Sei90, Gre05]. energetici [Fer38d]. Energia [Lom01, MF37]. énergie [Gol80, Mon66]. Energieinhalt [Flü39]. Energies [Fer53a, Fer54e, Fer52h]. energiyach [Fer52h]. Energy [AFM46, Ano39c, Ano39a, Ano39d, Buc83, Coc62, Fer40a, Fer41, Fer46a, Fer46f, Fer51a, Fer54g, Fis97, For01, Gow64b, GLR14, HA69, HDHA89, IAE97, Mai03, Mal03a, Mei62, Rho18a, Rub03, Seg85, Sei90, Sei03, Sei04, Smy45a, Stu93, Tur06a, Zin55, Bar87, Ber80, Cow09, Cra07, Esp08b, FP26, Fer38d, Fer50b, Fer54h, Fer55, Fer61d, Fer70c, Flü39, Fri54, Gol80, Gol82b, Gow64a, GA74a, GA74b, Hei55, Hei83, Lau46b, LW71, Lib85, Lom01, Sal02a, SW87, Smy45b, Fer57b, Man57, Mon66]. Engineer [Fer35a]. Engineering [EP08]. engineers [Jam11]. Enrico [All62a, Ano34c, Ano38a, Ano68c, Ano70, Ano95b, Ano95c, Bet55, Buc73b, Büh98a, CL03, CW71, Fri55b, Fri55a, Gar18, Hub13, Jam11, KGSO55, Lan18, Lan17, MR86, Pal69, Pie01, Rab63, Rab66, Sha67, Wes16, AGR04, All63, ASA55b, ASA55a, All57, Ama02a, And55, Ano47, Ano51, Ano56a, Ano58, Ano61, Ano62, Ano64, Ano68b, Ano73, Ano81, Ano82, Ano01a, Ano03, Ano07, Ano09d, Ano10, Ano12a, Ano15a, Bai55, Bal63, Ban03, Bas01, Bas02, Bat01, Ber01, BC03, BB04a, BB04b, BB02, Boc15, Bon02, BC55, Bro73c, Bru07, Bru11, Bru16, Büh98a, Cal75, Car01, Car05, CM02, Cat03, Cha73, Cif08, Com42, Coo99, Cor98, CS99, CDS01, DR01, De 05a, DE06a, De 06b, DE06b, DS07, DE07, DS09, DS10]. Enrico [DE15, Deb53, Der64, Dys04, ED83, EE70, EP08, EP10, FV02, FSA41, Fer46i, FH46, Fer49c, Fer54k, Fer65c, Fer82c, Fer87, Fis10b, Fis12b, Gam03, Gla02, Gol99b, Gol99a, Gri96, GLR06, GR09, GLR20, GS74, Hah62, Hol90c, Hol02, Hol07, JM02, Jam04a, Kon50, Kub09, Laf71, Mal03b, Mal03a, Mal10, Met54, Met55, Mil07, Min02, Mor63a, Nis18, Ore01a, Ore03a, Ore04, Peg39, Per55, Per02, PS07, Pie01, Pon55, PP72, Pon92, Pon93, Pon13, PZHC09, Pra04, Rab66, Ras55, Ras68, Ras02, Rho99, Ril70, Sal01, Sal02b, Sal03b, SS99, Sch17, Sch02, Sci01, Seg70a, Seg70b, Seg71, Seg87b, Seg96, Seg09, SH16, Sei03, Sel53, Ste01, Ste59, Stu06, Sum65, Tal17, Tel91, Tel01, Tel02, Tur06a, Ver01a, Ver01b, Ver02, Wig55]. Enrico [Wig96a, Wil71, Yan03, Yan04, Yan13a, dL63, dL64, dL65, Mal05, Ama73, Ano96b, Bad71, Buc73a, Dra72b, Gol70, Kle71, Wei71]. Enrico-Fermi [MR86]. Enriko [GS74, PP72]. enthalpies [SHL^{+}18]. entrance [LMR06]. Entropy [Fer23j]. entstehenden [HS39b]. Entstehung [HS39a]. epic [Rob12]. episode [Pen13]. episodes [Bra09, Pon92, Pon13]. episodi [Pon92, Pon13]. Epoch [Ano39d]. Equation [Bro61, GG51, Hil69, Oak18, Som32a, VD15, Esp02, GS89, Mir34, Som32b, ZZWK12]. Equations [BY90, FMT49a, FMT49b, Mar00, ABF14, Fer23g]. equazione [Mir34]. equazioni [Fer23g]. equilibrium [Fer24e]. equivalence [Bul81, Ber02c]. Equivalenza [Ber02c]. Era [Bia90, Dir77]. Erdalkalimetalle [HS39b]. Erforschung [Fis10a, Fis12a]. Ergodic [Gal03, Fer23b, Fer23d, Fer23e, Fer24f]. ergodico [Fer23d, Fer23e]. ergodisch [Fer23b]. Ergodischer [Fer24f]. erhielt [CSW97a]. Erinnerungen [Hah75a]. Erkenntnisse [Hah75b]. Erlebnisse [Hah75b]. Ermenc [Sei90]. Erratum [Ano83, Lie82, Zag13]. Error [Kow15]. erste [Lan01]. esistenza [Fer23g]. esperimenti [Car05]. Esposito [Jam11]. Essay [Bar05, Mot81]. Essays [APB96, ABP98, Ano94a, Ano02, FPLR59, Wig67, vB93]. Essen [KGSO55]. essenziale [Ano02]. est [Dem15]. Establishing [Com42]. establishment [Ols63]. Esther [Moo95]. Estimating [Oak18, VD15]. État [Fer32a]. Eternal [SAĆ16]. Eternity [AS13]. etica [GER^{+}92]. Ettore [Fer76, Mal10, ACM10, DE06a, DE06b, DE07, DE15, Esp08a, EAW15, GHH02, GR05, GR08, Mag09]. Eugene [HH04b, Wig96b]. Europe [Bra69, Cer69, DGS89, Dro20, Rei69, Smi68, Zuc70, Fer68b, Fer71c, Rod19]. European [AS11, Smi68, Zuc70]. Even [Wei94]. event [Wat82]. events [Fer50b]. Ever [Fra05]. Everybody [Jon85a, Jon85b, Tar03, Web15, Web02]. Everything [Gar18, Sch17, Lan18]. Evidence [FM44, Fri39, Hal28, HS39a, HS39a]. evidential [Fra04]. Evolution [Sol03]. Exact [SM67]. Examination [ACF^{+}42, Hal28]. Excellence [Bar05, Fri01]. Excerpt [Fer51d]. Excerpts [Sei90, Ore03a]. Exchange [FAL^{+}52, Tam34]. Exciting [Dir77]. Excursion [Gla02]. Exercise [Ste71]. Exiles [Fra05]. Exist [Sch81, Tip80, Gra16a, Tip81b, Tip81c]. Existence [Cha32b, HS89, Cha32a, Fer24f, HS39b]. Existenz [Fer24f]. Exoplanets [Bai17a]. expanding [Dea13]. Expansion [Rij66]. Expansions [Din58]. Experiences [Hah75b]. Experiment [AFF^{+}42b, Fer42s, F^{+}42b, Fer42q, Fra86, Fra90a, Wat92, CA36, DPR05, F^{+}42g, Lau46b, PZHC09]. Experimental [DuM28, Fer42d, F^{+}42a, F^{+}42c, F^{+}42d, F^{+}43b, F^{+}43c, F^{+}43d, F^{+}43e, F^{+}43f, F^{+}44c, F^{+}44d, F^{+}44e, F^{+}44f, F^{+}44g, F^{+}44a, F^{+}44b, Fer52a, Hei03, PZHC09, Rob83, Fer29c, F^{+}42f, F^{+}42e, F^{+}42g, F^{+}43h, Sal01, F^{+}43a, F^{+}43g]. experimentation [Dra76]. Experiments [And55, CW32a, CW32b, DRWZ34, F^{+}42d, FM47a, Car05, F^{+}42b, F^{+}42f, F^{+}42e, F^{+}43i]. explain [Kow15]. explanation [Har75]. Exploring [WH72]. Explosion [Ano39b]. Explosive [Wal04]. Exponential [AFF^{+}42b, Fer42c, F^{+}42a, Fer42q, F^{+}42d, Fer57a, Fer58b, F^{+}42b, F^{+}42c, F^{+}42f, F^{+}42e, F^{+}42g, Fer42j]. Exponents [BPP18]. extensions [GR86]. Extracting [Ano12c]. extraordinary [Mag09]. extraterrestres [Cha15]. Extraterrestrial [Bri83, Cro99, KB89, Oak18, Sag73, SD75, SN83, Sch81, Tar03, Tip80, Tip81a, Tip81b, Tip81c, VD15, Web15, AK17, Cro86, DS92, Nea14, Pap85, Sho09, Web02, Wes90, Wil01]. Extraterrestrials [HZ82, Hof90, Reg87, Cha15, Har75, ZH95].

F [Kai04, Oak18]. FA [KGSO55, Seg86b, KGSO55]. faces [Del17]. facing [Ano47]. facsimiles [Bey49]. facsimlie [Wri64, Wri65]. Fact [Pei86]. Factor [CFW42, Fer42b, Fer42g, Fer42h, F^{+}42a, Fer42s, F^{+}42b, MB00, F^{+}43h]. Factors [Din58, Ste77]. Fail [GLR14]. Failures [Bat03a]. Fall [Fer51d]. Family [Cat03, Fri55b, Wil04a, Ano95b, Fer54k, Fer82c, Fer87, Cam54]. Famous [Ano34d, HH04a, Wri64, Wri65]. Fano [Ano96a]. Fantastic [WD06]. Far [Ano39a, De 06a]. far-reaching [De 06a]. Farm [Ber96]. Fast [AFW42, AF42b, FB41, SZ39]. Father [Gar18, Pra04, Sch17, EE70, Lan18]. Favor [Hal28]. favorite [GS22]. Fe [Ano05]. Fear [Sei90]. Feasibility [Fer42k]. features [F^{+}42b]. February [MP77]. Feinberg [Mon66]. Feld [Ano93a]. Feldes [FR25d]. Feliu [GHMP87]. Fellowships [OBC^{+}49]. Fenomeni [Fer22c]. fenomeno [Fer29d]. Fermi [ASA55b, All57, All62a, Ama73, Ano58, Ano61, Ano62, Ano63a, Ano64, Ano68b, Ano68c, Ano70, Ano73, Ano81, Ano94a, Ano94b, Ano95b, Ano96b, Ano96a, Ano03, Ano05, Ano09d, Bad71, Bai55, Bal63, BY90, BHR19, Bet55, Blu01, BC55, Bro68, Bro73c, Buc73a, Buc73b, Büh98a, Cam54, CL03, Cre10, Dra72b, DuM28, Fal73, Fis10b, Fis12b, Fra95, Fra86, Fri55b, Fri55a, Gar18, Gla02, Gol70, Gol99a, GG51, Goo91, Gou59, Gra61, Gri96, Gro78, Ish19, Jam04a, Kle71, KGSO55, Laf71, Lan18, Lan17, Lie82, Mag61, Mal05, Mal10, Man57, MR86, Min02, Mon66, Moo95, Mor63a, Mot81, OVPL15, Pal69, PQQ94, Pie01, Rab63, Rab66, Rij66, Sei90, Sha67, Sol68, Ste93, Str98, Stu06, Tal17, Tar03, Tel02, Ver01a, Wei71, Wes16, Wig55]. Fermi [Wig96a, Wil71, Win86, Zuc70, AGR04, Agn03, Agn04, AG14, All63, ASA55a, Ama59, Ama87, Ama02a, AMS19, ABF14, And55, AA55, And74b, And74a, And99, Ano34b, Ano34c, Ano38a, Ano39a, Ano47, Ano50b, Ano51, Ano54d, Ano55a, Ano55b, Ano56a, Ano57, Ano59, Ano60, Ano63b, Ano65a, Ano66, Ano68a, Ano69, Ano74, Ano79, Ano82, Ano90, Ano93b, Ano94c, Ano01a, Ano01c, Ano06, Ano07, Ano09a, Ano09b, Ano09c, Ano10, Ano12a, Ano15a, Ano15b, Ano22, AYS95, AS13, BB10, BB15a, BB15b, BB15c, BB15d, Bai17a, Bai17b, Bai18, Ban03, BDM81, Bar87, Bas01, Bas02, Bat01, Bat03a, Bat03b, Bat06, Bax01, Bec95, Bel82, Bel87, Bel88, Bel94, Bel97, BPP18, Ber18, BI04, BI05, Ber01, Ber02a, BC03, Ber03, BB04a]. Fermi [BB04b, BRV84, Ber02c, Bet63, BB02, BGJR24, Bla57, BM35, Boc15, Bon02, BM20, Bra69, BP68, Bro73b, BR94, Bro61, Bru07, Bru11, Bru16, Büh98a, Bul81, Bun83, BP58, BP97, Bye02, Bye03, Cab03, Cal75, Cal01, Car01, Car05, CM02, Cat03, Cer69, CK08, Cha73, Cha15, CBH09, CA77, CIR04, Cif08, Cin02, Ćir09, Ćir16, Ćir18, CT67, Com42, Coo99, Cor98, CS99, CS00b, CS00a, CS00d, CDS01, Cow09, CW73, Cra88, Cro04, D’A05, DR01, DRWZ34, DPR05, Dau08, De 05a, DE06a, De 06b, DE06b, DS07, DE07, DS09, DS10, DS11, DE15, DJ99, De 03, Deb53, Del17, Der64, DF06, DE04, Din58, Döb22, Dug61, Dys04, ED83, EL07, EFGP99, EE70, Esp02, Esp05, EP08, Esp08b, EP10, FV02]. Fermi [Faz15, FU43b, Fer44d, Fer46i, FH46, Fer49c, Fer54k, Fer82c, Fer87, FMT49a, FMT49b, FG92, Fir05, Fol86, For19, Fra90b, Fra90c, Fre85, Fuk14, G.64, Gal02, Gal03, Gam03, Gar92, Gar04, GM85a, GM85b, Geo32, Gol99b, GR86, GR89, GS22, Gra15, Gra16b, Gra16a, GRJ03, Gri04, GA34, GR05, GLR06, GR08, GR09, GLR20, GS74, GP31, Hah62, Hal28, Hal69, HMB09, Har61, Hil69, Hoc83, Hol90a, Hol90c, Hol74, Hol02, Hol03, Hol07, HM07, Hub01, Hub13, JM02, JAH09, Jon85b, KS99, Kap40, Kat21, KHB12, Kon50, Kon55, KS42, Kot54, Kow15, KaKKZL05, Kub09, Lan12, Lan92, Lan01, Lau13b, Led04, LRS00, Lib85, LS77, Lie81, Lut66, Mac86, Mac87, Mad87, Mai91, Mal03b, Mal03a, Mal05, MP56, Mar75, Mar83, Mar00, Mar59, MS38, MS01]. Fermi [Met54, Met55, Met74, Mil07, Mir34, MRR^{+}02, Mor63b, Mos50, MŚ94, Nis18, Ore01a, Ore03a, Ore04, Par02, PG86, Peg39, Peo04, Per55, Per02, PS07, Pie01, Pin74, Pon55, PP72, Pon92, Pon93, Pon13, Pon72, Pon04, PZHC09, Pra04, Rab66, Ras55, Ras68, Ras02, Reg03, Rei69, Rho50, Rho99, Ric01, Ric02, Rig84, Ril70, RNKS04, Rob08, Rub03, Sal02a, Sal01, Sal02b, Sal03b, SAĆ16, SDO18, San81, SS99, Sch94, Sch89, Sch99a, SHL^{+}18, Sch17, Sch02, Sch03a, Sci01, SC03, Seg03, Seg55, Seg70a, Seg70b, Seg71, Seg87b, Seg88, Seg96, Seg09, SH16, Sei03, Sei04, Sei55, Sel53, Shi81, SS11, Sim03, SHM19, Smi68, Sol03, Som27, Som28a, Som28b, Som32a, Som32b, Som33, Spr91, Ste71, Ste01, SZ62, Ste59, SM67, Str63a]. Fermi [Sum65, Tam34, Tel91, Tel01, Tel04, Tre76, TSH^{+}19, Tur06a, Ula55, Ver01a, Ver01b, Ver02, Vol08, Wal64a, Wal74b, Wal74a, Wal64b, War61, Wat88, Wat04, Web02, Web15, Wei94, Wei63, WWL^{+}09, Wei97, WB64, WR63, Wes90, Whe93, Wig62, Wil68, Wil04a, Wil04b, Wil04c, Yan03, Yan04, Yan13a, Yan13b, Zag11, Zag13, ZZWK12, Zim62a, Zim62b, Zim62c, Zim62d, Zim62e, Zin55, dL63, dL64, dL65, dL74, vB93, vH01, Ano95b, Ano95c, Bad67, CW71, Jam11, Dra72a, Kai04, Mal05]. Fermi-Field [BR94]. Fermi-Point [Vol08]. FERMIAC [RC12]. Fermilab [BDH89]. Fermions [FK08, MZF^{+}12]. Fermis [Cor64, Fer80]. Fermi’schen [Som27, Som28a, Som28b, Som32a, Som33]. Ferretti [Shi81]. Festschrift [Mot81]. Feynman [Fer52d]. Fiction [Sol68]. Field [BR94, Fer21b, FR25a, FR25b, FR25c, Fer51h, Ste71, CF53b, CIR04, FR25d, FR33b, FR33c, Moo92b]. Fields [CF53a, Fer54b, Kae35, Fer26d, Mar75, Tho27]. Fifties [Pes03]. fiftieth [BRV84]. Fifty [BI05, BRV84, Kae48, Tar03, Rho18b, Web02]. fifty-seven [Rho18b]. figure [GS22]. File [McM94]. Film [CW71, FBL64, Pal69, AS96].

Filtering [AFM46]. Finding [SM09]. Fine [Bai17b]. fine-tuned [Bai17b]. finely [LBS17]. finely-tuned [LBS17]. Finite [Bro68, BP68]. First [AGR04, AT62a, AT62b, ATFF82, Ano46a, Ano52, Ano55b, Ano11, Fer46d, Fer70b, Fri55b, Ken54, ML48, Wat74, Cor98, Fer25a, Fis97, Lan01, Lau46b, SR92, Ste59, Wol80, Ber18, Stu18]. Fisica [AF38, Bas02, CL03, DE06a, DE15, Fer25a, Fer27a, Fer28c, Fer29b, Fer30c, Fer38a, FP38, Fer50a, Fer52b, Mal08, Ras02, Stu06, Ama02b, Ano01b, CM02, DS07, Dra72a, Dra76, Fer26b, Fer29e, JM02, PS07, Pon72, Pon04, Ras68, Seb97, Seg76, Seg83, Wei77]. fisiche [Fer29c]. fisico [CS99, Dra72b, Seg71, Seg87b, Seg09, Tur08, CL03]. fisicos [Seg87a]. fisikaim [Seg86a]. Fission [ABD^{+}39, AFG41, AF42b, AF44a, AF44b, AFN44, Ano41a, BHT86, BW39a, BW39b, FS41, FW43, FMM43, FHN44, FMM45, FW67, Gra64, Hah58, Hah62, HS89, Mei39, MF39c, NBDG40, Sim00, Sim03, Sim10, Sim12, Stu85, Whe62, Ama84, Bar00, CHW91, CSW97b, Fri74, FW85, GA71, HS39a, Ham02, Hol02, Lom01, MF39b, Pei86, Sim89, Whe09, WC84, Ano41b, CSW97a]. Fissionable [FA61]. fissione [Lom01]. Fissionist [Ano50a]. Five [Web15, LW71]. Flawed [Gra16b]. flee [JB83]. Fleming [Zuc70]. Flimsy [McM94]. Florence [Bec95]. flow [Som28a]. fluids [And99]. Focus [Wal04]. Focuses [Ken54]. Focusing [WWL^{+}09]. fondamenti [Fer29c, Fer30b, Kop23, Mai91]. Fondazione [Fer50a]. forbidden [Fer26d]. Force [Ano39d]. Forces [Bri65, BR94, BR96, Tam34, Pai86, Ree00]. Forest [Fer42s, F^{+}42b]. Forgan [Ish19]. Forgetting [Sim12]. Forging [Sei90]. Forgotten [GR08, GR05]. Formation [FM44, HS89, Fer23f, HS39a, SS99, SHL^{+}18]. Formazione [Fer23f, SS99]. Formed [Ano34c]. Former [Fir05, Ros04, Ano96a, Ano01b]. Formerly [Mon66]. formula [Fer26e]. formulation [CIR04, Kre71]. Forties [Hei03, Pes03]. fortunate [LBS17]. forty [Fis97]. forza [FP26]. fotoelettrico [Fer23n]. fotografica [Bat03b]. Found [Ano34d, Ano39a, Arb03]. Foundations [Bey49, GL23, Fer29c, Kop23]. Four [Ada72, Ber88, FB64, Cra92]. Four-Hundredth [FB64]. fra [FM03, Val06]. Fragments [HS89, HS39a]. francaise [Mon66]. France [Gol62]. Francesco [GB12]. Franck [Pri95]. Franco [Tur08, Goo01]. frange [Fer29i, Fer30n]. Frank [Hay16, Seg15]. Franklin [Ein39]. Frauen [Chi94]. freakishly [Bai17b]. fredda [Mal08]. freddo [Maf92]. Frédéric [Pin03]. Frederick [Sol68]. Free [AF36c, CRH^{+}56, Fer52c, RC53]. Freed [Ano39d]. Freeman [Bar05]. Frees [Ano39b]. French [Sol68, Cha15, Chw13, CCJ^{+}34, Dem15, Fer31a, Fer32a, Geo32, Gol80, Per55, Seg84, Six88, dL63]. frequency [KGSO55]. Fresh [GB12]. Friedman [Bar05]. Friedrich [ABES12]. friends [PP72, Pon93]. Fringes [Fer29i, Fer30n]. Frisch [CSW97a, Sol68, CSW97a]. Fritz [CSW97a, FH81, Hol90d, CSW97a]. front [Ano47]. Frontier [Cre10]. Frontiers [Nam85]. Fuchs [Sei90]. Full [Rho18b, Fer25a]. Function [Bla57, Rij66, SS11]. functional [SHL^{+}18, SS11]. Functions [Din58, GM85a, GM85b, MS38, Rho50, BDM81, Sch89]. Fundamental [Bha39, Fer51e, Bey49, SA79]. fundamentals [Mai91]. Funds [Bat03a].

Further [CW32a, FM47a, MF37, HS39a]. fusion [De 01, Rei72]. fusione [De 01]. Future [Fer46f, For01, Lan12, Pap85]. Futuro [For01].

G [Fer34p, Mon66]. G. [Fer34p]. gadgets [Moo95]. Galactic [Fer54b, Oli75]. Galaxy [Jon76, Cra00, Gri18, Kre71]. Galiaeana [Der64]. Galiaeana^{≫} [Ama59, Der64]. Galilaeana [LRS00]. Galilei [Büh98b, FB64]. Galileo [Büh98b, FB61, FBL64, FB64, FB69, FB03, Jam04b, Gra61]. galvano [Som28b]. galvano-magnetische [Som28b]. Gamba [Tur08]. Game [Win86]. Games [Tau63]. Gamma [Cow09, Din58, FR33a, Mor58]. Gamma-Ray [Cow09, Mor58]. Gamow [Wil71, AM23, MR86, Rei72]. Gas [DJ99, DuM28, KS42, LM83, MRR^{+}02, Fer23j, Fer26g, Fer26f, GRJ03, Str98]. Gases [Fer40a, And99, Fer26g]. Gate [Sei90]. Gauge [CIR04]. Gauge-invariant [CIR04]. Gaussian [GM85a, GM85b]. Geiger [LMR06]. geladenen [Fer24g]. gemacht [Flü39]. Gender [Hoo03, Ore03b]. General [Bad67, Sei90, EL07, Fer23b, Fer23d, Fer23e, F^{+}42b, Fuk14, Tre76, Reg03, Som28a]. generale [Fer23d, Fer23e]. Generalization [Fer23g]. Generalized [FMT49b, Mar59, FMT49a]. Generalizzazione [Fer23g]. generation [Gan02, RK95]. generations [Ada72, Kra99]. Generator [HS19, AFR37]. generatore [AFR37]. generazione [Gan02]. Genesis [Wei97, Fer55, Fer70c, Str98]. Geneva [Yor87]. genio [Bru07]. Genius [Bru16, LS92, Bat01, Bru07, Mag09, Tal17]. Geometrical [Lib85]. Georg [ABES12]. George [Bar05, Fri55b, Wil71, AM23, Ama56, Rei72]. German [KGSO55, Ano34d, Ano10, BF32, Bet00, BJ41, Büh98a, Büh98b, Chi94, CSW97a, Ern92, Fal28, FM03, Fer22f, Fer23b, Fer24f, Fer24g, Fer24h, FR25d, Fer26g, Fer26h, FR27a, Fer28i, Fer28e, Fer28b, Fer30o, Fer30p, Fer31e, FR31, FS33b, Fer34p, Fer34o, Fer38c, Fis10a, Fis12a, Flü39, GHH02, HS39a, HS39b, Hah75b, Hah75a, Lan01, Mal03b, Nod34, Ols63, Pau25, Sim12, Som27, Som28a, Som28b, Som32a, Som33]. Get [Wil69]. Gets [Ano55b, Ano58]. già [Ano01b]. Giacomini [Fer35a]. giant [Bat01]. Giants [MD67]. Gilbert [Sol68]. Gilberto [Gra61]. Ginestra [Bad67]. Gino [Lan17, Wes16]. giornata [Bas01]. giorni [Ama02b]. giovanili [Gal02]. girls [Kie13]. Giuseppe [Bru11, Tal17]. Given [Ano79, Ano95c, Cip94, Fer49e, Fer50d, Fer50e, Fer65c, Fer51d]. gives [Dug61]. glass [Dys10]. Glicksman [Gli04]. Glimpses [Gar04]. Go [Fra05]. godovschine [Pon55]. Gold [Hol74]. Golden [BM20, CBH09, DF06, RNKS04, Ste71, CIR04, Mar96, RK95, Seg03]. Goldman [Sei90]. Good [Met74]. Goodenough [Ano09b]. Gordon [Sei90]. got [Ols63]. Gottinga [CS00b]. Göttingen [CS00b]. Goudsmit [Gol89]. Government [Smy45b]. Grad [Ros04]. gradients [TSH^{+}19]. Grand [Wal74b]. grandes [Mon66]. Grant [Sei90]. Graphite [AFF^{+}42b, AFF^{+}42a, AFM46, FW41, Fer42a, Fer42d, Fer42i, Fer42n, Fer42o, FMM43, FHN44, PHK25, FSA41]. Gratitude [Fir05, Fer54c]. grave [Fer22a]. Gravitation [Tre76]. Gravitational [Fer21b, CF53b, Sch99b]. Gravitazionale [Fer21b]. Great [Ano49, Ćir18, Kra78, Sol68, Bak76, Büh98b, Kae48, Lau46b, Ols63, Ano50a, Car15, Wei70, Bri83]. Greatest [Ano39a]. Ground [Ken54]. Group [Bat03a, Far01, Hol74, Hol03, MS01, Fer44d]. Groups [Fer28a, Fer36c, Pau25]. Groves [Sei90]. Grund [Som28a, Som28b]. gruppi [Fer36c]. gruppo [MS01]. guerra [Mal08]. guerre [Mai91]. Guglielmo [Fer38b]. Guide [Ano09d, Dro20, Rod19, Sei90, Zim62a, Zim62b, Zim62c, Zim62d, Zim62e]. Guiding [Sal03b]. Guíxols [GHMP87]. Guy [Sol68].

H [Ano70, Bar05, FU43b, Hol90d, Ish19, Ano54a, FR76, Gon96b, Ken54, Wol09]. H-Bomb [Ano54a, Gon96b, Ken54, Wol09, FR76]. Habitable [Dol64, Dol70, Dol07]. Hafemeister [Sei90]. Hahn [CSW97a, Ern92, Ano39b, CSW97a, Ern92, FPLR59, Her79, Sim03, Sim12]. Hailed [Ano39a, Ano39d]. Half [Clo15, ML48, Seg15, Hay16]. Half-life [Clo15, ML48, Seg15, Hay16]. Hall [Ber96, CK08]. Hane [Sei90]. Hans [Ano61, Ano94a, Ste93, Ber80]. Hard [Sol68, Sol68]. Hard-Hit [Sol68]. Hard-Hitting [Sol68]. hardback [Cal13, Ish19]. Hargittai [Bar05]. Harmful [Gra16b]. Harold [Ano12b]. Harper [Wil71]. Harrison [Mag61]. Hart [Kow15, Wes90]. Hartree [Mar75]. harvest [Bra09]. Harvey [Sol68]. Hawking [BB10]. hbk [Lau13a]. He-cooled [ACF^{+}42]. head [Ano47]. head-and-shoulders [Ano47]. Hearings [Ano49]. heart [Mil05]. Heavy [Boh39, Fer40c, Fer40d, Fer42w, FU43b, FU43a, Fri39, HM07, Ano34e]. Heberling [Hol90d]. Hecker [Ano09b]. Heidelberg [KGSO55, Stu06]. Heidelberg/Der [KGSO55]. Heisenberg [MP77, Sch00, Sei90, Mac86, Mac87, Pau34]. Held [Bet55, Meh73, CCJ^{+}34, GHMP87, MR86, SA79]. Helium [Ano34b, Fer44c, Fer29g, Fer30k, LCM^{+}42]. helped [Kie13]. Henry [Lan17, Sei90, Wes16]. Her [Kra78, CSW97b]. Herewith [Ols63]. heritage [Esp08a, Kra15]. Hidden [JAH09, Döb22]. High [CK08, Cow09, Fer29b, Fer50b, Fer51a, Fer54g, HS19, Mai03, Mal03a, Rij66, Sei03, CIR04, CW32a, CW32b, Fer34m, Fer52b, Fer52h, Fer54h, Hei55, Hei83, KGSO55]. High-Energy [Mal03a, Sei03, Cow09]. high-field [CIR04]. high-frequency [KGSO55]. High-Temperature [CK08]. Higher [Ano34d, Ano01b, Fer34g, FRD34, Som33]. Highlights [Cif08, RK95]. Hintertreppe [Fis10a, Fis12a]. Hiroshima [Lau51, Yor87, Sei90]. histoire [Gol80, Mon66]. Historical [ABP98, Ano97, Cra02, GL23, Pie01, Rob83, Seg85, APB96, Bar05]. Histories [Car15]. History [Ano11, Bel87, Buc83, Fis97, GHMP87, Kev77, McM51, Sei90, Sim10, Swi45, Tip81a, WP85, Wei77, Ano98, Bat03b, CHW91, DPR05, Fra04, Gol80, Gol82b, GS22, GA71, HTS83, HA69, HDHA89, Hoc83, Kea10, Kev87, Kev95, Kra99, KHFA67, L’A07, Rho18a, WH72, WC84, Bon02, Seb97, Sei90]. Hit [Sol68]. Hitler [Ber96, JB83]. Hitting [Kow53, Sol68].

Hochfrequenzkinematographie [KGSO55]. Hoerlin [Lan17, Wes16]. höheren [Som33]. Holds [Ano55b]. Holes [Fal73]. Holt [Lan17, Wes16]. Homage [Ula55]. Home [Wal74a]. Honor [Bet55, Kra15]. Honored [Wal64a]. Honors [Wig62]. honour [Mot81]. honours [Ano96a]. Hotel [Ano05]. hours [AS13]. Houtermans [ABES12]. Hove [Whe93]. Howard [Sei90]. HTC [Whe93]. human [Rho18a]. Hundred [Bar05, FB64]. Hundredth [FB64]. Hungarian [Jon10, Mar96]. hunter [Sho09]. Hutchins [Mad87]. hybrid [MZF^{+}12]. Hydrodynamics [Tau63, Str98]. Hydrogen [AFNY52a, AF52, AFL^{+}52, AFLN52, AFMN53, Fer25d, Fer42e, FU43b, FU43a, FAL^{+}52, FMN53, FGMN53, FMA54, Ken54, LFA^{+}52, NAF^{+}52, Rho95]. hydrogenated [FAP^{+}34]. Hydrogenous [FPR34a, Fer36d]. Hyperfeinstruktur [FS33b]. Hyperfine [Gou32, FS33a, FS33b, Fer33c]. Hypothesis [Bal73, BHR19, SAĆ16, Fer26a, Fer27b, Fer49b, Bax01].

IAEA [Ano97]. Ice [AYS95]. Ich [Büh98a]. Ida [Hoo03, Ore03b]. idea [Ano17, Bro78, Cro86]. Ideal [Fer26f, Fer26g]. idealen [Fer26g]. Ideas [GHMP87, Lan92]. identical [Fer24b, Mil07]. identici [Fer24b]. idrogenate [FAP^{+}34, FPR34a, FPR34b, Fer36d]. idrogeno [Fer25d]. If [Cor64, Web02, Web15, Tar03]. ignited [Gre05]. Igniting [Fit99, Fit13]. Ignored [Hop90]. ihr [CSW97a]. ihre [Fer28b]. II [Fer65b, Pal69, Sha67, Som28b, Wig96b, ADF^{+}35d, ADF^{+}35a, AF35b, Bad05, BB15a, BB15b, CW32b, Dir03, Fer22a, FPR34a, FPR34b, Fer34l, Fer34k, Fer54m, Goe49, Kie13, Mal03a, PHK25, Sim12, Som28b, Yan13c]. III [ADF^{+}34b, AF36d, Din58]. ill [Bat06, Min02]. Illinois [Bro73c]. Illus [Gra61, Bro68]. Illustrious [Fer68b, Fer71c, Sol68, Zuc70, Bra69, Cer69, Rei69, Smi68]. im [Chi94, Fer23b, Pau25, Som33]. Image [Fer23f]. images [Ver01b, Sei90]. imagination [Hol78, Hol98]. Imaging [CBH09]. Imbedded [FW41, FSA41]. Immagini [Ver02, Fer23f, Ver01b]. Immigrants [Bra69, Cer69, Fer68b, Rei69, Smi68, Zuc70, Fer71c, Sol68]. Impact [Ano05, LW71]. Implications [Sag83, Döb22]. Important [CDFS53a, CDFS53b, CDFS54, Fer52c, Fer23a]. importanti [Fer23a]. impressions [Agn82]. improvement [GR05]. improvements [Spr91]. Impurities [Fer42x]. Incident [Moo92b]. including [Ano65b]. Incompressible [Fer51g, FvN51, FvN55]. Inconsistency [Zag11, Zag13]. Inconvenient [Sim10]. Independence [GA74a, GA74b]. indeterminazione [DS11]. Index [FZ44, Bar05]. Indium [FW43]. indotta [Car01, Fer34j]. Induced [AGR04, De 05b, Fer34h, GLR06, GR09, GLR20, ADF^{+}34c, ADF^{+}34d, ADF^{+}34a, ADF^{+}35e, ADF^{+}35b, ADF^{+}35c, Car01, De 05a, De 06a, De 06b, FAP^{+}34, Fer34j]. inediti [Ver02]. Inertia [Fer22a]. Inerzia [Fer22a]. Infamous [Boc15]. inference [Fer28f]. Influence [GR09, FV02, Fer22g, Fer22e, FR25d, FPR34a]. Influential [Rog10]. Infra [Fer25g]. Infra-red [Fer25g]. Infrared [Dys60, SW66]. infrarosse [Fer25g]. ing [Fer35a]. Inhomogeneous [LM83]. initiating [FS50]. Initiation [All62b]. Injustice [CSW97b, CSW96]. Innocence [Stu18]. Innsbruck [KGSO55]. Innsbruck/2 [KGSO55]. Inquiry [Ano49, Dar98]. Inquisitor [Mal10]. Inquisitore [Mal10]. insights [Hah75b]. Inspirations [All63]. inspire [PZHC09]. Instability [Fer51d, Fer51g, FvN51, FvN55]. Instantaneous [SZ39]. institut [CCJ^{+}34]. Institute [Ano01b, CCJ^{+}34, WH72, Ano01b]. Institutes [Fer38a]. Institutions [Buc80]. instructive [Pon92, Pon13]. instrumentation [Car05]. instruments [F^{+}42f]. Insulator [FK08]. integerali [Fer23g]. Integrable [BPP18]. Integral [Rho50, WR63, Fuk14]. integrale [Fer25a]. Integrals [Bla57, CT67, Din58, Fer23g]. Integration [Som32a, Mir34, Som32b]. Integrazione [Som32b, Mir34]. Intellectual [Bra69, Cer69, Fer68b, Rei69, Smi68, Zuc70, Fer71c]. Intellectuals [Smi68]. Intelligence [Reg87, Sag73, SD75, SN83, Sch94, Tip81a, DS92, Pap85, Sho09, Wes90, Wil01]. Intelligent [Bau85, Bri83, Gri11, Lem98, SS77, AS13, Gri18, Tip81b, Tip81c]. intensi [Fer26d]. intensità [Fer25c, Fer26d, Fer30m]. Intensitätsverhältnis [Fer30o]. intensiteiten [Fer24a]. Intensities [Fer24a, Fer25c]. intensity [Fer26d, Fer30o, Fer30m]. interacting [SW87]. Interaction [FM47c, FG92, LT05, SZ39, BRV84, BF32, Fra90c, SHL^{+}18, Ste77, TSH^{+}19]. Interactions [BP97, Cab03, Fra86, BP58, Cab02]. interazioni [Cab02]. Interference [Fer29i, FM47b, Mos50, Fer30n]. interferenza [Fer29i, Fer30n]. Intergalactic [AS13]. International [Ano03, Ano05, DGS89, GHMP87, KGSO55, Meh73, MR86, CCJ^{+}34, Fis97, IAE97]. Internationale [KGSO55]. Internationales [KGSO55]. internationalism [Cra92]. Interpretation [Fer27c, CSW97b, Fer29d, Fer30d, Fer30e, Kop23]. interpretazione [Kop23]. interpretazioni [Fer27c]. Interscience [Bro68]. Interstellar [CM59, Sch94, Ste77]. Interventi [Bas01]. Intervention [For01]. intriguing [San81]. Introduced [Bha39]. Introduction [Ekl62, Lar19, Seg62, D’A03, Dra76, Hol98, L’A07, Moo95, Fer28c]. Introduzione [Ber02b, Fer28c, Dra76]. Intuitive [Lar19]. invalidation [San81]. invariance [FP26]. invariant [CIR04]. invarianti [LC02]. invariants [LC02]. Inventing [Gol92]. Invention [Ano55b]. Inventions [Ano55c, Bos50]. Inventor [Wil71]. inventory [KHFA67]. Invisible [Sol68, Cha15, Tur06b]. invisibles [Cha15]. Involved [SZ62]. Involving [GM85a, GM85b, AG14]. Inward [Pai86]. Iodine [FR38]. iodio [FR38]. ioni [Fer30i, Fer30j, Fer31d]. Ionic [Fer30i, Fer30j]. Ionisierungsspannungen [Som33]. Ionization [Fer40a, Fer24e, Som33, SHL^{+}18]. ionizzazione [Fer24e]. ions [CW32a, CW32b, Fer31d, GP31, HM07]. iperfini [FS33a]. ipotesi [Fer26a, Fer27b]. Iridium [Arb03]. Irpinian [FM03]. irpino [FM03]. Irradiated [HS89]. irradiation [HS39a, HS39b]. Isaac [Rig00]. ISBN [Bat06, Cal13, Ish19, Lau13a, Mal08, Tur08, Wes16]. Isidor [Rig00]. Iskusstvennaya [ADF^{+}35a, FAD^{+}34b]. Isolation [ML48]. Isotope [PAF^{+}42, Fer42t]. Isotopes [Ama03, Ano53a, Arb00, Arb03, HS89, NBDG40, dH62, HS39a]. issue [Cif08].

Issued [Ano55b]. Issues [Sei90]. Istituti [Fer38a]. Istituto [Ano01b, Min02, Ano01b, CL03]. istruttivi [Pon92, Pon13]. István [Bar05]. Italia [PS07]. Italian [AFRS34, ADF^{+}34b, ADF^{+}34c, ADF^{+}34d, ADF^{+}34a, ADF^{+}35e, ADF^{+}35b, ADF^{+}35c, AF35a, AF35b, AF36b, AF36c, AF36d, AF36e, AFR37, AF38, Ama61, Ama02a, Ama02b, AFM88, AS96, Ano34c, Ano53a, Ano82, Ano95a, Ano01a, Ano01b, Ano02, Bas01, Bas02, Bat03b, Bel87, Bel88, Bel97, Ber01, Ber02a, Ber02c, Bia90, Bon02, BCR09, Bru07, Bru11, Buc80, Cab02, Cal75, Car01, Car05, CM02, Cin02, Cod88, CS99, CS00c, CS00b, CS00a, CS00d, CDS01, Cum02, DR01, DE06a, DS07, DS11, DE15, De 01, Der64, Dra72a, Dra72b, Dra76, FV02, Fer21b, Fer21a, Fer22a, Fer22g, Fer22b, Fer22c, Fer22d, Fer22e, Fer22h, Fer23a, Fer23c, Fer23d, Fer23e, Fer23f, Fer23g, Fer23h, FP23, Fer23k, Fer23m, Fer23l, Fer23j, Fer23n, Fer24b, Fer24d, Fer24e, FR25b, FR25c, Fer25a, Fer25g]. Italian [Fer25c, Fer25e, Fer25b, Fer25f, Fer25d, Fer26a, Fer27a, Fer26b, FP26, Fer26e, Fer26d, Fer26f, Fer27b, Fer27c, FR27b, Fer27e, Fer27d, Fer28c, Fer28d, Fer28f, Fer28g, Fer28h, Fer29a, Fer29b, Fer29c, Fer29d, Fer29e, Fer29g, Fer29f, Fer29h, Fer29i, Fer30a, Fer30c, Fer30d, Fer30e, Fer30k, Fer30i, Fer30j, Fer30g, Fer30h, Fer30l, Fer30m, Fer30n, Fer31b, Fer31d, Fer32b, Fer32c, Fer32e, Fer32f, FR33b, FR33c, FR33a, FU33b, FS33a, Fer33b, Fer33a, FAP^{+}34, FPR34a, Fer34f, FA34, Fer34d, Fer34j, Fer34l, Fer34k, FRD34, Fer34m, Fer34n, Fer34e, Fer35a, Fer35d, Fer35e, FR35, Fer36c, Fer36b, Fer36d, FR38, Fer38a, FP38, Fer38b, Fer38d, Fer38f, Fer49c, Fer52b, Fer52f, Fer58a, FB69, Fer72, Fer82a, Fer82b, Fer96, Ferxx, For01, Gal02, Gan02, JM02, Kop23, LMR06, LC02, Lom93, Lom01, Maf92, Mai91, Maj28]. Italian [Maj37, Mal10, MFS^{+}6x, MS01, Mir34, Par02, Per02, PS07, Pie01, PFM^{+}22, Pon92, Pon93, Pon13, Pon72, Pon04, RF26, Ras02, Ric02, Sag04, Sal02a, Sal01, Sal02b, SS99, Sci01, Seb97, Seg71, Seg76, Seg79, Seg87b, Seg09, Som32b, Tel01, Val06, Ver01a, Ver01b, Ver02, dL74]. Italiana [Min02, Ras02, Stu06, Bel87, DR01, Dra76, Mai91, Ras68, Min02, Goo91]. italiani [CDS01, MFS^{+}6x]. italiano [Ano82]. Italy [All62a, DGS89, Fer62a, Meh73, MR86, Pal69, Rab63, Cal01, Cip94, Hol74, Orl98, PS07]. iv [Cal13, ADF^{+}34c, Far01, HH04a]. IX [ADF^{+}35b].

J [Mon66, Sei90, Sol68]. J. [Ano63a, Ano63b, Bet63, PC06]. Jack [Sol68]. Jahre [GHH02]. Jahren [Ern92]. James [Kai04, Mal05, Sol68, Wil71, Bat06]. January [Rig00]. Jefferson [Sei90]. Jewish [Orl98]. Jianwei [Fra95]. Joanna [Cre10]. John [Ano09b, Ano68b, Ano68c, PZHC09, Tau63, vN96]. Joliot [Six88, Pin03, Six88]. Jones [Sei90]. Jordan [San81, Sch99b]. Joseph [Sei90]. Journal [Rom91]. journals [Bey49]. Journey [FR13]. journeys [WD06]. July [Gla02, MR86, Rig00]. June [Fer33c, F^{+}43g]. Just [Ree00, Rob08, Sol68]. Justice [Sei90].

Kann [Flü39]. Kant [Cro86]. karne [Seg86a]. keepers [Van03]. kept [Pur63]. Kernspaltung [CSW97a]. Key [Ano97, Ken54, Hoc83]. kinetic [FP26]. Klaus [Sei90]. kleinsten [Fis10a, Fis12a]. Knew [Gar18, Sch17, Lan18]. Known [Arb03, Arb00]. Kohlendioxyds [Fer31e]. Kommentierung [Ern92]. Komplexstruktur [Pau25]. Kongreß [KGSO55]. konspekt [Fer68a]. kotlov [Fer47a]. Kramers [MP56]. Kristalle [Fer38c]. Kurian [Bar05]. Kurzman [Sei90]. Kurzzeitphotographie [KGSO55]. Kvantovaya [Fer68a, Dir03]. kvarkim [Seg86a].

L [FSA41, Sei90]. Lab [Met74, Pri95, Wal74b, Wal74a]. Label [Ćir16]. Laboratories [Pin03]. Laboratory [AGR04, Tel87, Cre10, Fer51d, Fer54c, Hug03]. Laborde [Mon66]. lady [Dau08]. Landed [Ano95a]. Landmarks [Wal04]. Lanfranco [Goo91]. Laplace [KGSO55]. Laplace-Transformation [KGSO55]. Large [Pes03, AG14]. Larry [Win86]. Last [Ćir09, Gar18, Lan12, Sch17, Sol68, Ber18, Lan18, Ric02, Sal01]. late [MF37]. later [Esp08a]. l’atome [Mon66]. L’atomo [Bru11]. Lattice [AFF^{+}42b, AFF^{+}42a, Fer42d, Fer42n, Fer42o, Str98]. Lattice-gas [Str98]. Laue [FPLR59, Her79]. Laura [Ano95b, Bra69, Cam54, Cer69, Fri55b, Gra61, Mon66, Rei69, Smi68, Sol68, Gro78, Man57, Zuc70]. laureate [Ano65b]. Laureates [Bar05, Dar04]. Laves [Hol90d]. lavori [CS99, CS00c, Gal02]. lavoro [Pon92, Pon13]. Law [BM20, Fer35c, Gri04]. Lawren [Sei90]. laws [Bai17a]. Leader [Gol70]. Leaders [Lau51, Ols63]. leading [AK17, SW87]. Leakage [Fer42y]. leap [Fis10a, Fis12a]. Learn [Ber02a]. Leason [Hol90d]. Lecture [Fer51d, Fer52g, Fer39b, Fer70a, Rut20]. Lectures [Fer34b, FF55, Fer83, FF08, Ano65b, Ric02, SR92, Fer27a, Ferxx, Cif08]. led [Bel82]. Lederman [Ano93b]. Lee [Ano64]. L’effetto [Fer32b]. Legacy [And74b, Ano05, BB04b, Lan17, Bat03b, BB04a, JM02, Mal05, San81, Mal05, Stu06]. Leida [CS00b]. Leiden [CS00b]. lektsii [Fer68a]. l’Enciclopedia [DR01]. Length [Bha39]. Lenti [FR35, AF35a, AF35b, AF36b, AF36d, AF36e, Dra76, Fer36c, Fer38d]. Leo [Fra05, LS92]. Leona [Fol86]. L’eredità [JM02, Bat03b]. Leslie [Sei90]. lesson [Fer25a]. Lessons [Fer25a]. Letter [BR57, Com42, Ein39, Fer52d, Fer54c, Fer57c, Pau30, Phi93, Wei94, Dem15]. Letters [Ern92, Kat75]. leurs [Seg84]. Level [War61, Ano01c, Jam11]. levels [Fer38d]. Leverett [ACF^{+}42]. lezione [Fer25a]. Lezioni [Fer25a, Fer27a, Fer49d, Fer54d, Ferxx, Ric02]. L’histoire [Mon66]. li [Seg86b]. Libby [Fol86]. Liberate [Ano39d]. Liberated [Ano39a]. libero [AF36c]. Licei [AF38, Fer29b, Fer52b]. Liceo [Ver01a]. Lie [SAĆ16]. Life [Agn03, Ano54d, Bau85, BS17, Bri83, Cro99, Fer04a, Fri55b, Gar18, Gri11, Hay16, Kra78, Lem98, Oak18, Sch17, Seg15, VD15, Web15, AK17, ABES12, Ano95b, AS13, Bai17a, Clo15, Cro86, Dra76, Dys10, Fer54k, Fer82c, Fer87, Gre05, Gri18, Lan18, LBS17, ML48, Mag09, PC06, Pon92, Pon13, Sal02b, SS77, Sim96, Web02, dL74, Seg15, Tar03, Hay16]. Light [CBH09, Fit99, GB89, Sal03b, Fer23l, Fer26a, Fer27b, Fit13, Lou83]. Lilienthal [CFU47]. limiting [Ste77]. Line [Fer22c, Fer66c, Fer34m, Wil71]. Linea [Fer22c]. lineari [FV02]. Lines [Fer24a, Fer25c, Fer26d]. L’influenza [FV02]. linger [Mad87]. L’interpretazione [Fer29d, Fer30d, Fer30e]. Liquid [Fer51g, HM07, JAH09]. Liquids [FvN51, FvN55, JAH09, Sch99a]. Lise [Büh98b, CSW97a, Ano67b, Ban04, Bar00, Büh98b, Cra69, CSW97a, CSW97b, Ern92, FPLR59, Ham02, Her79, Kra78, Kra88, Mei64a, Mei64b, Rif06, Sim89, Sim96]. L’Istituto [Ano01b]. Lists [Win86]. ’little [KS99, Lan17]. livelli [Fer38d]. Lives [Sei90, Wil71]. Lloyd [Sei90]. Logo [Win86]. Lomax [Sol68]. London [Buc73b, Cal13, Fri55b]. long [Maf92]. Longitudinal [FW42]. Look [GB12]. looks [Mei64a, Mei64b]. L’opera [PS07]. Lord [MF37]. loro [Dra76, Fer27c]. Loss [Fer40a, Fer42x]. Louise [Sol68]. love [CA77, Kea10]. Lovely [Wat92]. Low [AFM46, Dug61]. Löwdin [SHL^{+}18, TSH^{+}19]. Lowell [Cro86]. Ltd [Fri55b]. luce [FR25b, FR25c, Fer26a, Fer27b]. Luisa [Mal05, Stu06]. Lunar [Szi71, Tel87]. lungo [Maf92]. Lyapunov [BPP18].

M [Mag61, Mon66, Mot81, Sei90]. M. [Geo32, Pau34]. M8 [LFA^{+}52]. Machine [Sch94, Ano74]. MacPherson [Sei90]. Madden [Sol68]. Made [Kae35, Lew21, Flü39, MB00, Moo95, Sei90]. madness [Kea10]. Maestro [Bas01, Fer37b, Lee01, Sal01]. maggiore [FRD34]. Magnetic [CF53a, FR25a, FR25b, FR25c, Fer26c, Fer30f, Fer30h, Fer51h, Fer54b, Fer75, CF53b, FR25d, Fer26d, Fer30p, FR33b, FR33c, Som28b]. magnetici [Fer26d, Fer30h]. magnetico [FR25b, FR25c, FR33b, FR33c]. magnetische [Som28b]. magnetischen [FR25d, Fer30p]. Magnetooptical [Fer27c]. magnetoottici [Fer27c]. mai [FV02]. Majorana [Fer76, GHH02, Mal10, ACM10, DE06a, DE06b, DE07, DE15, DE04, Esp02, Esp05, Esp08a, EAW15, FK08, GHH02, GR05, GR08, Mag09, MZF^{+}12]. Making [Ano39d, Rho86, Rho95, Sei90, Wal74a, Wol09, Yor87, Gro67a, Gro67b]. Malcolm [Sei90]. Man [Ano38b, Ano99, Bor57, Bor63, Bro73a, Dol64, Dol70, Dol07, Gar18, Ore03a, Peg39, Sch17, Lan18, LS92, Ore01a, Ste77, Ste59, dL64, dL65]. Manhattan [Moo95, Sei90, Fer54l, Fer54m, FS95, Gro67a, Gro67b, Gro62, Gro83, HTS83, HHW99, Nie89, Sch15, Sza92, Sei90]. manoscritti [Der64]. manuscript [Fer25a]. manuscripts [Ama59, Der64]. Many [Dys10, Kae35, Del17]. Many-colored [Dys10]. Marc [Bar05]. March [Ano16]. Marconi [Fer38b]. Margherita [Ano01b]. Mario [Fer37b]. Marks [Sch00]. Marshall [Fol86, Fer44d]. Martians [Mar96]. Martin [Sei90, Sol68]. Mass [FP23, Stu93, Fer22f, Fer23c, Fer29h, Fer31b, FS50, Nie89]. Mass-Energy [Stu93]. massa [FP23, Fer29h]. Masse [Fer21b, Fer22a, Fer23c, Fer23i, Fer31b, Fer22f]. Masses [Fer21b, Fer22a]. Master [Ore04, Pal69, Sal01, Ste01]. Match [Lan17]. Matematica [Fer25a]. matematici [Fer28d]. Materia [Ama61, Fer33a, Fer34e]. Material [FA56, FA61]. Materials [All62b, Fer39a, Fer40a, Fer42u, FSS47, F^{+}42c, F^{+}42d, F^{+}42f, F^{+}42e]. Mathematical [Fer25a, EAW15, KGSO55, Fer28d]. Mathematicians [Ano15b, Rid84]. Mathematics [PZHC09]. Matter [Bad67, Fer36a, FT47, FTW47, FTW65, Ama66, Fer33a, Fer34e, Pai86, Spr91, Ama61]. matters [Rei72]. Matthew [Oak18]. Mauro [Bar05]. Max [Fis10a, Fis12a, KGSO55, Fis10a, Fis12a, FPLR59, Gre05, Her79]. Max-Planck-Medaille [KGSO55]. Maxwell [Faz15]. May [Ano05]. Maybe [Cra00]. McDougall [Fuk14]. McKay [Sei90]. Mean [AF36c]. Meaning [Ano82, Fer42s, Ano01a]. Means [FR27b]. Measure [FR27b]. Measurement [Fer42c, FMM47, FR27a]. Measurements [AF41b, BFMM44, Fer42e]. Measures [Ano34b, Lau13b]. Measuring [AFW^{+}47, AFW^{+}57]. Mécanique [Geo32]. Meccanica [Fer96, Fer23a, FP26, Fer27d, Fer28d, Fer30b, Fer30d, Fer30e, Fer30l, Fer36b, Gal02, Mai91]. meccanico [Fer23d, Fer23e]. meccanismo [Fer27d]. mechanical [Fer23b, Fer23d, Fer23e]. Mechanics [Fer26c, Fer30l, Fer54i, GL23, Ano95c, Dir26, Fer23a, FP26, Fer26h, Fer27d, Fer28d, Fer30d, Fer30e, Fer36b, Fer61a, Fer65c, Fer68a, Fer96, Gal02, Geo32, Mai91, Gri96]. mechanisches [Fer23b]. Mechanism [BW39b, FW67, Whe09, Fer27d]. Medaille [KGSO55]. Medal [KGSO55, Ano70]. Medicine [Ama03, Cra02, Bar05]. Medie [FP38]. medio [AF36c]. medium [Fer23l]. Meeting [Ano39a, Ano41a, Bet55, CCJ^{+}34, Fer33c, GHMP87, Dys04]. meets [Whe93]. Megavoltage [Rob95]. Meitner [CSW97a, Ern92, Kra88, Ano67b, Ban04, Bar00, Büh98b, Cra69, CSW97a, CSW97b, Ern92, FPLR59, Ham02, Her79, Kra78, Mei64a, Mei64b, Rif06, Sim89, Sim96, Sim03]. mekhanika [Fer68a]. Mem [Fri55a]. Memoirs [Sei90, Wil71, SSSS95]. Memorable [Rom91]. Memorandum [Fer42l, FU43b]. Memorial [Fer52g, Rei72, Bet55]. memoriam [Fre08]. Memorie [Mor63a, Fer62a, Fer65b]. Memories [McM94, Pon93, Sal01, Hah75a, Tel01]. Memory [AMS19, Her79, Per55]. Men [Str63b, Str63c, Van03]. mentor [AM23]. Mercury [FR25a]. Meson [And55, Bha39, Sel53]. Mesons [FY49, Hei55, Hei83]. Mesotron [Bha39, Kap40]. Mesotrons [Fer39a, FT47, FTW47, FTW65]. Messung [FR27a]. Métadier [Mon66]. Metal [Ano60, Fer42f, Fer42p, F^{+}42b, F^{+}42g]. Metalle [Som27, Som28a, Som28b]. metalli [Fer30m]. Metals [Hal69, CW73, Dug61, Fer30m, HS39b, Hoc83, Som27, Som28a, Som28b, Zim62a, Zim62b, Zim62c, Zim62d, Zim62e, Fal73]. Meteorology [Tau63]. Method [AFW^{+}47, AFW^{+}57, CW32a, Fer27e, Fer36a, FS57a, FL57, FA61, SM67, Cra88, Fer28b, Fuk14]. Methode [Fer28i, Fer28b]. Methods [Fer28i, Fer42m, Fer42l, Fer44c, Mir34]. metodi [Mir34]. metodo [CM02, Fer27e]. MeV [FMN53]. Mexico [Ano05]. mezzo [Fer23l, FR27b]. Mi [Seg86a]. Mi-karne [Seg86a]. Microcrystalline [FSS47]. Microns [Cor64]. MicroReviews [Hub13]. Migdal [Bro68]. might [SHM19]. Migration [Bra69, Cer69, Fer68b, Rei69, Smi68, Zuc70, Fer71c]. Milano [Mal08]. Milestones [Ano54b, Gon96a, Gon96b]. Military [Smy45a, Smy45b]. Milky [Gri18]. millennio [GER^{+}92]. Miller [Sei90]. Million [Ano34b]. Millions [MB00]. Millionths [Ano34b]. mind [Seg93, WD06]. ming [Seg86b]. Minimum [FM52]. Minority [FR76]. Minutes [Ano41a, Fer33c]. mio [Lee01]. miracle [Hol02]. Miramare [Meh73]. Mirrors [FZ46a, FZ46b]. miscellaneous [F^{+}42f, F^{+}42e]. Misleading [Gra16b]. Missing [Ano50a]. Mission [Fak83]. mistake [Gar92]. misura [FR27b]. mittels [HS39b]. Model [AMS19, Ano60, BP97, DE04, GR08, MŚ94, BP58, Esp05, GR05, GP31, SW87, Som33]. Modell [Som33]. models [Str98]. Moderation [Fer42i]. Modern [Gol99a, Kev77, Ama66, Bra09, Coo99, Dra72a, Kev87, Kev95, Pon72, Pon04, Seg80, Seg87a, Seg07, Fer30c, Seg84]. Moderna [Fer30c, Pon72, Pon04, Dra72a]. modernes [Seg84]. modernim [Seg86a]. modernos [Seg87a]. molecola [Fer29g, Fer30k]. Molecole [Fer34f, Fer82a, Fer25g, Fer32b, Lut66]. molecular [GRJ03]. molecule [Fer29g, Fer30k]. Molecules [Fer25g, Fer32b, Fer66a, LS77, LT05, Lie81, Fer34f, Fer38c, Fer82a, Lie82]. Moleküle [Fer38c]. Møller [Kra92, Roq92]. Moment [Fer26c, Dra76, Fer30p]. Momente [Fer30p]. momenti [Fer30h]. momento [Dra76]. Momentous [Lan17]. Moments [Fer30f, Fer30h, Fer75, Büh98b]. Monastero [GER^{+}92]. Monatomic [Fer23j, Fer26f, Fer26g]. mondiale [Bel87, Goo91]. mondo [Ano82, Cum02]. Money [G.64]. monoatomico [Fer23j, Fer26f]. monograph [Jam11]. Monographs [Bro68]. Monopole [Fer49a]. Monsters [Fal73]. Monte [Fer46g, FR48]. Monte-Carlo [FR48]. Month [Ano11, F^{+}43a, F^{+}43b, F^{+}43c, F^{+}43d, F^{+}43e, F^{+}43f, F^{+}44c, F^{+}44d, F^{+}44e, F^{+}44f, F^{+}44g, F^{+}43g]. moon [SHM19]. Moore [ACF^{+}42]. Moose [ED83]. Moral [Phi93]. Morrison [Fre08]. Morton [Sol68]. Most [Rog10, Ano51]. motion [Fer21a, Fer29h, Fer36d, Seg93]. Moto [Fer21a, Fer29h, Fer36d]. Motz [Sei90]. Movement [Sei90, Fer34m]. Mr. [Geo32]. Müller [LMR06]. Multiple [Car15, Fer53a, Fer54f, Fer54e, Fra05, Fer25c]. Muon [Led03]. muons [Fra90c]. Museo [Pie01, CL03]. Muses [JB83]. museum [Pie01, Sim10]. Mussolini [Fer61c, Fer68c]. My [Ano95b, Fer04a, Fri55b, Fer54k, Fer82c, Fer87, Ste01, Wil71, Wil71]. mysterious} [Dau08]. Mystery [FR13]. myth [Mar96]. Myths [PHK25].

N [Gar18, Lan18, Mon66]. N.Y [Bar05]. nach [GHH02, Som27]. Nachweis [HS39a, HS39b, HS39a]. Named [Ano63b, Bet63]. Names [Laf71]. naming [Stu86]. Nanoscale [WWL^{+}09]. nanowire [MZF^{+}12]. narrative [Com56]. Nascita [Sal02a, Bas01, Ber02a]. nation [Fer54c]. National [Ano41a, Cre10, GR63, Wal74b, Wal74a, GC99, Ano41b, FSA41, Fer49c]. Nationalism [Cra92]. Natural [FBE^{+}34]. Nature [Bad67, Dar98, Fer57c, SP48, Tel57, Ama66, Meh73]. Nauchnye [Dir03, Fer71a, GS74]. Nauchnykh [Dir03]. Navigator [Ano54c, Ano82, Ano95a, Sag04]. navigatore [Ano82]. Nazionale [Fer49c].

Nazism [Ols63]. NDCrc [FSA41]. NDCrc-121 [FSA41]. Need [Kae35, Büh98a]. Negative [AFNY52b, AFL^{+}52, FT47, FTW47, FAL^{+}52, FGMN53, FMA54, FTW65, LFA^{+}52, NAF^{+}52]. neglected [Roq92]. negli [Car05, PS07]. nei [Fer22e, Fer26d, Fer30m, Fer32b, Gal02]. Neighborhood [Fer22c, Fer66c]. Neighbors [Tar03]. Neither [Gra15, Six88]. neitronnoi [ADF^{+}35a, FAD^{+}34b]. nel [Ano82, Ber02a, Cum02, Pon92, Pon13]. nella [AF36c, Dra76, Fer23i, Fer26b, FP26, Fer27d, Fer29d, Fer30d, Fer30e, Fer30l, Fer31b, JM02, Mai91, Mal08, Pon92, Pon13, Seg76, Seg83]. nell’alta [Fer38b]. nelle [Fer32b, Fer36d]. nell’insegnamento [DE06a, DE15]. neptunium [Ano34e, Nod34]. Neumann [Kow15, Tau63, vN96]. neutral [KS99]. Neutrino [Ano56b, CRH^{+}56, GLR14, KS99, RC53, Sut92, Bro78, Rei96]. neutrinos [Bon05, Fra04]. Neutron [AGR04, AFS39, AF41b, AFF^{+}42b, AFF^{+}42a, AFW^{+}47, Ano34b, Ano38b, Cha32b, De 05c, EP10, Fer34a, FAD^{+}34a, Fer34h, Fer34k, Fer34i, Fer39b, FB41, Fer42e, Fer42p, Fer42n, Fer42o, FZ44, Fer46c, Fer46b, Fer46h, Fer46i, FH46, FM46a, FM46b, FMM47, FM49, Fer50c, Fer51f, FA61, Fer70a, Fri39, GLR06, GR09, GLR20, Kae39, Lau13b, Rob50, Rob83, Rog13, Seg55, Stu93, AFRS34, ADF^{+}34b, ADF^{+}34c, ADF^{+}34d, ADF^{+}34a, ADF^{+}35d, ADF^{+}35a, ADF^{+}35e, ADF^{+}35b, ADF^{+}35c, Ama84, AFW^{+}57, Bro73b, Cha32a, De 05a, De 06a, FAP^{+}34, FAD^{+}34b, Fer34j, Fer34l, Fer65a, GS22, HS39a, HS39b, MF39b, Six88, Jam11]. Neutron-Absorption [AFW^{+}47, AFW^{+}57]. Neutron-Induced [AGR04, GLR06, GR09, GLR20, De 05a, De 06a, FAP^{+}34]. Neutronen [BJ41, HS39b]. Neutronenbestrahlung [HS39a]. Neutroni [AFRS34, Fer34k, FR35, Fer38d, ADF^{+}34b, ADF^{+}34c, ADF^{+}34d, ADF^{+}34a, ADF^{+}35e, ADF^{+}35b, ADF^{+}35c, AF35a, AF35b, AF36b, AF36c, AF36d, AF36e, AFR37, Car01, Car05, Dra76, FAP^{+}34, FPR34a, FPR34b, Fer34j, Fer34l, Fer36c, Fer36d, FR38]. Neutronic [FS55, FZ55, FA56, FS57a, FL57, FS57b, FZ57, FS58a, FZA58, Fer60, FS61]. Neutrons [AF35a, AF35b, AF36a, AF36b, AF36c, AF36d, Ama03, AFH39, AF39, AF40, AF41a, AFW42, AF42a, AF42b, AFRW42, AFN44, AFM46, AF4x, Ano34c, BM35, BFMM44, De 05b, FF42, FF48, FPR34a, Fer35b, Fer35c, Fer36c, FR38, FAW38, Fer40c, Fer40d, FW41, FSA41, Fer42a, Fer42f, Fer42i, Fer42p, Fer42r, Fer42v, Fer42w, FT43, FW43, FMM43, FZ44, FHN44, FZ46a, FZ46b, FM47a, FM47c, FM47b, FM47d, FM47e, FM47f, FSS47, GLR06, HS89, MF39a, SZ39, Tam34, Tur06a, AF36e, AFR37, Bon05, BJ41, Car01, Car05, De 06b, Dra76, FR35, Fer36d, Fer38d, F^{+}42g]. never [FV02, Gra16a]. News [Ano66, Stu85]. Newton [Büh98b]. Newtons [Büh98b]. next [Gan02]. ni [Six88]. Nichols [Sei90]. nicht [CSW97a]. Nicole [Mon66]. nine [Van03]. Nineteen [FB64]. Nitrogen [Fer42g]. No [Fer34c, Fer42j, Fre85, Kow15]. Nobel [Bar05, MFS^{+}6x, Ano65b, Bar05, CHU87, Cra92, CSW96, CSW97a, CSW97b, Cra02, Dar04, Fer38e, Fer39b, Fer70a, Fri01, Gra16a, Gre05, Har02, Kur02, MFS^{+}6x, Pau45, Peg39, Wil69]. Nobelist [Ano50b]. Nobelpreis [CSW97a]. Noble [Hal69]. Noddack [Hoo03, Ore03b]. Nominators [Bar05, CHU87]. Nominees [Bar05, CHU87].

Non [Mai91, Sch99a, FV02, Fer23g, Fer23h]. Non-Fermi [Sch99a]. nonexistence [Fer23g]. Nonlinear [Ano05, FV02, FPU55, FPU74, PUF65]. n’ont [Six88]. Nor [Gra15, Six88]. normal [Fer23b, Fer23d, Fer23e, Fer23g]. normale [Fer23d, Fer23e]. normali [Fer23g]. Normalsystem [Fer23b]. nostri [Ama02b]. Notable [Kae35]. Note [FR48, Fer62a, Fer65b, Mor63a, Ano94a, Moo92a]. Notebook [AGR04]. Notebooks [AA55]. Notes [Fer61a, Fer65c, Fer66b, Ore58, Ore82, Ore01b, Bar05, Spr91, Bel82, Gri96, Ano95c]. Novel [Bun83, EP08]. Noyau [Pau34, Fer32a]. noyaux [CCJ^{+}34]. nozione [FP26]. Np [ML48]. nucléaire [Gol80]. Nuclear [Ama03, Ano07, BHT86, Bad05, BW39b, Bri65, BR94, BR96, Cal75, Cal13, De 03, EP08, Fer40b, FBR^{+}41, F^{+}42c, FU43b, F^{+}43b, F^{+}43c, F^{+}43d, F^{+}43e, F^{+}43f, FU43a, F^{+}44c, F^{+}44d, F^{+}44e, F^{+}44f, F^{+}44g, F^{+}44a, F^{+}44b, Fer45, Fer49e, Fer50d, Fer50e, Fer51a, Fer57c, Fer58b, Fer61b, Fer74, Fol86, For01, Gar18, Gra64, Hei03, Hug03, Jam11, Jen00, Lau13a, LF62, MF39a, Mei62, Mla98, MŚ94, NBDG40, Pal69, Pin03, Pra04, Ros72, Rut20, Sag83, Sch17, Seg85, Sei90, Sim00, Sim10, Sim12, Stu84, Stu18, Tel57, Tur06b, WS48, Wat93, Wat04, Wig62, Wig96b, vHEGM12, Ada72, Ama84, And73a, Ano11, Ban03, Bar00, BRV84, Bey49, Cah95, CHW91, Cra07, CSW97b, Cum02, De 01, Esp08b, EP10, Fer38d, F^{+}42d, F^{+}42f, F^{+}42e, Fer47a, Fer50b, Fer55, Fer70c, Fri74]. nuclear [Gol80, Gol82b, GA71, Gra16a, GS89, Ham02, Hol02, Kae48, Lan18, Mag09, PHK25, Rif06, Sal02a, Ste59, SF44, Tur12, WH72, WC84, Wil75, Wol80, F^{+}43a, F^{+}43g, Fer52h, Lom01, Stu79, Sei90]. Nuclear-Fission [Sim10, Sim12]. Nucleare [For01, De 01, Lom01, Mal08, Sal02a]. nucleari [Cal75, Cum02, Fer38d]. Nuclei [Boh39, Fer30f, Fer54g, Fer75, Fri39, Goe48, Goe49, Ama87, Ama02b, BP68, CCJ^{+}34, Fer25d, Fer30h, Fer54h, Flü39, Maj28, Bro68]. Nucleon [Fer53a, Fer53b, Fer54f, Fer54e, Fer54a]. Nucleon-Nucleon [Fer53a, Fer54e]. Nucleonics [FFH^{+}44]. Nucleons [Fer83, FF55, FF08, Cif08]. Nucleus [Ano34b, Fer52c, FR13, Kow53, Mei39, MF39c, Fer30p, Fer32a, Fer52e]. Number [AFRW42, Ano34d, Fer34g, FRD34, Fer42p, Kre71, KaKKZL05]. Numbers [Lew21, AG14, Ree00]. numerica [Mir34]. Numerical [FM52, DPR05, Mir34, PZHC09]. numero [FRD34]. nuova [FP26, Fer28d, Fer30b]. nuove [Fer29c]. Nuovi [AFRS34, Lom93]. nuovo [Ano82]. nur [CSW97a]. nutzbar [Flü39].

O [Jam11, Bel82]. O. [Ber01]. obbediente [Bru07]. Obedient [Bru16, Tal17, Bru07]. Obey [Kap40]. Obeying [Mos50]. Obituary [FH81, Ama56]. Observations [Blu01]. Obtaining [CW32a]. occasion [BRV84, FPLR59, SA79]. Occur [Fer22c, Fer66c]. October [Ano41a, Ano03, CCJ^{+}34, Far01, FR76]. octobre [CCJ^{+}34]. Odds [Hof90, Lau13b]. odyssey [Yor87]. offering [KS99]. Office [Sei90]. official [Smy45b]. offprints [Met54]. often [Gra16a]. Oh [Mil05]. Old [DS09, Gol89, Ano15b, Bai17a, DS10]. Oliphant [Ano34b]. once [Ols63]. once-great [Ols63]. onde [Fer38b]. ondulatoria [Fer27d, Fer30l, FP26]. One [Bar05, LT05, Bai17a, Bia90, KS99]. One-Electron [LT05]. Ongoing [Tar03]. only [Ano39d, CSW97a, Fer52g, Gri18, LFA^{+}52, Mai91, NAF^{+}52]. Onset [DJ99]. Open [DF06]. Opening [Sch00]. Opens [Cow09]. opera [Sal02b]. Operating [FS57a]. operations [Som28a]. Opinion [Gol92, FR49]. Oppenheimer [Ano63a, Ano63b, Bet63, Ken54, PC06, Val06]. Opportunity [Rid84]. Optical [Fer24c]. Optimum [Fer42u]. ora [Sal01]. Oracles [Sei90]. Oraria [Fer22c]. orbital [SHL^{+}18, TSH^{+}19]. orbite [FA34]. orbits [FA34]. Order [Rho50]. Orders [CT67]. Ordinary [FAL^{+}52]. Origin [BR96, Fer49f, Fer49g, Fer54b, Bel97, Fer49b, Fri74, Pen79]. Original [HM07, AS96, PHK25]. originale [AS96]. originally [Bey49]. Origins [Mar83, Pri95]. Orso [Fer37b]. Oscar [FM03]. oscillating [Bon05]. oscillation [Fer32c]. Oscillations [CK08, KS99]. Oscillator [Mar59]. oscillazione [Fer32c, Fer32e, Fer32f]. other [Kea10, Rei72, Wol80]. Otto [CSW97a, Ern92, CSW97a, Ern92, FPLR59, Her79, Sim12]. Our [Wat74, And74b, And74a, Wil75]. Outer [Bau85]. ouvrage [Mon66]. ove [Seg86a]. ove-ad [Seg86a]. ove-tagliyotehe [Seg86a]. Ownership [Ano55b]. Oxford [Bar05, Gol99a]. Oxford/New [Bar05]. Oxide [AFF^{+}42b, AFF^{+}42a].

P [Fal73, HH04b, Mon66, Str98, Wig67]. Pacific [Ken54]. Padé [EFGP99]. pages [Ano95b, Ano95c, Ano96b, Bro73c, Gol99a, Seg15, Stu06, Wil71, Wil71]. Paired [Sei04]. Paired-up [Sei04]. Pandora [Van03]. panel [Ano51]. Pang [Fra95]. Panisperna [AS96, Tur08, Ama03, AFM88, AS96, Ano01b, Bia90, BCR09, CL03, Car05, Cod88, FM03, Lom93, Mal10, Pie01]. Papa [Mal10]. paper [Pau34]. Papers [All62a, Dir03, Mor63a, Pal69, Rab63, Rab66, Sch58, Sch03b, Sei90, Sha67, vN96, DP97, Fer62a, Fer65b, Fer71a, Rom91, Wig96b, Wri64, Wri65, Yan13c, Bal63]. Paradigm [JAH09]. Paradox [For19, Gra15, Gra16a, Tar03, AS13, BB15a, BB15c, Bai17a, Bai17b, Bai18, Cha15, Döb22, Kow15, Web02, Wes90, BHR19, BB10, BB15b, BB15d, Bax01, Ber18, Ćir09, Ćir16, Ćir18, Fre85, Gra15, Gra16b, Gra16a, HMB09, Lan12, SAĆ16, SDO18, Sch94, SHM19, Web15, Ish19]. paradoxe [Cha15]. Paraffin [AF36c, Büh98a, Hol02]. Paraffin. [Büh98a]. paraffina [AF36c]. Paris [KGSO55]. Paris/FA [KGSO55]. park [Kra15]. Part [Fer42n, Fer42o, Fer46c, ADF^{+}35d, BB15a, BB15b, Som28a, Som28b, Wig96b]. parte [Fer23l]. Particelle [Fer52f, Fer33a, Fer34e, JM02]. participants [Pau30]. Participation [Man57, Fer57b]. Particle [BH83a, BH83b, Lar19, Mai03, Nam85, Ser12, Ano51, BH82, BDH89, JM02, Rei96, PQQ94, Fra95]. Particles [Ano34c, FS41, FY49, Mos50, Fer24g, Fer25f, Fer51b, Fer51c, Fer51e, Fer52f, Fer01, Fis10a, Fis12a, Mil07, Hub13]. Partisans [Sol68]. Past [F^{+}43a, F^{+}43b, F^{+}43c, F^{+}43d, F^{+}43e, F^{+}43f, F^{+}44c, F^{+}44d, F^{+}44e, F^{+}44f, F^{+}44g, F^{+}43g]. Pasta [AMS19, Ano05, Ano15b, BPP18, BI04, BI05, DPR05, OVPL15, Str98, Wei97, Dau08, PZHC09]. Patent [Ano50b, Ano53a, Ano53b, Ano55b, Tur06a]. patenting [Tur06b]. Patents [Ano55c]. Path [AF36c, Bel82, Fer80]. Pauli [Sch99b]. Pay [Tur06a]. Payoff [Ano53a]. Payot [Mon66]. Peace [Sei90, Yor87]. Peaceful [Fer57b, Man57]. Pedagogic [Spr91]. pedagogical [DPR05]. Peierls [Ano81, DP97]. penetrante [FR33b, FR33c]. penetrating [FR33b, FR33c]. People [Lib79, Wat82, Sol68]. Percentage [Fer42x]. percorso [CS00d]. Perfect [Fer23j]. perfetto [Fer23j, Fer26f]. perfezionamento [Fer22g]. performed [F^{+}43i]. Periodic [Fer28a, Fer28h, Fer28f, Fer28b, Kea10]. periodico [Fer28f, Fer28h]. periodischen [Fer28b]. Persico [DS11]. Personaggi [Seg76, Seg83]. Personal [Fer62b, Sei90, Com56, IAE97, Sal01]. personali [Sal01]. Personalities [Seg76]. Perspective [Fer04b, Seg85]. Perspectives [Mai03, DPR05, SA79]. persuader [Sci01]. persuasore [Sci01]. Perturbation [SM67]. Peso [Fer21b, Fer22a, Fer23k]. Peter [Bad67, Sei90]. Petite [Mon66]. Phase [FM46a, FM46b, FM47d, FMA54]. Phenomena [Fer22c, FM47b, Fer66c, Mos50]. phenomenological [EAW15]. Phenomenon [Ano39b, Fer29d]. Philip [Fre08]. Philosophy [Ćir18]. Phonon [RNKS04]. Photo [FF42]. photoelectric [Fer23n]. photographic [Bat03b]. Photographs [FS95, Moo95]. photography [KGSO55]. photoproduction [Fer54a]. Photosource [FF48]. Physical [Bad67, Bet55, DGS89, Fri39, GL23, Ama66, Dra76, Fer29c, Pai86, Wal04]. physiciens [Seg84]. Physicist [Buc73b, Dra72b, Fer04a, Gol99b, Hay16, Kai04, Rab63, Rho99, Seg70a, Seg15, Zuc70, ABES12, Clo15, CS99, Gra16a, Gre05, Meh73, Orl98, Seg70b, Seg71, Seg87b, Seg09, Wil70, Yor87, Wil71, Bro73c, Ama73, Bad71, Buc73a, Gol70, Kle71, Wei71, Ano96b]. Physicists [Ano39a, Bad05, Ber88, Buc80, Kev77, Ore58, Ore82, Ore01b, Rid84, Ser12, Stu84, Ada72, HH04a, Jam04b, Jam11, Jon10, Kev87, Kev95, Seg80, Seg84, Seg87a, Seg07, dR85]. Physics [ABP98, Ano20, Ano39b, Ano56a, Ano64, Ano03, Ano11, Bad67, Bar05, Bro68, BH83a, BH83b, Bye03, CSW97b, DE06b, DS09, De 03, EP08, Fal73, Fer25a, Fer26b, Fer29e, Fer30c, Fer38e, FSA41, F^{+}42c, F^{+}42d, F^{+}43a, F^{+}43b, F^{+}43g, F^{+}43c, F^{+}43d, F^{+}43e, F^{+}43f, F^{+}44c, F^{+}44d, F^{+}44e, F^{+}44f, F^{+}44g, F^{+}44a, F^{+}44b, Fer46c, Fer46b, Fer46h, Fer51f, Fer55, Fer70c, Fer85, Fra95, Gam85, Gol70, Hei03, Hol74, Hug03, Jam11, Kae35, Laf71, Lan92, Lan17, Lar19, Led03, Mai03, Mal03a, Meh73, MR86, MKR87, Mla98, Nam85, Nis18, Orl98, Pal69, Pau45, Peg39, Pei93, Pin03, Pra04, Seg55, Seg79, SH16, Sei03, Sel53, Stu18, Vol08, Wei70, Wes16, APB96, Ama02b, Ano65b, Bas02, BRV84, Bey49, Bra09, BH82, BDH89, Büh98b, Bul81, Chw13, Con62]. physics [Coo99, CW73, CHU87, Dar04, De 05c, DE06a, DS07, DE07, DS10, DE15, Dea13, EP10, EAW15, Fer27a, Fer28c, Fer32a, FBR^{+}41, F^{+}42f, F^{+}42e, Fer46i, FH46, Fer49e, Fer50d, Fer50e, Fer74, FPLR59, GHMP87, GS22, Her79, JM02, Kae48, KS99, Kra99, KHFA67, LMR06, Meh75, PQQ94, PS07, Pon72, Pon04, Ras02, RK95, SA79, Seb97, Seg76, Sim96, Sop76, Sop80, Sop88, Stu79, Tur06b, Tur12, WP85, WH72, Wei77, Wri64, Wri65, AF38, Ano01b, Bar05, Cal13, CCJ^{+}34, Fer29b, Fer38a, FP38, F^{+}42f, F^{+}42e, F^{+}42g, F^{+}43h, F^{+}43i, Fer52b, Gol99a, Esp05, Wig96b, Lau13a, Rom91, Sei90, Dra72a]. physics/0511222 [Esp05]. Physik [Büh98b]. physique [Chw13, CCJ^{+}34, Fer32a]. piano [Fer23l]. Pico [Win86]. Pictures [Ver02]. Picturing [FS95, Moo95]. piece [Büh98a, Hol02]. Pierces [Ano34b]. Pierre [Six88]. pila [Sal02a]. Pile [AT62a, AT62b, AFW^{+}47, Ano46a, EP08, Fer42c, F^{+}42a, Fer42j, Fer42u, FZ43, FWN^{+}45, Fer46d, Fer47b, Fer57a, Fer70b, Gan02, Wat74, AFW^{+}57, F^{+}42b, F^{+}42f, F^{+}42e, F^{+}43h, F^{+}43i, Fer46e, Sal02a, AFMW42, Ano22, Wei63, Fri55b]. Pile-1 [Ano22]. Piles [F^{+}42a, Fer42m, Fer42l, F^{+}42b, F^{+}42c]. Pilgrimages [Bet04]. Pion [Fer53b, Fer53c, Fer54f, Fer54a, Led03]. Pioneer [Ano55c, Cra69, Ham02]. pioneers [Wil75]. Pions [AFNY52a, AFNY52b, AF52, AFL^{+}52, AFLN52, AFMN53, Fer51a, FAL^{+}52, Fer53a, FMN53, FGMN53, Fer54f, Fer54e, FMA54, Fer83, Kot54, LFA^{+}52, NAF^{+}52, Cif08, Fer54a, FF55, FF08, Fra90c, BDH89]. Pisa [Gam03, LRS00, Ver01a]. Pisanti [Jam11]. Pitchblende [Ano34d]. Pitchblende1 [SP48]. Place [Hol74, Dys10]. Placzek [Ama56]. plan [AG14]. Planck [Fis10a, Fis12a, KGSO55, Cra88, Fis10a, Fis12a]. plane [Fer23l]. Planetarium [Bax01]. Planets [Dol64, Dol70, Dol07]. planned [F^{+}42b]. plans [Pap85]. Plant [ACF^{+}42, BF42, LCM^{+}42]. plates [Fri55b]. Platinum [Arb00]. Ploeger [Cre10]. plurality [Cro86]. Plus [AFW42, AF42b, LW71]. plutonium [MB00, vHEGM12]. Pockets [CK08]. Poetry [Sol68]. Poincaré [Fer23g]. Point [Vol08]. Polarisation [FR25a, FR25d]. Polarization [FR25b, FR25c, Fer53b, Fer54g, Fer54h, Fer23l, FR25d]. polarizzazione [Fer23l, FR25b, FR25c]. Policies [Sei90, CHW91, WC84]. Policy [Sag83]. politica [Bel87, Goo91, Mal08]. Political [Bat03a, Sei90, Gol80, Gol82b]. Politics [Bar05, Hab69, Sim12, Wil04b, Bel87, Fri01, LW71, Sch99b, Sei90]. politique [Gol80]. poltergeist [Rei96]. Polygamma [Din58]. Pontecorvo [Cal13, Hay16, Lau13a, Mal08, Seg15, B^{+}13, Bon05, Clo15, Dra72a, Maf92, Tur12]. Pope [Lan17, Wes16, Mal10, SH16]. popularizer [Sci01]. Population [Bar05, CHU87, Cra92]. portrait [Ano47]. portraits [Far01]. Portuguese [Seg87a]. Poses [Hub01]. position [F^{+}43h]. Positive [AFNY52b, AFLN52, And33, CW32a, CW32b, GP31]. positron [Maj37]. positrone [Maj37]. positroni [FU33b]. Positrons [FU33a, FU33b]. possibilità [FRD34]. Possibilities [Kae39]. Possibility [FRD34, FS50]. Possible [Cha32b, Fer34g, Lew21, San81]. Postage [Hub01]. Postulate [Hal28]. Postwar [CSW97b]. potentials [SHL^{+}18]. potere [Val06]. Pourquoi [Six88]. Power [Ano67a, Ano16, BF42, Bun83, Com53, Cra07, Fer46a, Kae39, Sei90, EE70, LCM^{+}42, Rho18b, Val06]. powerful [Ano51]. Pp [Buc73b, Cal13, Fri55b, Bar05, Bat06, Bro68, Gra61, Ish19, Lau13a, Mal08, Min02, Tur08]. pp. [Lan17]. practices [Tur06b]. precede [CS00c]. Preceded [Ken54]. preceding [CS00b]. precedono [CS00b]. precision [HS19]. Prediction [GLR14]. Preface [Cif08]. pregiudizio [CM02]. Preis [CSW97a]. Preliminary [FSA41, Fer42q]. preludono [CS00c]. Premi [MFS^{+}6x]. première [Mon66]. preparatory [F^{+}42f, F^{+}42e]. Preparing [Nea14]. Presence [Fer42x, SP48, CF53b]. Present [Gou32, Kat75, Ama02b, CHW91, WC84]. Presentation [Ber03, Ano65b]. preserved [Der64]. President [Ein39]. Press [Ano95b, Ano95c, Ano96b, Bar05, Bat06, Bro73c, Buc73b, Cal13, Cam54, Gol99a, Ish19, Lau13a, Wil71]. Pressione [Fer34m]. presso [Der64]. Pressure [AF44b, AYS95, Dug61, Fer34m]. Pressure-Tuned [AYS95]. Prevalence [Oak18, VD15]. pri [Fer52h]. price [CSW97a]. prima [Fer25a]. primer [SR92, Ree17]. primi [CS99]. Princeton [FSA41, Ano58]. principi [Fer25e]. principio [DS11, Fer23h, FP26, Fer30d, Fer30e, Fer30l, Ber02c]. Principle [Fer30l, Bai18, Fer23h, FP26, Fer30d, Fer30e, Ber02c, DS11]. Principles [Fer25e]. privedshem [Bel82]. Prize [Ano68a, Bar05, Fer38e, Pau45, Ano55a, Ano63b, Bet63, CSW97a, CSW97b, Fri01, G.64, MFS^{+}6x, Peg39, Wal64b]. Prizes [Bar05, Cra02, CHU87, Bar05, Har02]. Prizewinners [Wil69]. prizewinning [Gra16a]. pro [Fer26a, Fer27b]. probabilità [Fer22h, Fer23m, Fer26e]. Probability [Fer23m, Fer22h, Fer24h, Fer26e, Zag11, Zag13]. probably [Gri18]. probe [Kow15]. Problem [Ano05, Ano15b, BPP18, BGJR24, Blu01, Boc15, Ćir16, ED83, Fer42r, FM52, Gal03, Hub01, Nis18, Tar03, Web15, Wei97, BI04, BI05, EFGP99, Shi81, Str98, Web02]. Problem-Solving [Nis18]. Probleme [Fer28i]. Problemi [Fer26b, Fer28d, Fer29e]. Problems [CF53b, Fer26b, Fer28i, Fer29e, F^{+}42c, F^{+}42d, F^{+}42f, F^{+}42e, AG14, EL07, Fer28d, FPU55, FPU74, PUF65, Rob08]. Proceedings [Ano03, MR86, GHMP87, SA79, Stu79, WH72, DGS89, GS74]. Process [Ano39d, FAP^{+}40, CSW97b]. processes [Fer26h, Fer52h, Som28b]. processi [Fer52h]. Prodotta [Fer34k]. Prodotti [AFRS34]. Produced [Ano34b, Fer34a, FAD^{+}34a, FPR34a, Fer34k, Fer34i, Fer39b, Fer40c, Fer40d, Fer51a, Fer70a, HS89, AFRS34, ADF^{+}35d, ADF^{+}35a, FAD^{+}34b]. Produces [Ano39d, Ano34c]. Producing [FRD34, Fer36a]. Product [Fer44e]. Production [Ama03, AFS39, AFH39, AF40, AF41a, AFF^{+}42b, AFM46, BM35, Fer34g, FAP^{+}40, Fer41, Fer42n, Fer52a, Fer53a, Fer54f, Fer54e, Kot54, PAF^{+}42, dH62, Fer42t, F^{+}42g, Hei55, Hei83]. Products [Mei39, MF39c, MF39b]. produrre [FRD34]. Prof [Ano64]. Prof. [Ano38a, FU43b, Fri55a]. Professor [Mot81, DS07, AA55, Ano34c]. professore [DS07]. Professors [Ano39a]. Professorship [Ano57, Ano64]. Program [Gon96b, Sei90, BDM81, Tur06a]. Progress [Kae35, Wal74b, BI04, BI05, Pap85]. proibite [Fer26d]. project [De 05a, De 06a, Fer55, Fer70c, Bet00, FS95, Fit99, Fit13, Gro67a, Gro67b, Gro62, Gro83, HTS83, HHW99, Nie89, Sch15, Sim12, Sza92, Moo95, Sei90]. Projected [F^{+}42b]. Projects [Coc62]. Proletarian [Sol68]. proliferation [vHEGM12]. Prometheus [Bar05]. Proof [HS89, Fer23b]. propagation [Fer38b]. propagazione [Fer38b]. Properties [BY90, Fer27e, Fer28g, Fer45, KSSW46, SW48, WS48, CCJ^{+}34, Fer28f, Fer28b, F^{+}42c, F^{+}42d, F^{+}42f, F^{+}42e, Lib85]. property [AF36e]. prophet [Ber80]. proprietá [Fer27e, Fer28g, AF36e, Fer28f]. propriétés [CCJ^{+}34]. Prospective [For01]. Prospects [De 01, Fer38f]. Prospectus [FFH^{+}44]. Prospettive [De 01, Fer38f]. prossima [Gan02]. Protactinium [BW39a]. protégé [AM23]. Protest [AAF^{+}52, Sol68]. Proton [Ano74, Fer53b]. Protons [BM35, Fer35b, Fer54g, Tam34, CW32b, Fer54h]. Prototypes [Bun83]. provacata [FAP^{+}34]. Providence [GLR20]. Proving [Ken54]. provocata [ADF^{+}34b, ADF^{+}34c, ADF^{+}34d, ADF^{+}34a, ADF^{+}35e, ADF^{+}35b, ADF^{+}35c, FPR34b, Fer34l]. provocate [FPR34a]. Proximity [FK08, Oli75]. pseudopotential [GS22]. Psi [Din58]. pubblicato [FV02]. Publication [GR08, GR05]. published [FV02]. Pulsars [Cow09, Pen13]. Pure [All62b, ML48, Mil05]. Purity [FA61]. Purpose [Fer42s]. Purposes [Smy45a, Smy45b]. Pursuing [Str98, Wei97]. puti [Bel82].

QED [Sch03a]. Qiu [Fra95, Fra95]. Quadrature [GM85a, GM85b]. quanta [Fer26a, Fer27b]. Quantelung [Fer26g]. Quantensprung [Fis10a, Fis12a]. Quantentheorie [Fal28]. Quantenzustände [Fer24h]. quanti [Fer23a, Fer25e, Fer26a, Fer27b, Fer29d]. quantica [CS00c, CS00b, CS00d, Mai91]. quantici [Fer23m]. quantique [Geo32]. quantistica [Cin02, Fer29f, Fer29i, Fer30d, Fer30e, Fer30g, Fer30n, Fer31b, Gal02, Mai91, Maj28, PS07, Seb97]. quantistica. [Fer31c]. Quantization [Fer26f, Fer24b, Fer26g]. quantizzazione [Fer24b, Fer26f]. Quantum [And99, CK08, DS09, DF06, Dir03, Fer23m, Fer25e, Fer26c, Fer29f, Fer29i, Fer30g, Fer32d, Fer54i, Fer66a, Gam85, GL23, Gri96, Kra99, Sch02, Sch03a, Sch58, Sch03b, SC03, Sop76, Sop80, Sop88, Ste71, Ano95c, Cin02, Con62, CS00c, CS00b, CS00d, DS10, Dir26, Fer23a, Fer24h, Fer29d, Fer30d, Fer30e, Fer30n, Fer31b, Fer61a, Fer65c, Fis10a, Fis12a, Gal02, Gre05, KHFA67, Lou83, Mai91, Maj28, PS07, Roq92, RK95, SW72, Seb97, Zag11, Zag13, Fal28, Fer68a, Geo32]. quark [San81, Seg83]. Quarks [Nam85, BDH89, Seg80, Seg84, Seg87a, Seg07]. quasi [Fer23b, Fer23d, Fer23e, Fer24f]. quasi-ergodic [Fer23b, Fer23d, Fer23e, Fer24f]. quasi-ergodico [Fer23d, Fer23e]. quasi-ergodisch [Fer23b]. quasi-Ergodischer [Fer24f]. quella [Fer23c]. quest [Com56]. Question [Jon85b]. Questions [Mor63b, And73a, KS99]. questo [Ano02]. Qui [Dem15]. Quick [Fri50, KS99]. quick-change [KS99].

R [Fri55a, Sei90]. Ra [AFRW42, FF48]. Rabi [Rig00]. racconta [Bru11]. raccontate [Lom93]. Race [Ano95a, Lau51, Rob95, Sei90, Mac86, Mac87, Sei90]. Rachel [Moo95]. radiant [Mil05]. Radiation [Dys60, FP23, Fer24d, FR25a, FR25b, FR25c, Fer32d, Fer49f, Fer49g, Fer54b, FR25d, Fer29d, Fer31a, FR33b, FR33c, Fer49b]. Radiative [De 05b, De 06b]. radiazione [FP23, FR33b, FR33c]. Radio [Fer42t, PAF^{+}42]. Radio-Active [PAF^{+}42, Fer42t]. Radioactive [FAP^{+}40, MA40, Rob50, SMKW46, SWK46, Maj28]. Radioactivity [AGR04, De 05b, De 06b, Fer34a, FAD^{+}34a, FPR34a, Fer34h, Fer34i, Fer35d, Fer35e, Fer39b, Fer70a, GLR06, GR09, GLR20, L’A07, Mon66, Wig43, ADF^{+}35d, ADF^{+}35a, Car01, De 05a, De 06a, FAP^{+}34, Fer34d, Fer35a, Fer38f, Fer65a, ADF^{+}34b, ADF^{+}34c, ADF^{+}34d, ADF^{+}34a, ADF^{+}35e, ADF^{+}35b, ADF^{+}35c, Fer34j, Fer34l, Fer34k]. radioaktivity [FAD^{+}34b]. radioaktivnoct [ADF^{+}35a, FAD^{+}34b]. radioattivi [Maj28]. Radioattività [ADF^{+}34b, ADF^{+}34c, ADF^{+}34d, ADF^{+}34a, ADF^{+}35e, ADF^{+}35b, ADF^{+}35c, Fer34j, Fer34l, Fer34k, Car01, FAP^{+}34, FPR34a, FPR34b, Fer34d, Fer35a, Fer35d, Fer35e, Fer38f]. Radioelementi [AFRS34]. radioelements [AFRS34]. Radiological [dR85]. Radisson [Ano05]. Radium [Ano34c, FF42, Kae48]. Radium-Beryllium [FF42]. Radon [AF42b]. ragazzi [AFM88, AS96, Ano01b, Bia90, BCR09, Cod88, FM03, Lom93]. Ragazzo [Tur08]. Raggi [Fer34n, Fer22b, Fer22d, Fer22e, Fer23f, FR33a, Fer33b, Seg83]. raios [Seg87a]. Ralph [AM23]. Raman [Fer31e, FR31, Fer32b]. Ramaneffekt [Fer31e, FR31]. Ramsay [Mon66]. Range [AFN44, FT43, FW43]. Ranges [FHN44]. Ranotgen [Seg86a]. rapporto [FR27b, Fer30m]. Rapports [CCJ^{+}34]. Rasetti [Goo01, Tur08]. Rate [Fer42r]. Ratio [FR27b, Fer42d, Fer42f, FR27a, Fer30o, Fer30m]. Rational [CT67, SS11]. Ratios [AFG41]. Ray [Cow09, FR33a, Mor58]. Raymond [Swi45]. rayonnement [Fer31a]. rayons [Seg84]. Rays [Kae39, Rob95, CA36, Fer22b, Fer22d, Fer22e, Fer23f, Fer33b, FS34a, FS34b, Fer34n, Fer34o, Seg80, Seg84, Seg87a, Seg07]. reaching [De 06a]. Reacting [AFW^{+}47, F^{+}42a, Fer42m, Fer42l, Fer42y, Fer46d, Fer47b, FL57, FL58, Fer70b, AFW^{+}57, F^{+}43i]. Reaction [All62b, And73b, Ano68d, Ano95a, Ano07, BHT86, Fer41, Fer42k, Fer42r, Fer52a, Gol62, MF39a, Wat74, And73a, Ano11, F^{+}42g, FS50, Ste59]. Reactions [Fer40c, Fer40d, FB41, FU43a, FS58b, Fer47a, SF44, FU43b]. Reactivity [AFW^{+}47, AFW^{+}57]. Reactor [ATFF82, Ano55b, Cal75, FS55, FZ55, FA56, FS57a, FZ57, FS58a, FZA58, Fer58b, Fer60, FS61, Fer61b, Fol86, LF62, dH62, Ban03, FS57b, Lan01, Wol80]. Reactors [Bun83, Fer70d, Jam11, EP10, Fer71b, Gan02, Cum02]. read [Gra16a]. Reader [Bad67]. Ready [Fra05]. reaktsiyami [Fer47a]. Real [Ano56b, Rab63, WWL^{+}09, Moo95]. realistic [BP68]. realities [WD06]. realization [Orl98]. really [FB69, Fra04]. realm [Kae48]. reasoning [D’A03]. Reattori [Cum02, Gan02]. Recalls [Lau51]. Recapture [Hol74]. Receive [Ano55c, Ano63b, Bet63, Ano81, CSW97b]. received [CSW97a]. Receives [Ano62, Ano68c, Ano70, Ano94b, Ano73]. Recenti [Fer35d, Fer35e]. Recently [CSW97b]. Recognition [Ano55b]. Recollections [ABP98, Dir77, HS11, APB96, B^{+}13]. Recombination [FU33a, Fer35b, FU33b]. Reconstruction [Ama79]. record [Wol80]. recording [Ano75, Fri74]. Recordings [CDFS53a, CDFS53b, CDFS54, Fer52c, Ber96]. Recovery [AGR04]. red [Fer25g]. Redazione [Fer35a]. Reference [Fer42l]. refinement [Fer22g].

Reflection [Fer24d, FZ46a, FZ46b, Fer24c]. reflections [Dys10, IAE97, Wig67]. Reflectors [BF42]. Refraction [FZ44]. Refugee [Seg85, Jon10]. Regelungsmathematik [KGSO55]. Related [Sei90, Lie81, Lie82]. relating [Ano02]. Relation [Fer25g, Fer45, MP56, Pau25]. relativa [Ano02]. relativist [Fer22f]. Relativistic [KS42, Kra92, LT05, Fer23c]. relativistica [Fer23c]. relativistischen [Fer22f]. Relatività [Fer25a, CS99, Fer22a, Fer23i, Kop23]. Relativity [Fer25a, Fer22a, Kop23, Tre76, CS99, Reg03]. relazione [Fer25g]. Released [Ano39c, CSW97b, Lau46b]. Relics [Mad87]. remark [Fer34p]. Remarkable [Jam04b]. Remarks [Cha73, FB41, Fer41, Sal03a]. remembered [Cro04, PP72, Rig84, Bat06, Mal05, Kai04]. Remembering [Shi91]. Remembers [Fir05]. Reminiscences [Ano75, BHB80, Fer62b, Ros72, Nie89, Wil75]. ren [Seg86b]. renewal [DE06a, DE15]. Report [ACF^{+}42, Ano39a, Ano41b, BC03, FSA41, Fer42q, Fer44d, Fer44e, Fer53c, FR76, KHFA67, Smy45b, Kae45, Pri95]. reported [Bey49, PZHC09]. Reports [CCJ^{+}34]. representation [Seg03]. Reproduction [AFF^{+}42a, CFW42, Fer42b, Fer42g, Fer42h, F^{+}42a, Fer42o, Fer42s, F^{+}42b, F^{+}43h]. reproductions [Wri64, Wri65]. Research [Cif08, FSA41, F^{+}42c, F^{+}42d, F^{+}42f, F^{+}42e, Fer49c, Met74, MS01, Pie01, FR35]. researchers [Jam11]. researches [Cor98]. resistance [Dug61]. Resistivity [Dug61]. Resolution [Bax01, BGJR24, Wes90]. Resolving [BHR19, SAĆ16]. Resonance [AYS95, Fer24d, FR25a, FR25b, FR25c, FSA41, FW43, FM47e, Fer24c, FR25d]. Resonanzstrahlung [FR25d]. Response [Ore03b]. Responsibility [Bad05]. Restructuring [DGS89]. resulting [HS39b]. Results [Fer35d, Fer35e, GR89]. retirement [SA79]. Retrieval [EP08]. retrospect [Stu79]. Returning [Ano01c]. Reveal [Ano09a, CSW97b]. Revealed [Kae45]. Review [All62a, Ama73, Ano95b, Ano95c, Ano96b, Bad67, Bad71, Bal63, Bar05, Bat06, Bra69, Bro68, Bro73c, Buc73a, Buc73b, Cam54, Cer69, CW71, Cre10, Dra72a, Dra72b, Dro20, Fal73, Fra95, Fri55b, Gar18, Gol70, Gol99a, Goo91, Gra61, Gri96, Hay16, Hub13, Ish19, Kai04, Lan18, Lau13a, Lut66, Mag61, Mal08, Man57, Min02, Moo95, Mot81, Oak18, Rab63, Rab66, Rei69, Rob83, Seg15, Sha67, Smi68, Ste93, Str98, Stu06, Tal17, Tar03, Tur08, Wei71, Wes16, Win86, Zuc70, AM23, Kle71, Ols63, Wal04]. Reviews [Cal13, FBL64, Mal05, Pal69, Sei90, Sol68, Wil71, PFM^{+}22]. Revised [Ore58, Ore82, Ore01b]. Revisiting [Ree17]. Revolution [FBL64, Gol99a, Kae48, FB61, FB03, Gal02, Gre05, Gra61]. revolutions [Coo99]. Rey [Mon66]. Rhetoric [Cre10]. Rhetoric/Communication [Cre10]. Rhodes [Moo95, Sei90]. Ricerche [CL03, FR35, MS01, Pie01, dL74, Fer49c]. Richard [Moo95, Sei90]. Richardson [Fer23n]. Richtmyer [Fer52g]. Rickover [Wal64a]. ricombinazione [FU33b]. Ricordi [Lee01, Tel01, Pon93, Sal01]. riflessione [Fer24d]. Righe [Fer34m, Fer25c, Fer26d]. Right [Mei62, Sch15, Fra90a]. rigid [Fer21a]. Rigido [Fer21a]. rigorous [GR89]. Riley [CW71]. rinnovamento [DE06a, DE15]. Risk [Car15]. risonanza [Fer24d, FR25b, FR25c]. risultati [Fer35d, Fer35e]. Rita [Sol68]. Rivista [PFM^{+}22]. rivoluzione [Gal02]. Rn [AFW42]. Road [Bar05, Sei90, Har02]. Roads [Mei62]. Robert [Ano63a, Ano63b, Ano95c, Bar05, Bet63, Sol68, Ano65a]. rock [FR31]. rod [F^{+}43h]. Roger [Sei90]. Role [Ken54, Pin03]. Roma [Min02, Ver01a, CL03]. Rome [Ano03, DGS89, SA79, BB02, Bet04, Hol03, SA79, Seg88, Ver01a]. Romer [Mon66]. Röntgen [Fer22b, Fer22d, Fer22e, Fer23f]. Roosevelt [Dem15, Ein39]. Roots [Pri95]. Rossi [Min02, LMR06]. rotante [Fer23l, RF26]. rotating [Fer23l]. rotation [Fer23l, Fer32c, Fer32e, Fer32f]. rotazione [Fer32c, Fer32e, Fer32f]. Route [OVPL15, Bel94, Her79]. Row [Wil71]. Rubbia [Goo91, Bel87, Bel88]. rule [CIR04, Seg03, BM20, CBH09, DF06, RNKS04, Ste71]. runs [Lau46b]. Russian [Bro68, ADF^{+}35a, Bel82, Dir03, FAD^{+}34b, Fer47a, Fer52h, Fer68a, Fer71a, Fer98, GS74, Pon55, PP72]. Rutherford [Mon66, MF37]. Rydberg [Fer28e, Fer28g]. Rydbergkorrektionen [Fer28e].

S [Bro68, Cre10, Jam11]. S. [Dra72a, Fer42l, Fri55a]. saggi [Ano02]. salt [FR31]. Salvini [Ano01b]. Salvo [Min02]. Salzman [Sol68]. Same [Fer36a]. Samra [Moo95]. Sant [GHMP87]. Santa [Ano05]. Sapienza [DGS89]. save [Cra07]. Savoia [Ano01b]. Saw [Sol68]. say [FB69]. Says [Ano34b]. sbarcato [Ano82]. scale [PQQ94, Fra95]. Scar [Sol68]. Scattered [AFNY52a, AFMN53, Fer54h, Ols63]. Scattering [AF52, Fer42i, FM46a, FM46b, FM47d, FM47e, FM47f, FM49, FAL^{+}52, Fer53b, Fer53c, FMN53, FGMN53, Fer54g, FMA54, LFA^{+}52, NAF^{+}52, RNKS04, Fer54a, Roq92]. scelse [Maf92]. scelti [Fer96]. Scenario [Vol08]. sceneggiatura [AS96]. Schatten [Chi94]. Schenectady [Ano41a]. Schluter [Ano95c]. scholars [Shi91]. School [MR86, Seg88, Wat88, HS11, Ver01a]. Schools [AF38, Fer29b, FP38, Fer52b]. Schroeer [Sei90]. Schwartz [Gar18, Lan18]. Science [AK16, Ano34f, Ano38b, Ano46b, Ano52, Ano53b, Ano54c, Ano54d, Ano56b, Bar05, Ćir18, Cra02, Hab69, Hol74, Kae35, Kae39, Mon66, Reg87, Wig62, Bat01, Bel87, Cra92, Dra76, Fri01, Gre05, Har02, HH04a, LW71, Mar96, Val06, vB93, Bar05, Ano94a]. Sciences [Ano41b, DGS89, Kae35, WH72]. Scientific [AF38, Bon02, Dar98, Deb53, Dir03, Dro20, Fer51d, FBL64, GHMP87, Gra61, Kev77, Rod19, Bey49, B^{+}13, DP97, Dra72b, DS92, FB61, FB03, GH90, Hol90d, Hol78, Hol98, Kev87, Kev95, Mal03b, Ols63, SS99, Sci01, Seg71, Seg87b, Seg09, Wig67, Wig96b, Wri64, Wri65]. Scientifica [Bon02, Dra76, Mal03b, SS99, Seg71, Seg87b, Seg09, Dra72b]. Scientifici [AF38]. scientifico [Sci01]. scientifiques [Mon66]. Scientist [Kae45, Kra78, Ols63, Ore04, Maf92, Sho09]. Scientists [AAF^{+}52, Ano53a, Ano12c, Bar05, Bax46, Bax68, CFU47, GB89, GR63, Rog10, Seg85, Sei90, Wea76, AK17, Har02, Kur02, Rob12, Shi91, Sza92, Bar05, GR63]. scienza [Bel87, Goo91, Val06, GER^{+}92]. scienziato [Maf92]. Scope [Jam11]. scoperta [Ano82, Ano01a, Car01, Dra76]. scoperte [Seg76, Seg83]. SCRAPS [FB64]. screenplay [AS96]. scritti [Fer96]. Scuole [FP38]. Sea [Ano06]. Seaborg [Ano59]. Search [All62b, Bau85, Dys60, SD75, SP48, Sim00, AK17, DS92, Pap85, Sho09, Wil01]. Searching [CM59]. seated [Ano51]. Seawater [Ano12c]. secchi [CM02]. secolo [Sal01, Wei77]. Second [Ano34b, Ano95c, BM20, Gri04, Stu18]. Secondary [FP38]. Secret [Ano95a, FS95, Kae45, Moo95, Wal04, Ber96, Pur63, Wea76]. Section [AF4x, AFL^{+}52, F^{+}43b, F^{+}43c, F^{+}43d, F^{+}43e, F^{+}43f, F^{+}44c, F^{+}44d, F^{+}44e, F^{+}44f, F^{+}44g, F^{+}44a, F^{+}44b, FMM47, AF42a, AF42b, BFMM44, FMM45, F^{+}43a, F^{+}43g]. Sections [AFW^{+}47, AFNY52b, AFLN52, AFW42, AFW^{+}57, Fer42f]. Security [Ken54]. Seeds [BF42]. Seeing [WWL^{+}09]. Seeking [BB15b, BB15a]. Seems [GLR14]. Segrè [Ano96b, Buc73b, Lan17, Wes16, Wil71, Ama73, All62a, Bad71, Bro73c, Buc73a, Dra72b, Gol70, Kle71, Pal69, Rab63, Seg93, Wei71]. seinem [GHH02]. Selected [DP97, Sch58, Sch03b, Sei90, Yan13c, B^{+}13, Fer96]. Selection [ABP98, Wal64a, APB96]. Selector [FMM47, Fer50c]. Self [Mar75, Ano11, SHL^{+}18, TSH^{+}19]. Self-consistent [Mar75]. self-interaction [SHL^{+}18, TSH^{+}19]. self-sustained [Ano11]. Semiclassical [RNKS04]. semiconductor [MZF^{+}12]. semiconductors [CIR04]. Separated [NBDG40]. September [Ano03, Ber02a, DGS89, GHMP87, Meh73, SA79]. septième [CCJ^{+}34]. Serie [Fer34m]. Series [GM85a, GM85b, Fer34m]. serious [Fer22a]. servizi [Mal08]. Session [Ano34b]. SETI [Pen13]. settembre [Ber02a, GER^{+}92]. Seven [Arb00, Rho18b]. Seventh [CCJ^{+}34, Far01]. Seventy [Web15, Esp08a]. Seventy-Five [Web15]. Several [Wil71]. Severi [GB12]. shade [Chi94]. shadows [LS92]. shape [Ree00]. shapes [HS19]. sharpening [AS13]. Shells [Goe48, Goe49]. Shield [FZ43, FZ57, HA69, Sei90, HAD90]. Shift [FMA54]. Shoenberg [Mot81]. Shook [Gam85]. Short [Ber03, KGSO55, Zim62a, Zim62b, Zim62c, Zim62d, Zim62e]. short-time [KGSO55]. shoulders [Ano47]. Shows [Ano60]. sicurezza [Mal08]. Sig [Ano09b]. signal [Shi81]. Signals [Wal64a]. Signatures [MZF^{+}12]. significato [Ano82, Ano01a]. Silence [Ćir18, Bri83]. Silver [CBH09]. simmetrica [Maj37]. Simone [Lau13a, Cal13, Mal08]. Simons [Sol68]. Simple [AF39, Str98, Sch89]. Simplified [Fer42u]. Simulation [Wei97, Str98]. since [Meh75]. Sinisgalli [BCR09]. Sironi [Mal08]. Sistema [Fer21a, Fer23d, Fer23e, Fer23g, Fer28f, Fer28h]. sistemi [FV02, Fer23h, Fer24b]. Site [Com42]. Sitte [Fer34p]. Six [Arb03, Sei90, AS13, Ree00]. Sixty [FB64, FR13]. Sixty-Four [FB64]. Size [Car15, Fer42c]. Skills [Win86]. Slow [AF35a, AF35b, AF36a, AF36b, AF36d, Ama03, AF40, AF41b, Fer35c, Fer36c, FAW38, Fer42e, FM47a, FM47b, FM47e, FM47f, FSS47, FM49, SZ39, Tur06a, AF36e, Bon05, Dra76, FR35, Fer38d]. Slowing [Fer42v, Fer42w, FMM43]. smallest [Fis10a, Fis12a]. smasher [Ano51]. Smith [Sei90]. Smyth [Kae45]. So-Called [Gra16b]. Sobranie [Dir03]. Social [Bad05, Wil71]. Società [Stu06]. Society [Bet55]. Socio [Ama02a]. Socrates [Gri04]. Soddy [Mon66]. soglie [GER^{+}92]. Sole [Fol86]. Solid [Fer25b, Bas02, Bul81]. solid-state [Bas02, Bul81]. solidi [Fer25b]. Solido [Bas02]. Solids [LS77]. Solipsist [SN83]. soliton [Bul81]. solo [Mai91]. Solution [Ano94a, FM52, GG51, HMB09, Ber18, Esp02, vB93, Ste93]. Solutions [BY90, Tar03, Web15, ABF14, Döb22, Web02]. Solvay [CCJ^{+}34, Far01, Meh75, CCJ^{+}34]. Solve [Ano15b, AG14]. Solved [Boc15]. Solving [ED83, For19, Nis18, Ish19]. Some [Fer27e, Fer28g, Fer41, F^{+}42a, Fer62b, GR86, Nie89, Ros04, Sag83, Wal64a, Dug61, Fer22h, Fer28f, Fer28b, Ols63, Ore03a, Spr91, Fer23a, Pon92, Pon13, Sal01]. Sommario [For01]. Sommerfeld [DuM28, Nis18]. Songs [Sol68]. Sopra [AF36b, Fer22c, Fer23j, Fer24d, Fer25c, Fer25b, Fer25d, Fer26e, Fer26d, Fer29f, Fer30g, Fer34m, RF26, Fer22g, Fer22d, Fer22e, Fer23g, FR25b, FR25c, Fer31c]. sora [CM02]. Sorry [BR57]. Sort [Ser12]. sostanze [FAP^{+}34, FPR34a, FPR34b, Fer36d]. Sound [Fri74, Ano75]. Source [AFRW42, FF42, FW43, Kae39]. Sources [Dys60, GLR06, KHFA67]. sous [CCJ^{+}34]. Souvenir [Per55]. Soviet [Gon96b, SSSS95]. sozdavaemaya [ADF^{+}35a, FAD^{+}34b]. Space [Bau85, FP23, WWL^{+}09, O’N74]. Spaceship [Sut92]. Spain [GHMP87]. spazio [FP23]. Special [Cif08, SSSS95]. Spectra [Fer30i, Fer30j, Fer31d, Pau25]. Spectral [Fer24a, Fer34m]. spectral-line [Fer34m]. spectrograph [FR33a]. spectrometry [Nie89]. spectroscopic [MS01]. Spectrum [HS19]. speculation [Pei86]. Speeches [Bas01, CDFS53a, CDFS53b, CDFS54, Fer52c, Ano65b]. Speed [Ano34b, Lau13b, Sei03]. spektraallijnen [Fer24a]. Spektren [Pau25]. Spencer [Sei90]. sperimentale [Fer29c, Fer30b, Sal01]. sperimentazione [Dra76]. Spettrali [Fer34m]. spettri [Fer30i, Fer30j, Fer31d]. spettrografo [FR33a]. spettroscopiche [MS01]. Sphere [FW41, FSA41]. Spies [Lan94]. Spin [FM47f, FM49, Gou71]. Spinning [RF26]. Spiral [CF53a]. Split [Ano39d]. Splits [Kae39]. spoon [Kea10]. Spostamento [Fer34m]. spreading [AS13]. Spriel [Mon66]. Springer [Stu06]. Springford [Mot81]. spy [Clo15, Seg15, Hay16, Sei90]. spymaster [SSSS95]. stability [CF53b, Spr91]. Stage [BP97, BP58]. Stagg [Moo92b]. staircase [Fis10a, Fis12a]. Stallion [Sei90]. Stamp [Hub01]. stand [Rob08]. Standardization [AFMW42]. Standards [AF41b]. Stanisław [PZHC09]. Stanley [Sol68]. Stars [Bau85, Fer30a, Spr91]. start [Wol80]. start-up [Wol80]. Started [Ano68d]. State [FMT49b, Bas02, Bul81, Fer32a, Fer32c, FMT49a, GS89]. Statement [OBC^{+}49]. States [DGS89, Fer23m, Fer65b, HA69, HDHA89, Pal69, Smy45b, Fer24h, Zag11, Zag13, Fer57b, Rid84, Wat88]. Stat’i [Dir03, Fer23m]. Statistica [Fer36b, Bel97, CS00c, CS00b, CS00a, CS00d, Fer23n, Fer28f, Fer28g, Fer29a, Fer96, Par02]. Statistical [DE04, Fer27e, Fer28i, Fer28g, Hal28, Sol03, Fer23n, Fer28f, Fer28b, Fer29a, GR05, SW87, Fer28e, Fer36b, Fer96]. statistico [Fer27e]. Statistics [Fer66a, Fer66b, Kap40, Mos50, Ore58, Ore82, Ore01b, SC03, Sei55, Ama87, Bel82, Bel94, Bel97, CS00c, CS00b, CS00a, CS00d, KaKKZL05, Par02, Sch89, Som27, Som28a, Som28b]. Statistik [Som28a, Som28b, Som27]. statistike [Bel82]. statistische [Fer28b, Fer28e]. statistischen [Fer28i]. Stato [Bas02, Fer32c]. Status [F^{+}42c, F^{+}42d, F^{+}42f, F^{+}42e]. Steinsalzes [FR31]. Stellar [Dys60]. stelle [Fer30a]. Stephen [Mon66, Tar03, BB10]. Steps [G.64]. Stern [Fer23j]. Sternstunden [Büh98b]. still [Kae48]. Stockholm [Bar05, Fer39b, Fer70a, Har02, WD06]. Stoner [Fuk14]. Stood [vH01]. Storia [Bel87, Goo91, Seb97, Wei77, Bat03b, Maf92, Val06]. Storico [CL03, Pie01]. Story [Gam85, Kae45, Kra18, Lau46b, Mon66, Fer61d, Gro67a, Gro67b, Gro62, Gro83, HTS83, Kie13, Lom93, Maf92, Pur63, Val06, Mon66]. Stoßes [Fer24g]. Stoßvorganges [Fer26h]. Strahlen [Fer34o]. Straßmann [CSW97a, CSW97a, FH81, Sim03]. Striking [Hol74]. Strömungs [Som28a]. Strömungs- [Som28a]. strong [Fer26d]. Strontium [FM47d]. Structure [CCJ^{+}34, Fer28i, Gou32, Pau34, FS33a, FS33b, Fer33c, Pau25, CCJ^{+}34]. strumentazione [Car05]. strutture [FS33a]. Stück [Büh98a]. Student [Fir05, Ano96a]. Students [Ros04, PP72, Pon93]. Studi [CL03, Pie01]. Studies [Ano05, Bar05, Cra02, Cre10, FPU55, FPU74, PUF65, Rob83, Bey49, Cra92, Hol78, dL74, Fer22d]. Studii [Fer22d]. study [Fis10a, Fis12a, Pie01]. sua [Sal02b]. suas [Seg87a]. Substance [Ano34d]. Substances [FPR34a, FAP^{+}40, FAP^{+}34, Fer36d]. success [Bel87]. Successful [Ken54]. successo [Bel87, Goo91]. Sudoplatov [McM94]. sue [Fer22h]. Suffer [Met74]. suggested [Gra16a]. Sugli [LC02]. Sui [Fer25e, Fer30h, Fer36c, FR35, FR38]. Sul [AF36c, CS00d, Fer21b, Fer23k, Fer23l, Fer27d, Fer29g, Fer29h, Fer30k, Fer30i, Fer30j, Fer30l, Fer30m, Fer31d, Fer36d]. Sulla [Fer21a, Fer22e, FP23, Fer23m, Fer23n, Fer25f, Fer26f, Fer28f, Fer28g, Fer28h, Fer29i, Fer30n, FU33b, FS33a, FRD34, CS99, Fer24b, FR33b, FR33c, FAP^{+}34, FPR34a, FPR34b]. Sull’assorbimento [AF35a, AF35b, AF36d]. Sulle [AF36e, Fer32e, Fer32f]. Sull’Elettrostatica [Fer21b]. Sull’equilibrio [Fer24e]. Sull’origine [Bel97]. Summary [F^{+}43a, F^{+}43b, F^{+}43g, F^{+}43c, F^{+}43d, F^{+}43e, F^{+}43f, F^{+}44c, F^{+}44d, F^{+}44e, F^{+}44f, F^{+}44g, F^{+}44a, F^{+}44b, For01, Sel53]. Summation [GM85a, GM85b]. Summer [Gol82a]. Sun [Rho95, HHW99]. Super [FR49]. Superbomb [GB89]. Supercomets [Gri11]. Superconducting [FK08]. superconductor [MZF^{+}12]. Superconductors [CK08]. Superheavy [Kra18]. superiore [Ano01b]. Superiori [FP38]. Supplement [GM85b]. Supply [Rub03]. Surface [Ano60, Fal73, FK08, Har61, KHB12, WWL^{+}09, Ano09c, CW73, Hoc83, Mot81, Zim62a, Zim62b, Zim62c, Zim62d, Zim62e, Mag61]. Surfaces [Ano09a, Hal69, Ano94c, Dug61, Pin74]. survey [GHH02]. Sustainability [HMB09]. sustained [And73a, Ano11]. Sustaining [FL57]. sviluppo [Cum02]. svolta [PS07]. Swados [Sol68]. Swedish [CSW97b]. Swing [Swi45]. Sword [Sei90]. Sylves [Sei90]. Symmetric [Fer51h]. Symmetrical [Maj37]. Symmetries [GHMP87, Wig67]. Symmetry [BY90]. Symposium [Bet55, KGSO55, Meh73, Stu79, Sch00, BDH89]. System [Fer28a, Fer28h, FL57, FL58, Fer21a, Fer23b, Fer23d, Fer23e, Fer23g, Fer28f, Fer28b, OVPL15]. Systeme [Fer24f]. Systems [Bro68, DF06, Fer28b, PG86, FV02, Fer23h, Fer24b, Fer24f]. Szilard [Lan01, And74b, And74a, Ano55b, Bye02, Fra05, LS92, Lan92, Lan01].

T [Ano93a, Sei90, Dir03]. Table [Fer42x, Kea10]. tables [Hol02]. tagliyotehe [Seg86a]. Take [Ano06]. taken [CSW97a]. Takes [G.64]. Taking [FR48]. Tale [CSW97b, CSW96]. tales [Kea10]. talk [Ano12b]. talking [Yor87]. talks [Bru11]. tallio [FR27b]. Tanner [Sol68]. Tape [CDFS53a, CDFS53b, CDFS54, Fer52c]. task [Orl98]. tasks [SSSS95]. Taylor [Fer51d, Fer51g, FvN51, FvN55]. teach [Rob08]. Teacher [Bas01, Ste01]. teachers [Shi91]. Teaching [DE06b, DE07, DE06a, DE15, SA79]. Technical [Ano01b, Fer38a]. technically [Flü39]. technisch [Flü39]. technologies [Cum02]. Technology [Peo04, Sei90, LW71]. Tecnici [Fer38a]. tecnico [Ano01b]. tecnologie [Cum02]. Teeming [Tar03, Web15, Web02]. Teil [Som28a, Som28b]. Teilchen [Fis10a, Fis12a, Fer24g]. Telegamma [Rob95]. Telescope [Cow09]. Tell [Ano39a, Wei94]. Teller [Ano62, Fir05, Sei90, Wig62]. Tells [Ano34b]. temperatura [Fer22e]. Temperature [CK08, CFW42, Fer42g, Fer42y, Fer22e, F^{+}42g, F^{+}43h]. Temperatures [Dug61]. tense [Lan12]. Tentativo [Cab03, Fer33b, Fer34n]. Tenth [Fer52g]. tenu [CCJ^{+}34]. teorema [Fer22h, Fer23g]. Teoremi [Mir34, Fer23a]. Teoria [Fer25a, Fer34n, Maj37, Fer22a, Fer22g, Fer22e, Fer23a, Fer23c, Fer23i, Fer23j, Fer23n, Fer25e, Fer25b, Fer25f, Fer28f, Fer28h, Fer29a, Fer29d, Fer29i, Fer30n, FS33a, Fer33b, Fer47a, Kop23, Maj28]. teorica [DE06a, DS07, DE15, Fer27a]. teorico [CS99, Sal01]. teorie [Fer29c]. Teoriya [Dir03]. Terme [Fer28e, Fer28e]. termica [Fer22g]. termico [Fer24e]. termini [Fer28g]. Termodinamica [Fer58a, Fer72, Fer82b]. Termodinamika [Fer98]. Terms [Fer28g, Fer29g, Fer30k]. terrestre [FR33b, FR33c]. terzo [GER^{+}92]. Test [Fer57a, Fer58b, Ken54]. testimonials [dL74]. testimonianze [dL74]. Testimony [Ano49]. Testing [FA56, FA61, Sei90, Sei90]. Tests [FZ43]. Texts [Bro68]. Thales [Bad67, Ama66]. Thallium [FR27b, FR27a]. Thalliumdampfes [FR27a]. Their [Buc80, Fer27c, HHW99, AM23, And99, Dra76, Dug61, FPLR59, Seg80, Seg84, Seg87a, Seg07]. them [Ols63]. theorem [Fer22h, Fer23g, D’A05]. theorems [Fer23a, Mir34]. Theoretical [DE06b, Fer42n, Fer42o, Meh73, Cor98, CS99, CSW97b, DE06a, DS07, DE07, DE15, EAW15, Fer27a, Sal01]. Theorie [Fer22f, Fer24g, Fer28b, FS33b, Fer34o, Fer31a]. Theories [BR94, Fer22f, Fer29c, Lie81, Lie82, dL64, dL65]. Theory [Bad67, Bha39, BGJR24, BM35, BR94, Dir03, DuM28, Fer23j, Fer25a, Fer25e, Fer25b, Fer28h, Fer29i, Fer32d, Fer33c, FS34a, FS34b, Fer47b, FMT49b, FG92, Fra86, Gam85, Gou32, GR09, Kon55, Kot54, LS77, LM83, Mon66, PG86, Sol03, Ste71, SZ62, SM67, Tam34, Wil68, Yan13b, Ama66, BRV84, Bul81, CA36, Dir26, Fal28, Fer22a, Fer22g, Fer22e, Fer23a, Fer23c, Fer23n, Fer24g, Fer25f, Fer28f, Fer28b, Fer29a, Fer29d, Fer30n, Fer31a, FS33a, FS33b, Fer33b, Fer34n, Fer34o, Fer46e, Fer47a, Fer01, FMT49a, Fra90b, GR86, GR89, Hoc83, Kop23, KaKKZL05, Lou83, Maj28, Maj37, Mar83, SW72, SHL^{+}18, SS11, Som27, Som28a, Som28b, Spr91, Whe93, Tau63, Bro68]. There [AK16, Fre85, Lem98, DS92, Fra04, Kow15, Ols63, Rob08]. Thermal [AF42a, AF4x, BFMM44, FW41, Fer42a, Fer42f, Fer42p, FZ44, FM47d, FMM47, BJ41, Fer22g, Fer24e, F^{+}42g]. thermalization [OVPL15]. Therman [FA61]. thermischer [BJ41]. Thermo [Som28b, Som28b]. Thermo-electric [Som28b]. Thermo-elektrische [Som28b]. thermo-magnetic [Som28b]. thermo-magnetische [Som28b]. Thermodynamic [De 05b, Kot54, De 06b]. Thermodynamics [Fer37a, Fer56, Fer66b, Fer10, KS42, BM20, Fer58a, Fer72, Fer82b, Fer98, Gri04]. Thermonuclear [Fit99, Fit13, Gon96a, Gon96b, FS50]. Things [Sol68]. Thinking [Win86]. Third [Ano58]. thirteen [Bey49]. Thirties [Hei03, Hug03, Pin03, Sol68, Dra76, Ama02b, Sol68, Stu93]. Thirty [Arb00, Arb03, Gam85]. Thirty-Six [Arb03]. Thomas [Bar05, Som32a, Som33, ABF14, BY90, BP68, Bro61, EFGP99, Esp02, Esp05, FMT49a, FMT49b, GR86, GR89, GG51, GR05, GR08, GP31, Hil69, LS77, Lie81, Lie82, MP56, Mar75, Mar83, Mar00, MŚ94, PG86, Rij66, Som32a, Som32b, Som33, Spr91, SZ62, SM67, Str98, ZZWK12]. Thomas-Fermischen [Som32a]. Thorium [HS39a, HS89, Mei39, MF39b]. Thousand [Ano34b]. Three [And73a]. Thus [Ano39a]. Time [And74b, And74a, Ano16, Bye03, Fer42r, Fer66c, Mac86, Mac87, Rog10, Seg79, Wat74, vHEGM12, Bax46, Bax68, KGSO55, Wil75, Sei90]. Times [Gar18, Kra78, Sch17, Ama66, Lan18, Dea13]. Titanic [Jon10]. title [Fer52g]. titled [Mon66]. Titus [Sei90]. Today [Ano98]. Tokyo [Bar05]. told [Gro62, Gro83]. too [Boc15]. Topological [FK08]. Torino [Tur08]. Total [AFNY52b, AFL^{+}52, AFLN52]. Tough [Sol68]. tra [Fer22a, Fer23c, Fer25g, Fer25f, Fer25d, Mai91]. Trackdown [Ols63]. trad [Dra72a]. trading [Tur06b]. traduction [Mon66]. tragedies [RK95]. trail [CS00d]. Trainer [Bas01]. Traité [Chw13]. Trajectories [Fer51h]. transfer [JB83]. transform [KGSO55]. Transformation [KGSO55]. Translated [Bro68]. translation [Dra72a]. translational [Fer21a]. translations [PHK25]. Transmission [FSS47]. Transmutation [Mon66]. Transplantation [Smi68]. transport [CIR04]. Transuranic [Kra18]. Transuranium [McM51, Sim00]. Trapped [DJ99]. trascinamento [Fer23l]. trascorso [Sal01]. Traslatorio [Fer21a]. trattato [Ber01]. Travel [Sch94]. Traveling [Dro20]. Travelling [Rod19]. treatise [Ber01, Chw13]. Trends [FPLR59, Met74]. trenta [Ama02b, Dra76]. Triatomic [Fer25g]. triatomiche [Fer25g]. tribute [Kub09, MF37]. Tributes [MF37]. Trieste [Meh73]. Trinity [Sei90, And74a, Kat21, Lam65, Lam85, Lew21]. trip [SHM19, WD06]. trolley [Fer46g]. troubled [Mag09]. Trudov [Dir03]. trudy [Fer71a, GS74]. true [Kea10]. Truman [Lef95]. truth [Cra07]. tsepnimi [Fer47a]. Tsung [Ano64]. Tsung-Dao [Ano64]. Tübingen [Pau30]. Tuned [AYS95, Bai17b, LBS17]. tuning [Bai17b]. Turchetti [Cal13, Lau13a, Mal08]. Turns [Ano97, BB15b, BB15d, BB15a, BB15c, KS99]. twelve [Wil75]. Twenties [PS07]. Twentieth [Bar05, Dar04, Kra99, Ste77, Wei77, Buc73b]. Twentieth-Century [Bar05, Dar04]. twenty [LW71]. twenty-five [LW71]. Two [BP58, BP97, FvN51, Lau51, BF32, CS00c, Hol02, Mai91]. Two-Stage [BP97, BP58]. Type [BY90, MF39a, Ser12, Mar00].

U.S. [Tur06a]. uchenikov [PP72]. Ugo [Ano96a]. Ulam [Ano05, OVPL15, Str98, AMS19, Ano15b, BPP18, BI04, BI05, DPR05, Dau08, PZHC09, Wei97]. ultimate [Fer33a, Fer34e]. ultime [Fer33a, Fer34e, Ric02]. Umberto [Ver01a, Ver01a]. unbounded [Kae48]. Uncertainty [DS11]. Underlying [BPP18]. undulatory [Fer27d]. Uniform [Fer21b, Fer23g]. Uniforme [Fer21b]. uniformi [Fer23g]. Unit [Fer42y, Fer44c]. United [DGS89, Fer65b, Pal69, Fer57b, HA69, HDHA89, Rid84, Smy45b, Wat88]. Universal [Bar05, Fra90c, Tre76]. Universe [Ano03, Cow09, Dav82, Gri11, Tar03, Web15, Bai17b, Dea13, Dys10, Lau46b, LBS17, Ree00, SS77, Web02, BS17, Lan12]. Università [DGS89, Ver01a]. University [Ano64, Ano95b, Ano95c, Ano96b, Bar05, Bat06, Bro73c, Buc73b, Cal13, Cam54, Fer65c, Gol99a, Ish19, Lau13a, SA79, Wil71, AAF^{+}52, Ano07, FSA41, Fer42q, Fer49e, Fer50d, Fer50e, Fer55, Fer70c, Fer85, Lau51, Shi91, Ver01a]. unknown [Fer54j]. unmasked [Ano09c]. Unpublished [Jam11, Ver02, EP10]. Unravelling [FR13]. Unseparated [AF42b]. untold [Gro67a, Gro67b, Kie13]. unwanted [SSSS95]. Unwin [Fri55b]. Upper [FP38, Fer38b]. Uran [HS39a]. Uranium [ABD^{+}39, AFS39, AFH39, AF39, AFG41, AF41a, AF42b, AFF^{+}42b, AFF^{+}42a, Ano46a, Ano34d, Ano39a, Ano39d, Ano12c, Bet00, FW41, FS41, FSA41, Fer41, Fer42d, Fer42n, Fer42o, Hah62, HS89, Kae39, Lib79, MF39a, MF39c, NBDG40, SMKW46, SWK46, SZ39, Ano34c, Ber96, HS39a, HS39b, MF39b, Ano41a]. Urans [HS39b]. Uranspaltung [HS39a]. Urey [FU43b]. URSS [Maf92]. urto [Fer25d]. USA [Ano05]. Use [BF42, Car15, Fal73, KGSO55, Mor57, CW73, Fer52b, Sti47, Sti76]. Users [Wal74a]. Uses [Ano34c, Ano39d, Fer57b, Man57]. Using [Cra88, Win86, F^{+}43i]. uso [Fer52b]. USSR [Maf92]. Utilization [FU43b, FU43a].

v [PP72, ADF^{+}34d]. Vakoch [Oak18]. Valeria [Tur08]. valore [Kop23]. Value [FMM45, Kop23]. Values [Fer44e]. Vannevar [Gol92, Hol90b]. vapor [FR27a]. Vapour [FR25a]. Varenna [GER^{+}92, MR86, Fer54d, HS11, Ric01]. variabile [Fer29h]. variable [Fer29h, Sch99b]. Variety [Ano34b]. various [F^{+}42c, F^{+}42d, F^{+}42f, F^{+}42e]. Vast [Ano39d]. Velocity [Fer35c, FMM47, CW32a, CW32b, Fer50c, Shi81]. Venti [PS07]. veramento [FB69]. Verhalten [HS39b]. Verhältnisses [FR27a]. Verlag [Stu06]. Verleihung [KGSO55]. Verschwinden [GHH02]. version [Mon66]. verso [CS00d]. Versuch [Fer34o]. Versus [Rob95]. very [Hei55, Hei83]. vi [Bar05, Tau63]. Via [Tur08, AS96, Ano01b, Bia90, BCR09, Lom93, Mal10, Ama03, AFM88, AS96, Bia90, BCR09, CL03, Car05, Cod88, Lom93, Ano01b, FM03, Pie01]. Vibration [Fer32e, Fer32f]. Vicinanza [Fer22c]. View [Fer54l, Fer54m, Mil07]. Views [Nis18]. VII [ADF^{+}34a, AYS95]. viii [Bar05, Bro68, ADF^{+}35e]. Viking [Wil71]. Villa [GER^{+}92]. Vincent [Sei90]. visit [Fer49c, Fer76]. visita [Fer49c]. Visualization [CBH09]. vita [Dra76, Pon92, Pon13, Sal02b, Tur08, dL74]. viva [FP26]. Voci [MFS^{+}6x]. voices [MFS^{+}6x]. Vol [Fer62a, Fer65b, Mor63a, Wig96b, All62a, Bro68, Mon66, Rab63, Rab66, Sha67]. Volt [Ano39d]. Volta [Som27, Som27]. Volta-Effektes [Som27]. voltages [Som33]. Volts [Ano39b, Ano39d]. Volume [Pal69, Ano02, Rei72, MKR87, Tau63]. Vorgänge [Som28b]. vospominanijakh [PP72]. vuoto [FP23].

W [Bar05, Bat06, Mag61, Mal05, Sei90, FZ43]. Wahrscheinlichkeit [Fer24h]. Walter [Ano70]. Walther [Six88]. War [Bad05, Cal13, Kie13, Lau13a, Mal03a, PHK25, Sag83, Sim12, Tur12, De 03]. Warfare [Fer57c, Tel57]. Warren [Sol68]. Wars [Mai91, Stu18]. wartime [CSW96]. Warum [CSW97a]. Was [Büh98a, Sei90, Bia90, Fer54l, Fer54m, Hop90]. Washington [Bet55]. Water [AF44b, AFN44, Bun83, Fer42w, Wig43]. Wave [Fer30l, FP26, Fer26h]. Wave-Mechanics [Fer30l]. waves [Fer38b]. Way [Wil69, Esp08b, Hol02, Gri18]. Weak [Cab03, Fra86, BRV84, Cab02]. weak-interaction [BRV84]. Weapon [Fit99, Fit13]. Weapons [Bad05, Yor87]. Weart [Sei90]. Weaver [Sei90]. Webb [Mag61, Tar03]. wechselnden [FR25d]. Wechselwirkung [BF32]. Week [Kae39]. Weight [Ano34b, Fer21b, Fer23k, Fer22a]. Weinberg [Ano81]. Weintraub [Sol68]. Weissert [Str98]. weiterer [HS39a]. Weizsäcker [BY90]. Welcome [Cor04]. well [KS99]. Wellenmechanik [Fer26h]. went [Ols63]. Wentzel [MP56]. Wer [Ano10]. werden [Flü39, CSW97a]. We’re [BR57, Cor64, Sei90, Bey49, Sol68]. Werner [MP77]. Whatever [Ćir16]. Wheeler [Ano68a, Ano68b, Ano68c]. Where [BB15c, BB15d, Cra00, Döb22, Jon85a, Tar03, Web15, vH01, HZ82, Ols63, ZH95, Jon85b]. Which [SAĆ16, Lau46b]. Who [Ano53a, Gar18, Sch17, Dem15, Gre05, Kie13, Lan18, Maf92, Ano10, Dem15]. wide [Bel87]. Widely [Ano39a]. Widerspruch [Fer22f]. Wife [Fer04a, Wei94]. Wigner [Ano58, HH04b, Seg03, Wig67, Wig96b]. Wild [Pei86]. Wiley [Bro68, Win86]. will [Döb22, Kra15, Rig84]. William [Sei90]. Williams [Sei90]. Wilson [Ano65a]. win [Kie13]. Window [Cow09]. winners [MFS^{+}6x]. Wins [Ano59, Ano93b]. witness [SSSS95]. Wolfenstein [Wol04]. Woman [Fol86]. Women [Bye03, AS11, HHW99, Kie13, Chi94]. Won [Ano95a]. Wong [Fal73]. Words [Win86]. Work [BB04b, Bon02, Deb53, F^{+}42g, F^{+}43h, F^{+}43i, Fol86, Lau51, Mal05, Stu06, BB04a, CS99, F^{+}42f, F^{+}42e, Gal02, Kra92, PS07, Pon92, Pon13, Sal02b]. Working [Wil04c, Fer71a]. Works [Dir03, Tau63, B^{+}13, CS00c, Wig96b]. World [Ano46a, CW71, Fer22c, FS95, GR63, HA69, Hol07, Man57, Moo95, Pes03, Ril70, Rub03, Sei90, Stu84, Stu18, Wil71, AK17, Ano51, Bel87, Cip94, Cra07, Cum02, Fer57b, Gre05, HA62, HA90, Kea10, LW71, Pai86, Ste59, Ano82, Ano95a, Bad05, Kie13, Mal03a, PHK25, Sim12]. world-wide [Bel87]. worlds [Cro86]. worldwide [Gol82b]. Writer [Sol68]. Writers [Sol68, Sol68]. Writings [ABP98, Jam11, APB96, EP10, Fer96]. Wrong [Mei62, Fra90a]. Wu [Seg86b].

x [Bar05, Gra61, ADF^{+}35c, CA36, Com42, Fer22b, Fer22d, Fer22e, Fer23f, Kae39, Rob95, Seg80, Seg83, Seg84, Seg87a, Seg07]. X-Rays [Kae39, Rob95, CA36, Fer22b, Fer22d, Fer22e, Fer23f, Seg80, Seg84, Seg87a, Seg07]. xian [Seg86b]. xii [Bat06, Buc73b, Stu06]. xix [Seg15]. xv [Bar05]. xvii [Bar05]. XX [Wei77].

Yadernie [Fer52h]. yadernimi [Fer47a]. Yang [Fra95]. Year [FR13, Kae35, Ano15b, Bai17a]. Years [Ama79, Ano05, Bar05, Ber88, Gam85, Mla98, Rog13, BI04, BI05, BRV84, Con62, CS00b, CDS01, Dea13, Ern92, Esp08a, Fis97, GHH02, Kae48, LW71, PZHC09, Sop88, Sza92, Sol68]. Yield [Ano34d]. York [Bar05, Bro68, Dea13, Gol99a, Gra61, Lan17, Seg15, Stu06, Wil71]. you’ll [Gra16a]. Yukawa [Jam04b].

Zametki [Bel82]. Zeilinger [Fis10a, Fis12a]. zero [Bai17a]. Zhaoming [Fra95]. Zinn [Ano70]. Zip [Ano46a, Ano46b]. Zoo [Bal73]. zum [Fis10a, Fis12a]. zur [Fer28b, Fer26g, Fer26h, FS33b, Fer34p, Som28a, Som28b]. Zusammenhang [Pau25]. zwei [BF32]. zwischen [Fer22f, Fer24g].

References

Anderson:1955:PFN

[AA55] H. L. Anderson and Samuel K. Allison. From Professor Fermi’s notebooks. Reviews of Modern Physics, 27(3):273–275, July 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

            (electronic), 1539-0756. URL http://link.aps.org/doi/10.
            1103/RevModPhys.27.273;   http://rmp.aps.org/abstract/
            RMP/v27/i3/p273_1.

Anderson:1952:PUC

[AAF^{+}52] Herbert L. Anderson, Samuel K. Allison, Enrico Fermi, Willard F. Libby, Joseph E. Mayer, Edward Teller, and Harold C. Urey. Protest from University of Chicago scientists. Bulletin of the Atomic Scientists, 8(7):255–256, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders.

Anderson:1939:FU

[ABD^{+}39] Herbert L. Anderson, E. T. Booth, J. R. Dunning, Enrico Fermi, G. N. Glasoe, and F. G. Slack. The fission of uranium. Physical Review, 55(5):511–512, March 1, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v55/i5/p511_2.

Amaldi:2012:ALF

[ABES12] Edoardo Amaldi, S. (Saverio) Braccini, Antonio Ereditato, and Paola Scampoli, editors. The adventurous life of Friedrich Georg Houtermans, physicist (1903–1966). Springerbriefs in Physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-32854-7 (paperback), 3-642-32855-5 (e-book). ISSN 2191-5423. xiv + 152 pp. LCCN Q124.6-127.2; QC16.H688 A43 2012. URL http://www.sif.it/

attivita/saggiatore/recensioni/e_amaldi.

Amore:2014:ACS

[ABF14] Paolo Amore, John P. Boyd, and Francisco M. Fernández. Accurate calculation of the solutions to the Thomas–Fermi equations. Applied Mathematics and Computation, 232(??):929–943, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300314001829.

Amaldi:1998:CPE

[ABP98] Edoardo Amaldi, Giovanni Battimelli, and Giovanni Paoloni, editors. 20th Century Physics: Essays and Recollections: a Selection of Historical Writings, volume 3 of Edoardo Amaldi Foundation series. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1998. ISBN 981-02-2369-2. xiii + 747 pp. LCCN Q127.I8 A53 1998. URL http://www.

worldscientific.com/worldscibooks/10.1142/2852.

Allison:1942:RCE

[ACF^{+}42] Samuel K. Allison, C. M. Cooper, Enrico Fermi, Eugene P. Wigner, and Leo Szilard. Report of the committee for the examination of the Moore–Leverett design of a He-cooled plant. Report CE-324, US Atomic Energy Commission, Washington, DC, USA, 1942. Included are four pages that seem to have been written by Enrico Fermi, chairman of the committee. Undated: year chosen according to report number.

Arimondo:2010:CEM

[ACM10] E. Arimondo, Charles W. Clark, and W. C. Martin. Colloquium: Ettore Majorana and the birth of autoionization. Reviews of Modern Physics, 82(3):1947–1958, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://adsabs.harvard.edu/
abs/2010RvMP...82.1947A;   http://link.aps.org/doi/10.
1103/RevModPhys.82.1947;   http://rmp.aps.org/abstract/
RMP/v82/i3/p1947_1.

Adams:1972:FGN

[Ada72] J. B. Adams, F.R.S. Four generations of nuclear physicists. Notes and Records of the Royal Society of London, 27(1):75–94, August 1972. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Amaldi:1934:RPBe

[ADF^{+}34a] Edoardo Amaldi, Oscar D’Agostino, Enrico Fermi, Franco Rasetti, Bruno Pontecorvo, and Emilio Segré. Radioattività provocata da bombardamento di neutroni. VII. (Italian) [Radioactivity induced by neutron bombardment. ]. La Ricerca Scientifica, 5(2):467–470, ???? 1934. CODEN RISCAZ. ISSN 0035-5011.

Amaldi:1934:RPBa

[ADF^{+}34b] Edoardo Amaldi, Oscar D’Agostino, Enrico Fermi, Franco Rasetti, and Emilio Segré. Radioattività provocata da bombardamento di neutroni. III. (Italian) [Radioactivity caused by neutron bombardment. III]. La Ricerca Scientifica, 5(1):452–453, ???? 1934. CODEN RISCAZ. ISSN 0035-5011.

Amaldi:1934:RPBb

[ADF^{+}34c] Edoardo Amaldi, Oscar D’Agostino, Enrico Fermi, Franco Rasetti, and Emilio Segré. Radioattività provocata da bombardamento di neutroni. IV. (Italian) [Radioactivity induced by neutron bombardment. IV]. La Ricerca Scientifica, 5(1):652–653, ???? 1934. CODEN RISCAZ. ISSN 0035-5011.

Amaldi:1934:RPBc

[ADF^{+}34d] Edoardo Amaldi, Oscar D’Agostino, Enrico Fermi, Franco Rasetti, and Emilio Segré. Radioattività provocata da bombardamento di neutroni. V. (Italian) [Radioactivity induced by neutron bombardment. V]. La Ricerca Scientifica, 5(2):21–22, ???? 1934. CODEN RISCAZ. ISSN 0035-5011.

Amaldi:1935:IRS

[ADF^{+}35a] Edoardo Amaldi, Oscar D’Agostino, Enrico Fermi, B. M. Pontecorvo, Franco Rasetti, and Emilio Segrè. Iskusstvennaya radioaktivnoct’, sozdavaemaya neitronnoi bombardirovkoi — II. (Russian) [Artificial radioactivity produced by neutron bombardment — II]. Uspekhi Fizicheskikh Nauk, 15(7):838–870, July 1935. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (elec-

tronic). URL http://ufn.ru/ru/articles/1935/7/d/. Russian

translation of [ADF^{+}35d].

Amaldi:1935:RPBb

[ADF^{+}35b] Edoardo Amaldi, Oscar D’Agostino, Enrico Fermi, Bruno Pontecorvo, and Emilio Segré. Radioattività provocata da bombardamento di neutroni. IX. (Italian) [Radioactivity induced by neutron bombardment. IX]. La Ricerca Scientifica, 6(1):435–437, ???? 1935. CODEN RISCAZ. ISSN 0035-5011.

Amaldi:1935:RPBc

[ADF^{+}35c] Edoardo Amaldi, Oscar D’Agostino, Enrico Fermi, Bruno Pontecorvo, and Emilio Segré. Radioattività provocata da bombardamento di neutroni. X. (Italian) [Radioactivity induced by neutron bombardment. X]. La Ricerca Scientifica, 6(1):581–584, ???? 1935. CODEN RISCAZ. ISSN 0035-5011.

Amaldi:1935:ARP

[ADF^{+}35d] Edoardo Amaldi, Oscar D’Agostino, Enrico Fermi, Bruno M. Pontecorvo, Franco Rasetti, and Emilio Segrè. Artificial radioactivity produced by neutron bombardment. Part II. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 149(868):522–558, April 10, 1935. ISSN 0950-1207 (print), 2053-9150 (electronic). URL http://

www.jstor.org/stable/96379.

Amaldi:1935:RPBa

[ADF^{+}35e] Edoardo Amaldi, Oscar D’Agostino, Enrico Fermi, Franco Rasetti, Bruno Pontecorvo, and Emilio Segré. Radioattività provocata da bombardamento di neutroni. VIII. (Italian) [Radioactivity induced by neutron bombardment. VIII]. La Ricerca Scientifica, 6(1):123–125, ???? 1935. CODEN RISCAZ. ISSN 0035-5011.

Amaldi:1935:SNLa

[AF35a] Edoardo Amaldi and Enrico Fermi. Sull’assorbimento dei neutroni lenti. I (Italian) [On the absorption of slow neutrons. I]. La Ricerca Scientifica, 6(2):344–347, ???? 1935. CODEN RISCAZ. ISSN 0035-5011.

Amaldi:1935:SNLb

[AF35b] Edoardo Amaldi and Enrico Fermi. Sull’assorbimento dei neutroni lenti. II (Italian) [On the absorption of slow neutrons. II]. La Ricerca Scientifica, 6(2):443–447, ???? 1935. CODEN RISCAZ. ISSN 0035-5011.

Amaldi:1936:ADS

[AF36a] Edoardo Amaldi and Enrico Fermi. On the absorption and the diffusion of slow neutrons. Physical Review, 50(10):899–928, November 15, 1936. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/

v50/i10/p899_1. English translation of [AF36b].

Amaldi:1936:SAD

[AF36b] Edoardo Amaldi and Enrico Fermi. Sopra l’assorbimento e la diffusione dei neutroni lenti. (Italian) [On the absorption and diffusion of slow neutrons]. La Ricerca Scientifica, 7(1):454–503, ???? 1936. CODEN RISCAZ. ISSN 0035-5011. See also English translation in [AF36a].

Amaldi:1936:SCL

[AF36c] Edoardo Amaldi and Enrico Fermi. Sul cammino libero medio dei neutroni nella paraffina. (Italian) [On the mean free path of neutrons in paraffin]. La Ricerca Scientifica, 7(1):223–225, ???? 1936. CODEN RISCAZ. ISSN 0035-5011.

Amaldi:1936:SNL

[AF36d] Edoardo Amaldi and Enrico Fermi. Sull’assorbimento dei neutroni lenti. III. (Italian) [On the absorption of slow neutrons. III]. La Ricerca Scientifica, 7(7):56–59, ???? 1936. CODEN RISCAZ. ISSN 0035-5011.

Amaldi:1936:SPD

[AF36e] Edoardo Amaldi and Enrico Fermi. Sulle proprietà di diffusione dei neutroni lenti. (Italian) [On a property of diffusion of slow neutrons]. La Ricerca Scientifica, 7(1):393–395, ???? 1936. CODEN RISCAZ. ISSN 0035-5011.

Amaldi:1938:FPL

[AF38] Edoardo Amaldi and Enrico Fermi. Fisica per i Licei Scientifici. (Italian) [Physics for Scientific Schools]. N. Zanichelli, Bologna, Italy, 1938. ???? pp. LCCN ????

Anderson:1939:SCN

[AF39] Herbert L. Anderson and Enrico Fermi. Simple capture of neutrons by uranium. Physical Review, 55(11):1106–1107, June 1, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v55/
i11/p1106_2.

Anderson:1940:PAS

[AF40] Herbert L. Anderson and Enrico Fermi. Production and absorption of slow neutrons by carbon. Report A-21, US Atomic Energy Commission, Washington, DC, USA, September 25, 1940.

Anderson:1941:PNU

[AF41a] Herbert L. Anderson and Enrico Fermi. Production of neutrons by uranium. Report A-6 (CPA-6), US Atomic Energy Commission, Washington, DC, USA, January 17, 1941.

Anderson:1941:SSN

[AF41b] Herbert L. Anderson and Enrico Fermi. Standards in slow neutron measurements. Report A-2, US Atomic Energy Commission, Washington, DC, USA, June 5, 1941.

Anderson:1942:ACSb

[AF42a] Herbert L. Anderson and Enrico Fermi. Absorption cross-section of boron for thermal neutrons. Report C-74, US Atomic Energy Commission, Washington, DC, USA, 1942. Undated: year chosen according to report number.

Anderson:1942:FCS

[AF42b] Herbert L. Anderson and Enrico Fermi. Fission cross-section of unseparated uranium for fast radon plus beryllium neutrons. Report C-83, US Atomic Energy Commission, Washington, DC, USA, 1942. Experiment in November, 1941. Undated: year chosen according to report number.

Anderson:1944:BAF

[AF44a] Herbert L. Anderson and Enrico Fermi. Boron absorption of fission activation. Report CP-2161, US Atomic Energy Commission, Washington, DC, USA, 1944. 3 pp. In: Fermi, E., et al. Argonne Laboratory. Report for month ending September 23, 1944.

Anderson:1944:DPW

[AF44b] Herbert L. Anderson and Enrico Fermi. Dissociation pressure of water due to fission. Report CP-1729, US Atomic Energy Commission, Washington, DC, USA, May 1944. 1 pp. In: Fermi, E., et al. Nuclear Physics Division. Part I of report for month ending May 25, 1944.

Anderson:194x:ACS

[AF4x] Herbert L. Anderson and Enrico Fermi. The absorption cross section of boron for thermal neutrons. Report CP-74, US Atomic Energy Commission, Washington, DC, USA, 194x.

Anderson:1952:SCP

[AF52] Herbert L. Anderson and Enrico Fermi. Scattering and capture of pions by hydrogen. Physical Review, 86(5):794, June 1, 1952. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v86/i5/p794_1.

Anderson:1942:NRL

[AFF^{+}42a] Herbert L. Anderson, B. T. Feld, Enrico Fermi, George Weil, and Walter H. Zinn. Neutron reproduction in a lattice of uranium oxide and graphite. Report C-20, US Atomic Energy Commission, Washington, DC, USA, March 26, 1942.

Anderson:1942:NPL

[AFF^{+}42b] Herbert L. Anderson, B. T. Feld, Enrico Fermi, George L. Weil, and Walter H. Zinn. Neutron production in a lattice of uranium oxide and graphite (exponential experiment). Report CP-20, US Atomic Energy Commission, Washington, DC, USA, March 26, 1942.

Anderson:1941:BRF

[AFG41] Herbert L. Anderson, Enrico Fermi, and A. V. Grosse. Branching ratios in the fission of uranium (235). Physical Review, 59(1):52– 56, January 1, 1941. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/
PR/v59/i1/p52_1.

Anderson:1939:PNU

[AFH39] Herbert L. Anderson, Enrico Fermi, and H. B. Hanstein. Production of neutrons in uranium bombarded by neutrons. Physical Review, 55(8):797–798, April 15, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v55/i8/p797_2.

Anderson:1952:TCSa

[AFL^{+}52] Herbert L. Anderson, Enrico Fermi, E. A. Long, R. Martin, and Darragh E. Nagle. Total cross section of negative pions in hydrogen. Physical Review, 85(5):934–935, March 1, 1952. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v85/i5/p934_2.

Anderson:1952:TCSb

[AFLN52] Herbert L. Anderson, Enrico Fermi, E. A. Long, and Darragh E. Nagle. Total cross sections of positive pions in hydrogen. Physical Review, 85(5):936, March 1, 1952. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http:

//inspirehep.net/record/28195;   http://prola.aps.org/
abstract/PR/v85/i5/p936_1.

Anderson:1946:PLE

[AFM46] Herbert L. Anderson, Enrico Fermi, and Leona Marshall. Production of low energy neutrons by filtering through graphite. Physical Review, 70(11–12):815–817, December 1, 1946. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v70/i11-12/p815_1.

Amelio:1988:RPI

[AFM88] Gianni Amelio, Ennio Fantastichini, and Laura Morante. I ragazzi di Via Panisperna. (Italian) [The boys of Via Panisperna]. 125 minute film from San Paolo audiovisivi, Italy., 1988. Screenplay by Gianni Amelio and Alessandro Sermoneta, photography Tonino Nardi, directed by Gianni Amelio, Andrea Prodan, and Ennio Fantastichini. Starring Virna Lisi, Andrea Prodan, and Alberto Gimignani.

Anderson:1953:ADP

[AFMN53] Herbert L. Anderson, Enrico Fermi, R. Martin, and Darragh E. Nagle. Angular distribution of pions scattered by hydrogen. Physical Review, 91(1):155–168, July 1, 1953. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v91/i1/p155_1.

Anderson:1942:SAP

[AFMW42] Herbert L. Anderson, Enrico Fermi, John Marshall, and Leona Woods. Standardization of the Argonne Pile. Report, US Atomic Energy Commission, Washington, DC, USA, 1942. Excerpt from Report CP-641 for month Ending May 10, 1943.

Anderson:1944:RFN

[AFN44] Herbert L. Anderson, Enrico Fermi, and Darragh E. Nagle. Range of fission neutrons in water. Report CP-1531, US Atomic Energy Commission, Washington, DC, USA, 1944. Excerpt from Metallurgical Laboratory Report CP-1531 for month ending March 25, 1944.

Anderson:1952:ADP

[AFNY52a] Herbert L. Anderson, Enrico Fermi, Darragh E. Nagle, and Gaurang B. Yodh. Angular distribution of pions scattered by hydrogen. Physical Review, 86(5):793, June 1, 1952. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v86/i5/p793_2.

Anderson:1952:DTC

[AFNY52b] Herbert L. Anderson, Enrico Fermi, Darragh E. Nagle, and Gaurang B. Yodh. Deuterium total cross sections for positive and negative pions. Physical Review, 86(3):413, May 1, 1952. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://inspirehep.net/record/47341;
http://prola.aps.org/abstract/PR/v86/i3/p413_1.

Amaldi:1937:GAN

[AFR37] Edoardo Amaldi, Enrico Fermi, and Franco Rasetti. Un generatore artificiale di neutroni. (Italian) [An artificial generator of neutrons]. La Ricerca Scientifica, 8(2):40–43, ???? 1937. CODEN RISCAZ. ISSN 0035-5011.

Amaldi:1934:NRP

[AFRS34] Edoardo Amaldi, Enrico Fermi, Franco Rasetti, and Emilio Segrè. Nuovi radioelementi prodotti con bombardamento di neutroni. (Italian) [New radioelements produced with neutron bombardment]. Il Nuovo Cimento (8), 11(7):442–447, ???? 1934. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL

http://www.springerlink.com/content/g14637jq4r7r54g8/.

Anderson:1942:NNE

[AFRW42] Herbert L. Anderson, Enrico Fermi, J. H. Roberts, and M. D. Whitaker. The number of neutrons emitted by a Ra + Be source (source I). Report C-21 (MDDC-880), US Atomic Energy Commission, Washington, DC, USA, March 21, 1942.

Anderson:1939:NPA

[AFS39] Herbert L. Anderson, Enrico Fermi, and Leo Szilard. Neutron production and absorption in uranium. Physical Review, 56(3):284– 286, August 1, 1939. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://adsabs.harvard.edu/abs/
1939PhRv...56..284A;   http://prola.aps.org/abstract/
PR/v56/i3/p284_1.

Anderson:1942:ACSa

[AFW42] Herbert L. Anderson, Enrico Fermi, and George L. Weil. Absorption cross-sections for Rn plus Be fast neutrons. Report C-72, US Atomic Energy Commission, Washington, DC, USA, 1942. Undated: year chosen according to report number.

Anderson:1947:MMN

[AFW^{+}47] Herbert L. Anderson, Enrico Fermi, A. Wattenberg, George L. Weil, and Walter H. Zinn. Method for measuring neutronabsorption cross sections by the effect on the reactivity of a chainreacting pile. Physical Review, 72(1):16–23, July 1, 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v72/i1/p16_1.

Anderson:1957:MMN

[AFW^{+}57] Herbert L. Anderson, Enrico Fermi, Albedo Wattenberg, George L. Weil, and Walter H. Zinn. Method for measuring neutron-absorption cross sections by the effect on the reactivity of a chain-reacting pile. Physical Review, 72(1):16–23, July 1957. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v72/
i1/p16_1.

Albarracin:2014:DPS

[AG14] Lluís Albarracín and Núria Gorgorió. Devising a plan to solve Fermi problems involving large numbers. Educational Studies in Mathematics, 86(1):79–96, May 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:

//link.springer.com/accesspage/article/10.1007/s10649-
013-9528-9.

Agnew:1982:EI

[Agn82] Harold M. Agnew. Early impressions. Bulletin of the Atomic Scientists, 38(10):20–21, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See erratum [Ano83] that corrects the photo caption: it shows the first nuclear pile after reassembly at Argonne National Laboratory, not the original pile at Stagg Field.

Agnew:2003:DFL

[Agn03] Harold Agnew. Documents on Fermi’s life. In Anonymous [Ano03], pages 289–294. ISBN 88-8286-032-9. LCCN

???? URL http:old.enea.it/produzione_scientifica/pdf.
../V2003_AttiFermi.pdf.

Agnew:2004:FCA

[Agn04] Harold Agnew. Fermi at Columbia, Los Alamos, and Chicago. In Orear [Ore04], chapter 21, pages 101–106. ISBN ???? LCCN ????

URL http://dspace.library.cornell.edu/handle/1813/62;
 http://hdl.handle.net/1813/74.

Acocella:2004:EFD

[AGR04] Giovanni Acocella, Francesco Guerra, and Nadia Robotti. Enrico Fermi’s discovery of neutron-induced artificial radioactivity: The recovery of his first laboratory notebook. Physics in Perspective (PIP), 6(1):29–41, April 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://www.

springerlink.com/content/5qgmvqy95d3t33ma/.

Al-Khalili:2016:ASA

[AK16] Jim Al-Khalili, editor. Aliens: Science Asks: Is There Anyone Out There? Profile Books, London, UK, 2016. ISBN 1-78125-681-0 (paperback), 1-78283-271-8 (e-book). 148 pp. LCCN QB54 .A45 2016.

Al-Khalili:2017:AWL

[AK17] Jim Al-Khalili, editor. Aliens: the world’s leading scientists on the search for extraterrestrial life. Picador, New York, NY, USA, 2017. ISBN 1-250-10963-9 (hardcover), 1-250-10965-5. 232 pp. LCCN QB54 .A395 2017.

Allison:1957:EF

[All57] Samuel K. Allison. Enrico Fermi 1901–1954. Biographical memoirs — National Academy of Sciences of the United States of America, 30(6):123–155, ???? 1957. CODEN BMNSAC.

ISSN 0077-2933. URL http://books.nap.edu/html/biomems/
efermi.pdf;   http://lccn.loc.gov/57002459.

Allison:1962:BRC

[All62a] Samuel K. Allison. Book review: The Collected Papers of Enrico Fermi. vol. 1. Italy, 1921–1938 by Enrico Fermi and Emilo Segrè. American Scientist, 50(4):424A–424A, ???? 1962. CO-DEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic).

URL http://www.jstor.org/stable/27838557. See [Fer62a].

Allison:1962:ICR

[All62b] Samuel K. Allison. Initiation of the chain reaction: The search for pure materials. International Atomic Energy Agency Bulletin, 4(0):12–14, ???? 1962. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-2690 (electronic). URL http://

en.wikipedia.org/wiki/Atlas_Uranium_Mill;   http://en.
wikipedia.org/wiki/Moab,_Utah;   http://en.wikipedia.
org/wiki/Moab_Tailings;   http://www.iaea.org/Publications/
Magazines/Bulletin/Bull040su/04004701214su.pdf.

Allibone:1963:IEF

[All63] T. E. Allibone. Inspirations of Enrico Fermi. Nature, 200(4910): 961–962, December 7, 1963. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v200/n4910/pdf/200961a0.pdf.

Alpher:2023:GGR

[AM23] Victor S. Alpher and Simon A. Mitton. George Gamow and Ralph Alpher: a review of their cosmological collaboration as mentor and protégé 1942–1955. European Physical Journal H, 48(1):??, December 2023. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL https://link.springer.com/article/
10.1140/epjh/s13129-023-00057-0.

Amaldi:1956:GPO

[Ama56] Edoardo Amaldi. George Placzek [obituary]. La Ricerca Scientifica, 26(??):2038–??, ???? 1956. CODEN RISCAZ. ISSN 0035-5011. Placzek had worked closely with Fermi’s group in Rome, especially on experiments on the Raman effect; see [Seg70b, page 60].

Amaldi:1959:FMF

[Ama59] Edoardo Amaldi. The Fermi manuscripts at the ^{≪}Domus Galiaeana^{≫}. Physis: Rivista Internazionale di Storia della Scienza, 1(2):69–72, 1959. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Amaldi:1961:MAI

[Ama61] Ginestra (Giovene) Amaldi. Materia e antimateria. (Italian) [Matter and antimatter]. A. Mondadori, Milano, Italy, 1961. 439 pp. LCCN ????

Amaldi:1966:NMP

[Ama66] Ginestra (Giovene) Amaldi. The nature of matter: physical theory from Thales to modern times. Allen and Unwin, London, UK, 1966. 332 pp. LCCN QC173 .A413.

Amaldi:1973:BRE

[Ama73] Edoardo Amaldi. Book review: Enrico Fermi, Physicist, by Emilio Segré. Physics Today, 26(6):58–60, July 1973. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://www.physicstoday.org/resource/1/
phtoad/v26/i6/p58_s1.

Amaldi:1979:YR

[Ama79] Edoardo Amaldi. The years of reconstruction. In Schaerf and Amaldi [SA79], pages 379–461. ISBN 3-7186-0007-2. LCCN QC770 .C748 1979.

Amaldi:1984:DND

[Ama84] Edoardo Amaldi. From the discovery of the neutron to the discovery of nuclear fission. Physics Reports, 111(1–4):1–331, September 1984. CODEN PRPLCM. ISSN 0370-1573 (print), 1873-6270

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037015738490214X.

Amaldi:1987:FDS

[Ama87] E. Amaldi. The Fermi–Dirac statistics and the statistics of nuclei. In García Doncel et al. [GHMP87], pages 253–277. ISBN 84-7488-148-X. LCCN QC174.17.S9 I57 1983.

Amaldi:2002:CDS

[Ama02a] Edoardo Amaldi. Commemorazione del socio Enrico Fermi. (Italian) [Commemoration of the associate Enrico Fermi ]. In Conoscere Fermi nel centenario della nascita: 29 settembre 1901– 2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901–2001] [Ber02a], pages 23–36. ISBN 88-7438-000-3. LCCN QC774.F4. 10.33 EUR. URL http://

prometeo.sif.it:8080/libri/fermi/17.pdf. English transla-

tion in [BB04b].

Amaldi:2002:FND

[Ama02b] Ugo Amaldi. La fisica dei nuclei dagli anni trenta ai giorni nostri. (Italian) [The physics of nuclei from the Thirties to the present day]. In Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901–2001] [Ber02a], pages 152–177. ISBN 88-7438-000-3. LCCN QC774.F4. 10.33 EUR. URL http:

//prometeo.sif.it:8080/libri/fermi/10.pdf. English trans-

lation in [BB04b].

Amaldi:2003:SNP

[Ama03] Ugo Amaldi. Slow neutrons at Via Panisperna: the discovery, the production of isotopes and the birth of nuclear medicine. In Anonymous [Ano03], pages 145–168. ISBN 88-8286-032-9. LCCN

???? URL http:old.enea.it/produzione_scientifica/pdf.
../V2003_AttiFermi.pdf.

Amati:2019:MEF

[AMS19] Graziano Amati, Hugues Meyer, and Tanja Schilling. Memory effects in the Fermi–Pasta–Ulam model. Journal of Statistical Physics, 174(1):219–257, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Anderson:1933:PE

[And33] Carl D. Anderson. The positive electron. Physical Review, 43(6): 491–494, March 15, 1933. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/ abstract/PR/v43/i6/p491_1. This paper reports the discovery of the positron, for which Anderson received the 1936 Nobel Prize in Physics.

Anderson:1955:MEE

[And55] Herbert L. Anderson. Meson experiments with Enrico Fermi. Reviews of Modern Physics, 27(3):269–272, July 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
27.269;   http://rmp.aps.org/abstract/RMP/v27/i3/p269_
1.

Anderson:1973:TQA

[And73a] Herbert Anderson. Three questions about the sustained nuclear chain reaction. The University of Chicago Magazine, 65(5):3– 7, March/April 1973. ISSN 0041-9508. URL http://library.

ucsd.edu/dc/object/bb3960604v.

Anderson:1973:EDC

[And73b] Herbert L. Anderson. Early days of the chain reaction. Bulletin of the Atomic Scientists, 29(4):8–12, April 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anderson:1974:AOTb

[And74a] Herbert L. Anderson. “All in our time”: Fermi, Szilard, and Trinity. Bulletin of the Atomic Scientists, 30(8):40–47, October 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also part 1 [And74b].

Anderson:1974:AOTa

[And74b] Herbert L. Anderson. “All in our time”: The legacy of Fermi and Szilard. Bulletin of the Atomic Scientists, 30(7):56–62, September 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Continued in [And74a].

Andrews:1999:QFB

[And99] Michael R. Andrews. Quantum fluids: Bose gases and their Fermi cousins. Nature, 398(6724):195–198, March 18, 1999. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v398/
n6724/full/398195a0.html.

Anonymous:1920:PBA

[Ano20] Anonymous. Physics at the British Association. Nature, 106 (2663):357–358, November 11, 1920. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v106/n2663/pdf/106357a0.pdf.

From this meeting report: “The results thus show that the elements may be considered as being composed of these hydrogen nuclei, or ‘protons’ as Sir Ernest Rutherford would have us call them, . . ..” It is believed that this is the first published mention of the word proton.

Anonymous:1934:DAE

[Ano34a] Anonymous. Dozen artificial elements. New York Times, ??(??): 25, June 5, 1934. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/101252432/.

Anonymous:1934:FMS

[Ano34b] Anonymous. Fermi measures speed of neutron: Says it pierces atom nucleus in a thousand million million millionths of a second. New helium produced. Oliphant tells the Cambridge session of variety with atomic weight of 6. New York Times, ??(??):25, October 5, 1934. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/101010661/.

Anonymous:1934:IPE

[Ano34c] Anonymous. Italian produces 93d element by bombarding uranium: Professor Enrico Fermi, academician, uses neutrons formed by decomposition of berylium under the action of alpha particles of radium. New York Times, ??(??):??, June 5, 1934. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL https://search.proquest.com/hnpnewyorktimes/
docview/101250532/.

Anonymous:1934:NEF

[Ano34d] Anonymous. New element found, has atomic count 93: Famous pitchblende deposits yield substance with higher number than uranium. (German) []. New York Times, ??(??):19, July 13, 1934. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL https://search.proquest.com/hnpnewyorktimes/
docview/101123324/.

Anonymous:1934:NHE

[Ano34e] Anonymous. New heavy element created [element 93: neptunium]. The Times, ??(??):??, June 4, 1934.

Anonymous:1934:SE

[Ano34f] Anonymous. Science: 93rd element? Time, ??(??):??, June 18, 1934. CODEN TYMEA9. ISSN 0040-781x (print), 2169-

1665 (electronic). URL http://www.time.com/time/magazine/
article/0,9171,747508,00.html. News story on the possible

discovery of the first transuranic element, later known as neptunium (atomic symbol Np).

Anonymous:1938:PEF

[Ano38a] Anonymous. Prof. Enrico Fermi. Nature, 142(3603):906–907, November 19, 1938. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v142/n3603/pdf/142906c0.pdf.

Anonymous:1938:SNM

[Ano38b] Anonymous. Science: Neutron man. Time, ??(??):??, November 21, 1938. CODEN TYMEA9. ISSN 0040-781x (print), 2169-

1665 (electronic). URL http://www.time.com/time/magazine/
article/0,9171,760318,00.html. News of the award of the

1938 Nobel Prize for Physics to Enrico Fermi.

Anonymous:1939:EFU

[Ano39a] Anonymous. 6 elements found in uranium atom: Physicists bare discovery of greatest amount of energy liberated thus far: Report widely hailed: Professors Bohr and Fermi, at Columbia meeting, tell of atomic ‘cannon ball’. New York Times, ??(??): 17, February 25, 1939. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/102952660/.

Anonymous:1939:AEF

[Ano39b] Anonymous. Atom explosion frees 200,000,000 volts; new physics phenomenon credited to Hahn. New York Times, ?? (??):2, January 29, 1939. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/102763891. From the article:

“American scientists heard today of a new phenomenon in physics — explosion of atoms with a discharge of 200,000,000 volts of energy. … Dr. Enrico Fermi of the University of Rome told yesterday that this had been accomplished by Dr. G. [sic] Hahn of Berlin. … Scientists at the meeting said the discovery was comparable in significance to the original discovery of radioactivity thirty years ago.”.

Anonymous:1939:AER

[Ano39c] Anonymous. Atomic energy released. The Science News-Letter, 35(6):??, February 11, 1939. CODEN SNLEAI. ISSN 0096-4018 (print), 2326-1285 (electronic). URL http://www.sciencenews.

org/view/feature/id/199036/title/Atomic_Energy_Released.

According to [BHT86, page 212, column 1], this news story “appears to be the first connection between fission and the possibility of explosives made in print.” Flügge’s scientific article on that topic in Naturwissenschaften [Flü39] did appear until June 1939.

Anonymous:1939:VEF

[Ano39d] Anonymous. Vast energy freed by uranium atom: Split, it produces 2 ‘cannonballs’, each of 100,000,000 electron volts: Hailed as epoch making; new process, announced at Columbia, uses only 1–30 volt to liberate big force. New York Times, ??(??): 18, January 31, 1939. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/102759255/.

Anonymous:1941:MMU

[Ano41a] Anonymous. Minutes of the meeting of the Uranium Advisory Committee of the National Academy on Atomic Fission at Schenectady on October 21, 1941. Report, US National Academy of Sciences, Washington, DC, USA, 1941.

Anonymous:1941:RNA

[Ano41b] Anonymous. Report of the National Academy of Sciences Committee on Atomic Fission. Report, US National Academy of Sciences, Washington, DC, USA, July 11, 1941.

Anomalous:1946:ZWF

[Ano46a] Anomalous. Zip out: World’s first uranium pile. Time, 48(??): 67–??, December 9, 1946. CODEN TYMEA9. ISSN 0040-781x (print), 2169-1665 (electronic).

Anonymous:1946:SZ

[Ano46b] Anonymous. Science: Zip out. Time, ??(??):??, December 9, 1946. CODEN TYMEA9. ISSN 0040-781x (print), 2169-1665 (elec-

tronic). URL http://www.time.com/time/magazine/article/

0,9171,793263,00.html. News story of the first successful controlled nuclear reaction in Chicago under the direction of Enrico Fermi on 2 December 1942.

Anonymous:1947:EFH

[Ano47] Anonymous. Enrico Fermi, head-and-shoulders portrait, facing

front, 1947. URL http://hdl.loc.gov/loc.pnp/cph.3c22125.

Anonymous:1949:GIT

[Ano49] Anonymous. The great inquiry: Testimony at AEC hearings. Bulletin of the Atomic Scientists, 5(8–9):221–250, 254, August/ September 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). This report is a condensation of the 2000-page transcript.

Anonymous:1950:GBM

[Ano50a] Anonymous. Great Britain: Missing fissionist. Time, ??(??):??, November 6, 1950. CODEN TYMEA9. ISSN 0040-781x (print),

2169-1665 (electronic). URL http://en.wikipedia.org/wiki/
Bruno_Pontecorvo;   http://www.time.com/time/magazine/
article/0,9171,813688,00.html. News story of the disappear-

ance on 31 August 1950 of Enrico Fermi’s former student, Bruno Pontecorvo. Several months later, Pontecorvo appeared publicly in the USSR.

Anonymous:1950:NFA

[Ano50b] Anonymous. Nobelist Fermi awarded patent on atomic device. The Science News-Letter, 58(20):311, November 11, 1950. CO-DEN SNLEAI. ISSN 0096-4018 (print), 2326-1285 (electronic).

URL http://www.jstor.org/stable/3927761.

Anonymous:1951:EFS

[Ano51] Anonymous. Enrico Fermi seated at control panel of a particle accelerator, the “world’s most powerful atom smasher”, 1951. URL

http://hdl.loc.gov/loc.pnp/cph.3c20917.

Anonymous:1952:SFD

[Ano52] Anonymous. Science: First decade. Time, ??(??):??, December 15, 1952. CODEN TYMEA9. ISSN 0040-781x (print), 2169-

1665 (electronic). URL http://www.time.com/time/magazine/
article/0,9171,820510,00.html. Commentary on the 10th

anniversary of the first successful controlled nuclear reaction in Chicago under the direction of Enrico Fermi on 2 December 1942.

Anonymous:1953:APP

[Ano53a] Anonymous. Atomic patent payoff: Italian scientists who derived isotopes are compensated. New York Times, ??(??):E7, August 9, 1953. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/112618167/.

Anonymous:1953:SAP

[Ano53b] Anonymous. Science: Atomic patent. Time, ??(??):??, August 10, 1953. CODEN TYMEA9. ISSN 0040-781x (print), 2169-

1665 (electronic). URL http://www.time.com/time/magazine/
article/0,9171,818653,00.html. News story on a 1935 patent

application by Enrico Fermi and his colleagues.

Anonymous:1954:AHB

[Ano54a] Anonymous. The atom: The H-bomb delay. Time, ??(??): ??, February 8, 1954. CODEN TYMEA9. ISSN 0040-781x (print), 2169-1665 (electronic). URL http://www.time.com/

time/magazine/article/0,9171,857625,00.html. News story

on the history of the development of the first thermonuclear weapon.

Anonymous:1954:MD

[Ano54b] Anonymous. Milestones, Dec. 6, 1954. Time, ??(??):??, December 6, 1954. CODEN TYMEA9. ISSN 0040-781x (print), 2169-

1665 (electronic). URL http://www.time.com/time/magazine/
article/0,9171,820990,00.html. From the article: “Died. En-

rico Fermi, 53, world-famed Italian-born nuclear physicist who supervised the building of the first successful nuclear reactor; of cancer; in Chicago.”.

Anonymous:1954:SDN

[Ano54c] Anonymous. Science: Death of a navigator. Time, ??(??): ??, December 6, 1954. CODEN TYMEA9. ISSN 0040-781x (print), 2169-1665 (electronic). URL http://www.time.com/

time/magazine/article/0,9171,820969,00.html.

Anonymous:1954:SLF

[Ano54d] Anonymous. Science: Life with Fermi. Time, ??(??):??, October 18, 1954. CODEN TYMEA9. ISSN 0040-781x (print), 2169-

1665 (electronic). URL http://www.time.com/time/magazine/
article/0,9171,936557,00.html. Book review of [Fer54k].

Anonymous:1955:FP

[Ano55a] Anonymous. Fermi Prize. Bulletin of the Atomic Scientists, 11 (2):70, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). The announcement begins: “The first AEC award of $25,000 for especially meritorious contribution to the development of atomic energy, provided for in the Atomic Energy Act of 1954, was given to Enrico Fermi, November 16, 1954, twelve days before his death.” It end with “Subsequent awards will be known as the “Fermi Prize”.

Anonymous:1955:PIF

[Ano55b] Anonymous. Patent is issued on first reactor; Fermi–Szilard invention gets recognition — A. E. C. holds ownership. New York Times, page 19, May 19, 1955. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/hnpnewyorktimes/docview/113334584/. See en-

try [FS55] for the patent.

Anonymous:1955:PAI

[Ano55c] Anonymous. Pioneer atom inventions receive patents. The Science News-Letter, 68(9):134, August 27, 1955. CODEN SNLEAI. ISSN 0096-4018 (print), 2326-1285 (electronic). URL http://www. jstor.org/stable/3935856. Discusses US patents 2,714,577 and 2,714,668, applied for in 1945, but not awarded for a decade, and after Fermi’s death.

Anonymous:1956:EFH

[Ano56a] Anonymous. Enrico Fermi and his contributions to physics. Nature, 177(4513):781, April 28, 1956. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v177/n4513/pdf/177781a0.pdf.

Anonymous:1956:SRN

[Ano56b] Anonymous. Science: The real neutrino. Time, ??(??):??, July 2, 1956. CODEN TYMEA9. ISSN 0040-781x (print), 2169-

1665 (electronic). URL http://www.time.com/time/magazine/
article/0,9171,891295,00.html. From the story: “[Neutrinos]

were reasoned into existence by Nobel Prizewinners Enrico Fermi and Wolfgang Pauli to fill a theoretical need, and the gnawing suspicion has long persisted that they do not exist. Last week from the Atomic Energy Commission came big news. Neutrinos do exist”.

Anonymous:1957:FP

[Ano57] Anonymous. Fermi professorship. Science, 125(3244):391, March 1, 1957. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1753499.

Anonymous:1958:PWG

[Ano58] Anonymous. Princeton’s Wigner gets third Enrico Fermi Award. The Science News-Letter, 74(24):372, December 13, 1958. CO-DEN SNLEAI. ISSN 0096-4018 (print), 2326-1285 (electronic).

URL http://www.jstor.org/stable/3939817.

Anonymous:1959:SWF

[Ano59] Anonymous. Seaborg wins Fermi Award. Science, 130(3387): 1465, November 27, 1959. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/

stable/1757520.

Anonymous:1960:MMS

[Ano60] Anonymous. Metal model shows Fermi surface. The Science News-Letter, 77(26):410, June 25, 1960. CODEN SNLEAI. ISSN 0096-4018 (print), 2326-1285 (electronic). URL http://www.

jstor.org/stable/3941940.

Anonymous:1961:EFA

[Ano61] Anonymous. Enrico Fermi Award [to Hans A. Bethe]. Physics Today, 14(10):74, October 1961. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v14/i10/p74_s1.

Anonymous:1962:DTR

[Ano62] Anonymous. Dr. Teller receives $50,000 Enrico Fermi Award. New York Times, ??(??):??, December 4, 1962. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Anonymous:1963:FAJ

[Ano63a] Anonymous. The [1963] Fermi Award [to J. Robert Oppenheimer]. Physics Today, 16(6):21–23, June 1963. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:1963:FPJ

[Ano63b] Anonymous. Fermi Prize: J. Robert Oppenheimer named to receive annual AEC Award. Science, 140(3563):161–163, April 12, 1963. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.jstor.org/stable/1710958.

Anonymous:1964:EFP

[Ano64] Anonymous. Enrico Fermi Professorship of Physics, Columbia University: Prof. Tsung-Dao Lee. Nature, 204(4961):827, November 28, 1964. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v204/n4961/pdf/204827b0.pdf.

Anonymous:1965:FAR

[Ano65a] Anonymous. Fermi Award to Robert Wilson. Bulletin of the Atomic Scientists, 21(2):57, February 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1965:NLI

[Ano65b] Anonymous, editor. Nobel lectures: including presentation speeches and laureates’ biographies: physics 1922–1941. Elsevier, Amsterdam, The Netherlands, 1965. x + 456 pp.

LCCN QC71 .P455 1965. URL http://nobelprize.org/
nobel_prizes/physics/laureates/1922/bohr-lecture.html;
 http://nobelprize.org/nobel_prizes/physics/laureates/
1922/bohr-lecture.pdf.

Anonymous:1966:NBF

[Ano66] Anonymous. News in brief: Fermi Award. Science, 153(3737): 723, August 12, 1966. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Anonymous:1967:APC

[Ano67a] Anonymous. Atomic power: Coming of age. Time, ??(??): ??, December 8, 1967. CODEN TYMEA9. ISSN 0040-781x (print), 2169-1665 (electronic). URL http://www.time.com/

time/magazine/article/0,9171,844231,00.html. Commen-

tary on the 25th anniversary of the first successful controlled nuclear reaction in Chicago under the direction of Enrico Fermi on 2 December 1942.

Anonymous:1967:ALM

[Ano67b] Anonymous. Award for Lise Meitner. International Atomic Energy Agency Bulletin, 9(2):17, ???? 1967. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-2690 (electronic). URL http:/

/www.iaea.org/Publications/Magazines/Bulletin/Bull092/

09201401717.pdf. Enrico Fermi Award for 1966.

Anonymous:1967:WAA

[Ano67c] Anonymous. When the Atomic Age began. International Atomic Energy Agency Bulletin, 9(6):3, ???? 1967. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-2690 (electronic). URL http:/

/www.iaea.org/Publications/Magazines/Bulletin/Bull096/
09605000303.pdf.

Anonymous:1968:DFP

[Ano68a] Anonymous. Distinctions: Fermi Prize for Wheeler. Nature, 220 (5162):4, October 5, 1968. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v220/n5162/pdf/220004b0.pdf.

Anonymous:1968:EFA

[Ano68b] Anonymous. The Enrico Fermi Award, 1968 [to John Archibald Wheeler]. Bulletin of the Atomic Scientists, 24(9):40, November 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1968:JWR

[Ano68c] Anonymous. John Wheeler receives AEC 1968 Enrico Fermi Award. Physics Today, 21(12):105, December 1968. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://www.physicstoday.org/resource/1/
phtoad/v21/i12/p105_s1.

Anonymous:1968:WCR

[Ano68d] Anonymous. When the chain reaction started. International Atomic Energy Agency Bulletin, 10(6):32–35, ???? 1968. CO-DEN IAEBAB. ISSN 0020-6067 (print), 1564-2690 (electronic).

Anonymous:1969:FA

[Ano69] Anonymous. Fermi Accelerator. Nature, 222(5195):718, May 24, 1969. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v222/n5195/pdf/222718a0.pdf.

Anonymous:1970:WHZ

[Ano70] Anonymous. Walter H. Zinn receives AEC Enrico Fermi Medal. Physics Today, 23(2):99, February 1970. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

www.physicstoday.org/resource/1/phtoad/v23/i2/p99_s3.

Anonymous:1973:BRE

[Ano73] Anonymous. Benedict receives Enrico Fermi Award. Physics Today, 26(1):103, January 1973. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v26/i1/p103_s2.

Anonymous:1974:PBF

[Ano74] Anonymous. Proton beams and the Fermi Machine. Science News (Washington, DC), 105(19):300, May 11, 1974. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic). URL

http://www.jstor.org/stable/3958907.

Anonymous:1975:RAS

[Ano75] Anonymous. Reminiscences of Los Alamos, 1943–1945 [sound recording]. Los Alamos National Laboratory. Presented at Santa Barbara Lectures on Science and Society. 1975. 1st annual, University of California at Santa Barbara, January 16–March 13, 1975, 1975.

Anonymous:1979:FAG

[Ano79] Anonymous. Fermi Award given. Science News (Washington, DC), 116(8):132, August 25, 1979. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic). URL http://www.jstor.

org/stable/3964302.

Anonymous:1981:WPR

[Ano81] Anonymous. Weinberg and Peierls receive Enrico Fermi Award. Physics Today, 34(3):85, March 1981. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v34/i3/p85_s1.

Anonymous:1982:EFS

[Ano82] Anonymous. Enrico Fermi: significato di una scoperta: il navigatore italiano sbarcato nel nuovo mondo. (Italian) [Enrico Fermi: The Meaning of a Discovery. A Disembarked Italian Navigator in the New World]. Forum italiano dell’energia nucleare, Roma, Italy, 1982. 121 pp. LCCN QC16.F46 E57 1982.

Anonymous:1983:E

[Ano83] Anonymous. Erratum. Bulletin of the Atomic Scientists, 39(2): 41, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Agn82].

Anonymous:1990:FA

[Ano90] Anonymous. Fermi Award [to ????]. Science, 250(4980):503, October 26, 1990. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.sciencemag.org/content/
250/4980/503.3.full.pdf.

Anonymous:1993:BBF

[Ano93a] Anonymous. Bulletin: Bernard T. Feld. Bulletin of the Atomic Scientists, 49(3):5, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:BLW

[Ano93b] Anonymous. Bulletin: Lederman wins Fermi Award. Bulletin of the Atomic Scientists, 49(7):7, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:BNF

[Ano94a] Anonymous. Book note: The Fermi Solution: Essays on Science, by Hans Christian von Baeyer. Bulletin of the Atomic Scientists, 50(1):61, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:DRF

[Ano94b] Anonymous. Dyson receives 1994 Fermi Award. Physics Today, 47(11):104, 1994. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://link.aip.org/link/?PTO/
47/104/1.

Anonymous:1994:FS

[Ano94c] Anonymous. Fermi surfaces. Physics World, 7(8):10, August 1994. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (elec-

tronic). URL http://physicsworldarchive.iop.org/full/
pwa-pdf/7/8/phwv7i8a11.pdf.

Anonymous:1995:AAI

[Ano95a] Anonymous. 50th anniversary article: “The Italian Navigator has landed in the New World.” secret race won with Chicago’s chain reaction. Los Alamos Science, ??(??):??, ???? 1995. CO-DEN LASCDI. ISSN 0273-7116. URL http://www.lanl.gov/

history/road/chicago-reactor.shtml.

Anonymous:1995:BRAg

[Ano95b] Anonymous. Book review: Atoms in the family: My life with Enrico Fermi: By Laura Fermi. The University of Chicago Press, Chicago. (1954). 267 pages. $13.95. Computers and Mathematics with Applications, 30(9):124, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
089812219590221X.

Anonymous:1995:BRN

[Ano95c] Anonymous. Book review: Notes on quantum mechanics: a course given by Enrico Fermi (second edition): Compiled by Robert A. Schluter. University of Chicago Press, Chicago. (1995). 188 pages. $14.95. Computers and Mathematics with Applications, 30(9):122, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0898122195901957.

Anonymous:1996:FAH

[Ano96a] Anonymous. 1996 Fermi Award honours his former student [Ugo Fano]. Nature, 379(6560):3, January 4, 1996. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v379/
n6560/pdf/379003b0.pdf.

Anonymous:1996:BREf

[Ano96b] Anonymous. Book review: Enrico Fermi, physicist: By Emilio Segrè. University of Chicago Press, Chicago. (1970). 276 pages. $13.95, £11.25. Computers and Mathematics with Applications, 31(2):132, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0898122196901157.

Anonymous:1997:ITK

[Ano97] Anonymous. The IAEA turns 40: Key dates and historical developments. International Atomic Energy Agency Bulletin, 39 (3S):1–24, September 1997. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-2690 (electronic). URL http://www.iaea.org/

Publications/Magazines/Bulletin/Bull393/39301205578su.
pdf.

Anonymous:1998:THD

[Ano98] Anonymous, editor. Today in history. December 2. Library of Congress, Washington, DC, USA, 1998. LCCN QC173. URL

http://memory.loc.gov/ammem/today/dec02.html.

Anonymous:1999:CM

[Ano99] Anonymous. The common man. Physics in Perspective (PIP), 1(3):336, October 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050024;   http://
www.springerlink.com/content/y1h02gr3rv7wurrl/.

Anonymous:2001:EFS

[Ano01a] Anonymous. Enrico Fermi: significato di una scoperta. (Italian) [Enrico Fermi: the meaning of a discovery]. ENEA, Roma, Italy, revised and expanded edition, 2001. ISBN 88-8286-022-1. 184 pp. LCCN QC16.F46 E57 2001.

Anonymous:2001:LFR

[Ano01b] Anonymous. L’Istituto di fisica e i ragazzi di via Panisperna / Istituto d’istruzione superiore di via Salvini, già Istituto tecnico Margherita di Savoia. (Italian) [The Institute of Physics and Via Panisperna boys / Institute of Higher Education Via Salvini, former Technical Institute Margherita di Savoia], volume 104 of La scuola adotta un monumento. F. lli Palombi, Roma, Italy, 2001. ISBN ???? 39 pp. LCCN ????

Anonymous:2001:RFL

[Ano01c] Anonymous. Returning to the Fermi level. Physics World, 14(9):15, September 2001. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/14/9/phwv14i9a15.
pdf.

Anonymous:2002:BER

[Ano02] Anonymous. Bibliografia essenziale relativa ai saggi contenuti in questo volume. (Italian) [Bibliography relating to the essays in this volume]. In Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901–2001] [Ber02a], pages 378– ?? ISBN 88-7438-000-3. LCCN QC774.F4. 10.33 EUR. URL

http://prometeo.sif.it:8080/libri/fermi/biblio.pdf. En-

glish translation in [BB04b].

Anonymous:2003:PIC

[Ano03] Anonymous, editor. Proceedings of the International Conference “Enrico Fermi and the Universe of Physics”, Rome, September 29–October 2, 2001. Accademia Nazionale dei Lincei and Instituto Nazionale di Fisica Nucleare, Rome, Italy, 2003. ISBN 88-

8286-032-9. LCCN ???? URL http:old.enea.it/produzione_
scientifica/pdf.../V2003_AttiFermi.pdf.

Anonymous:2005:CNS

[Ano05] Anonymous, editor. The Center for Nonlinear Studies 25th Annual International Conference: 50 Years of the Fermi–Pasta– Ulam Problem: Legacy, Impact, and Beyond: May 16–20, 2005, Radisson Santa Fe Hotel, Santa Fe, New Mexico USA. ????, ????, 2005. ISBN ???? LCCN ???? URL http://cnls.lanl.gov/

Conferences/annual25/agenda.htm.

Anonymous:2006:TDF

[Ano06] Anonymous. Take a dip in the Fermi sea. Science, 311(5765):1223, March 3, 2006. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.sciencemag.org/content/
311/5765/1223.2.full.pdf.

Anonymous:2007:EFN

[Ano07] Anonymous. Enrico Fermi and the nuclear chain reaction — University of Chicago. Web archive, November 5, 2007. URL http:

//digitalcollections.wordpress.com/2007/11/05/enrico-
fermi-and-the-nuclear-chain-reaction-university-of-chicago/
;   http://fermi.lib.uchicago.edu/.

Anonymous:2009:BDR

[Ano09a] Anonymous. Buried defects reveal Fermi surfaces. Science, 323(5918):1143, February 27, 2009. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/323/5918/1143.13.full.pdf.

Anonymous:2009:FAS

[Ano09b] Anonymous. Fermi Award [to Sig Hecker and John Goodenough]. Science, 325(5948):1605, September 25, 2009. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/325/5948/1605.3.full.pdf.

Anonymous:2009:FSU

[Ano09c] Anonymous. Fermi surface unmasked. Physics World, 22(4): 4, April 2009. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/22/04/phwv22i04a6.pdf.

Anonymous:2009:GEF

[Ano09d] Anonymous. Guide to the Enrico Fermi Collection 1918–1974. University of Chicago Library Web site, 2009. URL http://www.

lib.uchicago.edu/e/scrc/findingaids/view.php?eadid=ICU.
SPCL.FERMI.

Anonymous:2010:WIE

[Ano10] Anonymous. Wer ist’s? — Enrico Fermi. (German) [Who is he? — Enrico Fermi]. Nachrichten aus Chemie und Technik, 3(8):84–85, April 21, 2010. CODEN NCHTAD. ISSN 0027-

738X. URL http://onlinelibrary.wiley.com/doi/10.1002/
nadc.19550030804/abstract.

Anonymous:2011:MPH

[Ano11] Anonymous. This month in physics history: December 2, 1942: First self-sustained nuclear chain reaction. APS Physics, 20(11):

??, December 2011. URL http://www.aps.org/publications/
apsnews/201112/physicshistory.cfm.

Anonymous:2012:EF

[Ano12a] Anonymous. Enrico Fermi. In Encyclopædia Britannica Online Academic Edition. Encyclopædia Britannica, Chicago, IL,

USA, 2012. LCCN ???? URL http://www.britannica.com/
EBchecked/topic/204747/Enrico-Fermi.

Anonymous:2012:HAT

[Ano12b] Anonymous. Harold Agnew talk delights audience. Los Alamos National Laboratory News Center, May 11, 2012. URL http:/

/www.lanl.gov/news/stories/agnew-colloquium.html. From

the article: “Agnew, 92, fondly recounted his long association with Enrico Fermi, starting with his work on the team that created the first controlled nuclear reaction in a graphite pile at the University of Chicago. He remembers Fermi as, “a wonderful person, but just a regular guy,” who, in a very low-key manner, advanced key research in the development of the first atomic weapons and many other breakthrough scientific innovations.

He also told a story of Leo Szilard’s quirky behavior.

“At Chicago, Fermi used to give weekly lectures on Thursday evenings. Outside the lecture hall there was a sign up sheet, on a yellow legal pad. Szilard, who I think wasn’t sure we were going to win the war, refused to put his name on the sign up sheet, so Fermi had him listen to the lectures from the hallway, through an open door.”.

Anonymous:2012:SCE

[Ano12c] Anonymous. Scientists closer to extracting uranium from seawater. Web news report., September 7, 2012. URL http://

www.scientificcomputing.com/news-DS-Scientists-Closer-
to-Extracting-Uranium-from-Seawater-082912.aspx.

Anonymous:2015:EF

[Ano15a] Anonymous. Enrico Fermi. Web site, 2015. URL https://www.

aip.org/history/acap/biographies/bio.jsp?fermie.

Anonymous:2015:MSY

[Ano15b] Anonymous. Mathematicians solve 60-year-old Fermi–Pasta– Ulam problem. Scientific Computing, ??(??):??, March 24, 2015. CODEN SCHRCU. ISSN 1930-5753 (print), 1930-6156 (elec-

tronic). URL http://www.scientificcomputing.com/news/
2015/03/mathematicians-solve-60-year-old-fermi-pasta-

ulam-problem. See [OVPL15] for the research paper.

Anonymous:2016:MTA

[Ano16] Anonymous. The march of time: Atomic power! Periscope Film LLC archive, April 17, 2016. URL https://www.youtube.com/

watch?v=iTJJ7tkD5L4.

Anonymous:2017:CBI

[Ano17] Anonymous. Churchill’s big idea. Physics World, 30(4):3, 2017. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (elec-

tronic). URL http://stacks.iop.org/2058-7058/30/i=4/a=1.

Anonymous:2022:FCP

[Ano22] Anonymous. Fermi on Chicago Pile-1. Atomic Heritage Founda-

tion Web site., 2022. URL https://www.atomicheritage.org/
key-documents/fermi-chicago-pile-1.

Amaldi:1996:ERC

[APB96] Edoardo Amaldi, Giovanni Paoloni, and Giovanni Battimelli, editors. Essays and recollections on 20th century physics: a selection of historical writings, volume 2 of Edoardo Amaldi Foundation series. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1996. ISBN 981-02-2369-2. xiii + 747 pp. LCCN Q127.I8 A53 1996.

Arblaster:2000:DPI

[Arb00] J. W. Arblaster. The discoverers of the platinum isotopes: The discovery of the thirty seven known platinum isotopes between 1935 and 1996. Platinum Metals Review, 44(4):173–178, October 2000. CODEN PTMRA3. ISSN 0032-1400. URL http://www.

technology.matthey.com/article/44/4/173-178/.

Arblaster:2003:DII

[Arb03] J. W. Arblaster. The discoverers of the iridium isotopes: The thirty-six known iridium isotopes found between 1934 and 2001. Platinum Metals Review, 47(4):167–174, ???? 2003. CODEN PTMRA3. ISSN 0032-1400. URL http://www.

technology.matthey.com/article/47/4/167-174/;   http:/
/www.technology.matthey.com/pdf/167-174-pmr-oct03.pdf.

Amelio:1996:RPS

[AS96] Gianni Amelio and Alessandro Sermoneta. I ragazzi di via Panisperna: sceneggiatura originale dell’omonimo film di Amelio. (Italian) [The boys of Via Panisperna: original screenplay of the film by Amelio], volume 34 of Cinema. Circolo del cinema, Mantova, Italy, 1996. ISBN ???? 199 pp. LCCN ????

Apotheker:2011:EWC

[AS11] Jan Apotheker and Livia Simon Sarkadi, editors. European women in chemistry. Wiley-VCH Verlag and Co. KGaA, Weinheim, Germany, 2011. ISBN 3-527-63645-5 (e-book). 239 pp. LCCN QD20

.E976 2011. URL http://site.ebrary.com/id/10484792.

Armstrong:2013:ESH

[AS13] Stuart Armstrong and Anders Sandberg. Eternity in six hours: Intergalactic spreading of intelligent life and sharpening the Fermi paradox. Acta Astronautica, 89:1–13, August 2013. CODEN AASTCF. ISSN 0094-5765 (print), 1879-2030

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0094576513001148.

Allison:1955:EFb

[ASA55a] S. K. Allison, E. Segrè, and Herbert L. Anderson. Enrico Fermi. Bulletin of the Atomic Scientists, 11(1):2, 40, January 1955. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1955:EFa

[ASA55b] S. K. Allison, Emilio Segré, and Herbert L. Anderson. Enrico Fermi 1901–1954. Physics Today, 8(1):9–13, January 1955. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://www.physicstoday.org/resource/1/phtoad/v8/
i1/p9_s1.

Allardice:1962:FPa

[AT62a] Corbin Allardice and Edward R. Trapnell. The first pile. International Atomic Energy Agency Bulletin, 4(0):41–47, ???? 1962. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-

2690 (electronic). URL http://www.iaea.org/Publications/
Magazines/Bulletin/Bull040su/04005004147su.pdf.

Allardice:1962:FPb

[AT62b] Corbin Allardice and Edward R. Trapnell. The first pile. Bulletin of the Atomic Scientists, 18(10):19–24, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allardice:1982:AFR

[ATFF82] Corbin Allardice, Edward R. Trapnell, Enrico Fermi, and Laura Fermi. 40th anniversary: The first reactor. Report DOE/NE-0046, United States Department of Energy, Washington, DC,

USA, December 1982.  URL http://local.ans.org/mi/
Teacher_CD/Historical%20Info/DOE-NE-0046.pdf;   http:/
/www.osti.gov/accomplishments/documents/fullText/ACC0044.
pdf.

Aoki:1995:PTF

[AYS95] K. Aoki, H. Yamawaki, and M. Sakashita. Pressure-tuned Fermi resonance in ice VII. Science, 268(5215):1322–1324, June 2, 1995. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.jstor.org/stable/2888771.

Bilenkij:2013:BPS

[B^{+}13] Samoil Mihelevic Bilenkij et al., editors. Bruno Pontecorvo selected scientific works, recollections on Bruno Pontecorvo. Società italiana di fisica, Bologna, Italia, 2013. ISBN 88-7438-080-1.

xxxi + 615 pp. LCCN ???? URL http://www.sif.it/libri/
pontecorvo_2nd/contents.

Badash:1967:BRP

[Bad67] Lawrence Badash. Book review: Physics for the general reader: The Nature of Matter. Physical Theory from Thales to Fermi, by Ginestra Amaldi and Peter Astbury. Science, 155(3764):818, February 17, 1967. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/

1720247. See [Bad67].

Badash:1971:BRE

[Bad71] Lawrence Badash. Book review: Enrico Fermi, Physicist by Emilio Segrè. Isis, 62(4):561–562, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229863.

Badash:2005:APN

[Bad05] Lawrence Badash. American physicists, nuclear weapons in World War II, and social responsibility. Physics in Perspective (PIP), 7(2):138–149, June 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://www.

springerlink.com/content/u2638016081u076h/.

Bain:1955:EF

[Bai55] Read Bain. Enrico Fermi: 1901–1954. The Scientific Monthly, 80(2):92, February 1955. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL http://www.jstor.org/

stable/21365.

Bailey:2017:EBY

[Bai17a] David H. Bailey. Exoplanets, 4 billion-year-old life, Fermi’s paradox and zero–one laws. Web essay., March 3, 2017.

URL http://experimentalmath.org/2017/03/exoplanets-4-
billion-year-old-life-fermis-paradox-and-zero-one-laws/

.

Bailey:2017:FTF

[Bai17b] David H. Bailey. Fine tuning and Fermi’s paradox: A “freakishly” fine-tuned universe. Web essay., November 22, 2017. URL http:

//mathscholar.org/fine-tuning-and-fermis-paradox.

Bailey:2018:FPC

[Bai18] David H. Bailey. Fermi’s paradox and the Copernican principle. MathScholar blog., June 28, 2018. URL http://mathscholar.

org/2018/06/fermis-paradox-and-the-copernican-principle/

.

Baker:1976:ABG

[Bak76] Paul R. Baker, editor. The atomic bomb: the great decision. American problem studies. Dryden Press, Hinsdale, IL, USA, second revised edition, 1976. ISBN 0-03-089873-0. viii + 193 pp. LCCN D842 .B34 1976.

Balazs:1963:BRE

[Bal63] Nandor L. Balazs. Book review: Enrico Fermi: Collected Papers. Physics Today, 16(4):74–76, April 1963. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/?PTO/16/74/1;   http://link.aip.org/link/
phtoad/v16/i4/p74/s1.

Ball:1973:ZH

[Bal73] John A. Ball. The zoo hypothesis. Icarus: International Journal of Solar System Studies, 19(??):347–349, July 1973. CODEN ICRSA5. ISSN 0019-1035 (print), 1090-2643 (electronic). URL

http://en.wikipedia.org/wiki/Zoo_hypothesis.

Bankston:2003:EFN

[Ban03] John Bankston, editor. Enrico Fermi and the nuclear reactor. Unlocking the secrets of science. Mitchell Lane Publishers, Bear, DE, USA, 2003. ISBN 1-58415-184-6. 48 pp. LCCN QC16.F46 B36 2003.

Bankston:2004:LMA

[Ban04] John Bankston. Lise Meitner and the Atomic Age. Unlocking the secrets of science. Mitchell Lane Publishers, Bear, DE, USA, 2004. ISBN 1-58415-206-0. 48 pp. LCCN QC774.M4 B36 2004.

Barinaga:1987:FAA

[Bar87] Marcia Barinaga. Fermi Awards for atomic energy advances. Nature, 330(6145):201, November 19, 1987. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v330/n6145/pdf/330201b0.pdf.

Barron:2000:LMD

[Bar00] Rachel Stiffler Barron. Lise Meitner: discoverer of nuclear fission. Great scientists. Morgan Reynolds, Greensboro, NC, USA, 2000. ISBN 1-883846-52-8. 112 pp. LCCN QC774.M4 B37 2000.

Bartholomew:2005:ERO

[Bar05] James R. Bartholomew. Essay review: One hundred years of the Nobel Science Prizes: Elisabeth Crawford (Editor). Historical Studies in the Nobel Archives: The Prizes in Science and Medicine. viii + 161 pp., Tokyo: Universal Academy Press, 2002. Elisabeth Crawford. The Nobel Population, 1901–1950: A Census of the Nominators and Nominees for the Prizes in Physics and Chemistry. vi + 420 pp., Tokyo: Universal Academy Press, 2002. Mauro Dardo. Nobel Laureates and Twentieth-Century Physics. x + 515 pp., Cambridge: Cambridge University Press, 2004. Robert Marc Friedman. The Politics of Excellence: Behind the Nobel Prize in Science. xv + 400 pp., notes, index. New York: W. H. Freeman, 2001. István Hargittai. The Road to Stockholm: Nobel Prizes, Science, and Scientists. xvii + 342 pp., Oxford/New York: Oxford University Press, 2002. George Thomas Kurian. The Nobel Scientists: A Biographical Encyclopedia. 675 pp., Amherst, N.Y.: Prometheus Books, 2002. Isis, 96(4):625–632, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/10.
1086/498605;   http://www.jstor.org/stable/3652242. See

[CHU87, Cra92, Fri01, Cra02, Har02, Kur02, Dar04].

Bassani:2001:CCD

[Bas01] G. Franco Bassani. Celebrazioni del centenario della nascita di Enrico Fermi: Interventi della giornata dedicata a: ≪Fermi, maestro e didatta≫. (Italian) [celebrations of the centenary of the birth of Enrico Fermi: Speeches of the day dedicated to: ≪Fermi, teacher and trainer≫ ]. Il Nuovo Saggiatore, 17(5–6):5–47, September/December 2001. ISSN 0393-4578 (print), 1827-6148 (elec-

tronic). URL http://prometeo.sif.it:8080/papers/online/

sag/2001/05-06/pdf/05.pdf. Special issue in honor of the centennial of the birth of Enrico Fermi.

Bassani:2002:EFF

[Bas02] Franco Bassani. Enrico Fermi e la fisica dello stato solido. (Italian) [Enrico Fermi and solid-state physics]. In Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901–2001] [Ber02a], pages 57–67. ISBN 88-7438-000-3. LCCN

QC774.F4. 10.33 EUR. URL http://prometeo.sif.it:8080/

libri/fermi/03.pdf. English translation in [BB04b].

Battimelli:2001:EFG

[Bat01] Gianni Battimelli. Enrico Fermi: genius and giant of science. CERN Courier: International Journal of High-Energy Physics, 41(7):26–29, September 2001. CODEN CECOA2. ISSN 0304-288X (print), 2077-9550 (electronic). URL https://cds.cern.

ch/record/1733227.

Battimelli:2003:FFP

[Bat03a] Giovanni Battimelli. Funds and failures: the political economy of Fermi’s group. In Anonymous [Ano03], pages 169–184. ISBN 88-

8286-032-9. LCCN ???? URL http:old.enea.it/produzione_
scientifica/pdf.../V2003_AttiFermi.pdf.

Battimelli:2003:LFS

[Bat03b] Giovanni Battimelli. L’eredità di Fermi: storia fotografica dal 1927 al 1959 dagli archivi di Edoardo Amaldi. (Italian) [The legacy of Fermi: photographic history from 1927 to 1959 from the archives of Edoardo Amaldi]. Editori riuniti, Roma, Italia, 2003. ISBN 88-359-5428-2. 220 pp. LCCN Q127.I8 B24 2003.

Battimelli:2006:BRJ

[Bat06] Giovanni Battimelli. Book review: James W. Cronin (ed.), Fermi Remembered. Chicago: University of Chicago Press, 2004. xii + 287 pp., ill., ISBN 0-226-12111-9. Nuncius, 21(1):185–186, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539106x00357.

Baugher:1985:CSS

[Bau85] Joseph F. Baugher. On Civilized Stars: the Search for Intelligent Life in Outer Space. A spectrum book; Frontiers of science. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1985. ISBN 0-13-634411-9 (paperback), 0-13-634429-1 (hardcover). xi + 260 pp. LCCN QB54 .B38 1985.

Baxter:1946:SAT

[Bax46] James Phinney Baxter, 3rd. Scientists against time. Little, Brown and company, Boston, 1946. xv + 473 pp. LCCN Q127.U6 B3.

Baxter:1968:SAT

[Bax68] James Phinney Baxter, 3rd. Scientists against time. MIT Press, Cambridge, MA, USA, 1968. xxi + 473 pp. LCCN Q127.U6 B3 1968.

Baxter:2001:PHR

[Bax01] Stephen Baxter. The Planetarium Hypothesis: a resolution of the Fermi Paradox. Journal of the British Interplanetary Society, 54 (5–6):210–216, ???? 2001. CODEN JBISAW. ISSN 0007-084X.

Bethe:2002:EFR

[BB02] Hans A. Bethe and Henry Bethe. Enrico Fermi in Rome, 1931– 32. Physics Today, 55(6):28–29, June 2002. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

www.physicstoday.org/resource/1/phtoad/v55/i6/p28_s1.

Bernardini:2004:CFE

[BB04a] C. (Carlo) Bernardini and Luisa Bonolis, editors. Conoscere Fermi: Enrico Fermi: his work and legacy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 88-7438-015-1 (SIF, Bologna, Italy), 3-540-22141-7 (Springer-Verlag, Berlin, Heidelberg, New York). xii + 410 pp. LCCN QC774.F4 C6613 2004. URL http://www.loc.gov/

catdir/enhancements/fy0818/2004108212-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0818/2004108212-t.
html.

Bernardini:2004:EFH

[BB04b] C. (Carlo) Bernardini and Luisa Bonolis, editors. Enrico Fermi: His Work and Legacy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 88-7438-015-1 (SIF, Bologna, Italy), 3-540-22141-7 (Springer-Verlag, Berlin, Heidelberg, New York). xii + 410 pp. LCCN QC774.F4

C6613 2004. URL http://www.loc.gov/catdir/enhancements/
fy0818/2004108212-d.html;   http://www.loc.gov/catdir/
enhancements/fy0818/2004108212-t.html.

Bailey:2010:FPS

[BB10] David H. Bailey and Jonathan M. Borwein. Fermi’s Paradox and Stephen Hawking. Math Drudge, April 30, 2010. URL https://

experimentalmath.info/blog/2010/04/fermis-paradox-and-
stephen-hawking/.

Bailey:2015:DSFa

[BB15a] David H. Bailey and Jonathan M. Borwein. Desperately seeking ET: Fermi’s paradox turns 65 (Part II). Math Drudge, April 16,

2015. URL https://experimentalmath.info/blog/2015/04/
desperately-seeking-et-fermis-paradox-turns-65-part-ii/

.

Bailey:2015:DSFb

[BB15b] David H. Bailey and Jonathan M. Borwein. Desperately seeking ET: Fermi’s Paradox turns 65 (Part II). Huffington Post, ??(??):??, April 17, 2015. URL http://www.

huffingtonpost.com/david-h-bailey/where-is-et-fermis-

parado_b_7014044.html. See Part I [BB15d] and response [?].

Bailey:2015:WFPa

[BB15c] David H. Bailey and Jonathan M. Borwein. Where is ET? Fermi’s paradox turns 65. Math Drudge, April 6, 2015.

URL https://experimentalmath.info/blog/2015/04/where-
is-et-fermis-paradox-turns-2015/.

Bailey:2015:WFPb

[BB15d] David H. Bailey and Jonathan M. Borwein. Where is ET? Fermi’s Paradox turns 65. Huffington Post, ??(??):??, April 10, 2015. URL

http://www.huffingtonpost.com/david-h-bailey/where-is-
et-fermis-parado_b_7014044.html. See also Part II [?].

Bretscher:1955:EF

[BC55] E. Bretscher and John Douglas Cockcroft. Enrico Fermi, 1901– 1954. Biographical Memoirs of Fellows of the Royal Society, 1:68– 78, November 1955. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL http://www.jstor.org/stable/
769243;   https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1955.0006.

Bernardini:2003:RCC

[BC03] Carlo Bernardini and Rocco Capasso. Report on the celebrations for the centenary of Enrico Fermi’s birth. In Anonymous [Ano03], pages 399–405. ISBN 88-8286-032-9. LCCN

???? URL http:old.enea.it/produzione_scientifica/pdf.
../V2003_AttiFermi.pdf.

Bonolis:2009:SRP

[BCR09] Luisa Bonolis, Decio Cocolicchio, and Biagio Russo. Sinisgalli e i ragazzi di via Panisperna. (Italian) [Sinisgalli and the boys of Via Panisperna]. Pristem/Storia: Note di Mathematica, Storia, Cultura, 23/24:1–59, February 2009. ISSN 1825-5221. URL http:/

/matematica.unibocconi.it/pubblicazioni/pristemstoria-

23-24. Special issue: Un “Leonardo” del Novecento: Leonardo Sinisgalli (1908–1981).

Brown:1989:PQP

[BDH89] Laurie M. Brown, Max Dresden, and Lillian Hoddeson, editors. Pions to quarks: particle physics in the 1950s: based on a Fermilab symposium. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-30984-0. LCCN QC793.16 .P56 1989. URL

http://www.loc.gov/catdir/description/cam023/88025644.
html;   http://www.loc.gov/catdir/toc/cam028/88025644.
html.

Banuelos:1981:PCF

[BDM81] Alicia Bañuelos, Ricardo Angel Depine, and Roberto Claudio Mancini. A program for computing the Fermi–Dirac functions. Computer Physics Communications, 21(3):315–322, January 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900126.

Beck:1995:FF

[Bec95] Harold Beck. Fermi in Florence. Physics World, 9(2):18, February 1995. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-

7058 (electronic). URL http://physicsworldarchive.iop.
org/full/pwa-pdf/9/2/phwv9i2a19.pdf.

Belloni:1982:ZPP

[Bel82] L. Belloni. Zametki o puti, privedshem E. Fermi k statistike Fermi–Diraka. (Russian) [Notes on the path that led Fermi to Fermi–Dirac statistics]. Uspekhi Fizicheskikh Nauk, 136(1): 167–175, January 1982. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.ru/ru/

articles/1982/1/f/.

Belloni:1987:FRS

[Bel87] Lanfranco Belloni. Da Fermi a Rubbia: Storia e politica di un successo mondiale della scienza italiana. (Italian) [From Fermi to Rubbia: History and politics of a world-wide success of Italian science]. Rizzoli, Milano, Italy, 1987. ISBN 88-17-53095-6. 217 pp. LCCN QC9.I8 B44 1988. L22000.

Belloni:1988:FRI

[Bel88] Lanfranco Belloni. Da Fermi a Rubbia. (Italian) [From Fermi to Rubbia]. Rizzoli, Milano, Italia, 1988. ISBN 88-17-53095-6. 217 pp. LCCN QC9.I8 B44 1988. L22000.

Belloni:1994:FRF

[Bel94] Lanfranco Belloni. On Fermi’s route to Fermi–Dirac statistics. European Journal of Physics, 15(3):102–109, May 1994. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic). URL

http://iopscience.iop.org/0143-0807/15/3/002.

Belloni:1997:SDS

[Bel97] L. Belloni. Sull’origine della statistica di Fermi. (Italian) [On the origin of Fermi statistics]. Quaderni di Storia della fisica, 1:275– 286, 1997. CODEN ???? ISSN 1594-9974 (print), 1827-6164 (elec-

tronic). URL http://en.sif.it/journals/qsf/econtents.

Bernstein:1980:HBP

[Ber80] Jeremy Bernstein. Hans Bethe, prophet of energy. Basic Books, New York, NY, USA, 1980. ISBN 0-465-02903-5. xii + 212 pp. LCCN QC16.B46 B47. US$11.95.

Bernstein:1988:FPB

[Ber88] Barton J. Bernstein. Four physicists and the bomb: The early years, 1945–1950. Historical Studies in the Physical and Biological Sciences, 18(2):231–263, ???? 1988. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://

www.jstor.org/stable/27757603.

Bernstein:1996:HUC

[Ber96] Jeremy Bernstein. Hitler’s uranium club: the secret recordings at Farm Hall. American Institute of Physics, Woodbury, NY, USA, 1996. ISBN 1-56396-258-6. xxx + 427 + 4 pp. LCCN QC773.3.G3 B47 1995. US$34.95. Introduction by David Cassidy.

Bernardini:2001:EFT

[Ber01] Carlo Bernardini. Enrico Fermi e il trattato di O. D. Chwolson. (Italian) [Enrico Fermi and the treatise of O. D. Chwolson]. Il Nuovo Saggiatore, 17(5–6):15–19, September/December 2, 2001. ISSN 0393-4578 (print), 1827-6148 (electronic). URL http://

prometeo.sif.it:8080/papers/online/sag/2001/05-06/pdf/

05.pdf. Special issue in honor of the centennial of the birth of Enrico Fermi.

Bernardini:2002:CFN

[Ber02a] Carlo Bernardini. Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901–2001]. Edizioni scientifiche SIF, Bologna, Italy, second edition, 2002. ISBN 88-7438-000-3. viii + 383 pp. LCCN QC774.F4. 10.33 EUR.

URL http://www.gbv.de/dms/casalini/02/02220733.pdf;
http://www.sif.it/fermiindex.html; http://www.sif.it/
libri/conoscere_fermi;   http://www.zentralblatt-math.
org/zmath/en/search/?an=1073.01512. English translation in

[BB04b].

Bernardini:2002:I

[Ber02b] Carlo Bernardini. Introduzione. In Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901– 2001] [Ber02a], pages vii–?? ISBN 88-7438-000-3. LCCN

QC774.F4. 10.33 EUR. URL http://prometeo.sif.it:
8080/libri/fermi/introduzione.pdf. English translation in

[BB04b].

Bertotti:2002:CFP

[Ber02c] Bruno Bertotti. Le coordinate di Fermi e il Principio di Equivalenza. (Italian) [Fermi coordinates and the Equivalence Principle]. In Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901–2001] [Ber02a], pages 114–125. ISBN 88-7438-000-3. LCCN QC774.F4. 10.33 EUR. URL http:

//prometeo.sif.it:8080/libri/fermi/07.pdf. English trans-

lation in [BB04b].

Bernardini:2003:SPF

[Ber03] Carlo Bernardini. A short presentation of the Fermi Centennial Conference. In Anonymous [Ano03], pages 9–12. ISBN 88-

8286-032-9. LCCN ???? URL http:old.enea.it/produzione_
scientifica/pdf.../V2003_AttiFermi.pdf.

Berezin:2018:FLS

[Ber18] Alexander Berezin. “First in, last out” solution to the Fermi Paradox. arXiv.org, ??(??):1–3, March 20, 2018. URL https://

arxiv.org/abs/1803.08425.

Bethe:1955:MSH

[Bet55] Hans A. Bethe. Memorial Symposium Held in Honor of Enrico Fermi at the Washington Meeting of the American Physical Society, April 29, 1955. Reviews of Modern Physics, 27(3):249, July 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.27.249;   http://rmp.aps.org/abstract/
RMP/v27/i3/p249_1.

Bethe:1963:FPJ

[Bet63] Hans Bethe. Fermi Prize: J. Robert Oppenheimer named to receive annual AEC Award. Science, 140(3563):161–163, April 12, 1963. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Bethe:2000:GUP

[Bet00] Hans A. Bethe. The German Uranium Project. Physics Today, 53 (7):34–36, July 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Bethe:2004:PR

[Bet04] Hans Bethe. Pilgrimages to Rome. In Orear [Ore04], chapter 18, pages 75–82. ISBN ???? LCCN ???? URL http:

//dspace.library.cornell.edu/handle/1813/62;   http://
hdl.handle.net/1813/74.

Beyer:1949:FNP

[Bey49] Robert T. (Robert Thomas) Beyer, editor. Foundations of nuclear physics: facsimiles of thirteen fundamental studies as they were originally reported in the scientific journals. Dover, New York, NY, USA, 1949. 272 pp. LCCN QC173 .B485. See original paper [HS39b] and later translation [Gra64].

Bethe:1932:WZE

[BF32] Hans Bethe and Enrico Fermi. Über die Wechselwirkung von zwei Elektronen. (German) [On the interaction of two electrons]. Zeitschrift für Physik, 77(5–6):296–306, ???? 1932. CODEN

ZEPYAA. ISSN ???? URL http://www.springerlink.com/

content/l561955527vv4513/. See [Ber80, page 31] for Bethe’s account of how the research in this paper took two days. The writing took one more day, with Fermi proposing each sentence in German, and the native speaker, Bethe, occasionally revising it.

Breit:1942:URS

[BF42] Gregory Breit and Enrico Fermi. The use of reflectors and seeds in a power plant. Report C-11, US Atomic Energy Commission, Washington, DC, USA, March 9, 1942.

Bragdon:1944:MCS

[BFMM44] E. Bragdon, Enrico Fermi, John Marshall, and Leona Marshall. Measurements of the cross-section of boron for thermal neutrons. Report CP-1098, US Atomic Energy Commission, Washington, DC, USA, January 11, 1944.

Bini:2024:FRP

[BGJR24] Donato Bini, Andrea Geralico, Robert T. Jantzen, and Remo Ruffini. On Fermi’s resolution of the “4/3 problem” in the classical theory of the electron. Foundations of Physics, 54(3):??, June 2024. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s10701-024-00770-w.

Brown:1982:BEP

[BH82] Laurie M. Brown and Lillian Hoddeson. The birth of elementaryparticle physics. Physics Today, 35(4):36–43, April 1982. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://www.physicstoday.org/resource/1/

phtoad/v35/i4/p36_s1. This is a much-abridged version of [BH83a].

Brown:1983:BEP

[BH83a] Laurie M. Brown and Lillian Hoddeson. The birth of elementary particle physics. In The Birth of Particle Physics [BH83b], pages 3–36. ISBN 0-521-24005-0 (hardcover), 0-521-33837-9 (paperback). LCCN QC793 .B57 1983.

Brown:1983:BPP

[BH83b] Laurie M. Brown and Lillian Hoddeson, editors. The Birth of Particle Physics. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-24005-0 (hardcover), 0-521-33837-9 (paperback). LCCN QC793 .B57 1983.

Bhabha:1939:FLI

[Bha39] H. J. Bhabha. The fundamental length introduced by the theory of the mesotron (meson). Nature, 143(3616):276–277, February 18, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/

v143/n3616/pdf/143276b0.pdf. This paper immediately follows [Fri39], and proposes replacement of the name ‘mesotron’ used by Fermi [FTW47] with the shorter ‘meson’, the form that is now conventional.

Badash:1980:RA

[BHB80] Lawrence Badash, Joseph Oakland Hirschfelder, and Herbert P. Broida, editors. Reminiscences of Los Alamos, 1943–1945, volume 5 of Studies in the history of modern science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1980. ISBN 90-277-1097-X, 90-277-1098-8 (paperback). xxi + 188 pp. LCCN QC791.96 .R44.

Bennett:2019:CAH

[BHR19] Charles H. Bennett, Robin Hanson, and C. Jess Riedel. Comment on ‘The Aestivation Hypothesis for Resolving Fermi’s Paradox’. Foundations of Physics, 49(8):820–829, August 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). See [SAĆ16].

Badash:1986:NFR

[BHT86] Lawrence Badash, Elizabeth Hodes, and Adolph Tiddens. Nuclear fission: Reaction to the discovery in 1939. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 130(2):196–231, June 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://

www.jstor.org/stable/987181.

Berman:2004:FPU

[BI04] G. P. Berman and F. M. Izrailev. The Fermi–Pasta–Ulam problem: 50 years of progress. arxiv.org, ??(??):1–48, November 29,

2004. URL http://arxiv.org/abs/nlin/0411062v3.

Berman:2005:FPU

[BI05] G. P. Berman and F. M. Izrailev. The Fermi–Pasta–Ulam problem: Fifty years of progress. Chaos (Woodbury, NY), 15(1): 015104, 2005. CODEN CHAOEH. ISSN 1054-1500. URL http:

//chaos.aip.org/resource/1/chaoeh/v15/i1/p015104_s1.

Bianucci:1990:ERP

[Bia90] Piero Bianucci. Era uno dei ragazzi di via Panisperna. (Italian) [it was one of the boys of Via Panisperna]. Vita italiana. Cultura e scienza, A5(1):132–133, January/March 1990.

Bothe:1941:ATN

[BJ41] W. Bothe and P. Jensen. Die Absorption thermischer Neutronen in Elektrographit. (German) [The absorption of thermal neutrons in electrographite]. Report G-71, ????, ????, January 20, 1941. Captured German report cited in [Wea76, page 29] that contained incorrect results on the neutron capture cross-section of carbon. Those results caused the German Uranium Project scientists to switch from graphite to the hard-to-get deuterium (in heavy water obtained from the Rjukan Plant in Norway) as a neutron moderator, and likely, substantially delayed their progress in achieving a working nuclear reactor as a precursor to an atomic bomb.

Blankenbecler:1957:IFF

[Bla57] R. Blankenbecler. Integrals over the Fermi function. American Journal of Physics, 25(5):279–280, May 1957. CODEN AJPIAS.

ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://

ajp.aapt.org/resource/1/ajpias/v25/i5/p279_s1.

Blume:2001:EOF

[Blu01] Martin Blume. Earlier observations of the “Fermi Problem”. Science, 294(554):53, October 9, 2001. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/294/5540/53.2/reply. See [Hub01].

Bloch:1935:PNA

[BM35] F. Bloch and C. Møller. Production of neutrons by annihilation of protons and electrons according to Fermi’s theory. Nature, 136 (3451):987, December 21, 1935. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v136/n3451/pdf/136987a0.pdf.

Braak:2020:FGR

[BM20] D. Braak and J. Mannhart. Fermi’s Golden Rule and the Second Law of Thermodynamics. Foundations of Physics, 50(11): 1509–1540, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-020-00380-2.

Boccaletti:2015:WPS

[Boc15] D. Boccaletti. When a problem is solved too early: Enrico Fermi and the infamous 4/3 problem. Physis: Rivista Internazionale di Storia della Scienza. Nuova Serie, 50(1–2):??, ???? 2015. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Bohr:1939:DHN

[Boh39] Niels Bohr. Disintegration of heavy nuclei. Nature, 143 (3617):330, February 25, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://

www.chemteam.info/Chem-History/Bohr-Fission-1939.html;
 http://www.nature.com/nature/journal/v143/n3617/pdf/
143330a0.pdf.

Bonolis:2002:CDS

[Bon02] Luisa Bonolis. Cronologia dell’Opera scientifica di Enrico Fermi. (Italian) [History of the scientific work of Enrico Fermi]. In Conoscere Fermi nel centenario della nascita: 29 settembre 1901– 2001. (Italian) [Learn about Fermi in the centenary of his birth:

29 September 1901–2001] [Ber02a], pages 319–378. ISBN 88-7438-000-3. LCCN QC774.F4. 10.33 EUR. URL http://prometeo.

sif.it:8080/libri/fermi/crono.pdf. English translation in

[BB04b].

Bonolis:2005:BPS

[Bon05] Luisa Bonolis. Bruno Pontecorvo: From slow neutrons to oscillating neutrinos. American Journal of Physics, 73(6):487–499, 2005. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://scitation.aip.org/content/aapt/
journal/ajp/73/6/10.1119/1.1852540.

Born:1957:MA

[Bor57] Max Born. Man and the atom. Bulletin of the Atomic Scientists, 13(6):186–194, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Bor63]. See note in [BR57].

Born:1963:MA

[Bor63] Max Born. Man and the atom. In Grodzins and Rabinowitch [GR63], pages 590–601. LCCN D842 .B78. Reprint of [Bor57].

Boskey:1950:IA

[Bos50] Bennett Boskey. Inventions and the atom. Columbia Law Review, 50(4):433–447, April 1950. CODEN ???? ISSN 0010-1958 (print), 1945-2268 (electronic). URL http://www.jstor.org/

stable/1118964.

Byers:1958:TSM

[BP58] N. Byers and R. E. Peierls. Two-stage model of Fermi interactions. Il Nuovo Cimento, 10(3):520–524, November 1958. URL https:

//ui.adsabs.harvard.edu/#abs/1958NCim...10..520B.

Brink:1968:RTF

[BP68] D. M. Brink and Rudolf Peierls. A realistic Thomas–Fermi approach to finite nuclei. Comments on Nuclear and Particle Physics, 2:28–32, January/February 1968. CODEN CNPPAV. ISSN 0010-2709.

Byers:1997:TSM

[BP97] N. Byers and R. E. Peierls. Two-stage model of Fermi interactions. In Dalitz and Peierls [DP97], pages 449–453. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book).

LCCN QC21.2 .P42 1997. URL https://ui.adsabs.harvard.
edu/#abs/1997sspr.book..449B.

Benettin:2018:FPU

[BPP18] G. Benettin, S. Pasquali, and A. Ponno. The Fermi–Pasta–Ulam problem and its underlying integrable dynamics: An approach through Lyapunov exponents. Journal of Statistical Physics, 171(4):521–542, May 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Born:1957:LEW

[BR57] Max Born and Eugene I. Rabinowitch. Letter to the Editor: We’re sorry. Bulletin of the Atomic Scientists, 13(7):273, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1994:FTN

[BR94] Laurie M. Brown and Helmut Rechenberg. Field theories of nuclear forces in the 1930s: The Fermi-field theory. Historical Studies in the Physical and Biological Sciences, 25(1):1–24, ???? 1994. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.jstor.org/stable/27757733.

Brown:1996:OCN

[BR96] Laurie M. Brown and Helmut Rechenberg. The Origin of the Concept of Nuclear Forces. IOP Publishing, Bristol, UK, 1996. ISBN 0-7503-0373-5. xii + 392 pp. LCCN QC793.3.B5 B76 1996.

URL http://catdir.loc.gov/catdir/enhancements/fy0701/
96031758-d.html.

Braun:1969:BRI

[Bra69] Elisabeth Esser Braun. Book review: Illustrious Immigrants: The Intellectual Migration From Europe 1930–41, by Laura Fermi. Foro Internacional, 10(2):216–218, October/December

1969. ISSN 0185-013X. URL http://www.jstor.org/stable/
27737536.

Brandt:2009:HCD

[Bra09] Siegmund Brandt. The harvest of a century: discoveries of modern physics in 100 episodes. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-954469-7 (hardcover). xiv + 500 pp. LCCN QC7 .B64 2009.

Brink:1965:NF

[Bri65] David Maurice Brink. Nuclear Forces, volume 354 of Commonwealth and international library: Selected readings in physics. Pergamon, New York, NY, USA, 1965. ISBN 0-08-011034-7. viii + 232 pp. LCCN QC173 .B8513 1965.

Brin:1983:GSC

[Bri83] Glen David Brin. The ‘Great Silence’: the controversy concerning extraterrestrial intelligent life. Quarterly Journal of the Royal Astronomical Society, 24(??):283–309, September 1983. CODEN

QJRAAK. ISSN 0035-8738. URL http://adsabs.harvard.edu/
abs/1983QJRAS..24..283B.

Broyles:1961:DFT

[Bro61] Arthur A. Broyles. Derivation of the Fermi–Thomas equation. American Journal of Physics, 29(2):81–83, February 1961. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://ajp.aapt.org/resource/1/ajpias/v29/i2/p81_
s1.

Brown:1968:BRT

[Bro68] G. E. Brown. Book review: Theory of Finite Fermi Systems and Applications to Atomic Nuclei. A. B. Migdal. Translated from the Russian by S. Chomet. Interscience (Wiley), New York, 1967. viii + 319 pp., illus. $17.50. Interscience Monographs and Texts in Physics and Astronomy, vol. 19. Science, 159(3815):621, February 9, 1968. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/159/
3815/621.full.pdf.

Bronowski:1973:AM

[Bro73a] Jacob Bronowski. The Ascent of Man. British Broadcasting Corporation, London, UK, 1973. ISBN 0-563-10498-8. 448 pp. LCCN Q175 .B7918 1973; CB151.

Bronowski:1973:NCF

[Bro73b] Jacob Bronowski. The neutron: Chadwick and Fermi. In The Ascent of Man [Bro73a], pages 341–347. ISBN 0-563-10498-8. LCCN Q175 .B7918 1973; CB151.

Brown:1973:BRS

[Bro73c] Willard Brown. Book review: Segrè, Emilio. Enrico Fermi, physicist. Chicago, Illinois: The University of Chicago Press, 1970 (276 pages), $6.95. Science Education, 57(1):101, January/March 1973. CODEN SEDUAV. ISSN 0036-8326 (print), 1098-237X

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/sce.3730570130/abstract.

Brown:1978:IN

[Bro78] Laurie M. Brown. The idea of the neutrino. Physics Today, 31(9):23–28, September 1978. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

nobelprize.org/nobel_prizes/physics/laureates/1935/;
http://www.physicstoday.org/resource/1/phtoad/v31/i9/

p23_s1. Page 27 contains an English translation of Pauli’s famous letter proposing the existence of an as-yet-undiscovered particle, which he called the ‘neutron’. Fermi later renamed it ‘neutrino’ because of its small (and possibly zero) mass. The real neutron was first discovered by James Chadwick in 1932, a result for which he received the 1935 Nobel Prize in Physics.

Bruzzaniti:2007:EFG

[Bru07] Giuseppe Bruzzaniti. Enrico Fermi: il genio obbediente. (Italian) [Enrico Fermi: the obedient genius], volume 882 of Saggi. G. Einaudi, Torino, Italy, 2007. ISBN 88-06-16682-4 (paperback). xiii + 386 pp. LCCN QC774.F4 B78

2007. 24.50 EUR. URL http://matematica.unibocconi.it/
libri/enrico-fermi-il-genio-obbediente;   http://www.
loc.gov/catdir/toc/casalini06/06541291.pdf.

Bruzzaniti:2011:GBR

[Bru11] Giuseppe Bruzzaniti. Giuseppe Bruzzaniti racconta Enrico Fermi: L’atomo e la bomba atomica. (Italian) [Giuseppe Bruzzaniti talks about Enrico Fermi: the atom and the atomic bomb]. Video DVD, 2011. Directed by Michele Calvano.

Bruzzaniti:2016:EFO

[Bru16] Giuseppe Bruzzaniti. Enrico Fermi: the Obedient Genius. Springer Biographies. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 1-4939-3531-3 (hardcover), 1-4939-3533-X (e-book). xiii + 348 + 54 pp.

LCCN QC774.F4. URL http://link.springer.com/book/10.
1007/978-1-4939-3533-8.

Bertin:1984:FYW

[BRV84] Antonio Bertin, R. A. Ricci, and A. (Antonio) Vitale, editors. Fifty years of weak-interaction physics: on the occasion of the fiftieth anniversary of Fermi’s theory on nuclear beta-decay. Italian Physical Society, Bologna, Italy, 1984. 31 + xl + 797 + 3 pp. LCCN QC794.8.W4 F54 1984.

Bennett:2017:LU

[BS17] Jeffrey O. Bennett and G. Seth Shostak. Life in the Universe. Pearson, Boston, MA, USA, fourth edition, 2017. ISBN 0-13-408908-1. xvi + 510 + 34 pp. LCCN QH327 .B45 2017.

Buck:1973:BRE

[Buc73a] Barbara Reeves Buck. Book review: Enrico Fermi, Physicist by Emilio Segrè. British Journal for the History of Science, 6(3): 331–332, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print),

1474-001X (electronic). URL http://www.jstor.org/stable/

4025464. See [Seg70b].

Buck:1973:BRT

[Buc73b] Barbara Reeves Buck. Book review: Twentieth Century Enrico Fermi Physicist. By Emilio Segrè. Chicago and London: University of Chicago Press, 1970. Pp. xii + 276. £3.15. British Journal for the History of Science, 6(3):331–332, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

http://www.jstor.org/stable/4025464.

Buck:1980:IPT

[Buc80] Barbara Reeves Buck. Italian Physicists and Their Institutions, 1861–1911. Ph.D. thesis, Harvard University, Cam-

bridge, MA, USA, 1980. URL http://hollis.harvard.
edu/?itemid=%7Clibrary/m/aleph%7C003907650;   http://
search.proquest.com/docview/302964545.

Buck:1983:HAE

[Buc83] Alice L. Buck. A history of the Atomic Energy Commission. Report, US Department of Energy, Washington, DC, USA,

July 1983. 28 pp. URL http://www.atomictraveler.com/
HistoryofAEC.pdf.

Buhrke:1998:LWI

[Büh98a] Thomas Bührke. “Was Ich brauche, ist ein Stück Paraffin.” Enrico Fermi (1901–1954). (German) [“What I need is a piece of paraffin.” Enrico Fermi (1901–1954)]. In Newtons Apfel: Sternstunden der Physik; von Galilei bis Lise Meitner. (German) [Newton’s apple: great moments of physics; from Galileo to Lise Meitner] [Büh98b], pages 207–230. ISBN 3-406-44402-4, 3-406-42002-8 (paperback). LCCN ????

Buhrke:1998:NAS

[Büh98b] Thomas Bührke. Newtons Apfel: Sternstunden der Physik; von Galilei bis Lise Meitner. (German) [Newton’s apple: great moments of physics; from Galileo to Lise Meitner], volume 1202 of Beck’sche Reihe. Verlag C. H. Beck, München, Germany, 1998. ISBN 3-406-44402-4, 3-406-42002-8 (paperback). 258 pp. LCCN ????

Bullough:1981:BFE

[Bul81] R. K. Bullough. Bose–Fermi equivalence and soliton theory in solid-state physics. Nature, 292(5822):411–412, July 30, 1981. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v292/
n5822/pdf/292411a0.pdf.

Bunker:1983:ERF

[Bun83] Merle E. Bunker. Early reactors — from Fermi’s water boiler to novel power prototypes. Los Alamos Science, 7:124–131, Winter/ Spring 1983. CODEN LASCDI. ISSN 0273-7116. URL http://

library.lanl.gov/cgi-bin/getfile?07-14.pdf.

Bohr:1939:FP

[BW39a] Niels Bohr and John A. Wheeler. The fission of protactinium. Physical Review, 56(10):1065–1066, November 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v56/i10/p1065_2.

Bohr:1939:MNF

[BW39b] Niels Bohr and John Archibald Wheeler. The mechanism of nuclear fission. Physical Review (2), 56(5):426–450, September 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v56/
i5/p426_1.

Benguria:1990:SPS

[BY90] Rafael D. Benguria and Cecilia Yarur. Symmetry properties of the solutions to Thomas–Fermi–Dirac–von Weizsäcker type equations. Transactions of the American Mathematical Society, 320 (2):665–675, August 1990. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic). URL http://www.jstor.org/

stable/2001695.

Byers:2002:FS

[Bye02] N. Byers. Fermi and Szilard. ArXiv Physics e-prints, July 2002.

URL http://adsabs.harvard.edu/abs/2002physics...7094B.

Byers:2003:WPF

[Bye03] Nina Byers. Women in physics in Fermi’s time. In Anonymous [Ano03], pages 269–288. ISBN 88-8286-032-9. LCCN ???? URL

https://arxiv.org/html/physics/0302035.

Compton:1936:XRT

[CA36] Arthur Holly Compton and Samuel King Allison. X-rays in theory and experiment. D. Van Nostrand Company, Inc., New York, NY, USA, 1936. xiv + 828 pp. LCCN QC481 .C65 1936.

Chimbidis:1977:FLS

[CA77] James Chimbidis and Jay Andre, editors. To Fermi — with love. Argonne National Laboratory, Argonne, IL, USA, 1977. Sound recording (33 1/3 rpm).

Cabibbo:2002:IDI

[Cab02] Nicola Cabibbo. Le interazioni deboli. (Italian) [Weak interactions]. In Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901–2001] [Ber02a], pages 139–151. ISBN 88-7438-000-3. LCCN QC774.F4. 10.33 EUR. URL http:

//prometeo.sif.it:8080/libri/fermi/09.pdf. English trans-

lation in [BB04b].

Cabibbo:2003:FTW

[Cab03] Nicola Cabibbo. Fermi’s tentativo and weak interactions. In Anonymous [Ano03], pages 305–316. ISBN 88-8286-032-9. LCCN

???? URL http:old.enea.it/produzione_scientifica/pdf.
../V2003_AttiFermi.pdf.

Cahn:1995:DND

[Cah95] Robert Cahn. Over the desert, a nuclear dawn. The Christian Science Monitor, ??(??):??, July 13, 1995. ISSN 0882-7729 (print), 1540-4617 (electronic). URL http://search.proquest.

com/docview/405745975.

Calo:1975:EFC

[Cal75] Vincenzo Calò. Enrico Fermi e le centrali nucleari. (Italian) [Enrico Fermi and the Nuclear Reactor]. Mezzina, Molfetta, Italia, 1975. 64 pp. LCCN TK1078 .C34. L2500.

Calvani:2001:IAF

[Cal01] Paolo Calvani. Italy after Fermi. Physics World, 14(11):17, November 2001. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/14/11/phwv14i11a20.pdf.

Calver:2013:BRS

[Cal13] Neil Calver. Book reviews: Simone Turchetti, The Pontecorvo Affair: A Cold War Defection and Nuclear Physics. Chicago and London: The University of Chicago Press, 2012. Pp. iv + 292. ISBN 978-0-226-81664-7. £29.00 (hardback). British Journal for the History of Science, 46(1):173–174, March 2013. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Campbell:1954:BRA

[Cam54] Kathleen Campbell. Book review: Atoms in the Family, by Laura Fermi, Chicago: University of Chicago Press, 1954. Bulletin of the Atomic Scientists, 10(10):393–394, December 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carbonari:2001:EFS

[Car01] L. Carbonari. Enrico Fermi e la scoperta della radioattività indotta dai neutroni. (Italian) [Enrico Fermi and the discovery of radioactivity induced by neutrons]. Ph.D. thesis, Department of Physics, Università ‘La Sapienza’, Roma, Italy, 2001.

Carbonari:2005:SEF

[Car05] L. Carbonari. La strumentazione di Enrico Fermi negli esperimenti con i neutroni a Via Panisperna. (Italian) [The instrumentation in the experiments of Enrico Fermi with neutrons on Via Panisperna]. Quaderni di Storia della fisica, 13:93–123, 2005. CO-DEN ???? ISSN 1594-9974 (print), 1827-6164 (electronic). URL

http://en.sif.it/journals/qsf/econtents.

CarrChilders:2015:SRH

[Car15] Leisl Carr Childers. The Size of the Risk: Histories of Multiple Use in the Great Basin. University of Oklahoma Press, Norman, OK, USA, 2015. ISBN 0-8061-4927-2 (hardcover), 0-8061-5253-2 (e-book). x + 314 pp. LCCN HD210.G7 C55 2015. URL http:/

/www.h-net.org/reviews/showrev.php?id=45878.

Caton:2003:EFH

[Cat03] Alice Caton. Enrico Fermi and his family. In Anonymous [Ano03], pages 43–52. ISBN 88-8286-032-9. LCCN

???? URL http:old.enea.it/produzione_scientifica/pdf.
../V2003_AttiFermi.pdf.

Chen:2009:VFG

[CBH09] Chi Chen, C. A. Bobisch, and W. Ho. Visualization of Fermi’s Golden Rule through imaging of light emission from atomic silver chains. Science, 325(5943):981–985, August 21, 2009. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/325/
5943/981.full.pdf.

Cockcroft:1934:SPN

[CCJ^{+}34] J. Cockcroft, J. Chadwick, F. Joliot, J. Joliot, N. Bohr, G. Gamov, P. A. M. Dirac, and W. Heisenberg, editors. Structure et propriétés des noyaux atomiques. Rapports et discussions du septième conseil de physique tenu à Bruxelles du 22 au 29 octobre 1933 sous les auspices de l’institut international de physique Solvay. (French) [Structure and properties of atomic nuclei. Reports and discussions of the Seventh Meeting on Physics held in Brussels from 22 to 29 October 1933 under the auspices of the Solvay International Institute of Physics]. Gauthier-Villars et cie, Paris, France, 1934. LCCN ???? Publiés par la commission administrative de l’institut.

Condon:1953:TRIa

[CDFS53a] Edward U. Condon, K. K. Darrow, Enrico Fermi, and John C. Slater. The tape recordings of important speeches. American Journal of Physics, 21(2):150, February 1953. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://link.aip.org/link/?AJP/21/150/2;   http://
link.aip.org/link/ajpias/v21/i2/p150/s2.

Condon:1953:TRIb

[CDFS53b] Edward U. Condon, K. K. Darrow, Enrico Fermi, and John C. Slater. The tape recordings of important speeches. American Journal of Physics, 21(3):227, March 1953. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://link.aip.org/link/?AJP/21/227/1;   http://link.
aip.org/link/ajpias/v21/i3/p227/s1.

Condon:1954:TRIc

[CDFS54] Edward U. Condon, K. K. Darrow, Enrico Fermi, and John C. Slater. The tape recordings of important speeches. American Journal of Physics, 22(4):246, April 1954. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http:

//link.aip.org/link/?AJP/22/246/2;   http://link.aip.
org/link/ajpias/v22/i4/p246/s2.

Cordella:2001:EFA

[CDS01] Francesco Cordella, Alberto De Gregorio, and Fabio Sebastiani. Enrico Fermi: gli anni italiani. (Italian) [Enrico Fermi: the Italian years]. Nuova biblioteca di cultura. Editori riuniti, Roma, Italy, 2001. ISBN 88-359-5097-X. 332 pp. LCCN QC16.F46 C67 2001.

Cerase:1969:BRI

[Cer69] F. Cerase. Book review: Illustrious Immigrants: The Intellectual Migration From Europe 1930–41, by Laura Fermi. Genus, 25(1/4): 377, 1969. CODEN GNUSA7. ISSN 0016-6987 (print), 2035-5556

(electronic). URL http://www.jstor.org/stable/29787897.

Chandrasekhar:1953:MFS

[CF53a] Subrahmanyan Chandrasekhar and Enrico Fermi. Magnetic fields in spiral arms. Astrophysical Journal, 118:113–115, 1953. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Chandrasekhar:1953:PGS

[CF53b] Subrahmanyan Chandrasekhar and Enrico Fermi. Problems of gravitational stability in the presence of a magnetic field. Astrophysical Journal, 118:116–141, 1953. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic). URL http://

inspirehep.net/record/894281.

Compton:1947:SCL

[CFU47] Karl T. Compton, Enrico Fermi, and Harold Urey. Scientists comment on Lilienthal. Bulletin of the Atomic Scientists, 3(3):75, March 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Christy:1942:ETC

[CFW42] Robert F. Christy, Enrico Fermi, and Alvin M. Weinberg. Effect of temperature changes on reproduction factor. Report CP-254, US Atomic Energy Commission, Washington, DC, USA, September 14, 1942.

Chadwick:1932:EN

[Cha32a] James Chadwick. The existence of a neutron. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 136(830):692–708, June 1, 1932. ISSN 0950-1207 (print), 2053-9150 (electronic). URL http:// www.jstor.org/stable/95816. This paper reports the discovery of the neutron, for which Chadwick received the Nobel Prize in Physics in 1935.

Chadwick:1932:PEN

[Cha32b] James Chadwick. Possible existence of a neutron. Nature, 129 (3252):312, February 27, 1932. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v129/n3252/pdf/129312a0.pdf.

Chandrasekhar:1973:REF

[Cha73] S. Chandrasekhar. Remarks on Enrico Fermi. In Mehra [Meh73], chapter 43, pages 800–802. ISBN 90-277-0345-0, 90-277-2536-5.

LCCN QC173.96 .S95 1972. URL http://www.springer.com/
us/book/9789027703453.

Charin:2015:PFE

[Cha15] Gabriel Charin. Le paradoxe de Fermi et les extraterrestres invisibles. (French) [Fermi’s paradox and the invisible extraterrestrials]. Libération, ??(??):??, February 5, 2015. URL

http://www.liberation.fr/terre/2015/02/05/le-paradoxe-
de-fermi-et-les-extraterrestres-invisibles_1196456.

Chiu:1994:FIS

[Chi94] Ch. S. Chiu. Frauen im Schatten. (German) [Women in the shade]. J and V, Wien, Austria, 1994. ISBN 3-224-17669-5. 228 pp. LCCN CT3310 .C45 1994.

Crawford:1987:NPC

[CHU87] Elisabeth T. Crawford, J. L. Heilbron, and Rebecca Ullrich. The Nobel population 1901–1937: a census of the nominators and nominees for the prizes in physics and chemistry, volume 11; 4 of Berkeley papers in history of science; Uppsala studies in history of science. Office for History of Science and Technology, University of California at Berkeley, Berkeley, CA, USA, 1987. ISBN 0-918102-15-4 (paperback). 337 pp. LCCN QC28 .C74 1987.

Chwolson:1906:TPF

[Chw13] Orest Danilovich Chwolson. Traité de physique. (French) [Treatise on physics]. A. Hermann, Paris, France, 1906–1913. 4500 (est.: 5 volumes) pp. LCCN ???? Four volumes. Translated by E. Davaux from the Russian and German editions. Considerably expanded by the author. Followed by notes on theoretical physics by E. Cosserat and F. Cosserat. With a preface by E.-H. Amagat.

Cantelon:1991:AAD

[CHW91] Philip L. (Philip Louis) Cantelon, Richard G. Hewlett, and Robert Chadwell Williams, editors. The American atom: a documentary history of nuclear policies from the discovery of fission to the present. University of Pennsylvania Press, Philadelphia, second edition, 1991. ISBN 0-8122-3096-5 (hardcover), 0-8122-1354-8 (paper). xviii + 369 pp. LCCN UA23 .A597 1991.

Cifarelli:2008:PSI

[Cif08] L. Cifarelli. Preface: Special issue dedicated to Enrico Fermi: Lectures on pions and nucleons: Research highlights. Rivista del Nuovo Cimento, 31(1):1–7, January 2008. CODEN RNUCAC. ISSN 0393-697X (print), 1826-9850 (electronic).

Cini:2002:FEQ

[Cin02] Marcello Cini. Fermi e l’elettrodinamica quantistica. (Italian) [Fermi and quantum electrodynamics]. In Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901–2001] [Ber02a], pages 126–138. ISBN 88-7438-000-3. LCCN

QC774.F4. 10.33 EUR. URL http://prometeo.sif.it:8080/

libri/fermi/08.pdf. English translation in [BB04b].

Cipolla:1994:WIG

[Cip94] Gaetano Cipolla. What Italy has given the world. Legas, Brooklyn, NY, USA, 1994. ISBN 1-4619-0457-9 (e-book), 1-881901-04-1. 32 pp. LCCN DG441 .C56 1994.

Ciancio:2004:GIF

[CIR04] E. Ciancio, R. C. Iotti, and F. Rossi. Gauge-invariant formulation of Fermi’s golden rule: Application to high-field transport in semiconductors. Europhysics Letters, 65(2):242–247, January 2004. CODEN EULEEJ. ISSN 0295-5075 (print), 1286-4854

(electronic). URL http://iopscience.iop.org/0295-5075/65/
2/242/.

Cirkovic:2009:FPL

[Ćir09] Milan M. Ćirković. Fermi’s Paradox — the last challenge for Copernicanism? Serbian Astronomical Journal, 178(??):1–20, ???? 2009. CODEN SAJOFK. ISSN 1450-698X (print), 1820-9289 (electronic). arXiv:0907.3432.

Cirkovic:2016:FPD

[Ćir16] Milan M. Ćirković. Fermi’s Paradox is a daunting problem — under whatever label. Astrobiology, 16(10):737–740, October 2016. CODEN ASTRC4. ISSN 1531-1074 (print), 1557-8070

(electronic). URL https://www.liebertpub.com/doi/10.1089/

ast.2016.1498. See reply [Gra16b].

Cirkovic:2018:GSS

[Ćir18] Milan M. Ćirković. The Great Silence: Science and Philosophy of Fermi’s Paradox. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2018. ISBN 0-19-964630-9 (hardcover). xxvii + 395 pp. LCCN QB54 .C57 2018.

Chakravarty:2008:FPQ

[CK08] Sudip Chakravarty and Hae-Young Kee. Fermi pockets and quantum oscillations of the Hall coefficient in high-temperature superconductors. Proceedings of the National Academy of Sciences of the United States of America, 105(26):8835–8839, July 1, 2008. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (elec-

tronic). URL http://www.jstor.org/stable/25462878.

Carbonari:2003:IFP

[CL03] Luca Carbonari and Gilda Leoni. The Istituto Fisico on Via Panisperna: the new Museo Storico della Fisica e Centro Studi e Ricerche ‘Enrico Fermi’ di Roma. High Energy Physics Libraries Webzine., April 2003. URL http://library.web.cern.

ch/library/Webzine/7/papers/3/.

Close:2015:HLD

[Clo15] Frank E. Close. Half-life: the divided life of Bruno Pontecorvo, physicist or spy. Basic Books, New York, NY, USA, 2015. ISBN 0-465-06998-3 (hardcover), 0-465-04487-5 (e-book), 1-78074-582-6 (e-book). xix + 378 pp. LCCN QC774.P66 C56 2014. US$29.99.

Cocconi:1959:SIC

[CM59] Giuseppe Cocconi and Philip Morrison. Searching for interstellar communications. Nature, 184(4690):844–846, September 19, 1959. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v184/n4690/pdf/184844a0.pdf.

Cardone:2002:EFS

[CM02] Fabio Cardone and Roberto Mignani. Enrico Fermi e i secchi della sora Cesarina: metodo, pregiudizio e caso in fisica. (Italian) [????], volume 14 of Collana Arcobaleno. Di Renzo, Roma, Italy, nuova edition, 2002. ISBN 88-8323-050-7. 119 pp. LCCN QC16.F46 C37 2002.

Cockcroft:1962:EDC

[Coc62] Sir John Cockcroft. The early days of the Canadian and British atomic energy projects. International Atomic Energy Agency Bulletin, 4(0):18–20, ???? 1962. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-2690 (electronic). URL http://

en.wikipedia.org/wiki/AERE;   http://en.wikipedia.org/
wiki/Battle_of_the_Netherlands;   http://en.wikipedia.
org/wiki/Chalk_River_Laboratories;   http://en.wikipedia.
org/wiki/MAUD_Committee;   http://en.wikipedia.org/wiki/
Operation_Weser%C3%BCbung;   http://en.wikipedia.org/
wiki/Tube_Alloys;   http://www.iaea.org/Publications/
Magazines/Bulletin/Bull040su/04004701820su.pdf.

Codelli:1988:RPI

[Cod88] Lorenzo Codelli. I ragazzi di Via Panisperna. (Italian) [The Boys of Via Panisperna]. Comune, Gorizia, Italy, 1988. ISBN ???? iv + 81 pp. LCCN ???? For Sergio Amidei and Milko Tebaldi. With an introduction by Antonello Trombadori.

Compton:1942:LEF

[Com42] Arthur Compton. Letter to Enrico Fermi: Establishing Site X. US DOE report, September 14, 1942. URL http://www.osti.

gov/accomplishments/documents/fullText/ACC0038.pdf.

Compton:1953:BAP

[Com53] Arthur H. Compton. The birth of atomic power. Bulletin of the Atomic Scientists, 9(2):10–12, February 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Compton:1956:AQP

[Com56] Arthur Holly Compton. Atomic quest, a personal narrative. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1956. xix + 370 pp. LCCN QC773.A1 C65.

Condon:1962:YQP

[Con62] Edward U. Condon. 60 years of quantum physics. Physics Today, 15(10):37–49, October 1962. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/PHTOAD/v15/i10. Delayed 1951

Presidential address at the 1500th regular meeting of the American Philosophical Society of Washington, 2 December 1962, at the Natural History Museum Auditorium of the Smithsonian Institution, on the 60th anniversary of Planck’s constant, h. Reprinted in [WP85, pages 310–318].

Cooper:1999:EFR

[Coo99] Dan Cooper. Enrico Fermi and the revolutions in modern physics. Oxford portraits in science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-511762-X. 117 pp.

LCCN QC16.F46 C66 1999. URL http://www.loc.gov/catdir/
enhancements/fy0603/98034471-d.html.

Corben:1964:IMW

[Cor64] H. C. Corben. If microns were fermis. Journal of Mathematical Physics, 5(12):1664–1668, December 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i12/p1664_s1.

Cordella:1998:FTR

[Cor98] Francesco Cordella. The first theoretical researches of Enrico Fermi (1921–1926). Ph.D. thesis, Museum of Physics at the University “La Sapienza”, Rome, Italy, 1998. Awarded with the 1998 prize of the [Italian] National Academy of Sciences.

Corson:2004:WC

[Cor04] Dale Corson. Welcome to Cornell. In Orear [Ore04], chapter 16, pages 69–70. ISBN ???? LCCN ???? URL http:

//dspace.library.cornell.edu/handle/1813/62;   http://
hdl.handle.net/1813/74.

Cowen:2009:FON

[Cow09] Ron Cowen. Fermi opens new window on high-energy universe: Gamma-ray telescope detects bursts and pulsars. Science News (Washington, DC), 175(2):5–6, January 17, 2009. CODEN SC-NEBK. ISSN 0036-8423 (print), 1943-0930 (electronic). URL

http://www.jstor.org/stable/20465818.

Crawford:1969:LMA

[Cra69] Deborah Crawford. Lise Meitner, atomic pioneer. Crown Publishers, New York, NY, USA, 1969. 192 pp. LCCN QC774.M4 C7 1969.

Crawford:1988:UEM

[Cra88] Frank S. Crawford. Using Einstein’s method to derive both the Planck and Fermi–Dirac distributions. American Journal of Physics, 56(10):883–885, October 1988. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://

ajp.aapt.org/resource/1/ajpias/v56/i10/p883_s1.

Crawford:1992:NIS

[Cra92] Elisabeth T. Crawford. Nationalism and internationalism in science, 1880–1939: four studies of the Nobel population. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-40386-3 (hardcover). xii + 157 pp. LCCN Q126.9 .C73 1992. URL http://

www.loc.gov/catdir/description/cam024/91033702.html;
http://www.loc.gov/catdir/samples/cam031/91033702.html;
 http://www.loc.gov/catdir/toc/cam028/91033702.html.

Crawford:2000:WTM

[Cra00] Ian A. Crawford. Where are they? Maybe we are alone in the galaxy after all. Scientific American, 283(1):38–43, July 2000. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v283/n1/pdf/scientificamerican0700-38.pdf.

Crawford:2002:HSN

[Cra02] Elisabeth T. Crawford, editor. Historical Studies in the Nobel Archives: the Prizes in Science and Medicine, volume 31 of Uppsala studies in the history of science. Universal Academy Press, Tokyo, Japan, 2002. ISBN 4-946443-69-X. ISSN 0282-1036. vii + 161 pp. LCCN AS911.N9 .H57 2002.

Cravens:2007:PSW

[Cra07] Gwyneth Cravens. Power to save the world: the truth about nuclear energy. Alfred A. Knopf, New York, NY, USA, 2007. ISBN 0-307-26656-7. xv + 439 pp. LCCN

TK9146 .C65 2007.  URL http://www.loc.gov/catdir/
enhancements/fy0804/2007017611-b.ht;   http://www.loc.
gov/catdir/enhancements/fy0804/2007017611-d.ht;   http:
//www.loc.gov/catdir/enhancements/fy0804/2007017611-s.
ht;   http://www.loc.gov/catdir/toc/ecip0716/2007017611.
html.

Crease:2010:BRJ

[Cre10] Robert P. Crease. Book review: Joanna S. Ploeger, The Boundaries of the New Frontier: Rhetoric and Communication at Fermi National Accelerator Laboratory (Studies in Rhetoric/ Communication). Isis, 101(1):263–264, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/653917;   http://
www.jstor.org/stable/pdfplus/10.1086/653917.pdf.

Cowan:1956:DFN

[CRH^{+}56] C. L. Cowan, Jr., F. Reines, F. B. Harrison, H. W. Kruse, and A. D. McGuire. Detection of the free neutrino: a confirmation. Science, 1224(3212):103–104, July 20, 1956. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:

//www.sciencemag.org/content/124/3212/103.extract. See

also the first report of detection of the neutrino [RC53].

Crowe:1986:ELD

[Cro86] Michael J. Crowe. The extraterrestrial life debate 1750–1900: the idea of a plurality of worlds from Kant to Lowell. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-26305-0 (hardcover), 0-521-35986-4 (paperback). 680 pp. LCCN QB54 .C76 1986.

Crowe:1999:ELD

[Cro99] Michael J. Crowe. The Extraterrestrial Life Debate, 1750–1900. Dover, New York, NY, USA, 1999. ISBN 0-486-40675-X (paperback), 0-486-14501-8 (e-book). xxiv + 680 pp. LCCN QB54 .C76

Cronin:2004:FR

[Cro04] James W. Cronin, editor. Fermi remembered. University of Chicago Press, Chicago, IL, USA, 2004. ISBN 0-226-12111-9 (hardcover). xi + 287 pp. LCCN QC16.F46

F49 2004. URL http://www.loc.gov/catdir/description/
uchi052/2003020524.html;   http://www.loc.gov/catdir/
enhancements/fy0614/2003020524-b.html;   http://www.loc.
gov/catdir/toc/ecip049/2003020524.html.

Cordella:1999:DEF

[CS99] F. Cordella and F. Sebastiani. Il debutto di Enrico Fermi come fisico teorico: i primi lavori sulla relatività (1921–1922/23). (Italian) [The debut of Enrico Fermi as a theoretical physicist: the early work on Relativity (1921–1922/23)]. Quaderni di Storia della fisica, 5:69–88, 1999. CODEN ???? ISSN 1594-9974 (print),

1827-6164 (electronic). URL http://en.sif.it/journals/qsf/
econtents.

Cordella:2000:SFI

[CS00a] F. Cordella and F. Sebastiani. La statistica di Fermi. (Italian) [Fermi’ statistics]. Giornale di Fisica, 40(??):131–556, ???? 2000. CODEN GFSIAD. ISSN 0017-0283 (print), 1827-6156 (electronic).

Cordella:2000:FGL

[CS00b] Francesco Cordella and Fabio Sebastiani. Fermi a Gottinga e a Leida: gli anni che precedono la statistica quantica. (Italian) [Fermi in Göttingen and Leiden: years preceding quantum statistics]. Quaderni di Storia della fisica, 6:17–45, 2000. CODEN ???? ISSN 1594-9974 (print), 1827-6164 (electronic). URL http://en.

sif.it/journals/qsf/econtents.

Cordella:2000:DLC

[CS00c] Francesco Cordella and Fabio Sebastiani. I due lavori che preludono alla statistica quantica. (Italian) [the two works that precede quantum statistics]. Giornale di Fisica, 41(??):83–101, ???? 2000. CODEN GFSIAD. ISSN 0017-0283 (print), 1827-6156 (electronic).

Cordella:2000:SPF

[CS00d] Francesco Cordella and Fabio Sebastiani. Sul percorso di Fermi verso la statistica quantica. (Italian) [On Fermi’s trail to quantum statistics]. Il Nuovo Saggiatore, 16(??):11–22, ???? 2000. ISSN 0393-4578 (print), 1827-6148 (electronic).

Crawford:1996:NTW

[CSW96] Elisabeth Crawford, Ruth Lewin Sime, and Mark Walker. A Nobel tale of wartime injustice. Nature, 382(6590):393–395, August 1, 1996. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v382/n6590/abs/382393a0.html.

Crawford:1997:KIP

[CSW97a] Elisabeth Crawford, Ruth Lewin Sime, and Mark Walker. Die Kernspaltung und ihr Preis. Warum nur Otto Hahn den Nobelpreis erhielt, Otto Frisch, Lise Meitner und Fritz Straßmann dagegen nicht berücksichtigt werden. (German) [Fission and its price. Why only Otto Hahn received the Nobel Prize: Otto Frisch, Lise Meitner and Fritz Strassmann are not taken into consideration]. Kultur & Technik. Zeitschrift des Deutschen Museums München, 21(2):30–35, ???? 1997. ISSN 0344-5690. URL http:/

/www.deutsches-museum.de/fileadmin/Content/data/Insel/
Information/KT/heftarchiv/1997/21-2-30.pdf. Translation

from English by Dieter Beisel of [CSW96].

Crawford:1997:NTP

[CSW97b] Elisabeth Crawford, Ruth Lewin Sime, and Mark Walker. A Nobel tale of postwar injustice: Recently released Swedish documents reveal why Lise Meitner, codiscoverer of nuclear fission, did not receive the 1946 Physics Prize for her theoretical interpretation of the process. Physics Today, 50(9):26–32, September 1997. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Cody:1967:RCA

[CT67] W. J. Cody and Henry C. Thacher, Jr. Rational Chebyshev approximations for Fermi–Dirac integrals of orders −1/2, 1/2 and 3/2. Mathematics of Computation, 21(97):30–40, January 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.jstor.org/stable/2003468.

Cumo:2002:RTN

[Cum02] Maurizio Cumo. Reattori e tecnologie nucleari: Lo sviluppo nel mondo. (Italian) [Reactors and nuclear technologies: The development in the world]. In Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901–2001] [Ber02a], pages 223–241. ISBN 88-7438-000-3. LCCN QC774.F4. 10.33

EUR. URL http://prometeo.sif.it:8080/libri/fermi/13.

pdf. English translation in [BB04b].

Cockcroft:1932:EHVa

[CW32a] John D. Cockcroft and Ernest T. S. Walton. Experiments with high velocity positive ions: I. further developments in the method of obtaining high velocity positive ions. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 136(830):619–630, June 1, 1932. ISSN 0950-1207 (print), 2053-9150 (electronic). URL http://www. jstor.org/stable/95811. The work in this paper, and its companion [CW32b], won the Nobel Prize in Physics in 1951 for the authors “for their pioneer work on the transmutation of atomic nuclei by artificially accelerated atomic particles”.

Cockcroft:1932:EHVb

[CW32b] John D. Cockcroft and Ernest T. S. Walton. Experiments with high velocity positive ions: II. The disintegration of elements by high velocity protons. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 137(831):229–242, July 1, 1932. ISSN 0950-1207 (print), 2053-9150 (electronic). URL http://www.jstor.org/ stable/95941. The work in this paper, and its companion [CW32a], won the Nobel Prize in Physics in 1951 for the authors “for their pioneer work on the transmutation of atomic nuclei by artificially accelerated atomic particles”.

Cory:1971:FRW

[CW71] William B. Cory and Donald Wilke. Film review: The World of Enrico Fermi, edited by William Riley. American Journal of Physics, 39(7):849, July 1971. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.aapt.

org/resource/1/ajpias/v39/i7/p849_s1.

Cracknell:1973:FSC

[CW73] Arthur P. Cracknell and K. C. Wong. The Fermi surface; its concept, determination, and use in the physics of metals. Monographs on the physics and chemistry of materials. Clarendon Press, Oxford, UK, 1973. ISBN 0-19-851330-5. xii + 565 pp. LCCN QC176.8.E4 C68.

DAgostini:2003:BRD

[D’A03] G. (Giulio) D’Agostini. Bayesian reasoning in data analysis: a critical introduction. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2003. ISBN 981-238-356-5. xix + 329 pp. LCCN QA279.5 .D28 2003. URL http://

www.worldscientific.com/worldscibooks/10.1142/5262.

DAgostini:2005:FBT

[D’A05] G. D’Agostini. The Fermi’s Bayes Theorem. arxiv.org, ??(??):1–

4, September 10, 2005. URL http://arxiv.org/abs/physics/
0509080.

Darden:1998:NSI

[Dar98] Lindley Darden. The nature of scientific inquiry. Web docu-

ment, 1998. URL http://www.philosophy.umd.edu/Faculty/
LDarden/sciinq/.

Dardo:2004:NLT

[Dar04] M. (Mauro) Dardo. Nobel laureates and twentieth-century physics. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-83247-0, 0-521-54008-9 (paperback). xi + 533 pp.

LCCN QC7 .D27 2004. URL http://www.loc.gov/catdir/
description/cam041/2004049240.html;   http://www.loc.
gov/catdir/toc/cam041/2004049240.html.

Dauxois:2008:FPU

[Dau08] Thierry Dauxois. Fermi, Pasta, Ulam and a mysterious lady. arxiv.org, ??(??):3, January 10, 2008. URL http://arxiv.org/

abs/0801.1590.

Davies:1982:AU

[Dav82] P. C. W. Davies. The Accidental Universe. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24212-6 (hardcover), 0-521-28692-1 (paperback). xii + 139 pp. LCCN

QB981 .D268 1982. URL http://catdir.loc.gov/catdir/
description/cam022/81021592.html.

DeMarco:2001:PDF

[De 01] F. De Marco. Prospettive della fusione nucleare. (Italian) [prospects of nuclear fusion]. Il Nuovo Saggiatore, 17(5–6): 61–65, September/December 2001. ISSN 0393-4578 (print),

1827-6148 (electronic). URL http://prometeo.sif.it:8080/
papers/online/sag/2001/05-06/pdf/05.pdf. Special issue in

honor of the centennial of the birth of Enrico Fermi.

DeMaria:2003:FAN

[De 03] Michelangelo De Maria. Fermi and applied nuclear physics during the War (1939–1945). In Anonymous [Ano03], pages 217– 218. ISBN 88-8286-032-9. LCCN ???? URL http:old.enea.it/

produzione_scientifica/pdf.../V2003_AttiFermi.pdf.

DeGregorio:2005:EFD

[De 05a] Alberto De Gregorio. Enrico Fermi and the discovery of neutroninduced radioactivity: a project being crowned. arxiv.org, ??(??):

14, January 27, 2005. URL http://arxiv.org/abs/physics/
0501140.

DeGregorio:2005:RIN

[De 05b] Alberto De Gregorio. Radioactivity induced by neutrons: a thermodynamic approach to radiative capture. arxiv.org, pages 1–14,

2005. URL http://arxiv.org/abs/physics/0508049.

DeGregorio:2005:NPE

[De 05c] Alberto G. De Gregorio. Neutron physics in the early 1930s. Historical Studies in the Physical and Biological Sciences, 35(2):293– 340, March 2005. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

DeGregorio:2006:FRP

[De 06a] Alberto De Gregorio. A far-reaching project behind the discovery of neutron-induced radioactivity. Studies in History and Philosophy of Modern Physics, 37(2):330–346, June 2006. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219806000098.

DeGregorio:2006:RIN

[De 06b] Alberto De Gregorio. Radioactivity induced by neutrons: Enrico Fermi and a thermodynamic approach to radiative capture. American Journal of Physics, 74(7):614–620, July 2006. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://ajp.aapt.org/resource/1/ajpias/v74/i7/p614_s1.

DiGrezia:2004:FMS

[DE04] E. Di Grezia and S. Esposito. Fermi, Majorana and the statistical model of atoms. Foundations of Physics, 34(9):1431– 1450, September 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/B:FOOP.0000044099.18859.1d.

DeGregorio:2006:EFE

[DE06a] Alberto De Gregorio and Salvatore Esposito. Enrico Fermi e Ettore Majorana: continuitá e rinnovamento nell’insegnamento della fisica teorica. (Italian) [Enrico Fermi, Ettore Majorana: continuity and renewal in the teaching of theoretical physics]. arxiv.org, ??(??):13, February 1, 2006. URL http://arxiv.org/ abs/physics/0602008. To be published in Sapere [(Bari) (ISSN 0036-4681)].

DeGregorio:2006:TTP

[DE06b] Alberto De Gregorio and Salvatore Esposito. Teaching theoretical physics: the cases of Enrico Fermi and Ettore Majorana. arxiv.org, ??(??):21, February 21, 2006. URL http://arxiv.org/

abs/physics/0602146.

DeGregorio:2007:TTP

[DE07] Alberto De Gregorio and Salvatore Esposito. Teaching theoretical physics: the cases of Enrico Fermi and Ettore Majorana. American Journal of Physics, 75(9):781–790, September 2007. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://ajp.aapt.org/resource/1/ajpias/v75/i9/p781_s1.

DeGregorio:2015:EFE

[DE15] Alberto De Gregorio and Salvatore Esposito. Enrico Fermi e Ettore Majorana: continuità e rinnovamento nell’insegnamento della fisica teorica. (Italian) [Enrico Fermi and Ettore Majorana: continuity and renewal in the teaching of theoretical physics]. Report, L’Università di Roma^{≪}La Sapienza^{≫}e l Università di Napoli ≪Federico II^{≫}, Roma, Italia e Napoli, Italia, July 27, 2015. 13 pp.

URL https://www.researchgate.net/publication/2175080.

Dean:2013:NYT

[Dea13] Cornelia Dean, editor. The New York Times book of physics and astronomy: more than 100 years of covering the expanding universe. Sterling, New York, NY, USA, 2013. ISBN 1-4027-9320-0 (hardcover). xviii + 557 pp. LCCN QC7 .D43 2013. Foreword by Neil deGrasse Tyson.

Debye:1953:SWE

[Deb53] Peter Debye. The scientific work of Enrico Fermi. Proceedings of the American Academy of Arts and Sciences, 82(7):290–293, December 1953. CODEN PAAAAV. ISSN 0065-6836. URL http:

//www.jstor.org/stable/20023728.

Delmastro:2017:MFF

[Del17] Marco Delmastro. The many faces of Fermi. Physics World, 30 (12):40–41, December 2017. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL https://iopscience.iop.

org/article/10.1088/2058-7058/30/12/39.

Demortier:2015:QEQ

[Dem15] Guy Demortier. Qui est qui dans ^{≪}the letter to Roosevelt ^{≫}?. (French) [Who is who in “The letter to Roosevelt”?]. Revue des Questions Scientifiques, 186(1):125–152, ???? 2015. CO-DEN RQSCAN. ISSN 0035-2160. URL https://www.unamur.

be/sciences/philosoc/revueqs/textes-en-ligne/rqs_186_
1_demortier.

Derenzini:1964:AME

[Der64] Tullio Derenzini. Analisi dei manoscritti de Enrico Fermi conservati presso la ^{≪}Domus Galiaeana^{≫}. (Italian) [Analysis of the manuscripts of Enrico Fermi preserved at the Domus Galiaeana]. Physis: Rivista Internazionale di Storia della Scienza, 6(??):75– 85, 1964. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Derezinski:2006:FGR

[DF06] Jan Dereziński and Rafal Früboes. Fermi Golden Rule and open quantum systems. Lecture Notes in Mathematics, 1882:67– 116, 2006. CODEN LNMAA2. ISBN 3-540-30993-4 (print), 3-540-33967-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/3-540-33967-1_2.pdf.

DeMaria:1989:PIC

[DGS89] Michelangelo De Maria, Mario Grilli, and Fabio Sebastiani, editors. Proceedings of the International Conference on the Restructuring of Physical Sciences in Europe and the United States, 1945– 1960: Università “La Sapienza,” Rome, Italy, 19–23 September 1988. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1989. ISBN 9971-5-0740-4. LCCN QC9.U5 I57 1988.

deHevesy:1962:RPI

[dH62] George de Hevesy. The reactor and the production of isotopes. International Atomic Energy Agency Bulletin, 4(0):37, ???? 1962. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-

2690 (electronic). URL http://www.iaea.org/Publications/
Magazines/Bulletin/Bull040su/04004703737su.pdf.

Dingle:1958:AEC

[Din58] R. B. Dingle. Asymptotic expansions and converging factors. III. gamma, psi and polygamma functions, and Fermi–Dirac and Bose–Einstein integrals. Proceedings of the Royal Society of London. Series A, Mathematical and Physical Sciences, 244(1239): 484–490, April 22, 1958. CODEN PRLAAZ. ISSN 0080-4630.

URL http://www.jstor.org/stable/100264.

Dirac:1926:TQM

[Dir26] P. A. M. Dirac. On the theory of quantum mechanics. Proceedings of the Royal Society of London. Series A, Mathematical and Physical Sciences, 112(762):661–667, October 1, 1926. CO-DEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor.org/ stable/94692. Russian translation in [Dir03, Vol. II, pp. 147ff].

Dirac:1977:REE

[Dir77] P. A. M. Dirac. Recollections of an exciting era. In Weiner [Wei77], pages 109–146. ISBN 0-12-368857-4. LCCN QC7 .V37 1977.

Dirac:2003:SNT

[Dir03] P. A. M. Dirac. Sobranie Nauchnykh Trudov T. II. Kvantovaya Teoriya (Nauchnye Stat’i: 1924–1947). (Russian) [Collected Scientific Works. Quantum Theory (Scientific Papers: 1924–1947)]. Klassiki Nauki. Fizmatlit, Moscow, Russia, 2003. ISBN 5-9221-0381-4. 846 pp. LCCN ???? Edited by A. D. Sukhanov.

DeMarco:1999:OFD

[DJ99] B. DeMarco and D. S. Jin. Onset of Fermi degeneracy in a trapped atomic gas. Science, 285(5434):1703–1706, September 10, 1999. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.jstor.org/stable/2898982.

deLatil:1963:EFO

[dL63] Pierre de Latil. Enrico Fermi, ou le Christophe Colomb de l’atome. (French) [ Enrico Fermi, or the Christopher Columbus of the atom], volume 18 of Savants du monde entier. Seghers, Paris, France, 1963. 222 pp. LCCN ????

deLatil:1964:EFMb

[dL64] Pierre de Latil. Enrico Fermi: the man and his theories. A Profile in science. Paul S. Eriksson, New York, NY, USA, 1964. 178 pp. LCCN QC16.F46 L3814 1964. Some library catalogs list the year as 1966.

deLatil:1965:EFM

[dL65] Pierre de Latil. Enrico Fermi: the man and his theories. A Profile in science. Souvenir Press, London, UK, 1965. 178 pp. LCCN

QC774.F4 L313 1965. URL http://books.google.com/books?
id=YCE6AQAAIAAJ;   http://catalog.hathitrust.org/api/
volumes/oclc/2311870.html.

deLatil:1974:FVR

[dL74] Pierre de Latil. Fermi: la vita, le ricerche, le testimonianze. (Italian) [Fermi: life, studies, testimonials]. Accademia, Milano, Italy, 1974. 227 pp. LCCN ???? Translation by Livia Zannini of [dL63] to Italian.

Dobler:2022:WWT

[Döb22] Niklas Alexander Döbler. Where will they be: hidden implications of solutions to the Fermi paradox. International Journal of Astrobiology, ??(??):1–5, April 2022. CODEN IJANFR. ISSN 1473-5504 (print), 1475-3006 (electronic). URL https://www.

cambridge.org/core/journals/international-journal-of-
astrobiology/article/where-will-they-be-hidden-implications-
of-solutions-to-the-fermi-paradox/38F67DD23A230F8532F85A552D6A447C.

Dole:1964:HPM

[Dol64] Stephen H. Dole. Habitable Planets for Man. A Blaisdell book on the pure and applied sciences. Blaisdell Publishing Company, New York, NY, USA, 1964. xiii + 158 pp. LCCN QB54.D63 1964.

Dole:1970:HPMh

[Dol70] Stephen H. Dole. Habitable Planets for Man. American Elsevier Publishing Company, New York, NY, USA, second edition, 1970. ISBN 0-444-00092-5. xiii + 158 pp. LCCN QB54.

Dole:2007:HPM

[Dol07] Stephen H. Dole. Habitable Planets for Man. RAND, Santa Monica, CA, USA, 2007. ISBN 0-8330-4227-0 (paperback), 0-8330-4813-9 (e-book). xiii + 158 pp. LCCN QB54 .D63

2007. URL http://bibpurl.oclc.org/web/63740;   http:/
/www.rand.org/pubs/commercial\_books/CB179-1/.

Dalitz:1997:SSP

[DP97] Richard Henry Dalitz and Sir Rudolf Peierls, editors. Selected scientific papers of Sir Rudolf Peierls: with commentary, volume 19 of World Scientific series in 20th century physics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1997. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). xxiii + 805 pp. LCCN QC21.2 .P42 1997. URL

https://ui.adsabs.harvard.edu/#abs/1997sspr.book.....
D;   https://ui.adsabs.harvard.edu/#abs/1997ssps.book.
....P;   https://ui.adsabs.harvard.edu/#abs/1997WSSP..
.19.....D;   https://www.worldscientific.com/worldscibooks/
10.1142/3128.

Dauxois:2005:FPU

[DPR05] Thierry Dauxois, Michel Peyrard, and Stefano Ruffo. The Fermi– Pasta–Ulam ‘numerical experiment’: history and pedagogical perspectives. European Journal of Physics, 26(5):S3–S11, 2005. CO-DEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic).

URL https://arxiv.org/pdf/nlin/0501053.pdf.

delRegato:1985:RP

[dR85] Juan A. del Regato. Radiological physicists. American Institute of Physics, Woodbury, NY, USA, 1985. ISBN 0-88318-469-9. 188 pp. LCCN QC774.A2 D44 1985.

DAgostino:2001:EFE

[DR01] Salvo D’Agostino and Arcangelo Rossi, editors. Enrico Fermi e l’Enciclopedia italiana . (Italian) [Enrico Fermi and the Italian Encyclopedia]. Istituto della Enciclopedia italiana, Roma, Italy, 2001. ISBN ???? xxvi + 191 pp. LCCN QC3 .E39154 2001. Documentary appendices by Roberto Vergara Caffarelli.

Dragoni:1972:BRB

[Dra72a] Giorgio Dragoni. [Book review:] B. Pontecorvo, Fermi e la fisica moderna, trad. it. di S. Amadesi. (Italian) [B. Pontecorvo, Fermi and modern physics, Italian translation by S. Amadesi]. Physis: Rivista Internazionale di Storia della Scienza, 14(??):312–315, 1972. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Dragoni:1972:BRS

[Dra72b] Giorgio Dragoni. [Book review:] E. Segrè, Enrico Fermi, fisico. Una bibliografia scientifica. (Italian) [E. Segrè, Enrico Fermi, Physicist. A scientific bibliography]. Physis: Rivista Internazionale di Storia della Scienza, 14(??):307–311, 1972. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Dragoni:1976:MDV

[Dra76] Giorgio Dragoni. Un momento della vita scientifica italiana degli anni trenta: La scoperta dei neutroni lenti e la loro introduzione nella sperimentazione fisica. (Italian) [A moment in the life of Italian science in the thirties: The discovery of slow neutrons and their introduction in physical experimentation]. Physis: Rivista Internazionale di Storia della Scienza, 18(??):131–164, 1976. CO-DEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Dronsfield:2020:BRT

[Dro20] Alan Dronsfield. Book review: Traveling with the Atom: a Scientific Guide to Europe and Beyond. Web site, March 20, 2020.

URL https://www.chemistryworld.com/review/traveling-
with-the-atom-a-scientific-guide-to-europe-and-beyond/
4011154.article.

Danysz:1934:EFE

[DRWZ34] M. Danysz, J. Rotblat, L. Wertenstein, and M. Zyw. Experiments on the Fermi effect. Nature, 134(3399):970–971, December 22, 1934. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v134/n3399/pdf/134970a0.pdf.

Drake:1992:ATSf

[DS92] Frank D. Drake and Dava Sobel. Is anyone out there? The scientific search for extraterrestrial intelligence. Delacorte Press, New York, NY, USA, 1992. ISBN 0-385-30532-X. xv + 272 + 16 pp. LCCN QB54 .D72 1992.

DeGregorio:2007:DEF

[DS07] A. De Gregorio and F. Sebastiani. Il debutto di Enrico Fermi come professore di fisica teorica. (Italian) [The debut of Enrico Fermi as a professor of theoretical physics]. Physis: Rivista Internazionale di Storia della Scienza. Nuova Serie, 44(2):469–499, 2007. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

DeGregorio:2009:EFO

[DS09] Alberto De Gregorio and Fabio Sebastiani. Enrico Fermi and the old quantum physics. arxiv.org, pages 1–12, April 1, 2009. URL

http://arxiv.org/abs/0904.0260.

DeGregorio:2010:EFO

[DS10] A. De Gregorio and F. Sebastiani. Enrico Fermi and the old quantum physics. Quaderni di Storia della fisica, 16:81–102, 2010. CODEN ???? ISSN 1594-9974 (print), 1827-6164 (electronic).

URL http://en.sif.it/journals/qsf/econtents.

DeGregorio:2011:FPP

[DS11] A. De Gregorio and F. Sebastiani. Fermi, Persico e il principio di indeterminazione. (Italian) [Fermi, Persico, and the Uncertainty Principle]. Quaderni di Storia della fisica, 17:83–106, 2011. CO-DEN ???? ISSN 1594-9974 (print), 1827-6164 (electronic). URL

http://en.sif.it/journals/qsf/econtents.

Dugdale:1961:ERL

[Dug61] J. S. Dugdale. Electrical resistivity at low temperatures: the pressure dependence of the electrical resistance of metals gives some clues about their Fermi surfaces. Science, 134(3472):77–86, July 14, 1961. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.sciencemag.org/content/
134/3472/77.full.pdf.

DuMond:1928:ECS

[DuM28] Jesse W. M. DuMond. Experimental confirmation for Sommerfeld– Fermi–Dirac degenerate gas theory of conduction electrons. Science, 68(1767):452, November 9, 1928. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

jstor.org/stable/1652417;   http://www.sciencemag.org/
content/68/1767/452.1.full.pdf.

Dyson:1960:SAS

[Dys60] Freeman J. Dyson. Search for artificial stellar sources of infrared radiation. Science, 131(3414):1667–1668, June 3, 1960. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1705101;
 http://www.sciencemag.org/content/131/3414/1667.full.

pdf. See comments [SW66].

Dyson:2004:MEF

[Dys04] Freeman Dyson. A meeting with Enrico Fermi. Nature, 427 (6972):297, January 22, 2004. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v427/n6972/full/427297a.html.

Dyson:2010:MCG

[Dys10] Freeman J. Dyson. Many-colored glass: reflections on the place of life in the universe. University of Virginia Press, Charlottesville, VA, USA, 2010. ISBN 0-8139-2973-3. xi + 162 pp. LCCN ???? Page–Barbour Lectures for 2004.

Esposito:2015:PEM

[EAW15] Salvatore Esposito, E. (Evgeny) Akhmedov, and Frank Wilczek. The physics of Ettore Majorana: phenomenological, theoretical, and mathematical. Cambridge University Press, Cambridge, UK, 2015. ISBN 1-107-04402-2 (hardcover), 1-316-19108-7 (PDF ebook). xi + 382 pp. LCCN QC19.6 .E87 2015.

Edge:1983:PSE

[ED83] Douglas R. M. Edge and Michael K. Dirks. Problem solving, Enrico Fermi and the bull moose. School Science and Mathematics, 83(7):601–608, November 1983. CODEN SSMAAC. ISSN 0036-6803 (print), 1949-8594 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.1111/j.1949-8594.1983.
tb10144.x/abstract.

Epstein:1970:EFF

[EE70] Sam Epstein and Beryl Williams Epstein. Enrico Fermi, father of atomic power. Americans all. Garrard Publishing Company, Champaign, IL, USA, 1970. ISBN 0-8116-4558-4. 95 pp. LCCN PZ10.E6 En. Illustrations by Raymond Burns.

Epele:1999:PAA

[EFGP99] L. N. Epele, H. Fanchiotti, C. A. García Canal, and J. A. Ponciano. Padé approximant approach to the Thomas–Fermi problem. Physical Review A (Atomic, Molecular, and Optical Physics), 60(1):280–283, July 1999. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL http://link.

aps.org/doi/10.1103/PhysRevA.60.280.

Einstein:1939:AEL

[Ein39] Albert Einstein. Albert Einstein’s letter to President Franklin Delano Roosevelt. World-Wide Web document, August 2, 1939.

URL http://hypertextbook.com/eworld/einstein.shtml;
http://www.anl.gov/Science_and_Technology/History/Anniversary_
Frontiers/aetofdr.html;   http://www.lanl.gov/history/
road/pdf/Einstein.pdf.

Eklund:1962:I

[Ekl62] Sigvard Eklund. Introduction. International Atomic Energy Agency Bulletin, 4(0):3–5, ???? 1962. CODEN IAE-BAB. ISSN 0020-6067 (print), 1564-2690 (electronic). URL

http://www.iaea.org/Publications/Magazines/Bulletin/..

./04005080305su.pdd. Special number to mark the 20th anniversary of the world’s first nuclear reactor (December 2, 1962).

Efthimiou:2007:CFP

[EL07] C. J. Efthimiou and R. A. Llewellyn. Cinema, Fermi problems and general education. Physics Education, 42(3):253–261, 2007. CO-DEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic).

URL http://stacks.iop.org/0031-9120/42/i=3/a=003.

Esposito:2008:EFP

[EP08] Salvatore Esposito and Ofelia Pisanti. Enrico Fermi and the physics and engineering of a nuclear pile: the retrieval of novel documents. arxiv.org, ??(??):78 + 7, March 7, 2008. URL http:

//arxiv.org/abs/0803.1145.

Esposito:2010:NPN

[EP10] Salvatore Esposito and O. (Ofelia) Pisanti, editors. Neutron physics for nuclear reactors: unpublished writings [of Enrico Fermi]. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2010. ISBN 981-4291-22-6 (hardcover). xxxv + 665 pp. LCCN QC793.5.N462 F47 2010.

Ernst:1992:LMO

[Ern92] Sabine Ernst. Lise Meitner an Otto Hahn: Briefe aus den Jahren 1912 bis 1924: Edition und Kommentierung. (German) [Lise Meitner, Otto Hahn: Letters from the years 1912 to 1924: Edition and Commentary], volume 65 of Quellen und Studien zur Geschichte der Pharmazie. Wissenschaftliche Verlagsgesellschaft mbH, Stuttgart, Germany, 1992. ISBN 3-8047-1254-1. vi + 267 pp. LCCN RS61 .Q8 Bd. 65. With a foreword by Fritz Krafft.

Esposito:2002:MST

[Esp02] Salvatore Esposito. Majorana solution of the Thomas–Fermi equation. American Journal of Physics, 70(8):852–856, August 2002. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://ajp.aapt.org/resource/1/ajpias/v70/
i8/p852_s1.

Esposito:2005:AMT

[Esp05] Salvatore Esposito. Again on Majorana and the Thomas–Fermi model: a comment to physics/0511222. arxiv.org, ??(??):

??, December 5, 2005. URL http://arxiv.org/abs/physics/

0512259. See [GR05].

Esposito:2008:EMH

[Esp08a] Salvatore Esposito. Ettore Majorana and his heritage seventy years later. Annalen der Physik (1900), 17(5):302–318, May 2008. ISSN 0003-3804 (print), 1521-3889 (electronic). URL http://

adsabs.harvard.edu/abs/2008AnP...520..302E.

Esposito:2008:FAE

[Esp08b] Salvatore Esposito. Fermi at Los Alamos and the early Britain’s way to nuclear energy. arxiv.org, ??(??):1–7, May 2, 2008. URL

http://arxiv.org/abs/0805.0229v1. Report number DSF-

31/2007.

Fermi:1942:ECR

[F^{+}42a] Enrico Fermi et al. The experimental chain reacting pile and reproduction factor in some exponential piles. Report CP-341, US Atomic Energy Commission, Washington, DC, USA, 1942. In: Fermi, E., et al. Physics Research. Report for month ending November 15, 1942.

Fermi:1942:PEAb

[F^{+}42b] Enrico Fermi et al. The projected experiment at Argonne Forest and the reproduction factor in metal piles (exponential experiments; general features of the pile planned at Argonne Forest). Report CP-297, US Atomic Energy Commission, Washington, DC, USA, 1942. 4 pp. In: Fermi, E., et al. Physics Research. Report for month ending October 15, 1942.

Fermi:1942:SRPa

[F^{+}42c] Enrico Fermi et al. Status of research problems in experimental nuclear physics (exponential piles; nuclear properties of various materials). Report CP-133, US Atomic Energy Commission, Washington, DC, USA, 1942. 2 pp. In: Fermi, E., et al. Metallurgical Project. Report for week ending June 20, 1942.

Fermi:1942:SRPb

[F^{+}42d] Enrico Fermi et al. Status of research problems in experimental physics (exponential experiments; nuclear properties of various materials). Report CP-207, US Atomic Energy Commission, Washington, DC, USA, 1942. 3 pp. In: Fermi, E., et al. Metallurgical Project. Report for week ending July 25, 1942.

Fermi:1942:SRPd

[F^{+}42e] Enrico Fermi et al. Status of research problems of the Physics Division (exponential experiments; nuclear properties of various materials; miscellaneous experiments in nuclear physics; preparatory work for the experimental pile). Report CP-257, US Atomic Energy Commission, Washington, DC, USA, 1942. 4 pp. In: Fermi, E., et al. Physics Research. Report for month ending September 15, 1942.

Fermi:1942:SRPc

[F^{+}42f] Enrico Fermi et al. Status of research problems of the Physics Division (exponential experiments; nuclear properties of various materials; miscellaneous experiments in nuclear physics; preparatory work for the experimental pile; instruments). Report CP-235, US Atomic Energy Commission, Washington, DC, USA, 1942. In: Fermi, E., et al. Physics Research. Report for month ending August 15, 1942.

Fermi:1942:WCP

[F^{+}42g] Enrico Fermi et al. Work carried out by the Physics Division (experimental production of a chain reaction; exponential experiment with metal; effective temperature of the thermal neutrons). Report CP-387, US Atomic Energy Commission, Washington, DC, USA, 1942. 1 pp. In: Fermi, E., et al. Report on Physics Research for month ending December 15, 1942.

Fermi:1943:SAEa

[F^{+}43a] Enrico Fermi et al. Summary of the activities of the Experimental Section of the Nuclear Physics Division in the past month. Report CP-570, US Atomic Energy Commission, Washington, DC, USA, 1943. 1 pp. In: Fermi, E., and Herbert L. Anderson. Physics Research. Report for month ending April 17, 1943.

Fermi:1943:SAEb

[F^{+}43b] Enrico Fermi et al. Summary of the activities of the experimental section of the nuclear physics division in the past month. Report CP-641, US Atomic Energy Commission, Washington, DC, USA, 1943. 2 pp. In: Fermi, E., et al. Physics Research. Report for month ending May 10, 1943.

Fermi:1943:SAEd

[F^{+}43c] Enrico Fermi et al. Summary of the activities of the experimental section of the nuclear physics division in the past month. Report CP-781, US Atomic Energy Commission, Washington, DC, USA, 1943. 3 pp. In: Fermi, E., et al. Physics Research. Report for month ending July 10, 1943.

Fermi:1943:SAEe

[F^{+}43d] Enrico Fermi et al. Summary of the activities of the experimental section of the nuclear physics division in the past month. Report CP-1016, US Atomic Energy Commission, Washington, DC, USA, October 1943. 2 pp. In: Fermi, E., et al. Physics Research. Report for month ending October 23, 1943.

Fermi:1943:SAEf

[F^{+}43e] Enrico Fermi et al. Summary of the activities of the experimental section of the nuclear physics division in the past month. Report CP-1088, US Atomic Energy Commission, Washington, DC, USA, November 1943. 3 pp. In: Fermi, E., et al. Nuclear Physics Research. Report for month ending November 23,1943.

Fermi:1943:SAEg

[F^{+}43f] Enrico Fermi et al. Summary of the activities of the experimental section of the nuclear physics division in the past month. Report CP-1175, US Atomic Energy Commission, Washington, DC, USA, 1943. 2 pp. In: Fermi, E., et al. Nuclear Physics Research. Report for month ending December 25,1943.

Fermi:1943:SAEc

[F^{+}43g] Enrico Fermi et al. Summary of the activities of the Experimental Section of the Nuclear Physics Division in the past month [June 15, 1943]. Report CP-718, United States Department of Energy, Washington, DC, USA, 1943. 2 pp. URL http://www.osti.

gov/accomplishments/documents/fullText/ACC0039.pdf. In:

Fermi, E., et al. Physics Research. Report for month ending June 12, 1943.

Fermi:1943:WCPa

[F^{+}43h] Enrico Fermi et al. Work carried out by the Physics Division (position of control rod in experimental pile; temperature coefficient of effective reproduction factor). Report CP-416, US Atomic Energy Commission, Washington, DC, USA, 1943. 3 pp. In: Fermi, E., et al. Report on physics research for month ending January 15, 1943.

Fermi:1943:WCPb

[F^{+}43i] Enrico Fermi et al. Work carried out in the Physics Division (experiments performed using the chain reacting pile). Report CP-455, US Atomic Energy Commission, Washington, DC, USA, 1943. 4 pp. In: Fermi, E., et al. Report on physics research for period ending February 6, 1943.

Fermi:1944:SAEf

[F^{+}44a] Enrico Fermi et al. Summary of the activities of the experimental section of the nuclear physics division. Report CP-1761, US Atomic Energy Commission, Washington, DC, USA, May 1944. 2 pp. In: Fermi, E., et al. Nuclear Physics Division. Part II of report for month ending May 25, 1944.

Fermi:1944:SAEg

[F^{+}44b] Enrico Fermi et al. Summary of the activities of the experimental section of the nuclear physics division. Report CP-1827, US Atomic Energy Commission, Washington, DC, USA, June 1944.

2 pp. In: Fermi, E., et al. Nuclear Physics Division. Report for month ending June 25, 1944.

Fermi:1944:SAEa

[F^{+}44c] Enrico Fermi et al. Summary of the activities of the experimental section of the nuclear physics division in the past month. Report CP-1255, US Atomic Energy Commission, Washington, DC, USA, January 1944. 5 pp. In: Fermi, E., et al. Nuclear Physics Research. Report for month ending January 24, 1944.

Fermi:1944:SAEb

[F^{+}44d] Enrico Fermi et al. Summary of the activities of the experimental section of the nuclear physics division in the past month. Report CP-1389, US Atomic Energy Commission, Washington, DC, USA, February 1944. 5 pp. In: Fermi, E., et al. Nuclear Physics Research. Report for month ending February 24, 1944.

Fermi:1944:SAEc

[F^{+}44e] Enrico Fermi et al. Summary of the activities of the experimental section of the nuclear physics division in the past month. Report CP-1531, US Atomic Energy Commission, Washington, DC, USA, 1944. 3 pp. In: Fermi, E., et al. Nuclear Physics Research. Report for month ending March 25, 1944.

Fermi:1944:SAEd

[F^{+}44f] Enrico Fermi et al. Summary of the activities of the experimental section of the nuclear physics division in the past month. Report CP-1592, US Atomic Energy Commission, Washington, DC, USA, April 1944. 3 pp. In: Fermi, E., et al. Nuclear Physics Research. Report for month ending April 24, 1944.

Fermi:1944:SAEe

[F^{+}44g] Enrico Fermi et al. Summary of the activities of the experimental section of the nuclear physics division in the past month. Report CP-1729, US Atomic Energy Commission, Washington, DC, USA, 1944. 2 pp. In: Fermi, E., et al. Nuclear Physics Division. Part I of report for month ending May 25, 1944.

Fermi:1934:ODE

[FA34] Enrico Fermi and Edoardo Amaldi. Le orbite 1s degli elementi. (Italian) [The orbits of 1s of the elements]. R. Accad. d’Italia. Memorie, 6(1 (Fis.)):119–149, 1934.

Fermi:1956:TMN

[FA56] Enrico Fermi and Herbert L. Anderson. Testing material in a neutronic reactor. US Patent 2,768,134., October 23, 1956. US Patent Application 613,156 filed August 28, 1945.

Fermi:1961:MTT

[FA61] Enrico Fermi and Herbert L. Anderson. Method of testing therman neutron fissionable material for purity. US Patent 2,969,307., January 24, 1961. US Patent Application 630,123 filed November 21, 1945.

Fermi:1934:ARPb

[FAD^{+}34a] Enrico Fermi, Edoardo Amaldi, Oscar D’Agostino, Franco Rasetti, and Emilio Segrè. Artificial radioactivity produced by neutron bombardment. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 146(857):483–500, September 1, 1934. ISSN 0950-1207 (print), 2053-9150 (electronic). URL http://www.jstor.org/

stable/2935604.

Fermi:1934:IRS

[FAD^{+}34b] Enrico Fermi, Edoardo Amaldi, Oscar D’Agostino, Franco Rasetti, and Emilio Segrè. Iskusstvennaya radioaktivnoct’, sozdavaemaya neitronnoi bombardirovkoi. (Russian) [Artificial radioaktivity produced by neutron bombardment]. Uspekhi Fizicheskikh Nauk, 14(8):933–952, August 1934. CODEN UF-NAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL

http://ufn.ru/ru/articles/1934/8/b/. Russian translation of

[FAD^{+}34a].

Fakley:1983:BM

[Fak83] Dennis C. Fakley. The British Mission. Los Alamos Science, 7: 186–189, Winter/Spring 1983. CODEN LASCDI. ISSN 0273-

7116. URL http://library.lanl.gov/cgi-bin/getfile?07-
21.pdf.

Falkenhagen:1928:QC

[Fal28] Hans Falkenhagen. Quantentheorie und Chemie. (German) [Quantum theory and chemistry], volume 1928 of Leipziger Vorträge. S. Hirzel, Leipzig, Germany, 1928. viii + 142 pp.

LCCN QD461 .F3. URL http://books.google.com/books?
id=x5xKAAAAMAAJ;   http://catalog.hathitrust.org/api/
volumes/oclc/4335846.html.

Fermi:1952:OES

[FAL^{+}52] Enrico Fermi, Herbert L. Anderson, Arné Lundby, Darragh E. Nagle, and Gaurang B. Yodh. Ordinary and exchange scattering of negative pions by hydrogen. Physical Review, 85(5):935–936, March 1, 1952. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://inspirehep.net/record/
47336;   http://prola.aps.org/abstract/PR/v85/i5/p935_
1.

Falicov:1973:BRE

[Fal73] L. M. Falicov. Book review: Electrons, holes, and monsters: The Fermi Surface. Its Concept, Determination, and Use in the Physics of Metals by A. P. Cracknell and K. C. Wong. Science, 182 (4118):1240, December 21, 1973. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.

org/stable/1737568.

Fermi:1934:ASIa

[FAP^{+}34] Enrico Fermi, Edoardo Amaldi, Bruno Pontecorvo, Franco Rasetti, and Emilio Segrè. Azione di sostanze idrogenate sulla radioattività provacata da neutroni. (Italian) [Action of hydrogenated substances by neutron-induced radioactivity]. La Ricerca Scientifica, 5(2):282–283, ???? 1934. CODEN RISCAZ. ISSN 0035-5011. See [Seg70b, pp. 81–82] for an English translation, and a description of how this short paper was written in just one evening, and given to the journal editor the next morning.

Fermi:1940:PPR

[FAP^{+}40] Enrico Fermi, Edoardo Amaldi, Bruno Pontecorvo, Franco Rasetti, and Emilio Segré. Process for the production of radioactive substances. US Patent 2,206,634., July 2, 1940. URL http:

//www.google.com/patents/US2206634. US Patent Application

43,462, filed October 3, 1935.

Fara:2001:GPI

[Far01] Patricia Fara. Group portraits IV — the Seventh Solvay Conference [Bruxelles, 22–29 October 1933]. Endeavour, 25(4):137–138, December 2001. CODEN ENDEAS. ISSN 0160-9327 (print), 1873-1929 (electronic).

Fermi:1938:ASN

[FAW38] Enrico Fermi, Edoardo Amaldi, and Gian Carlo Wick. On the albedo of slow neutrons. Physical Review, 53(6):493, March 15, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v53/
i6/p493_1.

Fazarinc:2015:FDB

[Faz15]  Zvonko Fazarinc.  Fermi–Dirac, Bose–Einstein, Maxwell–

Boltzmann, and computers. Computer Applications in Engineering Education, 23(5):746–759, September 2015. CO-DEN CAPEED. ISSN 1061-3773 (print), 1099-0542 (elec-

tronic). URL https://dblp.org/db/journals/caee/caee23.
html#Fazarinc15;   https://www.wikidata.org/entity/Q60032924.

Fermi:1941:RFN

[FB41] Enrico Fermi and Gregory Breit. Remarks on fast neutron reactions. Report A-46, US Atomic Energy Commission, Washington, DC, USA, October 6, 14, 1941.

Fermi:1961:GSR

[FB61] Laura Fermi and Gilberto Bernardini. Galileo and the scientific revolution. Science and discovery. Basic Books, New York, NY, USA, 1961. 150 pp. LCCN QB36.G2 F43.

Fermi:1964:SNH

[FB64] Laura Fermi and Gilberto Bernardini. SCRAPS: Nineteen hundred and sixty-four: The four-hundredth anniversary of the birth of Galileo Galilei. The Physics Teacher, 2:140–??, 1964. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic). URL

http://tpt.aapt.org/resource/1/phteah/v2/i3/p140_s1.

Fermi:1969:CCV

[FB69] Laura Fermi and Gilberto Bernardini. Che cosa ha veramento detto Galileo. (Italian) [What did Galileo really say?]. Ubaldini, Roma, Italy, 1969. 133 pp. LCCN QB36.G2 F42.

Fermi:2003:GSR

[FB03] Laura Fermi and Gilberto Bernardini. Galileo and the scientific revolution. Dover, New York, NY, USA, 2003. ISBN 0-486-43226-2 (paperback). vi + 128 pp. LCCN QB36.G2

F43 2003. URL http://www.loc.gov/catdir/enhancements/
fy0614/2003050231-d.html.

Fermi:1934:NDD

[FBE^{+}34] Enrico Fermi, Hans A. Bethe, Walter M. Elsasser, H. O. W. Richardson, and K. Sitte. Natural β-decay: Discussion. In International Conference on Physics, London, Vol. I. Nuclear Physics, pages 66–71. Physical Society, London, UK (??), 1934.

Fermi:1964:BFR

[FBL64] Laura Fermi, Gilberto Bernardini, and Noel C. Little. Book and film reviews: Galileo and the scientific revolution. The Physics Teacher, 2:180–??, 1964. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic). URL http://tpt.aapt.org/

resource/1/phteah/v2/i4/p180_s1.

Fermi:1941:NP

[FBR^{+}41] Enrico Fermi, Gregory Breit, I. I. Rabi, Eugene P. Wigner, and John H. van Vleck, editors. Nuclear physics. University Press of Pennsylvania, Philadelphia, PA, USA, 1941. v + 68 pp.

Fermi:1921:SDS

[Fer21a] Enrico Fermi. Sulla dinamica di un sistema rigido di cariche elettriche in moto traslatorio. (Italian) [On the dynamics of a rigid system of electric charges in translational motion]. Il Nuovo Cimento (6), 22(1):199–207, ???? 1921. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://www.

springerlink.com/content/03523mx3119g89m6/.

Fermi:1921:SCG

[Fer21b] Enrico Fermi. Sull’elettrostatica di un campo gravitazionale uniforme e sul peso delle masse elettromagnetiche. (Italian) [On the electrostatics of a uniform gravitational field and the weight of electromagnetic masses]. Il Nuovo Cimento (6), 22(1):176– 188, ???? 1921. CODEN NUCIAD. ISSN 0029-6341 (print),

1827-6121 (electronic). URL http://www.springerlink.com/
content/fn2lk82543252j6t/.

Fermi:1922:CGD

[Fer22a] Enrico Fermi. Correzione di una grave discrepanza tra la teoria delle masse elettromagnetiche e la teoria della relatività. Inerzia e peso dell’elettricità. I, II. (Italian) [Correction of a serious discrepancy between theory of the masses and the electromagnetic theory of relativity. Inertia and weight of electricity. I, II]. Rendiconti dell’Accademia Nazionale dei Lincei, 31(1):184–187, 306– 309, 1922. CODEN AANLAW. ISSN 0001-4435.

Fermi:1922:RRI

[Fer22b] Enrico Fermi. I raggi Röntgen. (Italian) [X-rays]. Il Nuovo Cimento (6), 24(1):133–163, ???? 1922. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://www.

springerlink.com/content/533p81w97245h921/.

Fermi:1922:SFC

[Fer22c] Enrico Fermi. Sopra i fenomeni che avvengono in vicinanza di una linea oraria. (Italian) [On the phenomena that occur in the neighborhood of a world line]. Rendiconti dell’Accademia Nazionale dei Lincei, 31(1):21–23, 51–52, 101–103, 1922. CODEN AANLAW. ISSN 0001-4435. Source unknown. See [Fer66c] for an English translation, and images of the original document in Italian.

Fermi:1922:SSR

[Fer22d] Enrico Fermi. Studii sopra i raggi Röntgen. (Italian) [Studies on X-rays]. Tesi di laurea, Reale Scuola Superiore Normale, Università di Pisa, Pisa, Italy, June 20, 1922. Thesis manuscript in the library of the University of Pisa. This thesis was lost for many years (a fact lamented in volume 1 of the Collected Papers of Enrico Fermi [Fer62a, page 35]), and only recently discovered by Roberto Vergara Caffarelli at the University of Pisa; the book spine was incorrectly labeled ‘Terni’ instead of ‘Fermi’ [DS09, page 5]. Chapter 1 was published in [Fer22b] and Chapter 4 in [Fer23f].

Fermi:1922:STD

[Fer22e] Enrico Fermi. Sulla teoria dell’influenza della temperatura sopra la diffrazione dei raggi Röntgen nei cristalli. (Italian) [On the theory of the influence of temperature on the diffraction of X-rays in crystals]. Manuscript in the Pisa archives [DS09, page 4]., 1922.

Fermi:1922:WZE

[Fer22f] Enrico Fermi. Über einen Widerspruch zwischen der elektrodynamischen und relativistischen Theorie der elektromagnetischen Masse. (German) [Discrepancy between electrodynamic relativist theories of electromagnetic mass]. Physikalische Zeitschrift, 23(17):340–344, September 1, 1922. CODEN PHZ-

TAO. ISSN 0369-982X. URL http://en.wikisource.org/wiki/
Electrodynamic_and_Relativistic_Theory_of_Electromagnetic_
Mass;   http://hdl.handle.net/2027/mdp.39015086723239.

Fermi:1922:PDT

[Fer22g] Enrico Fermi. Un perfezionamento della teoria di Debye dell’influenza dell’agitazione termica sopra la diffrazione. (Italian) [A refinement of the theory of Debye on the influence of thermal diffraction]. Manuscript in the Pisa archives [DS09, page 4]., 1922.

Fermi:1922:TCD

[Fer22h] Enrico Fermi. Un teorema di calcolo delle probabilità ed alcune sue applicazioni. (Italian) [A theorem on probability and some of its applications]. Tesi di abilitazione, Scuola Normale Superiore di Pisa, Pisa, Italy, July 1922. Diplome in mathematics (cum laude), published later [Fer26e].

Fermi:1923:ATM

[Fer23a] Enrico Fermi. Alcuni teoremi di meccanica analitica importanti per la teoria dei quanti. (Italian) [Some important theorems of analytical mechanics to quantum theory]. Il Nuovo Cimento (6), 25(1):271–285, ???? 1923. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://www.

springerlink.com/content/pr766v13267h1157/.

Fermi:1923:BDM

[Fer23b] Enrico Fermi. Beweis, daß ein mechanisches Normalsystem im allgemeinen quasi-ergodisch ist. (German) [Proof that a mechanical normal system is in general quasi-ergodic]. Physikalische Zeitschrift, 24(??):261–265, ???? 1923. CODEN PHZTAO. ISSN 0369-982X.

Fermi:1923:CCT

[Fer23c] Enrico Fermi. Correzione di una contraddizione tra la teoria elettrodinamica e quella relativistica delle masse elettromagnetiche. (Italian) [Correction of a contradiction between the theory of electrodynamics and relativistic electromagnetic mass]. Il Nuovo Cimento (6), 25(1):159–170, ???? 1923. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://www.

springerlink.com/content/6h82m8g485410636/.

Fermi:1923:DCGa

[Fer23d] Enrico Fermi. Dimostrazione che in generale un sistema meccanico normale è quasi-ergodico. (Italian) [Demonstration that in general a normal mechanical system is quasi-ergodic]. Il Nuovo Cimento (6), 25(6):1–5, ???? 1923. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://www.

springerlink.com/content/7223216276011616/.

Fermi:1923:DCGb

[Fer23e] Enrico Fermi. Dimostrazione che in generale un sistema meccanico normale è quasi-ergodico. (Italian) [Demonstration that in general a normal mechanical system is quasi-ergodic]. Il Nuovo Cimento (6), 25(7):267–269, ???? 1923. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://

www.springerlink.com/content/7223216276011616/.

Fermi:1923:FIC

[Fer23f] Enrico Fermi. Formazione di immagini coi raggi Röntgen. (Italian) [Image formation with X-rays]. Il Nuovo Cimento (6), 25(1): 63–68, ???? 1923. CODEN NUCIAD. ISSN 0029-6341 (print),

1827-6121 (electronic). URL http://www.springerlink.com/
content/3242726132792134/.

Fermi:1923:GDT

[Fer23g] Enrico Fermi. Generalizzazione del teorema di Poincaré sopra la non esistenza di integerali uniformi di un sistema di equazioni canoniche normali. (Italian) [Generalization of the theorem of Poincaré on the nonexistence of uniform integrals of a normal canonical system of equations]. Il Nuovo Cimento (6), 26(1):105– 115, ???? 1923. CODEN NUCIAD. ISSN 0029-6341 (print),

1827-6121 (electronic). URL http://www.springerlink.com/
content/c264617500h5k701/.

Fermi:1923:PDA

[Fer23h] Enrico Fermi. Il principio delle adiabatiche ed i sistemi che non ammettono coordinate angolari. (Italian) [The adiabatic principle and systems that do not allow angular coordinates]. Il Nuovo Cimento (6), 25(1):171–175, ???? 1923. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://

www.springerlink.com/content/7884315r31620004/.

Fermi:1923:MNT

[Fer23i] Enrico Fermi. Le masse nella teoria della relatività. In I fondamenti della relatività Einsteiniana: valore e interpretazione della teoria. (Italian) [The foundations of Einsteinian relativity: value and interpretation of the theory] [Kop23], pages 342–344. LCCN ???? Appendix.

Fermi:1923:STSa

[Fer23j] Enrico Fermi. Sopra la teoria di Stern della costante assoluta dell’entropia di un gas perfetto monoatomico. (Italian) [On Stern’s theory of the absolute constancy of the entropy of a perfect monatomic gas]. Rendiconti dell’Accademia Nazionale dei Lincei, 32(2):395–398, ???? 1923. CODEN AANLAW. ISSN 0001-4435.

Fermi:1923:SPC

[Fer23k] Enrico Fermi. Sul peso dei corpi elastici. (Italian) [On the weight of elastic bodies]. Memorie Lincei, 14(5):114–124, ???? 1923.

Fermi:1923:STD

[Fer23l] Enrico Fermi. Sul trascinamento del piano di polarizzazione da parte di un mezzo rotante. (Italian) [On the rotation of the plane of polarization of light in a rotating medium]. Rendiconti dell’Accademia Nazionale dei Lincei, 32(5):115–118, ???? 1923. CODEN AANLAW. ISSN 0001-4435.

Fermi:1923:SPD

[Fer23m] Enrico Fermi. Sulla probabilità degli stati quantici. (Italian) [On the probability of quantum states]. Rendiconti dell’Accademia Nazionale dei Lincei, 32(5, ii):493–495, ???? 1923. CODEN AAN-LAW. ISSN 0001-4435.

Fermi:1923:STSb

[Fer23n] Enrico Fermi. Sulla teoria statistica di Richardson dell’effetto fotoelettrico. (Italian) [On the Richardson statistical theory of the photoelectric effect]. Il Nuovo Cimento (6), 26(1):97–104, ???? 1923. CODEN NUCIAD. ISSN 0029-6341 (print),

1827-6121 (electronic). URL http://www.springerlink.com/
content/l0510q70l8367582/.

Fermi:1924:BIV

[Fer24a] Enrico Fermi. Berekingen over de intensiteiten van spektraallijnen. (Dutch) [Calculations of the intensities of spectral lines]. Physika, 4(??):340–343, ???? 1924.

Fermi:1924:CSQ

[Fer24b] Enrico Fermi. Considerazioni sulla quantizzazione dei sistemi che contengono degli elementi identici. (Italian) [Considerations on the quantization of systems that contain identical elements]. Il Nuovo Cimento (8), 1(1):145–152, December 1924. CODEN NU-CIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL

http://www.springerlink.com/content/97203r302q8w7unj/.

Fermi:1924:ORR

[Fer24c] Enrico Fermi. Optical resonance reflection and diffusion. Rendiconti dell’Accademia Nazionale dei Lincei, 33(1):90–93, ???? 1924. CODEN AANLAW. ISSN 0001-4435.

Fermi:1924:SRD

[Fer24d] Enrico Fermi. Sopra la riflessione e la diffusione di risonanza. (Italian) [On the reflection and diffusion of resonance radiation]. Rendiconti dell’Accademia Nazionale dei Lincei, 33(1):90–93, ???? 1924. CODEN AANLAW. ISSN 0001-4435.

Fermi:1924:STI

[Fer24e] Enrico Fermi. Sull’equilibrio termico di ionizzazione. (Italian) [On thermal equilibrium ionization]. Il Nuovo Cimento (8), 1(1):153– 158, ???? 1924. CODEN NUCIAD. ISSN 0029-6341 (print),

1827-6121 (electronic). URL http://www.springerlink.com/
content/ag34574751012h22/.

Fermi:1924:EQE

[Fer24f] Enrico Fermi. Úber die Existenz quasi-Ergodischer Systeme. (German) [On the existence of quasi-ergodic systems]. Physikalische Zeitschrift, 25(??):166–167, ???? 1924. CODEN PHZTAO. ISSN 0369-982X.

Fermi:1924:TSZ

[Fer24g] Enrico Fermi. Über die Theorie des Stoßes zwischen Atomen und elektrisch geladenen Teilchen. (German) [On the theory of the collisions between atoms and electrically-charged particles]. Zeitschrift für Physik, 29(1):315–327, December 1924. CODEN ZEPYAA. ISSN ???? URL http://

inspirehep.net/record/2807;   http://www.springerlink.
com/content/9302470h61p55636/.

Fermi:1924:WQG

[Fer24h] Enrico Fermi. Über die Wahrscheinlichkeit der Quantenzustände. (German) [On the probability of quantum states]. Zeitschrift für Physik, 26(1):54–56, ???? 1924. CODEN ZEPYAA.

ISSN ????  URL http://www.springerlink.com/content/
jq848v33515m123l/.

Fermi:1925:EPL

[Fer25a] Enrico Fermi. Elettrodinamica: prima lezione integrale del dattiloscritto del Corso di Fisica Matematica (Lezioni ei Eletrodinamica e Teoria della Relatività) del 1924–1925. (Italian) [Electrodynamics: the first lesson of the full manuscript of the Course of Mathematical Physics (Lessons and Theory of Relativity and Electrodynamics) of 1924–1925]. Presso l’Università di Firenze, Firenze, Italy, 1925. ???? pp. LCCN ????

Fermi:1925:STC

[Fer25b] Enrico Fermi. Sopra la teoria dei corpi solidi. (Italian) [On the theory of solid bodies]. Periodico di Matematiche, 5(4):264–274, ???? 1925.

Fermi:1925:SID

[Fer25c] Enrico Fermi. Sopra l’intensità delle righe multiple. (Italian) [On the intensities of multiple lines]. Rendiconti dell’Accademia Nazionale dei Lincei, 1(6):120–124, 1925. CODEN AANLAW. ISSN 0001-4435.

Fermi:1925:SUT

[Fer25d] Enrico Fermi. Sopra l’urto tra atomi e nuclei di idrogeno. (Italian) [On collisions between atoms and hydrogen nuclei]. Rendiconti Accad. d. L. Roma (6), 1(6):77–80, 1925.

Fermi:1925:SPD

[Fer25e] Enrico Fermi. Sui principi della teoria dei quanti. (Italian) [On the principles of the quantum theory]. Rendiconti del Seminario matematico dell’Università di Roma, 8:7–12, ???? 1925.

Fermi:1925:STD

[Fer25f] Enrico Fermi. Sulla teoria dell’urto tra atomi e corpuscoli elettrici. (Italian) [On the theory of collisions between atoms and electrically-charged particles]. Il Nuovo Cimento (8), 2(2):143– 158, ???? 1925. CODEN NUCIAD. ISSN 0029-6341 (print),

1827-6121 (electronic). URL http://www.springerlink.com/

content/d4k7020l17x76169/. English translation in [Fer01].

Fermi:1925:RTC

[Fer25g] Enrico Fermi. Una relazione tra le costanti delle bande infrarosse delle molecole triatomiche. (Italian) [A relation between the constants of the infra-red bands of triatomic molecules]. Rendiconti dell’Accademia Nazionale dei Lincei, 1(6):386–387, ???? 1925. CODEN AANLAW. ISSN 0001-4435.

Fermi:1926:APC

[Fer26a] Enrico Fermi. Argomenti pro e contro la ipotesi dei quanti di luce. (Italian) [Arguments for and against the hypothesis of light quanta]. Il Nuovo Cimento (8), 3(1–2):47–54, ???? 1926. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL

http://www.springerlink.com/content/lm022085605043uh/.

Fermi:1926:PCN

[Fer26b] Enrico Fermi. Problemi di chimica nella fisica dell’atomo. (Italian) [The problems of chemistry in atomic physics]. Periodico di Matematiche, 6(4):19–26, ???? 1926.

Fermi:1926:QMM

[Fer26c] Enrico Fermi. Quantum mechanics and the magnetic moment of atoms. Nature, 118(2981):876, December 18, 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v118/
n2981/pdf/118876a0.pdf.

Fermi:1926:SID

[Fer26d] Enrico Fermi. Sopra l’intensità delle righe proibite nei campi magnetici intensi. (Italian) [On the intensity of forbidden lines in strong magnetic fields]. Rendiconti dell’Accademia Nazionale dei Lincei, 3(6):478–483, ???? 1926. CODEN AANLAW. ISSN 0001-4435.

Fermi:1926:SFC

[Fer26e] Enrico Fermi. Sopra una formula di calcolo delle probabilità. (Italian) [On a formula for calculating the probability]. Il Nuovo Cimento (8), 3(1–2):313–318, ???? 1926. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://

www.springerlink.com/content/41428t2uk9671532/. This pa-

per is the publication of Fermi’s Diplome degree [Fer22h].

Fermi:1926:SQD

[Fer26f] Enrico Fermi. Sulla quantizzazione del gas perfetto monoatomico. (Italian) [On the quantization of the ideal monatomic gas]. Rendiconti dell’Accademia Nazionale dei Lincei, 3(6):145–149, ???? 1926. CODEN AANLAW. ISSN 0001-4435.

Fermi:1926:QIE

[Fer26g] Enrico Fermi. Zur Quantelung des idealen einatomigen Gases. (German) [On the quantization of an ideal monatomic gas]. Zeitschrift für Physik, 36(11–12):902–912, November 1926. CO-

DEN ZEPYAA. ISSN ???? URL http://www.springerlink.
com/content/k763270092273181/.

Fermi:1926:WSG

[Fer26h] Enrico Fermi. Zur Wellenmechanik des Stoßvorganges. (German) [On the wave mechanics of collision processes]. Zeitschrift für Physik, 40(5):399–402, May 1926. CODEN ZEPYAA.

ISSN ????  URL http://www.springerlink.com/content/
wg37207762047g85/.

Fermi:1926:LFT

[Fer27a] Enrico Fermi. Lezioni di Fisica teorica. (Italian) [Lectures on theoretical physics]. Stabilimento Tipolitografico del Genio Civile, Roma, Italia, 1926–1927. 60 pp. Notes by Dei and Martinozzi.

Fermi:1927:APC

[Fer27b] Enrico Fermi. Argomenti pro e contro la ipotesi dei quanti di luce. (Italian) [Arguments for and against the hypothesis of light quanta]. Rendiconti Seminario Mat. Roma (2), 4(??):31–32, ???? 1927.

Fermi:1927:EEM

[Fer27c] Enrico Fermi. Gli effetti elettro e magnetoottici e le loro interpretazioni. (Italian) [Electro- and magnetooptical effects and their interpretation]. L’Energia Elettrica, ??(??):109–120, ???? 1927. Special issue on the centenary of the death of A. Volta.

Fermi:1927:SMD

[Fer27d] Enrico Fermi. Sul meccanismo dell’emissione nella meccanica ondulatoria. (Italian) [Mechanism of emission according to undulatory mechanics]. Rendiconti dell’Accademia Nazionale dei Lincei, 5(6):795–800, ???? 1927. CODEN AANLAW. ISSN 0001-4435.

Fermi:1927:MSP

[Fer27e] Enrico Fermi. Un metodo statistico per la determinazione di alcune proprietá dell’atomo. (Italian) [A statistical method for the determination of some atomic properties]. Rendiconti dell’Accademia Nazionale dei Lincei, 6(6):602–607, ???? 1927.

CODEN AANLAW. ISSN 0001-4435. This paper is the Fermi side of the Thomas–Fermi model, which Fermi developed independently a year after the work of Llewellyn H. Thomas ([Tho27]). See [Seg70b, page 53], [GR08], and [DS09] for historical background.

Fermi:1928:AGP

[Fer28a] Enrico Fermi. Anomalous groups in the periodic system of elements. Nature, 121(3048):502, March 31, 1928. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v121/
n3048/pdf/121502b0.pdf.

Fermi:1928:SMB

[Fer28b] Enrico Fermi. Eine statistische Methode zur Bestimmung einiger Eigenschaften des Atoms und ihre Anwendung auf die Theorie des periodischen Systems der Elemente. (German) [A statistical method for the determination of some atomic properties and its application to the theory of the periodic system of elements]. Zeitschrift für Physik, 48(1–2):73–79, January 1928. CODEN

ZEPYAA. ISSN ???? URL http://www.springerlink.com/

content/v762582061464612/. See English translation in [Mar75, pp. 205–213].

Fermi:1928:IAF

[Fer28c] Enrico Fermi. Introduzione alla Fisica Atomica. (Italian) [Introduction to atomic physics]. N. Zanichelli, Bologna, Italy, 1928. 330 pp. LCCN ????

Fermi:1928:PMC

[Fer28d] Enrico Fermi. Problemi matematici connessi alla nuova meccanica. (Italian) [Mathematical problems associated with the new mechanics]. Rendiconti Seminario Mat. Roma, 5(??):23, ???? 1928.

Fermi:1928:SBR

[Fer28e] Enrico Fermi. Statistische Berechnung der Rydbergkorrektionen der s-Terme. (German) [Statistical calculation of the Rydberg correction s-terme]. Zeitschrift für Physik, 49(7–8):550–554, ???? 1928. CODEN ZEPYAA. ISSN ???? URL http://www.

springerlink.com/content/g467772270t34273/.

Fermi:1928:SDSa

[Fer28f] Enrico Fermi. Sulla deduzione statistica di alcune proprietà dell’atomo. Applicazione alla teoria del sistema periodico degli elementi. (Italian) [On statistical inference of some atomic properties. Applying the theory to the periodic system of elements]. Rendiconti dell’Accademia Nazionale dei Lincei, 7(6):342–346, ???? 1928. CODEN AANLAW. ISSN 0001-4435.

Fermi:1928:SDSb

[Fer28g] Enrico Fermi. Sulla deduzione statistica di alcune proprietá dell’atomo. Calcolo della correzione di Rydberg per i termini s. (Italian) [On the statistical deduction of some atomic properties. Calculation of the Rydberg corrections to s-terms]. Rendiconti dell’Accademia Nazionale dei Lincei, 7(6):726–730, ???? 1928. CODEN AANLAW. ISSN 0001-4435.

Fermi:1928:STD

[Fer28h] Enrico Fermi. Sulla teoria del sistema periodico. (Italian) [On the theory of the periodic system]. Rendiconti dell’Accademia Nazionale dei Lincei, 7(6):66–??, ???? 1928. CODEN AANLAW. ISSN 0001-4435.

Fermi:1928:ASM

[Fer28i] Enrico Fermi. Über die Anwendung der statistischen Methode auf die Probleme des Atombaues. (German) [On the application of statistical methods to the problems of atomic structure]. In Quantentheorie und Chemie. (German) [Quantum theory and chemistry] [Fal28], pages 95–111. LCCN QD461 .F3. URL

http://books.google.com/books?id=x5xKAAAAMAAJ;   http:
//catalog.hathitrust.org/api/volumes/oclc/4335846.html.

Leipziger Vorträge.

Fermi:1929:AET

[Fer29a] Enrico Fermi. Affinità elettronica e teoria statistica dell’atomo. (Italian) [Electron affinity and statistical theory of the atom]. Il Nuovo Cimento (8), ??(??):13–14, 1929. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Fermi:1929:FPL

[Fer29b] Enrico Fermi. Fisica per i Licei. (Italian) [Physics for High Schools]. N. Zanichelli, Bologna, Italy, 1929. 239 (volume 1) + 243 (volume 2) pp. LCCN ???? Two volumes.

Fermi:1929:FSD

[Fer29c] Enrico Fermi. I fondamenti sperimentale delle nuove teorie fisiche. (Italian) [The experimental foundations of the new physical theories]. Atti Soc. It. Progr. Sci. 18a Riunione, 1:365–371, 1929.

Fermi:1929:LDF

[Fer29d] Enrico Fermi. L’interpretazione del fenomeno dell’irradiazione e dell’assorbimento nella attuale teoria dei quanti. (Italian) [The interpretation of the phenomenon of radiation and absorption in the current quantum theory]. Il Nuovo Cimento (8), ??:16–17, 1929. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Fermi:1929:PAD

[Fer29e] Enrico Fermi. Problemi attuali della fisica. (Italian) [The contemporary problems of physics]. Annali dell’Istruzione media, 5(??): 424–428, ???? 1929.

Fermi:1929:SEQ

[Fer29f] Enrico Fermi. Sopra l’elettrodinamica quantistica. (Italian) [On quantum electrodynamics]. Rendiconti dell’Accademia Nazionale dei Lincei, 9(6):881–887, ???? 1929. CODEN AANLAW. ISSN 0001-4435.

Fermi:1929:SCD

[Fer29g] Enrico Fermi. Sul complesso 4d della molecola di elio. (Italian) [On complex 4d terms of the helium molecule]. Rendiconti dell’Accademia Nazionale dei Lincei, 10(??):515–517, ???? 1929. CODEN AANLAW. ISSN 0001-4435.

Fermi:1929:SMC

[Fer29h] Enrico Fermi. Sul moto di un corpo di massa variabile. (Italian) [On the motion of a body of variable mass]. Rendiconti Accad. d. L. Roma (6), 9(6):984–986, ???? 1929.

Fermi:1929:STQ

[Fer29i] Enrico Fermi. Sulla teoria quantistica delle frange di interferenza. (Italian) [On the quantum theory of interference fringes]. Rendiconti dell’Accademia Nazionale dei Lincei, 10(6):72–77, ???? 1929. CODEN AANLAW. ISSN 0001-4435.

Fermi:1930:ASI

[Fer30a] Enrico Fermi. Atomi e stelle. (Italian) [Atoms and stars]. Atti Società Italiana per il Progresso delle Scienze, 19a Riunione, 1: 228–235, 1930.

Fermi:1930:FSD

[Fer30b] Enrico Fermi. I fondamenti sperimentale della nuova meccanica atomica. Periodico di Matematiche, 10(??):71–84, ???? 1930.

Fermi:1930:FMI

[Fer30c] Enrico Fermi. La fisica moderna. (Italian) [Modern physics]. Nuova Antologia, 65(??):137–145, November 1930.

Fermi:1930:LDPa

[Fer30d] Enrico Fermi. L’interpretazione del principio di causalità nella meccanica quantistica. (Italian) [The interpretation of the principle of causality in quantum mechanics]. Rendiconti dell’Accademia Nazionale dei Lincei, 11(??):980–985, ???? 1930. CODEN AAN-LAW. ISSN 0001-4435.

Fermi:1930:LDPb

[Fer30e] Enrico Fermi. L’interpretazione del principio di causalità nella meccanica quantistica. (Italian) [The interpretation of the principle of causality in quantum mechanics]. Il Nuovo Cimento (8), 7(10):361–366, ???? 1930. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Fermi:1930:MMAb

[Fer30f] Enrico Fermi. Magnetic moments of atomic nuclei. Nature, 125 (3140):16, January 4, 1930. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v125/n3140/pdf/125016a0.pdf.

Fermi:1930:SEQ

[Fer30g] Enrico Fermi. Sopra l’elettrodinamica quantistica. (Italian) [On quantum electrodynamics]. Rendiconti dell’Accademia Nazionale dei Lincei, 12(6):431–435, ???? 1930. CODEN AANLAW. ISSN 0001-4435.

Fermi:1930:SMM

[Fer30h] Enrico Fermi. Sui momenti magnetici dei nuclei atomici. (Italian) [On the magnetic moments of atomic nuclei]. Mem. Accad. d’Italia, 1 (Fis.)(??):139–148, ???? 1930.

Fermi:1930:SCDb

[Fer30i] Enrico Fermi. Sul calcolo degli spettri degli ioni. (Italian) [On the calculation of ionic spectra]. Reale Accademia d’Italia. Memorie della Classe di Scienze Fisiche Matematische e Naturali, 1 (Fis.) (??):149–156, ???? 1930.

Fermi:1930:SCDc

[Fer30j] Enrico Fermi. Sul calcolo degli spettri degli ioni. (Italian) [On the calculation of ionic spectra]. Il Nuovo Cimento (8), 8(??):7–14, ???? 1930. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Fermi:1930:SCDa

[Fer30k] Enrico Fermi. Sul complesso 4d della molecola di elio. (Italian) [On the complex 4d terms of the helium molecule]. Il Nuovo Cimento (8), 7(1):159–161, ???? 1930. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://www.

springerlink.com/content/qu2v070619524651/.

Fermi:1930:SPC

[Fer30l] Enrico Fermi. Sul principio di causalità nella meccanica ondulatoria. (Italian) [On the principle of causality in wave-mechanics]. Rendiconti dell’Accademia Nazionale dei Lincei, 11(6):980–??, ???? 1930. CODEN AANLAW. ISSN 0001-4435.

Fermi:1930:SRD

[Fer30m] Enrico Fermi. Sul rapporto delle intensità nei doppietti dei metalli alcalini. (Italian) [On the intensity ratio of the doublets of alkali metals]. Il Nuovo Cimento (8), 7(2):201–207, ???? 1930. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL

http://www.springerlink.com/content/n634245424810826/.

Fermi:1930:STQ

[Fer30n] Enrico Fermi. Sulla teoria quantistica delle frange di interferenza. (Italian) [On the quantum theory of interference fringes]. Il Nuovo Cimento (8), 7(1):153–158, ???? 1930. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://www.

springerlink.com/content/11u3u0132r2167u8/.

Fermi:1930:IDA

[Fer30o] Enrico Fermi. Über das Intensitätsverhältnis der Dublettkomponenten der Alkalien. (German) [On the intensity ratio of the alkali doublet]. Zeitschrift für Physik, 59(9–10):680–686, ???? 1930.

CODEN ZEPYAA. ISSN ???? URL http://www.springerlink.
com/content/u61157208733h542/.

Fermi:1930:MMAa

[Fer30p] Enrico Fermi. Über die magnetischen Momente der Atomkerne. (German) [On the magnetic moment of the atomic nucleus]. Zeitschrift für Physik, 60(5–6):320–333, ???? 1930. CODEN

ZEPYAA. ISSN ???? URL http://www.springerlink.com/
content/r2xvj07l47qj3005/.

Fermi:1931:TDR

[Fer31a] Enrico Fermi. La théorie du rayonnement. (French) [The theory of radiation]. Annales de l’Institut Henri Poincaré, 1(1):53–74, 1931. CODEN AIHPA2. ISSN 0365-320x (print), 2400-4855 (elec-

tronic). URL http://www.numdam.org/item?id=AIHP_1930__
1_1_53_0.

Fermi:1931:MEN

[Fer31b] Enrico Fermi. Le masse elettromagnetiche nella elettrodinamica quantistica. (Italian) [The electromagnetic mass in quantum electrodynamics]. Il Nuovo Cimento (8), 8(1):121–132, ???? 1931. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121

(electronic). URL http://www.springerlink.com/content/
666766873474t135/.

Fermi:1931:SEQ

[Fer31c] Enrico Fermi. Sopra l’elettrodinamica quantistica. Il Nuovo Cimento (8), 8(??):207–219, ???? 1931. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Fermi:1931:SCD

[Fer31d] Enrico Fermi. Sul calcolo degli spettri degli ioni. (Italian) [On the calculation of spectra of ions]. Il Nuovo Cimento (8), 8(1): 7–14, ???? 1931. CODEN NUCIAD. ISSN 0029-6341 (print),

1827-6121 (electronic). URL http://www.springerlink.com/
content/q7655qwn747j9635/.

Fermi:1931:RKG

[Fer31e] Enrico Fermi. Über den Ramaneffekt des Kohlendioxyds. (German) [On the Raman effect in carbon dioxide]. Zeitschrift für Physik, 71(3–4):250–259, March 1931. CODEN ZEPYAA.

ISSN ????  URL http://www.springerlink.com/content/
m55p73766721t070/.

Fermi:1932:EAP

[Fer32a] Enrico Fermi. État actuel de la physique du noyau atomique. (French) [The current state of physics of the atomic nucleus]. In Robert de Valbreuze, editor, Comptes Rendus du Congrès International d’Electricité Paris 1932, pages 789–807. Gauthier-Villars et cie, Paris, France, 1932.

Fermi:1932:LRN

[Fer32b] Enrico Fermi. L’effetto Raman nelle molecole e nei cristalli. (Italian) [The Raman effect in molecules and crystals]. Mem. Accad. d’Italia, 3 (Fis.)(??):239–256, ???? 1932.

Fermi:1932:SAO

[Fer32c] Enrico Fermi. Lo stato attuale di oscillazione e rotazione dell’ammoniaca. (Italian) [The current state of oscillation and rotation in ammonia]. La Ricerca Scientifica, 3(2):101–113, ???? 1932. CODEN RISCAZ. ISSN 0035-5011.

Fermi:1932:QTR

[Fer32d] Enrico Fermi. Quantum theory of radiation. Reviews of Modern Physics, 4(1):87–132, January 1, 1932. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
4.87;   http://rmp.aps.org/abstract/RMP/v4/i1/p87_1.

Fermi:1932:SBOa

[Fer32e] Enrico Fermi. Sulle bande di oscillazione e rotazione dell’ammoniaca. (Italian) [Vibration and rotation bands of ammonia]. Rendiconti dell’Accademia Nazionale dei Lincei, 16(6):179–185, ???? 1932. CODEN AANLAW. ISSN 0001-4435.

Fermi:1932:SBOb

[Fer32f] Enrico Fermi. Sulle bande di oscillazione e rotazione dell’ammoniaca. (Italian) [Vibration and rotation bands of ammonia]. Il Nuovo Cimento (8), 9(9):277–283, ???? 1932. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://www.

springerlink.com/content/g67688w169u75n61/.

Fermi:1933:UPC

[Fer33a] Enrico Fermi. Le ultime particelle costitutive della materia. (Italian) [The ultimate constituents of matter]. Atti Società Italiana per il Progresso delle Scienze, 22a Riunione, 2(??):7–14, ???? 1933. 22nd meeting, Bari (October 12–18, 1933). See also [Fer34e].

Fermi:1933:TTD

[Fer33b] Enrico Fermi. Tentativo di una teoria dell’emissione dei raggi ‘beta’. (Italian) [Attempt at a theory of the emission of β rays]. Ricerca Scientifica ed il Progresso Tecnico, 4(2):491–495, 1933. CODEN RSPTB6. ISSN ???? See [Jen00] for a history of the theory of beta decay.

Fermi:1933:THS

[Fer33c] Enrico Fermi. Theory of hyperfine structure [minutes of the Chicago Meeting, June 19–24, 1933]. Physical Review, 44(4): 313–330, August 1933. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.org/

doi/10.1103/PhysRev.44.313. Only title appears in the min-

utes.

Fermi:1934:ARN

[Fer34a] Enrico Fermi. Artificial radioactivity produced by neutron bombardment. In International Conference on Physics, London, Vol. I. Nuclear Physics, pages 75–77. Physical Society, London, UK, 1934.

Fermi:1934:CL

[Fer34b] Enrico Fermi. Conferencias (lectures). Faculdad de Ciencias Exactas, Físicas y Naturales, Serie B, Publicación 15, Buenos Aires, Argentina, 1934.

Fermi:1934:EN

[Fer34c] Enrico Fermi. Element no. 93. Nature, 133(3371):863–864, June 9, 1934. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v133/n3371/pdf/133863e0.pdf.

Fermi:1934:RAI

[Fer34d] Enrico Fermi. La radioattività artificiale. (Italian) [Artificial radioactivity]. Atti Società Italiana per il Progresso delle Scienze, 23a Riunione, 1:34–39, 1934.

Fermi:1934:UPC

[Fer34e] Enrico Fermi. Le ultime particelle costitutive della materia. (Italian) [The ultimate constituents of matter]. Scientia (Milan), 55 (261):21–28, January 1934. CODEN SCIMAI. ISSN 0036-8687 (print), 1825-4373 (electronic). URL http://amshistorica. unibo.it/7. See also [Fer33a].

Fermi:1934:MCI

[Fer34f] Enrico Fermi. Molecole e cristalli. (Italian) [Molecules and crystals]. N. Zanichelli, Bologna, Italy, 1934. 303 pp. LCCN ????

Fermi:1934:PPE

[Fer34g] Enrico Fermi. Possible production of elements of atomic number higher than 92. Nature, 133(3372):898–899, June 16, 1934. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.chemteam.info/Chem-History/
Fermi-transuranics-1934.html;   http://www.nature.com/
nature/journal/v133/n3372/pdf/133898a0.pdf.

Fermi:1934:RIN

[Fer34h] Enrico Fermi. Radioactivity induced by neutron bombardment. Nature, 133(3368):757, May 19, 1934. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v133/n3368/pdf/133757a0.pdf.

Fermi:1934:RPN

[Fer34i] Enrico Fermi. Radioactivity produced by neutron bombardment. Nature, 134(??):668, ???? ??, 1934. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Fermi:1934:RPBa

[Fer34j] Enrico Fermi. Radioattività indotta da bombardamento di neutroni. I (Italian) [Radioactivity induced by neutron bombardment. I]. La Ricerca Scientifica, 5(1):283, ???? 1934. CODEN RISCAZ. ISSN 0035-5011.

Fermi:1934:RPBc

[Fer34k] Enrico Fermi. Radioattività prodotta da bombardamento di neutroni. II. (Italian) [Radioactivity produced by neutron bombardment. II]. Il Nuovo Cimento (8), 11(7):429–441, ???? 1934. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121

(electronic). URL http://www.springerlink.com/content/
q107q23kqk234806/.

Fermi:1934:RPBb

[Fer34l] Enrico Fermi. Radioattività provocata da bombardamento di neutroni. II. (Italian) [Radioactivity caused by neutron bombardment. II]. La Ricerca Scientifica, 5(1):330–331, ???? 1934. CO-DEN RISCAZ. ISSN 0035-5011.

Fermi:1934:SSP

[Fer34m] Enrico Fermi. Sopra lo spostamento per pressione delle righe elevate delle serie spettrali. (Italian) [On the movement under high pressure of the spectral-line series]. Il Nuovo Cimento (8), 11(3): 157–166, ???? 1934. CODEN NUCIAD. ISSN 0029-6341 (print),

1827-6121 (electronic). URL http://www.springerlink.com/
content/04267161t7608557/.

Fermi:1934:TTR

[Fer34n] Enrico Fermi. Tentativo di una teoria dei raggi β. (Italian) [Draft of a theory of β rays]. Il Nuovo Cimento (8), 11(1):1–19, January 1934. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121

(electronic). URL http://inspirehep.net/record/25382;
http://www.springerlink.com/content/2t3p38675v840182/.

Fermi:1934:VTS

[Fer34o] Enrico Fermi. Versuch einer Theorie der β-Strahlen. I. (German) [Attempt at a theory of β-rays. I]. Zeitschrift für Physik, 88(3–4): 161–177, ???? 1934. CODEN ZEPYAA. ISSN ???? URL http://

www.springerlink.com/content/wq2r06j058382226/. English

translation in [].

Fermi:1934:BGB

[Fer34p] Enrico Fermi. Zur Bemerkung von G. Beck und K. Sitte. (German) [On the remark of G. Beck und K. Sitte]. Zeitschrift für Physik, 89(7–8):522, ???? 1934. CODEN ZEPYAA.

ISSN ????  URL http://www.springerlink.com/content/
m106547200268615/.

Fermi:1935:RAR

[Fer35a] Enrico Fermi. La radioattività artificiale: Redazione a cura del dott. ing. A. Giacomini. (Italian) [Artificial radioactivity: edited by Dr. Engineer A. Giacomini]. Rendiconti del Seminario Matematico e Fisico di Milano, 9(1):169–173, ???? 1935. CODEN

RSMFAG. ISSN 0370-7377. URL http://www.springerlink.
com/content/1424-9294/;   http://www.springerlink.com/
content/q4w87277111q0226/.

Fermi:1935:RNP

[Fer35b] Enrico Fermi. On the recombination of neutrons and protons. Physical Review (2), 48(6):570, September 15, 1935. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v48/i6/p570_1.

Fermi:1935:VDL

[Fer35c] Enrico Fermi. On the velocity distribution law for the slow neutrons. In ????, editor, Zeeman Verhandelingen, pages 128–130. Martinus Nijhoff, The Hague, The Netherlands, 1935. LCCN ????

Fermi:1935:RRDa

[Fer35d] Enrico Fermi. Recenti risultati della radioattività artificiale. (Italian) [Recent results in artificial radioactivity.]. La Ricerca Scientifica, 6(2):399–402, ???? 1935. CODEN RISCAZ. ISSN 0035-5011.

Fermi:1935:RRDb

[Fer35e] Enrico Fermi. Recenti risultati della radioattività artificiale. (Italian) [Recent results in artificial radioactivity.]. Atti Società Italiana per il Progresso delle Scienze, 24a Riunione, 3:116–120, ???? 1935.

Fermi:1936:CMM

[Fer36a] Enrico Fermi. Composition of matter and method of producing the same. US Patent ????, January 2, 1936. US patent application number 57,325, filed January 2, 1936.

Fermi:1936:SMI

[Fer36b] Enrico Fermi. Statistica meccanica. (Italian) [Statistical mechanics]. In ????, editor, Enciclopedia Italiana di Scienze, Lettere ed Arti, volume 32, pages 518–523. Istituto G. Treccani, Roma, Italy, 1936. LCCN ????

Fermi:1936:SGN

[Fer36c] Enrico Fermi. Sui gruppi di neutroni lenti. (Italian) [On groups of slow neutrons]. La Ricerca Scientifica, 7(1):310–315, ???? 1936. CODEN RISCAZ. ISSN 0035-5011.

Fermi:1936:SMN

[Fer36d] Enrico Fermi. Sul moto dei neutroni nelle sostanze idrogenate. (Italian) [On the motion of neutrons in hydrogenous substances]. La Ricerca Scientifica, 7(2):13–52, ???? 1936. CODEN RISCAZ. ISSN 0035-5011. English translation by G. Temmer in Collected Papers.

Fermi:1937:T

[Fer37a] Enrico Fermi. Thermodynamics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1937. vii + 160 pp. LCCN ????

Fermi:1937:MOM

[Fer37b] Enrico Fermi. Un maestro: Orso Mario Corbino. Nuova Antologia, 72(??):313–316, February 1, 1937.

Fermi:1938:FPI

[Fer38a] Enrico Fermi. Fisica per Istituti Tecnici Commerciali. (Italian) [Physics for Commercial Technical Institutes]. N. Zanichelli, Bologna, Italy, 1938. ???? pp. LCCN ????

Fermi:1938:GMP

[Fer38b] Enrico Fermi. Guglielmo Marconi e la propagazione delle onde elettromagnetiche nell’alta atmosfera. (Italian) [Guglielmo Marconi and the propagation of electromagnetic waves in the upper atmosphere]. Atti della Società Italiana per il Progresso delle Scienze, Collectanea Marconiana, Roma, pages 1–5, 1938.

Fermi:1938:MKG

[Fer38c] Enrico Fermi. Moleküle und Kristalle. (German) [Molecules and crystals]. Johann Ambrosius Barth, Leipzig, Germany, 1938. Translation of [Fer34f] to German by M. Schön and K. Birus.

Fermi:1938:NLL

[Fer38d] Enrico Fermi. Neutroni lenti e livelli energetici nucleari. (Italian) [Slow neutrons and nuclear energy levels]. Il Nuovo Cimento (8), 15(1):41–42, ???? 1938. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://www.

springerlink.com/content/x673m748r7512344/.

Fermi:1938:NPP

[Fer38e] Enrico Fermi. The Nobel Prize in Physics 1938. Nobel-

prize.org, 1938. URL http://www.nobelprize.org/nobel_
prizes/physics/laureates/1938/. The Nobel Prize in Physics

1938 was awarded to Enrico Fermi “for his demonstrations of the existence of new radioactive elements produced by neutron irradiation, and for his related discovery of nuclear reactions brought about by slow neutrons”.

Fermi:1938:PAD

[Fer38f] Enrico Fermi. Prospettive di applicazioni della radioattività artificiale. (Italian) [Prospects for applications of artificial radioactivity]. Rendiconti dell’Istituto di Sanità Pubblica, 1:421–432, 1938.

Fermi:1939:AMA

[Fer39a] Enrico Fermi. The absorption of mesotrons in air and in condensed materials. Physical Review, 56(12):1242, December 15, 1939. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v56/i12/p1242_1.

Fermi:1939:ARP

[Fer39b] Enrico Fermi. Artificial radioactivity produced by neutron bombardment: Nobel lecture delivered at Stockholm, December 12, 1938. In Les Prix Nobel en 1938: Les Conférences Nobel, Stockholm, pages 1–8. Norstedt & Söner, Stockholm, Sweden, 1939. ISSN 0546-8175. Reprinted in [Fer70a].

Fermi:1940:ILE

[Fer40a] Enrico Fermi. The ionization loss of energy in gases and in condensed materials. Physical Review, 57(6):485–493, March 15, 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://inspirehep.net/record/47115;
http://prola.aps.org/abstract/PR/v57/i6/p485_1.

Fermi:1940:ND

[Fer40b] Enrico Fermi. Nuclear disintegrations. Electrical Engineering (American Institute of Electrical Engineers), 59(2):57–58, February 1940. CODEN ELENAC. ISSN 0095-9197.

Fermi:1940:RPNa

[Fer40c] Enrico Fermi. Reactions produced by neutrons in heavy elements. Nature, 146(3707):640–642, November 16, 1940. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v146/
n3707/pdf/146640a0.pdf.

Fermi:1940:RPNb

[Fer40d] Enrico Fermi. Reactions produced by neutrons in heavy elements. Science, 92(2387):269–271, September 27, 1940. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic).  URL http://www.jstor.org/stable/1667328;
http://www.sciencemag.org/content/92/2387/269.extract.

Fermi:1941:SRP

[Fer41] Enrico Fermi. Some remarks on the production of energy by a chain reaction in uranium. Report A-14, US Atomic Energy Commission, Washington, DC, USA, June 30, 1941.

Fermi:1942:AGT

[Fer42a] Enrico Fermi. The absorption of graphite for thermal neutrons. Report C-154, US Atomic Energy Commission, Washington, DC, USA, June 30, 1942. Notes on Lecture of June 30, 1942.

Fermi:1942:CRF

[Fer42b] Enrico Fermi. Calculation of the reproduction factor. Report CP-358, US Atomic Energy Commission, Washington, DC, USA, 1942. Notes on Lecture of November 10, 1942.

Fermi:1942:CSM

[Fer42c] Enrico Fermi. The critical size — measurement of k in the exponential pile. Report C-289, US Atomic Energy Commission, Washington, DC, USA, 1942.

Fermi:1942:DGU

[Fer42d] Enrico Fermi. Design of the graphite–uranium lattice: Experimental determination of f_{t} from the Cd ratio. Report CP-337, US Atomic Energy Commission, Washington, DC, USA, 1942. Notes on Lectures of October 27 and November 3, 1942.

Fermi:1942:DAM

[Fer42e] Enrico Fermi. Determination of albedo; measurements of neutron density; slow neutron capture by hydrogen. Report C-31, US Atomic Energy Commission, Washington, DC, USA, March 10, 1942. Notes on Lecture of March 17, 1942.

Fermi:1942:DRB

[Fer42f] Enrico Fermi. The determination of the ratio between the absorption cross-sections of metal and carbon for thermal neutrons. Report C-84, US Atomic Energy Commission, Washington, DC, USA, May 15, 1942.

Fermi:1942:EAN

[Fer42g] Enrico Fermi. Effect of atmospheric nitrogen and of changes of temperature on the reproduction factor. Report C-85, US Atomic Energy Commission, Washington, DC, USA, May 19, 1942.

Fermi:1942:EBR

[Fer42h] Enrico Fermi. The effect of bismuth on the reproduction factor. Report CA-320, US Atomic Energy Commission, Washington, DC, USA, 1942. 1 pp. In: Fermi, E., and R. F. Christy. Bulletin for week ending October 31, 1942.

Fermi:1942:ECB

[Fer42i] Enrico Fermi. Effect of chemical binding in scattering and moderation of neutrons by graphite. Report C-87, US Atomic Energy Commission, Washington, DC, USA, 1942. Undated: year chosen according to report number.

Fermi:1942:EPN

[Fer42j] Enrico Fermi. Exponential Pile No. 11. Report CA-247, US Atomic Energy Commission, Washington, DC, USA, 1942. 2 pp. In: Whitaker, M. D., Enrico Fermi, E. O. Woolan, and A. H. Snell. Bulletin for week ending August 29, 1942.

Fermi:1942:FCR

[Fer42k] Enrico Fermi. Feasibility of a chain reaction. Report CP-383, US Atomic Energy Commission, Washington, DC, USA, November 26, 1942.

Fermi:1942:MKA

[Fer42l] Enrico Fermi. Memorandum to S. K. Allison with reference to designs and methods of cooling of chain reacting piles. Report Memo-io, US Atomic Energy Commission, Washington, DC, USA, October 5, 1942. 6 pp.

Fermi:1942:MCC

[Fer42m] Enrico Fermi. Methods of cooling chain reacting piles. Memo CP-10, US Atomic Energy Commission, Washington, DC, USA, October 5, 1942.

Fermi:1942:NPL

[Fer42n] Enrico Fermi. Neutron production in a lattice of uranium and graphite (theoretical part). Report CP-12, US Atomic Energy Commission, Washington, DC, USA, March 17, 1942.

Fermi:1942:NRL

[Fer42o] Enrico Fermi. Neutron reproduction in a lattice of uranium and graphite. theoretical part. Report C-12, US Atomic Energy Commission, Washington, DC, USA, March 17, 1942.

Fermi:1942:NNE

[Fer42p] Enrico Fermi. Number of neutrons emitted by metal per thermal neutron absorbed. Report C-190, US Atomic Energy Commission, Washington, DC, USA, July 16, 1942.

Fermi:1942:PRE

[Fer42q] Enrico Fermi. Preliminary report on the exponential experiment at Columbia University. Report CP-26, US Atomic Energy Commission, Washington, DC, USA, March/April 1942.

Fermi:1942:PTD

[Fer42r] Enrico Fermi. Problem of time dependence of the reaction rate: Effect of delayed neutrons emission. Report CP-291, US Atomic Energy Commission, Washington, DC, USA, 1942. Notes on Lecture of October 7, 1942.

Fermi:1942:PEAa

[Fer42s] Enrico Fermi. Purpose of the experiment at the Argonne Forest: Meaning of the reproduction factor k. Report CP-283, US Atomic Energy Commission, Washington, DC, USA, 1942.

Fermi:1942:RAI

[Fer42t] Enrico Fermi. Radio-active isotope production. Canadian Patent 407559., October 22, 1942. URL http://brevets-patents.

ic.gc.ca/opic-cipo/cpd/eng/patent/407559/summary.html.

This patent was never issued in the USA; see [Tur06b, pages 165, 171].

Fermi:1942:SCO

[Fer42u] Enrico Fermi. A simplified control. Optimum distribution of materials in the pile. Report CP-314, US Atomic Energy Commission, Washington, DC, USA, 1942. Notes on Lecture of October 20, 1942.

Fermi:1942:SDN

[Fer42v] Enrico Fermi. Slowing down and diffusion of neutrons. Report C-29, US Atomic Energy Commission, Washington, DC, USA, March 10, 1942.

Fermi:1942:SNH

[Fer42w] Enrico Fermi. The slowing down of neutrons in heavy water. Report CP-530, US Atomic Energy Commission, Washington, DC, USA, March 19, 1942.

Fermi:1942:TCP

[Fer42x] Enrico Fermi. A table for calculating the percentage loss due to the presence of impurities in alloy. Report C-5, US Atomic Energy Commission, Washington, DC, USA, February 10, 1942.

Fermi:1942:TER

[Fer42y] Enrico Fermi. The temperature effect on a reacting unit. Effect of the change of leakage. Report C-8, US Atomic Energy Commission, Washington, DC, USA, February 25, 1942.

Fermi:1944:A

[Fer44a] Enrico Fermi. Absorption of 49. Report CP-1592, US Atomic Energy Commission, Washington, DC, USA, April 1944. Excerpt from Metallurgical Laboratory Report CP-1592 for month ending April 24, 1944.

Fermi:1944:DB

[Fer44b] Enrico Fermi. Discussion on breeding. Report N-1729, US Atomic Energy Commission, Washington, DC, USA, 1944. Excerpt from Metallurgical Laboratory Report N-1729, notes on meeting of April 26, 1944.

Fermi:1944:MAH

[Fer44c] Enrico Fermi. Methods for analysis of helium circulating in the 105 unit. Document HW 3-492, US Atomic Energy Commission, Washington, DC, USA, August 7, 1944.

Fermi:1944:RFA

[Fer44d] Enrico Fermi. Report of Fermi’s activities with the Marshall Group. Report CP-1389, US Atomic Energy Commission, Washington, DC, USA, February 1944. Excerpt from Metallurgical Laboratory Report CP-1389 for month ending February 24, 1944.

Fermi:1944:RRV

[Fer44e] Enrico Fermi. Report on recent values of constants of 25 and product. Metallurgical Laboratory Report CK-1788, US Atomic Energy Commission, Washington, DC, USA, May 19, 1944.

Fermi:1945:RBN

[Fer45] Enrico Fermi. Relation of breeding to nuclear properties. Report CF-3199, US Atomic Energy Commission, Washington, DC, USA, 1945. Excerpt from Metallurgical Laboratory Report CF-3199, discussion on breeding, June 19–20, 1945.

Fermi:1946:AEP

[Fer46a] Enrico Fermi. Atomic energy for power. In The George Westinghouse Centennial Forum, volume 1: Science and Civilization — The Future of Atomic Energy. McGraw-Hill, New York, NY, USA, May 1946. LCCN ???? Also in Report MDDC-1, Atomic Energy Commission.

Fermi:1946:CNPb

[Fer46b] Enrico Fermi. A course in neutron physics. Report LAMS-347, US Atomic Energy Commission, Washington, DC, USA, February 5, 1946.

Fermi:1946:CNPa

[Fer46c] Enrico Fermi. A course in neutron physics, part 1. Document LADC-255, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, February 5, 1946. Notes by I. Halpern. Part 2, declassified in 1962.

Fermi:1946:DFC

[Fer46d] Enrico Fermi. The development of the first chain reacting pile. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 90(1):20–24, January 1946. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://www.jstor.org/stable/

3301034. Reprinted in [Fer70b].

Fermi:1946:ETP

[Fer46e] Enrico Fermi. Elementary theory of the pile. Report LAMS-427, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1946.

URL http://lib-www.lanl.gov/cgi-bin/getfile?00349505.
pdf.

Fermi:1946:FAE

[Fer46f] Enrico Fermi. The future of atomic energy. Report to United States Atomic Energy Commission MDDC-1, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, May

27, 1946. URL http://www.osti.gov/accomplishments/
documents/fullText/ACC0043.pdf. Report undated, but as-

signed release date.

Fermi:1946:MCT

[Fer46g] Enrico Fermi. Monte Carlo trolley. Unpublished., 1946. URL

https://en.wikipedia.org/wiki/FERMIAC. This one-of-a-kind

device, also known as the FERMIAC analog computer, was built by L. D. P. King at the suggestion of Enrico Fermi to simulate neutron diffusion. It was used for about two years, and is on display at the Bradbury Science Museum in Los Alamos.

Fermi:1946:NP

[Fer46h] Enrico Fermi. Neutron physics. Report MDCC-320, US Atomic Energy Commission, Washington, DC, USA, February 5, 1946.

Fermi:1946:NPCa

[Fer46i] Enrico Fermi. Neutron physics: a course by Enrico Fermi. Technical report, Los Alamos Scientific Laboratory, Los Alamos, NM,

USA, 1946. 115 pp. URL http://lib-www.lanl.gov/cgi-bin/
getfile?00191399.pdf.

Fermi:1947:ETK

[Fer47a] Enrico Fermi. Elementarnaya teoria kotlov c tsepnimi yadernimi reaktsiyami. (Russian) [Elementary theory of boilers with nuclear chain reactions]. Uspekhi Fizicheskikh Nauk, 32(5):54–65, May 1947. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-

6652 (electronic). URL http://ufn.ru/ru/articles/1947/5/

d/. Russian translation of [Fer47b].

Fermi:1947:ETC

[Fer47b] Enrico Fermi. Elementary theory of the chain-reacting pile. Science, 105(2715):27–32, January 10, 1947. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

jstor.org/stable/1675136;   http://www.sciencemag.org/
content/105/2715/27.extract.

Fermi:1949:DM

[Fer49a] Enrico Fermi. The Dirac monopole. Report VPI-EPP-5-86, ????, ????, 1949.

Fermi:1949:HOC

[Fer49b] Enrico Fermi. An hypothesis on the origin of the cosmic radiation. Il Nuovo Cimento (9), 6(S3):317–323, ???? 1949. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL

http://www.springerlink.com/content/a27kn2l157686t47/.

Fermi:1949:VEF

[Fer49c] Enrico Fermi. La visita di Enrico Fermi al Consiglio Nazionale delle Ricerche. (Italian) [The visit of Enrico Fermi to the National Research Council]. La Ricerca Scientifica, 19(??):1113–1118, ???? 1949. CODEN RISCAZ. ISSN 0035-5011.

Fermi:1949:LD

[Fer49d] Enrico Fermi. Lezioni donegani, 1949.

Fermi:1949:NPC

[Fer49e] Enrico Fermi. Nuclear physics. A course given at the University of Chicago. University of Chicago Press, Chicago, IL, USA, 1949. vii + 246 pp. Notes compiled by Jay Orear, A. H. (Arthur Hinton) Rosenfeld, and R. A. (Robert Arvel) Schluter.

Fermi:1949:OCRa

[Fer49f] Enrico Fermi. On the origin of the cosmic radiation. Physical Review, 75(8):1169–1174, April 15, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http:

//inspirehep.net/record/43073;   http://prola.aps.org/
abstract/PR/v75/i8/p1169_1.

Fermi:1949:OCRb

[Fer49g] Enrico Fermi. Origin of cosmic radiation. Physics Today, 2(5):30, May 1949. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/?PTO/2/30/1;
http://link.aip.org/link/phtoad/v2/i5/p30/pdf.

Fermi:1950:CFA

[Fer50a] Enrico Fermi, editor. Conferenze di Fisica Atomica (Fondazione Donegani). Accademia Nazionale dei Lincei, Roma, Italy, 1950. LCCN ???? Reprinted in full in [Fer65b, pages 684–788].

Fermi:1950:HEN

[Fer50b] Enrico Fermi. High energy nuclear events. Progress of Theoretical Physics, 5:570–583, 1950. CODEN PTPKAV. ISSN 1347-4081.

Fermi:1950:NVS

[Fer50c] Enrico Fermi. Neutron velocity selector. US Patent 2,524,379., October 3, 1950. US Patent Application 617,121 filed September 18, 1945.

Fermi:1950:NPCa

[Fer50d] Enrico Fermi. Nuclear physics. A course given at the University of Chicago. University of Chicago Press, Chicago, IL, USA, revised edition, 1950. ix + 246 pp. Notes compiled by Jay Orear, A. H. (Arthur Hinton) Rosenfeld, and R. A. (Robert Arvel) Schluter.

Fermi:1950:NPCb

[Fer50e] Enrico Fermi. Nuclear physics. A course given at the University of Chicago. Cambridge University Press, Cambridge, UK, revised edition, 1950. ix + 246 pp. Notes compiled by Jay Orear, A. H. (Arthur Hinton) Rosenfeld, and R. A. (Robert Arvel) Schluter.

Fermi:1951:ADP

[Fer51a] Enrico Fermi. Angular distribution of the pions produced in high energy nuclear collisions. Physical Review, 81(5):683–687, March 1, 1951. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://inspirehep.net/record/45476;
http://prola.aps.org/abstract/PR/v81/i5/p683_1.

Fermi:1951:EPa

[Fer51b] Enrico Fermi. Elementary particles. Mrs. Hepsa Ely Silliman memorial lectures: 1950. Yale University Press, New Haven, CT, USA, 1951. xii + 110 pp. LCCN ????

Fermi:1951:EPb

[Fer51c] Enrico Fermi. Elementary particles. Geoffrey Cumberlege, London, UK, 1951. xii + 110 pp. LCCN ????

Fermi:1951:ELT

[Fer51d] Enrico Fermi. Excerpt from a lecture on Taylor instability. Given during the fall of 1951 at Los Alamos Scientific Laboratory. Report, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1951.

Fermi:1951:FP

[Fer51e] Enrico Fermi. Fundamental particles. In Proceedings of the International Conference on Nuclear Physics and the Physics of Fundamental Parti des. The University of Chicago, September 17–22,

1951. ????, ????, 1951. URL http://inspirehep.net/record/

42971. Reprinted in [Fer65b, pages 825–828].

Fermi:1951:NP

[Fer51f] Enrico Fermi. Neutron physics. Report AECD-2664, US Atomic Energy Commission, Washington, DC, USA, October 16,

1951. 102 pp. URL http://www.osti.gov/accomplishments/
documents/fullText/ACC0040.pdf.  Technical Information

Branch, Oak Ridge, Tennessee. Notes by I. Halperin, revised by J. C. Beckerley, from Enrico Fermi’s lectures in 1945.

Fermi:1951:TIIa

[Fer51g] Enrico Fermi. Taylor instability of an incompressible liquid. Document AECU-2979, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, September 4, 1951. Part 1.

Fermi:1951:TCS

[Fer51h] Enrico Fermi. Trajectories in a cylindrically symmetric magnetic field. Notes and program for Maniac computer. Box 22, Folder 1 of the Enrico Fermi Collection in the University of Chicago Library, Special Collections Research Center., 1951. URL http://

fermi.lib.uchicago.edu/maniac.pdf. This undated document

contains handwritten notes, a typewritten machine-language program for the Los Alamos MANIAC computer, and a typewritten computed table of trajectories.

Fermi:1952:EPD

[Fer52a] Enrico Fermi. Experimental production of a divergent chain reaction. American Journal of Physics, 20(9):536–558, December 1952. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://link.aip.org/link/?AJP/20/536/1;
 http://link.aip.org/link/ajpias/v20/i9/p536/s1. The

American Journal of Physics is pleased to have the opportunity of publishing this paper on the tenth anniversary of the achievement of a nuclear chain reaction, December 2, 1942.

Fermi:1952:FAU

[Fer52b] Enrico Fermi. Fisica: ad uso dei licei. (Italian) [Physics: for use by high schools]. N. Zanichelli, Bologna, Italy, 1952. ???? pp. LCCN ????

Fermi:1952:FTR

[Fer52c] Enrico Fermi. Free tape recordings of important speeches: The nucleus. American Journal of Physics, 20(9):558, December 1952. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Fermi:1952:LF

[Fer52d] Enrico Fermi. Letter to Feynman, January 18, 1952. Reprinted in [Fer65b, pages 844–846].

Fermi:1952:N

[Fer52e] Enrico Fermi. The nucleus. Physics Today, 5(3):6–9, March 1952. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/?PTO/5/6/1;
http://link.aip.org/link/phtoad/v5/i3/p6/s1.

Fermi:1952:PEI

[Fer52f] Enrico Fermi. Particelle elementari. (Italian) [Elementary particles]. Edizioni scientifiche Einaudi, Torino, Italy, 1952. 194 pp. LCCN ???? Translated and extended by Piero Caldirola.

Fermi:1952:TRM

[Fer52g] Enrico Fermi. Tenth Richtmyer Memorial Lecture of the AAPT [title only]. Physical Review, 86(4):611, May 1952. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Fermi:1952:YPP

[Fer52h] Enrico Fermi. Yadernie processi pri bolschich energiyach. (Russian) [Nuclear processes at high energies]. Uspekhi Fizicheskikh Nauk, 46(1):71–95, January 1952. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.ru/

ru/articles/1952/1/c/.

Fermi:1953:MPP

[Fer53a] Enrico Fermi. Multiple production of pions in nucleon-nucleon collisions at cosmotron energies. Physical Review, 92(2):452–453, October 15, 1953. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://inspirehep.net/record/
47364;   http://prola.aps.org/abstract/PR/v92/i2/p452_
1.

Fermi:1953:NPP

[Fer53b] Enrico Fermi. Nucleon polarization in pion–proton scattering. Physical Review (2), 91(4):947–948, August 15, 1953. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v91/i4/p947_1.

Fermi:1953:RPS

[Fer53c] Enrico Fermi. Report on pion scattering. In ????, editor, Proceedings of the Third Annual Rochester Conference, December 18–20, 1952, page ?? ????, ????, 1953. LCCN ???? Reprinted in [Fer65b, pages 855–860].

Fermi:1954:PNS

[Fer54a] E. Fermi. Pion nucleon scattering and photoproduction of pions. In H. P. Noyes, E. M. Hafner, J. Klarmann, and A. E. Woodruff, editors, 4th Annual Rochester Conference on High-Energy and Nuclear Physics, pages 92–120. Rochester University, Rochester, NY, USA, 1954.

Fermi:1954:GMF

[Fer54b] Enrico Fermi. Galactic magnetic fields and the origin of cosmic radiation. Astrophysical Journal, 119(??):1–6, ???? 1954. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Fermi:1954:LAL

[Fer54c] Enrico Fermi. Letter: Los Alamos Laboratory has deserved the gratitude of this nation. Bulletin of the Atomic Scientists, 10(9): 359, November 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fermi:1954:LV

[Fer54d] Enrico Fermi. Lezioni Varenna, 1954.

Fermi:1954:MPPb

[Fer54e] Enrico Fermi. Multiple production of pions in nucleon-nucleon collisions at cosmotron energies. Physical Review, 93(6):1434–1435, March 15, 1954. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/
PR/v93/i6/p1434_9.

Fermi:1954:MPPa

[Fer54f] Enrico Fermi. Multiple production of pions in pion–nucleon collisions. Anais da Academia Brasileira de Ciências, 26(??):61–63, ???? 1954. CODEN AABCAD. ISSN 0001-3765 (print), 1678-2690 (electronic).

Fermi:1954:PES

[Fer54g] Enrico Fermi. Polarization in the elastic scattering of high energy protons by nuclei. Private Communication, March 24, 1954. Reprinted in [Fer65b, pages 994–995].

Fermi:1954:PHE

[Fer54h] Enrico Fermi. Polarization of high energy protons scattered by nuclei. Il Nuovo Cimento (9), 11(4):407–411, April 1, 1954. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL

http://www.springerlink.com/content/84074721xu54h801/.

Fermi:1954:QM

[Fer54i] Enrico Fermi. Quantum mechanics. Technical report, Enrico Fermi Institute for Nuclear Studies, University of Chicago, Chicago, IL, USA, 1954. Lectures to students at the University of Chicago, 1954. A few copies on file at the Enrico Fermi Institute.

Fermi:1954:U

[Fer54j] Enrico Fermi. [unknown]. Article for U. S. Information Service, May 1954.

Fermi:1954:AFM

[Fer54k] Laura Fermi. Atoms in the family: my life with Enrico Fermi. University of Chicago Press, Chicago, IL, USA, 1954. 267 pp.

LCCN QC774.F4 F4 1954. URL http://www.press.uchicago.
edu/ucp/books/book/chicago/A/bo3628811.html.

Fermi:1954:WMDa

[Fer54l] Laura Fermi. That was the Manhattan District: A domestic view — I. The New Yorker, 30(??):25–??, July 24, 1954. ISSN 0028-

792X. URL http://www.newyorker.com/archive/1954/07/24/
1954_07_24_025_TNY_CARDS_000244162.

Fermi:1954:WMDb

[Fer54m] Laura Fermi. That was the Manhattan District: A domestic view — II. The New Yorker, 30(??):27–??, July 31, 1954. ISSN 0028-

792X. URL http://www.newyorker.com/archive/1954/07/31/
1954_07_31_027_TNY_CARDS_000242787.

Fermi:1955:PCU

[Fer55] Enrico Fermi. Physics at Columbia University: The genesis of the nuclear energy project. Physics Today, 8(11):12–16, November 1955. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/?PTO/8/
12/1;   http://link.aip.org/link/phtoad/v8/i11/p12/s1.

Reprinted in [Fer70a] and [WP85, pages 282–286].

Fermi:1956:T

[Fer56] Enrico Fermi. Thermodynamics. Dover, New York, NY, USA, 1956. ISBN 0-486-60361-X. x + 160 pp. LCCN QC311 .F47

Fermi:1957:TEP

[Fer57a] Enrico Fermi. Test exponential pile. US Patent 2,780,595., February 5, 1957. US Patent Application 534,129 filed May 4, 1944.

Fermi:1957:AWU

[Fer57b] Laura Fermi. Atoms for the world: United States participation in the Conference on the Peaceful Uses of Atomic Energy. University of Chicago Press, Chicago, IL, USA, 1957. xi + 227 pp. LCCN TK9006 .I5 1955j.

Fermi:1957:LEN

[Fer57c] Laura Fermi. Letter to the Editor: The nature of nuclear warfare. Bulletin of the Atomic Scientists, 13(6):232, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tel57].

Fermi:1958:TIT

[Fer58a] Enrico Fermi. Termodinamica. (Italian) [Thermodynamics]. Boringhieri, Torino, Italy, 1958. 179 pp. LCCN ???? Translation of [Fer37a] to Italian by A. Scotti.

Fermi:1958:TEN

[Fer58b] Enrico Fermi. Test exponential nuclear reactor. Canadian Patent 552693., February 04, 1958. URL http://brevets-patents.ic.

gc.ca/opic-cipo/cpd/eng/patent/552693/summary.html.

Fermi:1960:NR

[Fer60] Enrico Fermi. Neutronic reactor. US Patent 2,931,762., April 5,

1960. URL http://www.google.com/patents/US2931762. US

Patent Application 593,510 filed May 12, 1945.

Fermi:1961:NQM

[Fer61a] Enrico Fermi. Notes on quantum mechanics. University of Chicago Press, Chicago, IL, USA, 1961. vii + 171 pp.

Fermi:1961:NR

[Fer61b] Enrico Fermi. Nuclear reactor. Canadian Patent 619065.,

March 25, 1961. URL http://brevets-patents.ic.gc.ca/
opic-cipo/cpd/eng/patent/619065/summary.html.

Fermi:1961:M

[Fer61c] Laura Fermi. Mussolini. University of Chicago Press, Chicago, IL, USA, 1961. vii + 477 pp. LCCN DG575.M8 F42.

Fermi:1961:SAE

[Fer61d] Laura Fermi. The story of atomic energy. Landmark books. Random House, New York, NY, USA, 1961. 184 pp.

LCCN QC778 .F47. URL http://www.archive.org/details/
storyofatomicene00ferm.

Fermi:1962:CPN

[Fer62a] Enrico Fermi. Collected papers (Note e memorie). Vol. I: Italy, 1921–1938. University of Chicago Press, Chicago, IL, USA, 1962. xlii + 1043 + 8 pp.

Fermi:1962:SPR

[Fer62b] Laura Fermi. Some personal reminiscences. International Atomic Energy Agency Bulletin, 4(0):38–40, ???? 1962. CODEN IAE-BAB. ISSN 0020-6067 (print), 1564-2690 (electronic). URL http:

//www.iaea.org/Publications/Magazines/Bulletin/Bull040su/
04005003840su.pdf.

Fermi:1965:ARN

[Fer65a] Enrico Fermi. Artificial radioactivity by neutron bombardment. In Anonymous [Ano65b], pages 414–421. LCCN QC71 .P455

1965. URL http://nobelprize.org/nobel_prizes/physics/
laureates/1922/bohr-lecture.html;   http://nobelprize.
org/nobel_prizes/physics/laureates/1922/bohr-lecture.

pdf. Nobel Lecture given in December 1938 in Stockholm.

Fermi:1965:CPN

[Fer65b] Enrico Fermi. Collected papers (Note e memorie). Vol. II: United States, 1939–1954. University of Chicago Press, Chicago, IL, USA, 1965. xvi + 1085 + 10 pp. Edited in the decade after the death of Fermi in 1954 by Edoardo Amaldi, Herbert L. Anderson, Enrico Persico, Franco Rasetti, C. S. Smith, A. Wattenberg, and Emilio Segrè (chairman).

Fermi:1965:NQM

[Fer65c] Enrico Fermi. Notes on quantum mechanics: a course given by Enrico Fermi at the University of Chicago. University of Chicago Press, Chicago, IL, USA, second edition, 1965. ISBN 0-226-24381-8. vii + 188 pp. LCCN ???? With problems compiled by Robert A. Schluter.

Fermi:1966:MCQ

[Fer66a] Enrico Fermi, editor. Molecules, Crystals, and Quantum Statistics. W. A. Benjamin, Inc., New York, NY, USA, 1966. xiv + 300 pp. LCCN QC173 .F4213. Translation of Molecole e cristalli [Fer34f] by M. Ferro-Luzzi. Edited by Lloyd Motz.

Fermi:1966:NTS

[Fer66b] Enrico Fermi. Notes on Thermodynamics and Statistics. University of Chicago Press, Chicago, IL, USA, 1966. viii + 182 pp. LCCN QC311.5 .F4 1966. Manuscript prepared for a course on thermodynamics held in 1951–1952.

Fermi:1966:PON

[Fer66c] Enrico Fermi. The phenomena that occur in the neighborhood of a time line. Report N67 15554, National Aeronautics and Space Administration, Washington, DC, USA, Octo-

ber 1966. 18 pp. URL http://archive.org/details/nasa_

techdoc_19670006225. English translation of [Fer22c]. Includes pages images of Italian original.

Fermi:1968:KMK

[Fer68a] Ènriko Fermi. Kvantovaya mekhanika (konspekt lektsii). (Russian) [Quantum mechanics (konspekt lektsii)]. Izdat. “Mir”, Moscow, USSR, 1968. 367 pp.

Fermi:1968:III

[Fer68b] Laura Fermi. Illustrious Immigrants; the Intellectual Migration from Europe, 1930–41. University of Chicago Press, Chicago, IL, USA, 1968. xi + 440 pp. LCCN E184.A1 F47.

Fermi:1968:M

[Fer68c] Laura Fermi. Mussolini. University of Chicago Press, Chicago, IL, USA, 1968. vii + 477 pp. LCCN DG575.M8 F42.

Fermi:1970:ARP

[Fer70a] Enrico Fermi. Artificial radioactivity produced by neutron bombardment: Nobel lecture delivered at Stockholm, December 12, 1938. In Enrico Fermi: physicist [Seg70b], chapter A-2, pages 214–221. ISBN 0-226-74472-8. LCCN QC16.F46 S4.

Fermi:1970:DFC

[Fer70b] Enrico Fermi. The development of the first chain reacting pile. In Enrico Fermi: physicist [Seg70b], chapter A-4, pages 231–239. ISBN 0-226-74472-8. LCCN QC16.F46 S4. Reprint of [Fer46d].

Fermi:1970:PCU

[Fer70c] Enrico Fermi. Physics at Columbia University: The genesis of the nuclear energy project. In Enrico Fermi: physicist [Seg70b], chapter A-3, pages 222–230. ISBN 0-226-74472-8. LCCN QC16.F46 S4. Reprint of [Fer55].

Fermi:1970:BR

[Fer70d] Laura Fermi. Bombs or reactors? Bulletin of the Atomic Scientists, 26(6):28–29, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fermi:1971:NTR

[Fer71a] Enrico Fermi. Nauchnye trudy (1921–1938). (Russian) [Working papers (1921–1938)]. Nauka, Moscow, USSR, 1971. 818 pp. LCCN ????

Fermi:1971:BR

[Fer71b] Laura Fermi. Bombs or reactors? In Lewis and Wilson [LW71], page ?? ISBN 0-670-11151-1. LCCN QC792 .A43 1971. With Eugene Rabinowitch.

Fermi:1971:III

[Fer71c] Laura Fermi. Illustrious immigrants: the intellectual migration from Europe, 1930–41. University of Chicago Press, Chicago, IL, USA, second edition, 1971. ISBN 0-226-24376-1 (hardcover), 0-226-24378-8 (paperback). xi + 431 pp. LCCN E184.A1 F47 1971.

Fermi:1972:TIT

[Fer72] Enrico Fermi. Termodinamica. (Italian) [Thermodynamics]. Boringhieri, Torino, Italy, second edition, 1972. ISBN 88-339-5182-0. 179 pp. LCCN ????

Fermi:1974:NP

[Fer74] Enrico Fermi. Nuclear physics. University of Chicago Press, Chicago, IL, USA, 1974. ISBN 0-226-24365-6. ix + 248 pp. LCCN ???? Notes compiled by Jay Orear, A. H. (Arthur Hinton) Rosenfeld, and Robert A. Schluter.

Fermi:1975:MMN

[Fer75] Enrico Fermi. 4. magnetic moments of nuclei. In Mehra [Meh75], chapter 7, pages 190–191. ISBN 90-277-0635-2. LCCN QC1.S792 M43.

Fermi:1976:VEM

[Fer76] Laura Fermi. A visit to the “Ettore Majorana” Centre. Physics Today, 29(8):9–13, August 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:/

/adsabs.harvard.edu/abs/1976PhT....29h...9F;   http://
www.physicstoday.org/resource/1/phtoad/v29/i8/p9_s1.

Fermi:1980:FPA

[Fer80] Laura Fermi. The Fermis’ path to Los Alamos. In Badash et al. [BHB80], pages 89–104. ISBN 90-277-1097-X, 90-277-1098-8 (paperback). LCCN QC791.96 .R44.

Fermi:1982:MCI

[Fer82a] Enrico Fermi. Molecole e cristalli. (Italian) [Molecules and crystals]. N. Zanichelli, Bologna, Italy, 1982. ISBN ???? 301 pp. LCCN ????

Fermi:1982:TIT

[Fer82b] Enrico Fermi. Termodinamica. (Italian) [Thermodynamics]. Boringhieri, Torino, Italy, 1982. ISSN ???? 178 pp. LCCN ????

Fermi:1982:AFM

[Fer82c] Laura Fermi. Atoms in the family: my life with Enrico Fermi. University of New Mexico Press, Albuquerque, NM, USA, 1982. ISBN 0-8263-1060-5 (paperback). ix + 267 + 24 pp. LCCN QC774.F4 F4 1988.

Fermi:1983:LPN

[Fer83] Enrico Fermi. Lectures on pions and nucleons. In Anonymous, editor, Celebrazioni del Trentennale della Scuola Internazionale di Fisica ≪Enrico Fermi≫, pages 17–95. Editrice Compositori e Società Italiana di Fisica, Bologna, Italia, 1983. ISBN ???? LCCN ???? Posthumous republication of lecture given in 1954 at the II Course of the International School of Physics at Villa Monastero, Varenna, Italy. Also available in [FF55, FF08].

Fermi:1985:PCU

[Fer85] Enrico Fermi. Physics at Columbia University. In Weart and Phillips [WP85], pages 282–286. ISBN 0-88318-468-0 (paperback). LCCN QC7 .H694 1985. Reprint of [Fer55].

Fermi:1987:AFM

[Fer87] Laura Fermi. Atoms in the family: my life with Enrico Fermi, volume 9 of The History of modern physics, 1800–1950. Tomash Publishers, Los Angeles, CA, USA, 1987. ISBN 0-88318-524-5. 267 + 24 pp. LCCN QC16.F46 F47 1987. US$35.00.

Fermi:1996:MSS

[Fer96] Enrico Fermi. Meccanica statistica: scritti scelti. (Italian) [Statistical mechanics: selected writings], volume 20 of I fondamenti della scienza. Teknos, Roma, Italy, 1996. ISBN ???? xxviii + 154 pp. LCCN ???? Edited by Guido Altarelli and Giorgio Capon.

Fermi:1998:TRT

[Fer98] Enrico Fermi. Termodinamika. (Russian) [Thermodynamics]. Izhevsk: Nauchno-Izdatel’skij Tsentr “Regulyarnaya i Khaoticheskaya Dinamika”, 1998.

Fermi:19xx:LEI

[Ferxx] Enrico Fermi. Lezioni di elettrodinamica. (Italian) [Lectures on electrodynamics]. Stabilimento Tipolitografico del Genio Civile, Roma, Italia, 19xx. 95 pp. Notes by A. Morelli.

Fermi:2001:TCB

[Fer01] Enrico Fermi. On the theory of collisions between atoms and electrically charged particles. In W. Marciano and S. White, editors, Workshop on Electromagnetic Probes of Fundamental Physics 16–21 Oct 2001. Erice, Sicily, Italy, pages 243–252. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2001. ISBN ???? LCCN

???? URL http://emcsc.ccsem.infn.it/ccsem01/White01.
html;   http://inspirehep.net/record/42723. Translated

from the Italian original [Fer25f] by Michele Gallinaro and Sebastian White.

Fermi:2004:MLP

[Fer04a] Laura Fermi. My life as a physicist’s wife. In Orear [Ore04], chapter 30, pages 145–154. ISBN ???? LCCN ????

URL http://dspace.library.cornell.edu/handle/1813/62;
 http://hdl.handle.net/1813/74.

Fermi:2004:DP

[Fer04b] Nelia Fermi. A different perspective. In Orear [Ore04], chapter 26, pages 129–138. ISBN ???? LCCN ???? URL http:

//dspace.library.cornell.edu/handle/1813/62;   http://
hdl.handle.net/1813/74.

Fermi:2010:T

[Fer10] Enrico Fermi. Thermodynamics. Dover books on physics. Dover, New York, NY, USA, 2010. ISBN 0-486-13485-7, 0-486-60361-X, 1-62198-581-4. 160 pp. LCCN QC311. URL http://app.knovel.

com/hotlink/toc/id:kpT0000001/thermodynamics.

Feynman:1980:LAB

[Fey80] Richard P. Feynman. Los Alamos from below. In Badash et al. [BHB80], pages 105–132. ISBN 90-277-1097-X, 90-277-1098-8 (paperback). LCCN QC791.96 .R44. URL https://www.chilton-

computing.org.uk/harwell/literature/FeynmanLosAlamos.
htm.

Feld:1942:NER

[FF42] B. T. Feld and Enrico Fermi. Neutrons emitted by a radiumberyllium photo source. Report C-89 (MDDC-1438), US Atomic Energy Commission, Washington, DC, USA, 1942. Undated: year chosen according to report number.

Feld:1948:NER

[FF48] B. T. Feld and Enrico Fermi. Neutrons emitted by a Ra + Be photosource. Report CP-89, US Atomic Energy Commission, Washington, DC, USA, November 5, 1948.

Fermi:1955:LPN

[FF55] Enrico Fermi and B. T. Feld. Lectures on pions and nucleons. Il Nuovo Cimento (10), 2(Supplement 1):1–73, January 1955. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121

(electronic). URL http://inspirehep.net/record/809073;
http://link.springer.com/article/10.1007/BF02746078;
http://www.springerlink.com/content/h5667333h80qn024/.

Reprinted in [FF08]. Issue also cited as volume 31, but publisher Web page says volume 2.

Fermi:2008:LPN

[FF08] Enrico Fermi and B. T. Feld. Lectures on pions and nucleons. Rivista del Nuovo Cimento, 31(1):1–73, January 2008. CODEN RNUCAC. ISSN 0393-697X (print), 1826-9850 (electronic). URL

http://inspirehep.net/record/809073. Reprint of [FF55].

Fermi:1944:PN

[FFH^{+}44] Enrico Fermi, James Franck, T. R. Hogness, Zay Jeffries, R. S. Mulliken, R. S. Stone, and C. A. Thomas. Prospectus on nucleonics. Report MUC-RSH-234, Metallurgical Project, University of Chicago, Chicago, IL, USA, November 18, 1944. 67

pp. URL http://www.marshallfoundation.org/library/wp-
content/uploads/sites/16/2015/05/xerox1482-45_opt.pdf.

Classified-secret report submitted to Arthur Compton on the military, scientific, and industrial implications of nuclear energy.

Feynman:1992:TFI

[FG92] R. P. Feynman and M. Gellmann. Theory of the Fermi interaction. Current Science, 63(2):76–81, July 25, 1992. CODEN CUSCAM. ISSN 0011-3891. Reprinted from Physical Review, 1958.

Fermi:1953:SNP

[FGMN53] Enrico Fermi, M. Glicksman, R. Martin, and D. Nagle. Scattering of negative pions by hydrogen. Physical Review, 92(1):161–163, October 1, 1953. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/
PR/v92/i1/p161_1.

Fermi:1946:NPCb

[FH46] Enrico Fermi and Isaac Halpern. Neutron physics: a course by Enrico Fermi. Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1946. iv + 74 pp.

Friedlander:1981:OFS

[FH81] Gerhart Friedlander and Günter Herrmann. [Obituary:] Fritz Strassmann. Physics Today, 34(4):84–86, April 1981. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Fermi:1944:CRG

[FHN44] Enrico Fermi, A. Heskett, and Darragh E. Nagle. Comparison of the ranges in graphite of fission neutrons from 49 and 25. Report CP-1592, US Atomic Energy Commission, Washington, DC, USA, April 1944. Excerpt from Metallurgical Laboratory Report CP-1592 for month ending April 24, 1944.

Firor:2005:FSR

[Fir05] John Firor. Former student remembers Teller and Fermi with gratitude. Physics Today, 58(2):10, February 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://physicstoday.org/resource/1/phtoad/v58/i2/
p10_s1.

Fischer:1997:HIA

[Fis97] David Fischer. History of the International Atomic Energy Agency: the first forty years. IAEA, Vienna, Austria, 1997. ISBN 92-0-102397-9. 550 pp. LCCN QC770 .F56 1997. URL

http://www-pub.iaea.org/mtcd/publications/pdf/pub1032_
web.pdf.

Fischer:2010:HQE

[Fis10a] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger]. Herbig, München, Germany, 2010. ISBN 3-7766-2643-7. 350 pp. LCCN ????

Fischer:2010:EF

[Fis10b] Ernst Peter Fischer. Enrico Fermi (1901–1954). In Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger] [Fis10a], pages 188–198. ISBN 3-7766-2643-7. LCCN ????

Fischer:2012:HQE

[Fis12a] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger], volume 19406 of Fischer. Fischer-Taschenbuch-Verlag, Frankfurt am Main, Germany, 2012. ISBN 3-596-19406-7. 350 pp. LCCN ????

Fischer:2012:EF

[Fis12b] Ernst Peter Fischer. Enrico Fermi (1901–1954). In Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger] [Fis12a], pages 188–198. ISBN 3-596-19406-7. LCCN ????

Fitzpatrick:1999:ILE

[Fit99] Anne Fitzpatrick. Igniting the light elements: The Los Alamos Thermonuclear Weapon Project, 1942–1952. Thesis LA-13577-T, Los Alamos National Laboratory, Los Alamos, NM, USA, July

1999. xi + 45 pp. URL https://fas.org/sgp/othergov/doe/
lanl/docs1/00460048.pdf.

Fitzpatrick:2013:ILE

[Fit13] Anne C. Fitzpatrick. Igniting the light elements: the Los Alamos Thermonuclear Weapon Project, 1942-1952. Biblioscholar, ????, 2013. ISBN 1-288-82498-X. x + 325 (est.) pp. LCCN ????

Fu:2008:SPE

[FK08] Liang Fu and C. L. Kane. Superconducting proximity effect and Majorana fermions at the surface of a topological insulator. Physical Review Letters, 100(9):096407, March 2008. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-

0145. URL http://link.aps.org/doi/10.1103/PhysRevLett.
100.096407.

Fermi:1957:MSN

[FL57] Enrico Fermi and Miles C. Leverett. Method of sustaining a neutronic chain reacting system. US Patent 2,813,070., November 12, 1957. US Patent Application 631,406 filed November 28, 1945.

Fermi:1958:CRS

[FL58] Enrico Fermi and Miles C. Leverett. Chain reacting system. US Patent 2,837,477., June 3, 1958. URL http://www.google.com/ patents/US2837477. US Patent Application 578,278 filed February 16, 1945.

Flugge:1939:KEA

[Flü39] Siegfried Flügge. Kann der Energieinhalt der Atomkerne technisch nutzbar gemacht werden?. (German) [Can the energy content of atomic nuclei be made technically available?]. Naturwissenschaften, 27(23–24):402–410, June 1939. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://

www.springerlink.com/content/g55p220623l6813n/.

Fermi:1944:EF

[FM44] Enrico Fermi and Leona Marshall. Evidence for the formation of 26. Report CP-1531, US Atomic Energy Commission, Washington, DC, USA, 1944. Excerpt from Metallurgical Laboratory Report CP-1531 for month ending March 25, 1944.

Fermi:1946:PNSa

[FM46a] Enrico Fermi and Leona Marshall. Phase of neutron scattering. Report MDDC-55 (CU-i), US Atomic Energy Commission, Washington, DC, USA, July 1946. Abstract published in Phys. Rev. 70:103 (1946). Physical Society Cambridge (England) Conference Report, 1947, pp. 94–97.

Fermi:1946:PNSb

[FM46b] Enrico Fermi and Leona Marshall. Phase of neutron scattering. Physical Review (2), 70(1–2):103, July 1946. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v70/i1-2/p99_1.

Fermi:1947:FES

[FM47a] Enrico Fermi and Leona Marshall. Further experiments with slow neutrons. Report, US Atomic Energy Commission, Washington, DC, USA, 1947. Excerpts from quarterly reports CF-3574, Argonne National Laboratory, July 26, 1946, and CP-3750 and CP-3801, Argonne National Laboratory, January 17 and April 14, 1947.

Fermi:1947:IPS

[FM47b] Enrico Fermi and Leona Marshall. Interference phenomena of slow neutrons. Physical Review, 71(10):666–677, May 15, 1947. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v71/i10/p666_1.

Fermi:1947:IBN

[FM47c] Enrico Fermi and Leona Marshall. On the interaction between neutrons and electrons. Physical Review, 72(12):1139– 1146, December 15, 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://inspirehep.net/

record/47899;   http://prola.aps.org/abstract/PR/v72/
i12/p1139_1.

Fermi:1947:PST

[FM47d] Enrico Fermi and Leona Marshall. Phase of scattering of thermal neutrons by aluminum and strontium. Physical Review, 71(12): 915, June 15, 1947. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/
PR/v71/i12/p915_1.

Fermi:1947:RSS

[FM47e] Enrico Fermi and Leona Marshall. Resonance scattering of slow neutrons. Report, US Atomic Energy Commission, Washington, DC, USA, June 1947.

Fermi:1947:SDS

[FM47f] Enrico Fermi and Leona Marshall. Spin dependence of scattering of slow neutrons by Be, Al, and Bi. Physical Review, 72(5):408– 410, September 1, 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v72/i5/p408_1.

Fermi:1949:SDS

[FM49] Enrico Fermi and Leona Marshall. Spin dependence of slow neutron scattering by deuterons. Physical Review, 75(4):578, February 15, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v75/i4/p578_1.

Fermi:1952:NSM

[FM52] Enrico Fermi and Nicholas Metropolis. Numerical solution of a minimum problem. Report LA-1492, US Atomic Energy Commission, Washington, DC, USA, November 19, 1952.

Feoli:2003:OAI

[FM03] Antonio Feoli and Giuliano Minichiello. Oscar D’Agostino: un irpino fra i ragazzi di via Panisperna. (German) [Oscar D’Agostino, an Irpinian among the boys of Via Panisperna], volume 7 of Studi meridionali. Edizioni del Centro Dorso, Avellino, Italy, 2003. ISBN ???? 122 pp. LCCN ????

Fermi:1954:PSA

[FMA54] Enrico Fermi, Nicholas Metropolis, and E. Felix Alei. Phase shift analysis of the scattering of negative pions by hydrogen. Physical Review, 95(6):1581–1585, September 15, 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://prola.aps.org/abstract/PR/v95/i6/
p1581_1;   http://prola.aps.org/pdf/PR/v95/i6/p1581_1.

Fermi:1943:SFN

[FMM43] Enrico Fermi, John Marshall, and Leona Marshall. Slowing down of fission neutrons in graphite. Report CP-1084, US Atomic Energy Commission, Washington, DC, USA, November 25, 1943.

Fermi:1945:FCS

[FMM45] Enrico Fermi, John Marshall, and Leona Marshall. Fission crosssection and ν-value for 25. Report CP-1186, US Atomic Energy Commission, Washington, DC, USA, December 31, 1945.

Fermi:1947:TNV

[FMM47] Enrico Fermi, John Marshall, and Leona Marshall. A thermal neutron velocity selector and its application to the measurement of the cross section of boron. Physical Review, 72(3):193–196, August 1, 1947. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/
PR/v72/i3/p193_1.

Fermi:1953:SMP

[FMN53] Enrico Fermi, R. L. Martin, and Darragh E. Nagle. Scattering of 169 and 192 MeV pions by hydrogen. Physical Review, 91(2):467, July 1953. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Feynman:1949:ESEa

[FMT49a] Richard P. Feynman, Nicholas Metropolis, and Edward Teller. Equations of state of elements based on the generalized Fermi– Thomas theory. Technical Report AECD-2448, Technical Information Branch, Oak Ridge Operations, AEC, Oak Ridge, TN, USA, January 20, 1949. 41 pp. URL http://www.osti.

gov/accomplishments/documents/fullText/ACC0107.pdf;
http://www.osti.gov/servlets/purl/4417654-BCgOtj/native/

.

Feynman:1949:ESEb

[FMT49b] Richard P. Feynman, Nicholas Metropolis, and Edward Teller. Equations of state of elements based on the generalized Fermi– Thomas theory. Physical Review (2), 75(10):1561–1573, May 15, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Folkart:1986:LML

[Fol86] Burt A. Folkart. Leona Marshall Libby dies; sole woman to work on Fermi’s 1st nuclear reactor. Los Angeles Times, ??(??):??, November 13, 1986. ISSN 0458-3035. URL

http://articles.latimes.com/1986-11-13/local/me-24930_
1_nuclear-reactor.

Fornaciari:2001:SDD

[For01] P. Fornaciari. Sommario dell’Intervento ≪Prospective dell’Energia nucleare in futuro≫. (Italian) [summary of the intervention ≪Prospective nuclear energy in the future≫ ]. Il Nuovo Saggiatore, 17(5–6):65–66, September/December 2001. ISSN 0393-4578 (print), 1827-6148 (electronic). URL http://prometeo.sif.it:

8080/papers/online/sag/2001/05-06/pdf/05.pdf. Special is-

sue in honor of the centennial of the birth of Enrico Fermi.

Forgan:2019:SFP

[For19] Duncan Forgan. Solving Fermi’s Paradox, volume 10 of Cambridge astrobiology. Cambridge University Press, Cambridge, UK, 2019. ISBN 1-107-16365-X (hardcover), 1-316-68151-3 (e-book). xvii + 413 pp. LCCN QB54 .F77945 2019.

Fermi:1923:SMD

[FP23] Enrico Fermi and A. Pontremoli. Sulla massa della radiazione in uno spazio vuoto. (Italian) [On the mass of radiation in empty space]. Rendiconti dell’Accademia Nazionale dei Lincei, 32(1): 162–165, ???? 1923. CODEN AANLAW. ISSN 0001-4435.

Fermi:1926:PDA

[FP26] Enrico Fermi and Enrico Persico. Il principio delle adiabatiche e la nozione di forza viva nella nuova meccanica ondulatoria. (Italian) [The principle of adiabatic invariance and the concept of kinetic energy in the new wave mechanics]. Rendiconti dell’Accademia Nazionale dei Lincei, 4 (II)(6):452–457, ???? 1926. CODEN AANLAW. ISSN 0001-4435.

Fermi:1938:FPS

[FP38] Enrico Fermi and Enrico Persico. Fisica per le Scuole Medie Superiori. (Italian) [Physics for Upper Secondary Schools]. N. Zanichelli, Bologna, Italy, 1938. 314 pp.

Frisch:1959:TAP

[FPLR59] O. R. Frisch, F. A. Paneth, F. Laves, and P. Rosbaud, editors. Trends in atomic physics; essays dedicated to Lise Meitner, Otto Hahn, Max von Laue on the occasion of their 80th birthday: Atomic physics. Interscience Publishers, New York, NY, USA, 1959. 285 pp. LCCN QC475 .B45 1959a. An identical edition is published simultaneously under the title Beiträge zur Physik und Chemie des 20. Jahrhunderts.

Fermi:1934:ASIb

[FPR34a] Enrico Fermi, Bruno Pontecorvo, and Franco Rasetti. Azione di sostanze idrogenate sulla radioattività provocate da neutroni. II. (Italian) [Influence of hydrogenous substances on the radioactivity produced by neutrons. II]. La Ricerca Scientifica, 5(2):282–283, ???? 1934. CODEN RISCAZ. ISSN 0035-5011. English translation by Franco Rasetti and Emilio Segré in Collected Papers.

Fermi:1934:ESI

[FPR34b] Enrico Fermi, Bruno Pontecorvo, and Franco Rasetti. Effetto di sostanze idrogenate sulla radioattività provocata da neutroni. II. La Ricerca Scientifica, 5(2):380–381, ???? 1934. CODEN RISCAZ. ISSN 0035-5011.

Fermi:1955:SNP

[FPU55] Enrico Fermi, John Pasta, and Stanisław M. Ulam. Studies of nonlinear problems. I. Technical Report LA-1940, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, May 1955. 22

pp. URL http://www.osti.gov/accomplishments/documents/

fullText/ACC0041.pdf. Also in Enrico Fermi: Collected Papers, volume 2, edited by Edoardo Amaldi, Herbert L. Anderson, Enrico Persico, Emilio Segré, and Albedo Wattenberg. Chicago: University of Chicago Press, 1965, pages 978–988.

Fermi:1974:SNP

[FPU74] Enrico Fermi, John Pasta, and Stanisław Ulam. Studies of nonlinear problems. I. In Nonlinear Wave Motion, Proceedings of the Summer Seminar, Potsdam (New York) 1972, pages 143–156. ????, ????, 1974.

Fermi:1925:EAM

[FR25a] Enrico Fermi and Franco Rasetti. Effect of an alternating magnetic field on the polarisation of the resonance radiation of mercury vapour. Nature, 115(2898):764, May 16, 1925. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v115/
n2898/pdf/115764b0.pdf.

Fermi:1925:ECMa

[FR25b] Enrico Fermi and Franco Rasetti. Effetto di un campo magnetico alternato sopra la polarizzazione della luce di risonanza. (Italian) [The effect of an alternating magnetic field on the polarization of resonance radiation]. Rendiconti dell’Accademia Nazionale dei Lincei, 1(6):716–722, ???? 1925. CODEN AANLAW. ISSN 0001-4435.

Fermi:1925:ECMb

[FR25c] Enrico Fermi and Franco Rasetti. Effetto di un campo magnetico alternato sopra la polarizzazione della luce di risonanza. (Italian) [The effect of an alternating magnetic field on the polarization of resonance radiation]. Rendiconti dell’Accademia Nazionale dei Lincei, 2(6):117–120, ???? 1925. CODEN AANLAW. ISSN 0001-4435.

Fermi:1925:EWM

[FR25d] Enrico Fermi and Franco Rasetti. Über den Einfluß eines wechselnden magnetischen Feldes auf die Polarisation der Resonanzstrahlung. (German) [On the influence of an alternating magnetic field on the polarization of resonance radiation]. Zeitschrift für Physik, 33(1):246–250, December 1925. CODEN ZEPYAA.

ISSN ????  URL http://www.springerlink.com/content/
p255870791167083/.

Fermi:1927:MVD

[FR27a] Enrico Fermi and Franco Rasetti. Eine Messung des Verhältnisses h/k durch die anomale Dispersion des Thalliumdampfes. (German) [A measurement of the ratio h/k in the anomalous dispersion of thallium vapor]. Zeitschrift für Physik, 43(5–6):379–383, May 1927. CODEN ZEPYAA. ISSN ???? URL http://www.

springerlink.com/content/j64786482l232041/.

Fermi:1927:MDR

[FR27b] Enrico Fermi and Franco Rasetti. Una misura del rapporto h/k per mezzo della dispersione anomala del tallio. (Italian) [A measure of the ratio h/k by means of the anomalous dispersion of thallium]. Rendiconti dell’Accademia Nazionale dei Lincei, 5(6): 566–570, ???? 1927. CODEN AANLAW. ISSN 0001-4435.

Fermi:1931:RSG

[FR31] Enrico Fermi and Franco Rasetti. Über den Ramaneffekt des Steinsalzes. (German) [On the Raman effect of rock salt]. Zeitschrift für Physik, 71(9–10):689–695, ???? 1931. CODEN

ZEPYAA. ISSN ???? URL http://www.springerlink.com/
content/h02441plp5389123/.

Fermi:1933:SPR

[FR33a] Enrico Fermi and Franco Rasetti. Uno spettrografo per raggi ‘gamma’ a cristello di bismuto. (Italian) [A gamma ray spectrograph for bismuth crystals]. La Ricerca Scientifica, 4(2):299–302, ???? 1933. CODEN RISCAZ. ISSN 0035-5011.

Fermi:1933:ADCa

[FR33b] Enrico Fermi and Bruno Rossi. Azione del campo magnetico terrestre sulla radiazione penetrante. (Italian) [Action of the Earth’s magnetic field on penetrating radiation]. Rendiconti dell’Accademia Nazionale dei Lincei, 17(??):346–350, ???? 1933. CODEN AANLAW. ISSN 0001-4435.

Fermi:1933:ADCb

[FR33c] Enrico Fermi and Bruno Rossi. Azione del campo magnetico terrestre sulla radiazione penetrante. (Italian) [Action of the Earth’s magnetic field on penetrating radiation]. Il Nuovo Cimento (8), 10(8):333–338, ???? 1933. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://www.

springerlink.com/content/h753rn58m8153231/.

Fermi:1935:RSN

[FR35] Enrico Fermi and Franco Rasetti. Ricerche sui neutroni lenti. (Italian) [Research on slow neutrons]. Il Nuovo Cimento (8), 12(4):201–210, April 1935. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://www.

springerlink.com/content/urq611j856217343/.

Fermi:1938:ADB

[FR38] Enrico Fermi and Franco Rasetti. Azione del boro sui neutroni caratteristici dello iodio. (Italian) [The action of boron on the neutrons characteristic of iodine]. La Ricerca Scientifica, 9(2): 472–473, ???? 1938. CODEN RISCAZ. ISSN 0035-5011.

Fermi:1948:NCT

[FR48] Enrico Fermi and R. D. Richtmyer. Note on census-taking in Monte-Carlo calculations. Report AECD-3164 (LADC-946; LAMS-805), US Atomic Energy Commission, Washington, DC, USA, July 11, 1948. 7 pp. URL http://scienze-como.

uninsubria.it/bressanini/montecarlo-history/fermi-1948.
pdf.

Fermi:1949:ODS

[FR49] Enrico Fermi and Isidor Isaac Rabi. An opinion on the development of the Super. This minority report to the Science Advisory Committee (SAC) opposed the development of the hydrogen bomb on ethical grounds. It is mentioned, but not properly cited, in [Rig00, Wol09]. It is reprinted in [WC84, pages 120–127] and [CHW91]., 1949.

Fermi:1976:MRH

[FR76] Enrico Fermi and Isadore I. Rabi. Minority report on the H-bomb: October 30, 1949. Bulletin of the Atomic Scientists, 32(10):58, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from Herbert F. York’s book, The Advisors/Oppenheimer, Teller, and the superbomb.

Fernandez:2013:UMA

[FR13] Bernard Fernandez and Georges Ripka. Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). xviii + 522 pp. LCCN QC773 .F47 2013.

Franklin:1986:EDT

[Fra86] Allan Franklin. Experiment and the development of the theory of weak interactions: Fermi’s theory. PSA: Proceedings of the Biennial Meeting of the Philosophy of Science Association, 1986 (1449):163–179, ???? 1986. ISSN 0270-8647. URL http://www.

jstor.org/stable/192798.

Franklin:1990:ERW

[Fra90a] Allan Franklin. Experiment, right or wrong. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-38207-6. x + 230 pp. LCCN Q175 .F785 1990. URL http://www.loc.gov/

catdir/description/cam023/89023848.html;   http://www.
loc.gov/catdir/toc/cam029/89023848.html.

Franklin:1990:FT

[Fra90b] Allan Franklin. Fermi’s theory. In Experiment, right or wrong [Fra90a], pages 9–24. ISBN 0-521-38207-6. LCCN Q175

.F785 1990. URL http://www.loc.gov/catdir/description/
cam023/89023848.html;   http://www.loc.gov/catdir/toc/
cam029/89023848.html.

Franklin:1990:TUF

[Fra90c] Allan Franklin. Toward a universal Fermi interaction: muons and pions. In Experiment, right or wrong [Fra90a], pages 25–41. ISBN 0-521-38207-6. LCCN Q175 .F785 1990. URL http://

www.loc.gov/catdir/description/cam023/89023848.html;
http://www.loc.gov/catdir/toc/cam029/89023848.html.

Frampton:1995:BRP

[Fra95] Paul H. Frampton. Book review: Particle Physics at the Fermi Scale by Yang Pang, Jianwei Qiu and Zhaoming Qiu. American Scientist, 83(3):274, ???? 1995. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://www.jstor.

org/stable/29775450. See [PQQ94].

Franklin:2004:TRN

[Fra04] Allan Franklin. Are there really neutrinos?: an evidential history. Perseus Publishers, Cambridge, MA, USA, 2004. ISBN 0-7382-0265-7, 0-8133-4128-0 (paperback). ix + 371 pp. LCCN QC793.5.N42 F73 2004. URL http://www.loc.gov/

catdir/enhancements/fy0837/2008530710-d.html;   http:/
/www.loc.gov/catdir/toc/fy0805/2008530710.html.

Frank:2005:ERG

[Fra05] Tibor Frank. Ever ready to go: The multiple exiles of Leo Szilard. Physics in Perspective (PIP), 7(2):204–252, June 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://adsabs.harvard.edu/abs/2005PhP...
..7..204F;   http://www.springerlink.com/content/1422-

Fermi:1934:SPP

[FRD34] Enrico Fermi, Franco Rasetti, and Oscar D’Agostino. Sulla possibilità di produrre elementi di numero atomico maggiore di 92. (Italian) [On the possibility of producing elements of atomic number higher than 92]. La Ricerca Scientifica, 5(1):536–537, ???? 1934. CODEN RISCAZ. ISSN 0035-5011.

Freitas:1985:TNF

[Fre85] Robert A. Freitas, Jr. There is no Fermi Paradox. Icarus: International Journal of Solar System Studies, 62(3):518–520, June 1985. CODEN ICRSA5. ISSN 0019-1035 (print), 1090-2643

(electronic). URL https://www.sciencedirect.com/science/
article/abs/pii/0019103585901927.

French:2008:MPM

[Fre08] A. P. French. In memoriam: Philip Morrison. Physics in Perspective (PIP), 10(1):110–122, March 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-007-0343-5.

Frisch:1939:PED

[Fri39] Otto Robert Frisch. Physical evidence for the division of heavy nuclei under neutron bombardment. Nature, 143(3616):276, February 18, 1939. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.chemteam.info/Chem-
History/Frisch-Fission-1939.html;   http://www.nature.
com/nature/journal/v143/n3616/pdf/143276a0.pdf.

Friendly:1950:QD

[Fri50] Fred Friendly. The quick and the dead. Four-part NBC radio miniseries on the creation of the atomic bomb, narrated by American entertainer Bob Hope, with Helen Hayes playing Lise Meitner and with Enrico Fermi and other leading scientists from the Manhattan Project receiving some mention., 1950. URL https://

www.discogs.com/National-Broadcasting-Company-The-The-
Quick-And-The-Dead-Volume-1-The-Atom-Bomb-/release/5350974;
 https://www.questia.com/library/journal/1P3-1718351551/
living-with-the-bomb-fred-friendly-s-the-quick-and.

Frisch:1954:AEH

[Fri54] Professor O. R. Frisch, O.B.E., F.R.S. Atomic energy — how it all began. British Journal of Applied Physics, 5(3):81–84, March 1954. CODEN BJAPAJ. ISSN 0508-3443 (print), 2057-7656 (elec-

tronic). URL http://iopscience.iop.org/0508-3443/5/3/
301;   http://stacks.iop.org/0508-3443/5/i=3/a=301. A

lecture delivered in London to the Education Group of The Institute of Physics on 20 October, 1953.

Frisch:1955:PEF

[Fri55a] O. R. Frisch. Prof. Enrico Fermi, For. Mem. R. S. Nature, 175 (4444):18–19, January 1, 1955. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v175/n4444/pdf/175018a0.pdf.

Frisch:1955:BRA

[Fri55b] Otto Robert Frisch. Book review: Atoms in the Family: My Life with Enrico Fermi — Designer of the First Atomic Pile, by Laura Fermi. Pp. 284 + 15 plates (London: George Allen and Unwin, Ltd., 1955). Nature, 176(4488):850, November 5, 1955. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v176/
n4488/pdf/176850b0.pdf.

Frisch:1974:ONF

[Fri74] Otto Robert Frisch. The origin of nuclear fission. [Sound recording]. Spring Green Multimedia, Washington, DC, USA, 1974. LCCN QC790; RYB 6354. 1 cassette tape.

Friedman:2001:PEB

[Fri01] Robert Marc Friedman. The politics of excellence: behind the Nobel Prize in science. Times Books, New York, NY, USA, 2001. ISBN 0-7167-3103-7. xv + 379 pp. LCCN QC49 .F75 2001.

Fermi:1933:STD

[FS33a] Enrico Fermi and Emilio Segrè. Sulla teoria delle strutture iperfini. (Italian) [On the theory of hyperfine structure]. Mem. Accad. d’Italia, 4 (Fis.):131–158, 1933.

Fermi:1933:THG

[FS33b] Enrico Fermi and Emilio Segrè. Zur Theorie der Hyperfeinstruktur. (German) [On the theory of hyperfine structure]. Zeitschrift für Physik, 82(11–12):729–749, November 1933. CO-

DEN ZEPYAA. ISSN ???? URL http://www.springerlink.
com/content/m585205865327k6g/.

Fermi:1934:TRa

[FS34a] Enrico Fermi and Emilio Segrè. Theory of p-rays. Il Nuovo Cimento (8), 11(??):1–19, ???? 1934. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Fermi:1934:TRb

[FS34b] Enrico Fermi and Emilio Segrè. Theory of p-rays. Zeitschrift für Physik, 88(??):761–771, ???? 1934. CODEN ZEPYAA. ISSN 0044-3328.

Fermi:1941:FUA

[FS41] Enrico Fermi and Emilio Segrè. Fission of uranium by alphaparticles. Physical Review, 59(8):680–681, April 15, 1941. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v59/i8/p680_2.

Fermi:1950:PIT

[FS50] Enrico Fermi and Leo Szilard. On the possibility of initiating a thermonuclear reaction in a mass of deuterium. Report, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, Summer 1950.

Fermi:1955:NRa

[FS55] Enrico Fermi and Leo Szilárd. Neutronic reactor. US Patent 2,708,656., May 17, 1955. URL http://www.google.com/ patents?vid=2708656. US Patent Application 568,904 filed December 19, 1944.

Fermi:1957:MON

[FS57a] Enrico Fermi and Leo Szilárd. Method of operating a neutronic reactor. US Patent 2,798,847., July 9, 1957. US Patent Application 323,452 filed December 1, 1952.

Fermi:1957:NR

[FS57b] Enrico Fermi and Leo Szilárd. Neutronic reactor. US Patent 2,807,581., September 24, 1957. URL http://www.google.com/ patents/US2807581.. US Patent Application 621,838 filed October 11, 1945.

Fermi:1958:ACN

[FS58a] Enrico Fermi and Leo Szilárd. Air cooled neutronic reactor. US Patent 2,836,554., May 27, 1958. URL http://www.google.com/ patents/US2836554. US Patent Application 596,465 filed May 29, 1945.

Fermi:1958:CR

[FS58b] Enrico Fermi and Leo Szilard. Chain reactions. Canadian Patent

552312., January 28, 1958. URL http://brevets-patents.ic.
gc.ca/opic-cipo/cpd/eng/patent/552312/summary.html.

Fermi:1961:ACN

[FS61] Enrico Fermi and Leo Szilard. Air-cooled neutronic reactor. Canadian Patent 620923., May 30, 1961. URL http://brevets-

patents.ic.gc.ca/opic-cipo/cpd/eng/patent/620923/summary.
html.

Fermi:1995:PBP

[FS95] Rachel Fermi and Esther Samra. Picturing the Bomb: Photographs from the Secret World of the Manhattan Project. H. N. Abrams, New York, NY, USA, 1995. ISBN 0-8109-3735-2. 232 pp. LCCN QC773.A1 F47 1995. Introduction by Richard Rhodes.

Fermi:1941:PRN

[FSA41] Enrico Fermi, Henry D. Smyth, and Herbert L. Anderson. A preliminary report to the National Defense Research Committee for the Department of Physics of Princeton University under Contract NDCrc-121. Appendix A: Enrico Fermi and Herbert L. Anderson: Capture of Resonance Neutrons by a Uranium Sphere Imbedded in Graphite. Report A-12, US Atomic Energy Commission, Washington, DC, USA, June 1, 1941. 10 pp.

Fermi:1947:TSN

[FSS47] Enrico Fermi, W. J. Sturm, and Robert G. Sachs. The transmission of slow neutrons through microcrystalline materials. Physical Review, 71(9):589–594, May 1, 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v71/i9/p589_1.

Fermi:1943:RDN

[FT43] Enrico Fermi and G. Thomas. The range of delayed neutrons. Report, US Atomic Energy Commission, Washington, DC, USA, November 23, 1943. Excerpt from Report CP-1088 for Month Ending November 23, 1943.

Fermi:1947:CNM

[FT47] Enrico Fermi and E. Teller. The capture of negative mesotrons in matter. Physical Review (2), 72(5):399–408, September 1, 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://inspirehep.net/record/44994;
http://prola.aps.org/abstract/PR/v72/i5/p399_1.

Fermi:1947:DNM

[FTW47] Enrico Fermi, Edward Teller, and Victor F. Weisskopf. The decay of negative mesotrons in matter. Physical Review, 71(5):314–315, March 1, 1947. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://inspirehep.net/record/
45186;   http://prola.aps.org/abstract/PR/v71/i5/p314_

1. Reprinted in [FTW65].

Fermi:1965:DNM

[FTW65] E. Fermi, E. Teller, and V. Weisskopf. The decay of negative mesotrons in matter. In Nuclear Forces [Bri65], pages 227–229. ISBN 0-08-011034-7. LCCN QC173 .B8513 1965. Reprint of [FTW47].

Fermi:1933:REP

[FU33a] Enrico Fermi and George E. Uhlenbeck. On the recombination of electrons and positrons. Physical Review, 44(6):510–511, September 15, 1933. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v44/i6/p510_2.

Fermi:1933:SRE

[FU33b] Enrico Fermi and George E. Uhlenbeck. Sulla ricombinazione di elettroni e positroni. (Italian) [On the recombination of electrons and positrons]. Ricerca Scientifica ed il Progresso Tecnico, 4(2): 157–160, 1933. CODEN RSPTB6.

Fermi:1943:UHH

[FU43a] Enrico Fermi and H. C. Urey. The utilization of heavy hydrogen in nuclear chain reactions. Technical Report A-544, US Atomic Energy Commission, Washington, DC, USA, March 1943. Memorandum of Conference on March 6–8, 1943.

Fermi:1943:MCB

[FU43b] Enrico Fermi and Harold C. Urey. Memorandum of conference between Prof. Enrico Fermi and Prof. H. C. Urey. (on the Utilization of Heavy Hydrogen in Nuclear Reactions). Report A-554, US Atomic Energy Commission, Washington, DC, USA, March 6–8, 1943.

Fukushima:2014:CGI

[Fuk14] Toshio Fukushima. Computation of a general integral of Fermi– Dirac distribution by McDougall–Stoner method. Applied Mathematics and Computation, 238(??):485–510, July 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S009630031400561X.

Falcioni:2002:CEF

[FV02] Massimo Falcioni and Angelo Vulpiani. Il contributo di Enrico Fermi ai sistemi non lineari: L’influenza di un articolo mai pubblicato. (Italian) [The contribution by Enrico Fermi to nonlinear systems: The influence of an article never published]. In Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901–2001] [Ber02a], pages 274–289. ISBN 88-7438-000-3. LCCN QC774.F4. 10.33 EUR. URL http://prometeo.sif.it: 8080/libri/fermi/15.pdf. English translation in [BB04b].

Fermi:1951:TIIb

[FvN51] Enrico Fermi and John von Neumann. Taylor instability at the boundary of two incompressible liquids. Document AECU-2979, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, August 19, 1951. Part 2.

Fermi:1955:TII

[FvN55] Enrico Fermi and John von Neumann. Taylor instability of incompressible liquids. U.S. Government Document AECU-2979, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, November

19, 1955. 17 pp. URL http://www.osti.gov/accomplishments/
documents/fullText/ACC0042.pdf. Reprinted in [Tau63, Paper

31].

Fermi:1941:ATN

[FW41] Enrico Fermi and George L. Weil. The absorption of thermal neutrons by a uranium sphere imbedded in graphite. Report A-1, US Atomic Energy Commission, Washington, DC, USA, July 3,

Fermi:1942:LDC

[FW42] Enrico Fermi and Alvin M. Weinberg. Longitudinal diffusion in cylindrical channels. Report C-170, US Atomic Energy Commission, Washington, DC, USA, July 7, 1942.

Fermi:1943:RIR

[FW43] Enrico Fermi and George L. Weil. Range of indium resonance neutrons from source of fission neutrons. Report CP-871, US Atomic Energy Commission, Washington, DC, USA, August 1943. 2 pp. In: Fermi, E., et al. Physics Research. Report for month ending August 14, 1943.

Frisch:1967:DFH

[FW67] Otto Robert Frisch and John A. Wheeler. The discovery of fission: How it all began and mechanism of fission. Physics Today, 20(11):43–52, November 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

www.physicstoday.org/resource/1/phtoad/v20/i11/p43_s1.

Reprinted in [WP85, pages 272–281].

Frisch:1985:DF

[FW85] Otto R. Frisch and John A. Wheeler. The discovery of fission. In Weart and Phillips [WP85], pages 272–281. ISBN 0-88318-468-0 (paperback). LCCN QC7 .H694 1985. Reprint of [FW67].

Fermi:1945:BPD

[FWN^{+}45] Enrico Fermi, Eugene P. Wigner, Lothar W. Nordheim, Alvin M. Weinberg, H. Soodak, H. S. Brown, Miles C. Leverett, F. Daniels, G. Young, and Glenn T. Seaborg. Breeder pile discussion. Report, US Atomic Energy Commission, Washington, DC, USA, June 19– 20, 1945.

Fermi:1949:MEP

[FY49] Enrico Fermi and Chen-Ning Franklin Yang. Are mesons elementary particles? Physical Review, 76(12):1739–1743, December 15, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://inspirehep.net/record/21797;
http://prola.aps.org/abstract/PR/v76/i12/p1739_1.

Fermi:1943:TSW

[FZ43] Enrico Fermi and Walter H. Zinn. Tests on a shield for the W pile. Report CP-684, US Atomic Energy Commission, Washington, DC, USA, May 25, 1943.

Fermi:1944:CNB

[FZ44] Enrico Fermi and Walter H. Zinn. Collimation of neutron beam from thermal column of CP-3 and the index of refraction for thermal neutrons. Report CP-1965., US Atomic Energy Commission, Washington, DC, USA, July 1944. 3 pp. In: Fermi, E., et al. Argonne Laboratory. Physics Research. Report for month ending July 29, 1944.

Fermi:1946:RNMa

[FZ46a] Enrico Fermi and Walter H. Zinn. Reflection of neutrons on mirrors. Report MDCC-56 (CU-10; MUC-AJD-166), US Atomic Energy Commission, Washington, DC, USA, July 1946. Abstract published in Phys. Rev. 70:103 (1946). Physical Society Cambridge (England) Conference Report 1947, page 92.

Fermi:1946:RNMb

[FZ46b] Enrico Fermi and Walter H. Zinn. Reflection of neutrons on mirrors. Physical Review (2), 70(1–2):103, July 1946. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v70/i1-2/p99_1.

Fermi:1955:NRb

[FZ55] Enrico Fermi and Walter H. Zinn. Neutronic reactor. US Patent 2,714,577., August 2, 1955. US Patent Application 626,383 filed November 2, 1945.

Fermi:1957:NRS

[FZ57] Enrico Fermi and Walter H. Zinn. Neutronic reactor shield. US Patent 2,807,727., September 24, 1957. US Patent Application 641,625 filed January 16, 1946.

Fermi:1958:NR

[FZA58] Enrico Fermi, Walter H. Zinn, and Herbert L. Anderson. Neutronic reactor. US Patent 2,852,461., September 16, 1958. US Patent Application 621,837 filed October 11, 1945.

G:1964:FPM

[G.64] D. S. G. Fermi Prize money: Congressional Committee takes steps to assume control of annual $50,000 award. Science, 143(3612): 1305, March 20, 1964. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
1713224;   http://www.sciencemag.org/content/143/3612/

1305.1.full.pdf. See also [Wal64b].

Grosse:1934:FE

[GA34] A. V. Grosse and M. S. Agruss. Fermi’s element 93. Nature, 134 (3394):773, November 17, 1934. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v134/n3394/pdf/134773b0.pdf.

Graetzer:1971:DNF

[GA71] Hans G. Graetzer and David L. Anderson. The discovery of nuclear fission: a documentary history, volume 20 of Van Nostrand Reinhold momentum books. Van Nostrand Reinhold, New York, NY, USA, 1971. viii + 120 pp. LCCN QC790 .G68.

Gowing:1974:IDBa

[GA74a] Margaret Gowing and Lorna Arnold. Independence and deterrence: Britain and atomic energy, 1945–1952. Macmillan Publishing Company, New York, NY, USA, 1974. ISBN 0-333-15781-8 (vol. 1). ???? pp. LCCN QC773.3.G7 G68 1974b.

Gowing:1974:IDBb

[GA74b] Margaret Gowing and Lorna Arnold. Independence and deterrence: Britain and atomic energy, 1945–1952. St. Martin’s Press, New York, NY, USA, 1974. ???? pp. LCCN QC773.3.G7 G68 1974.

Gallavotti:2002:MCR

[Gal02] Giovanni Gallavotti. La meccanica classica e la rivoluzione quantistica nei lavori giovanili di Fermi. (Italian) [Classical mechanics and the quantum revolution in Fermi’s early work]. In Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901–2001] [Ber02a], pages 76–84. ISBN 88-7438-000-3. LCCN QC774.F4. 10.33 EUR. URL http://prometeo.sif. it:8080/libri/fermi/05.pdf. English translation in [BB04b].

Gallavotti:2003:FEP

[Gal03] Giovanni Gallavotti. Fermi and the ergodic problem. In Anonymous [Ano03], pages 295–302. ISBN 88-8286-032-9. LCCN

???? URL http:old.enea.it/produzione_scientifica/pdf.
../V2003_AttiFermi.pdf.

Gamow:1985:TYS

[Gam85] George Gamow. Thirty Years That Shook Physics: the Story of Quantum Theory. Dover, New York, NY, USA, 1985. ISBN 0-486-24895-X (paperback). xiv + 224 + 9 pp. LCCN QC174.12

.G35 1985. US$4.95. URL http://www.loc.gov/catdir/
description/dover032/85006797.html.

Gambassi:2003:EFP

[Gam03] Andrea Gambassi. Enrico Fermi in Pisa. Physics in Perspective (PIP), 5(4):384–397, December 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

www.springerlink.com/content/542n12h2hmxk4qdp/.

Gandini:2002:DCP

[Gan02] Augusto Gandini. Dalla Chicago Pile 1 ai reattori della prossima generazione. (Italian) [From Chicago Pile 1 to the next generation of reactors]. In Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901–2001] [Ber02a], pages 205– 222. ISBN 88-7438-000-3. LCCN QC774.F4. 10.33 EUR. URL

http://prometeo.sif.it:8080/libri/fermi/12.pdf. English

translation in [BB04b].

Garwin:1992:FM

[Gar92] Richard L. Garwin. Fermi’s mistake? Nature, 355(6362):668, February 20, 1992. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v355/n6362/pdf/355668d0.pdf.

Garwin:2004:GFC

[Gar04] Dick Garwin. Glimpses of Fermi in Chicago and Los Alamos. In Orear [Ore04], chapter 24, pages 121–124. ISBN ???? LCCN ????

URL http://dspace.library.cornell.edu/handle/1813/62;
 http://hdl.handle.net/1813/74.

Gariboldi:2018:BRD

[Gar18] Leonardo Gariboldi. Book review: David N. Schwartz. The Last Man Who Knew Everything: The Life and Times of Enrico Fermi, Father of the Nuclear Age. Isis, 109(2):425–427, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Galison:1989:LSD

[GB89] Peter Galison and Barton Bernstein. In any light: Scientists and the decision to build the superbomb, 1952–1954. Historical Studies in the Physical and Biological Sciences, 19(2):267–347, ????

1989. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355

(electronic). URL http://www.jstor.org/stable/27757627.

Goodstein:2012:FLF

[GB12] Judith Goodstein and Donald Babbitt. A fresh look at Francesco Severi. Notices of the American Mathematical Society, 59(8): 1064–1075, September 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://www.ams.org/

notices/201208/rtx120801064p.pdf.

Garraty:1999:ANB

[GC99] John A. (John Arthur) Garraty and Mark C. (Mark Christopher) Carnes, editors. American national biography. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-520635-5 (set), 0-19-512780-3 (vol. 1), 0-19-512781-1 (vol. 2), 0-19-512782-X (vol. 3), 0-19-512783-8 (vol. 4), 0-19-512784-6 (vol. 5), 0-19-512785-4 (vol. 6), 0-19-512786-2 (vol. 7), 0-19-512787-0 (vol. 8), 0-19-512788-9 (vol. 9), 0-19-512789-7 (vol. 10), 0-19-512790-0 (vol. 11), 0-19-512791-9 (vol. 12), 0-19-512792-7 (vol. 13), 0-19-512793-5 (vol. 14), 0-19-512794-3 (vol. 15), 0-19-512795-1 (vol. 16), 0-19-512796-X (vol. 17), 0-19-512797-8 (vol. 18), 0-19-512798-6 (vol. 19), 0-19-512799-4 (Vol. 20), 0-19-512800-1 (vol. 21), 0-19-512801-X (vol. 22), 0-19-512802-8 (vol. 23), 0-19-512803-6 (vol. 24). 24 v. pp. LCCN CT213 .A68 1999. URL http://www.loc.

gov/catdir/enhancements/fy0603/98020826-d.html.

George:1932:MQCa

[Geo32] André George. Mécanique quantique et causalité d’après M. Fermi. (French) [Quantum mechanics and casuality, according to Mr. Fermi], volume 38 of Actualités scientifiques et industrielles. Hermann, Paris, France, 1932. 18 pp. LCCN ???? With remarks of M. Louis de Broglie.

Galbiata:1992:PPF

[GER^{+}92] Domenico Galbiata, Padre Eligio, R. A. Ricci, et al., editors. Scienza ed etica alle soglie del terzo millennio: Varenna, Villa Monastero, 28–30 settembre 1992, volume 36. Società italiana di fisica, Bologna, Italy, 1992. ISBN 88-7794-051-4. LCCN ????

Gombas:1951:STF

[GG51] P. Gombás and R. Gáspár. Solution of the Thomas–Fermi– Dirac equation. Nature, 168(4264):122, July 21, 1951. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v168/
n4264/pdf/168122a0.pdf.

Gillispie:1981:DSB

[GH90] Charles Coulston Gillispie and Frederic Lawrence Holmes, editors. Dictionary of scientific biography. Charles Scribner’s Sons, New York, NY, USA, 1981–1990. ISBN 0-684-16963-0 (v.1/2), 0-684-16964-9 (v.3/4), 0-684-16965-7 (v.5/6), 0-684-16966-5 (v.7/8), 0-684-16967-3 (v.9/10), 0-684-16968-1 (v.11/12), 0-684-16969-X (v.13/14), 0-684-16970-3 (v.15/16), 0-684-19177-6 (v.17), 0-684-19178-4 (v.18), 0-684-16962-2 (set). various pp. LCCN Q141 .D5 1981.

Giacomuzzi:2002:EMB

[GHH02] Salvatore Matteo Giacomuzzi, Gerhard Holzmüller, and Gerhard Huemer. Ettore Majorana (1906–1938): Eine Bestandsaufnahme 64 Jahre nach seinem Verschwinden. (German) [Ettore Majorana (1906–1938): A survey 64 years after his disappearance]. Berichte zur Wissenschaftsgeschichte, 25(2):137–148, June 2002. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

GarciaDoncel:1987:SPP

[GHMP87] Manuel García Doncel, A. Hermann, L. Michel, and A. Pais, editors. Symmetries in physics (1600-1980): proceedings of the 1st International Meeting on the History of Scientific Ideas held at Sant Feliu de Guíxols, Catalonia, Spain, September 20–26, 1983. Seminari d’História de les Ciéncies, Universitat Autónoma de Barcelona, Bellaterra, Barcelona, Spain, 1987. ISBN 84-7488-148-X. LCCN QC174.17.S9 I57 1983.

Golub:2023:HPF

[GL23] Robert Golub and Steve Keith Lamoreaux. The Historical and Physical Foundations of Quantum Mechanics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2023. ISBN 0-19-186123-5, 0-19-255536-7, 0-19-882218-9 (hardcover), 0-19-882219-7 (paperback). xiii + 747 pp. LCCN QC174.12 .G65 2023.

Glauber:2002:EEF

[Gla02] Roy Glauber. An excursion with Enrico Fermi, 14 July 1954. Physics Today, 55(6):44–46, June 2002. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v55/i6/p44_s1.

Glicksman:2004:GC

[Gli04] Maurice Glicksman. Glicksman comment. In Orear [Ore04], chapter 28, pages 141–142. ISBN ???? LCCN ???? URL http:

//dspace.library.cornell.edu/handle/1813/62;   http://
hdl.handle.net/1813/74.

Guerra:2006:EFD

[GLR06] Francesco Guerra, Matteo Leone, and Nadia Robotti. Enrico Fermi’s discovery of neutron-induced artificial radioactivity: Neutrons and neutron sources. Physics in Perspective (PIP), 8(3):255–281, September 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://www.

springerlink.com/content/w233203272t22206/.

Guerra:2014:WEC

[GLR14] Francesco Guerra, Matteo Leone, and Nadia Robotti. When energy conservation seems to fail: The prediction of the neutrino. Science & Education (Springer), 23(6):1339–1359, June 2014. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Guerra:2020:EFD

[GLR20] Francesco Guerra, Matteo Leone, and Nadia Robotti. Enrico Fermi’s discovery of neutron-induced artificial radioactivity: A case of “Emanation” from “Divine Providence”. Physics in Perspective (PIP), 22(3):129–161, September 2020. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Gautschi:1985:GQI

[GM85a] Walter Gautschi and Gradimir V. Milovanovi. Gaussian quadrature involving Einstein and Fermi functions with an application to summation of series. Mathematics of Computation, 44(169): 177–190, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.jstor.org/

stable/2007801.

Gautschi:1985:SGQ

[GM85b] Walter Gautschi and Gradimir V. Milovanovi. Supplement to Gaussian quadrature involving Einstein and Fermi functions with an application to summation of series. Mathematics of Computation, 44(169):S1–S11, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.jstor.org/stable/2007816.

GoeppertMayer:1948:CSN

[Goe48] Maria Goeppert Mayer. On closed shells in nuclei. Physical Review (2), 74(3):235–239, August 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRev.74.235.

GoeppertMayer:1949:CSN

[Goe49] Maria Goeppert Mayer. On closed shells in nuclei. II. Physical Review (2), 75(12):1969–1970, June 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http:

//link.aps.org/doi/10.1103/PhysRev.75.1969;   http://
prola.aps.org/abstract/PR/v75/i12/p1969_1;   http://www.
nobelprize.org/nobel_prizes/physics/laureates/1963/.

Goldschmidt:1962:FCD

[Gol62] Bertrand Goldschmidt. France’s contribution to the discovery of the chain reaction. International Atomic Energy Agency Bulletin, 4(0):21–24, ???? 1962. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-2690 (electronic). URL http://www.iaea.org/

Publications/Magazines/Bulletin/Bull040su/04004782124su.
pdf.

Goldberger:1970:BRL

[Gol70] M. L. Goldberger. Book review: A leader in physics: Enrico Fermi, Physicist by Emilio Segrè. Science, 169(3948):847, August 28, 1970. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1729724;
http://www.sciencemag.org/content/169/3948/847.1.full.

pdf. See [Seg70b].

Goldschmidt:1980:CAH

[Gol80] Bertrand Goldschmidt. Le complexe atomique: histoire politique de l’énergie nucléaire. (French) [The Atomic Complex: political history of nuclear energy]. Fayard, Paris, France, 1980. ISBN 2-213-00773-X. 493 pp. LCCN QC773 .G63. 69.00F.

Goldschmidt:1982:SC

[Gol82a] B. Goldschmidt. Summer 1942 in Chicago. International Atomic Energy Agency Bulletin, 24(4):3–6, ???? 1982. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-2690 (electronic). URL http:/

/www.iaea.org/Publications/Magazines/Bulletin/Bull244/
24404780306.pdf.

Goldschmidt:1982:ACW

[Gol82b] Bertrand Goldschmidt. The atomic complex: a worldwide political history of nuclear energy. American Nuclear Society, La Grange Park, IL, USA, 1982. ISBN 0-89448-550-4 (paperback). xiv + 479 pp. LCCN QC773 .G6313 1982. Translated from the French by Bruce M. Adkins. Revised and updated translation of Le complexe atomique [Gol80].

Goldberg:1989:BON

[Gol89] Stanley Goldberg. Between old and new: Goudsmit at Brookhaven. In De Maria et al. [DGS89], pages 125–129. ISBN 9971-5-0740-4. LCCN QC9.U5 I57 1988.

Goldberg:1992:ICO

[Gol92] Stanley Goldberg. Inventing a climate of opinion: Vannevar Bush and the decision to build the bomb. Isis, 83(3):429–452, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233904.

Goldberger:1999:BRD

[Gol99a] Marvin L. Goldberger. Book review: Dan Cooper, Enrico Fermi and the Revolution in Modern Physics. New York: Oxford University Press, 1999, 117 pages. $21.00 (cloth). Physics in Perspective (PIP), 1(2):226–227, June 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Goldberger:1999:AEF

[Gol99b] Marvin L. Goldberger. In appreciation: Enrico Fermi (1901– 1954): The complete physicist. Physics in Perspective (PIP), 1(3):328–336, October 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050024;   http://
www.springerlink.com/content/y1h02gr3rv7wurrl/.

Goncharov:1996:TMA

[Gon96a] German A. Goncharov. Thermonuclear milestones: (1) the American effort. Physics Today, 49(11):45–48, November 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Goncharov:1996:TMB

[Gon96b] German A. Goncharov. Thermonuclear milestones: (2) beginnings of the Soviet H-bomb program. Physics Today, 49(11):50– 54, November 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Goodstein:1991:BRL

[Goo91] Judith Goodstein. Book review: Lanfranco Belloni: Da Fermi a Rubbia: Storia e politica di un successo mondiale della scienza italiana. Isis, 82(4):770–771, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233387.

Goodstein:2001:CFR

[Goo01] Judith R. Goodstein. A conversation with Franco Rasetti. Physics in Perspective (PIP), 3(3):271–313, September 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/

PL00000534. This interview was conducted 4 February 1982 but was unpublished for 19 years.

Goudsmit:1932:PDT

[Gou32] Samuel Goudsmit. Present difficulties in the theory of hyperfine structure. In Convegno di Fisica Nucleare, Ottobre 1931, pages 33–49. Reale Accademia d’Italia, Rome, Italy, 1932.

Goudsmit:1959:FB

[Gou59] S. A. Goudsmit. The fermi and the barn. Physical Review Letters, 3(4):161–162, August 1959. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1971:DES

[Gou71] Samuel A. Goudsmit. The discovery of the electron spin. Web

document, April 1971. URL http://www.lorentz.leidenuniv.
nl/history/spin/goudsmit.html. English translation by J. H.

van der Waals of Goudsmit’s lecture in Dutch for the golden jubilee of the Dutch Physical Society in April 1971.

Gowing:1964:BAEa

[Gow64a] Margaret Gowing. Britain and atomic energy, 1939–1945. Macmillan Publishing Company, New York, NY, USA, 1964. xvi + 464 pp. LCCN QC773.A1 G6 1964a.

Gowing:1964:BAEb

[Gow64b] Margaret Gowing. Britain and Atomic Energy, 1939–1945. St. Martin’s Press, New York, NY, USA, 1964. xvi + 464 pp. LCCN QC773.A1 G6 1964. With an introductory chapter by Kenneth Jay.

Guth:1931:AFT

[GP31] Eugen Guth and Rudolf E. Peierls. Application of the Fermi– Thomas model to positive ions. Physical Review (2), 37:217, January 1931. CODEN PHRVAO. ISSN 0031-899x (print),

1536-6065 (electronic). URL https://ui.adsabs.harvard.edu/
#abs/1931PhRv...37..217G.

Grodzins:1963:AAS

[GR63] Morton Grodzins and Eugene I. Rabinowitch, editors. The Atomic Age: Scientists in National and World Affairs. Articles from the Bulletin of the Atomic Scientists 1945–1962. Basic Books, New York, NY, USA, 1963. xviii + 616 pp. LCCN D842 .B78. With the assistance of Harvey Flaumenhaft and Lois Gradner.

Goldstein:1986:SET

[GR86] Jerome A. Goldstein and Gisèle Ruíz Rieder. Some extensions of Thomas–Fermi theory. Lecture Notes in Mathematics, 1223: 110–121, 1986. CODEN LNMAA2. ISBN 3-540-17191-6 (print), 3-540-47350-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099187/.

Goldstein:1989:RRR

[GR89] Jerome A. Goldstein and Gisèle Ruíz Rieder. Recent rigorous results in Thomas–Fermi theory. Lecture Notes in Mathematics, 1394:68–82, 1989. CODEN LNMAA2. ISBN 3-540-51594-1 (print), 3-540-46679-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086753/.

Guerra:2005:FPE

[GR05] Francesco Guerra and Nadia Robotti. A forgotten publication of Ettore Majorana on the improvement of the Thomas–Fermi statistical model. arxiv.org, ??(??):??, November 28, 2005. URL http:

//arxiv.org/abs/physics/0511222. See rebuttal in [Esp05].

Guerra:2008:EMF

[GR08] Francesco Guerra and Nadia Robotti. Ettore Majorana’s forgotten publication on the Thomas–Fermi model. Physics in Perspective (PIP), 10(1):56–76, March 2008. CODEN PHPEF2.

ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-007-0340-8;
http://www.springerlink.com/content/53l26g8353l2421g/.

Guerra:2009:EFD

[GR09] Francesco Guerra and Nadia Robotti. Enrico Fermi’s discovery of neutron-induced artificial radioactivity: The influence of his theory of beta decay. Physics in Perspective (PIP), 11(4):379–404, December 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http:/

/adsabs.harvard.edu/abs/2009PhP....11..379G;   http://
www.springerlink.com/content/65w213132524072n/.

Graubard:1961:BRG

[Gra61] Mark Graubard. Book review: Galileo and the Scientific Revolution. Laura Fermi and Gilberto Bernardini. Basic Books, New York, 1961. x + 150 pp. Illus. $3.50. Science, 133(3463):1472, May 12, 1961. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/133/
3463/1472.3.full.pdf.

Graetzer:1964:DNF

[Gra64] Hans G. Graetzer. Discovery of nuclear fission. American Journal of Physics, 32(9):9–15, January 1964. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). See [HS39b, Bey49, HS89].

Gray:2015:FPN

[Gra15] Robert H. Gray. The Fermi Paradox is neither Fermi’s nor a paradox. Astrobiology, 15(3):195–199, March 2015. CODEN ASTRC4. ISSN 1531-1074 (print), 1557-8070 (electronic). URL https://

www.liebertpub.com/doi/10.1089/ast.2014.1247.

Gray:2016:FPF

[Gra16a] Robert H. Gray. The Fermi Paradox is not Fermi’s, and it is not a paradox: Despite what you’ll often read, the Nobel prizewinning nuclear physicist never suggested that aliens don’t exist. Scientific American blog Web site, January 29,

2016. URL https://blogs.scientificamerican.com/guest-
blog/the-fermi-paradox-is-not-fermi-s-and-it-is-not-a-
paradox/.

Gray:2016:CFP

[Gra16b] Robert H. Gray. The so-called Fermi Paradox is misleading, flawed, and harmful. Astrobiology, 16(10):741–743, October 2016. CODEN ASTRC4. ISSN 1531-1074 (print), 1557-8070 (elec-

tronic). URL https://www.liebertpub.com/doi/pdfplus/10.
1089/ast.2016.0823.rcm. Reply to [Ćir16].

Greenspan:2005:ECW

[Gre05] Nancy Thorndike Greenspan. The end of the certain world: the life and science of Max Born: the Nobel physicist who ignited the quantum revolution. Basic Books, New York, NY, USA, 2005. ISBN 0-7382-0693-8 (hardcover). x + 374 + 16 pp. LCCN QC16.B643 G74 2005.

Griffiths:1996:BRE

[Gri96] David J. Griffiths. Book review: Enrico Fermi, Notes on Quantum Mechanics, 2nd edition. American Journal of Physics, 64(4): 510–511, April 1996. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://ajp.aapt.org/resource/
1/ajpias/v64/i4/p510_s1;   http://link.aip.org/link/?
AJP/64/510/1.

Grimvall:2004:SFS

[Gri04] Göran Grimvall. Socrates, Fermi, and the Second Law of Thermodynamics. American Journal of Physics, 72(9):1145–1146, September 2004. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Gribbin:2011:AUS

[Gri11] John R. Gribbin. Alone in the Universe: Supercomets and the Emergence of Intelligent Life on Earth. Wiley, New York, NY, USA, 2011. ISBN 1-118-14797-9 (hardcover), 1-118-17539-5 (e-book). 237 pp. LCCN QB632 .G75 2011. URL http://

ebookcentral.proquest.com/lib/qut/detail.action?docID=
3058724.

Gribbin:2018:AMW

[Gri18] John Gribbin. Alone in the Milky Way: Why we are probably the only intelligent life in the galaxy. Scientific American, 319 (2):94–99, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL https://www.nature.com/

scientificamerican/journal/v319/n3/full/scientificamerican0918-
94.html.

Greiner:2003:EMB

[GRJ03] Markus Greiner, Cindy A. Regal, and Deborah S. Jin. Emergence of a molecular Bose–Einstein condensate from a Fermi gas. Nature, 426(6966):537–540, November 26, 2003. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v426/
n6966/full/nature02199.html.

Groves:1962:NIC

[Gro62] Leslie R. Groves. Now it can be told: the story of the Manhattan Project. Harper & Row, New York, NY, USA, 1962. xiv + 464 pp. LCCN QC773.A1 G7.

Groueff:1967:MPUa

[Gro67a] Stéphane Groueff. Manhattan Project: the untold story of the making of the atomic bomb. Collins, London, UK, 1967. 416 pp. LCCN QC773.3.U5 G7 1967b.

Groueff:1967:MPUb

[Gro67b] Stéphane Groueff. Manhattan Project; the untold story of the making of the atomic bomb. Little, Brown and Co., Boston, MA, USA, 1967. xii + 372 pp. LCCN QC773.3.U5 G7.

Grodzins:1978:LF

[Gro78] Ruth Grodzins. Laura Fermi, 1907–1977. Bulletin of the Atomic Scientists, 34(5):2–3, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Groves:1983:NIC

[Gro83] Leslie R. Groves. Now it can be told: the story of the Manhattan Project. Da Capo Press, New York, NY, USA and Cambridge, MA, USA, 1983. ISBN 0-306-80189-2, 0-7867-4822-2 (e-book). xiv + 464 + 8 pp. LCCN QC773.A1 G7.

Gurevich:1974:NTE

[GS74] I. I. Gurevich and J. A. Smorodinsky. Nauchnye trudy Enriko Fermi. (Russian) [Proceedings of Enrico Fermi]. Uspekhi Fizicheskikh Nauk, 114(10):385–387, October 1974. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://

ufn.ru/ru/articles/1974/10/k/.

Greiner:1989:NES

[GS89] Walter Greiner and Horst Stöcker, editors. The nuclear equation of state, volume 216A, 216B of NATO ASI series. Series B, Physics. Plenum Press, New York, NY, USA; London, UK, 1989. ISBN 0-306-43486-5 (part A), 0-306-43487-3 (part B). LCCN QC793.3.N8 N38 1989.

Gould:2022:FFF

[GS22] Christopher R. Gould and Eduard I. Sharapov. Fermi’s favorite figure: the history of the pseudopotential concept in atomic physics and neutron physics. European Physical Journal H, 47 (1):??, December 2022. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.springer.

com/article/10.1140/epjh/s13129-022-00042-z.

Hewlett:1962:NW

[HA62] Richard G. Hewlett and Oscar E. Anderson, Jr. The new world: 1939/1946, volume 1 of A History of the United States Atomic Energy Commission. Pennsylvania State University Press, University Park, PA, USA, 1962. xv + 766 pp. LCCN HD9698.U52 H48.

Hewlett:1969:HUS

[HA69] Richard G. Hewlett and Oscar E. Anderson. A history of the United States Atomic Energy Commission: The New World, 1939/1946: Atomic Shield, 1947. Pennsylvania State University Press, University Park, PA, USA, 1969. ???? pp. LCCN HD9698.U52 H48.

Hewlett:1990:NW

[HA90] Richard G. Hewlett and Oscar E. Anderson. The new world, 1939/1946, volume 1 of A history of the United States Atomic Energy Commission. University of California Press, Berkeley, CA, USA, 1990. ISBN 0-520-07186-7. xv + 766 pp. LCCN HD9698.U52 H4 1990 v.1.

Haberer:1969:PCS

[Hab69] Joseph Haberer. Politics and the Community of Science. Van Nostrand Reinhold, New York, NY, USA, 1969. vi + 337 pp. LCCN Q125 .H23 1969.

Hewlett:1990:AS

[HAD90] Richard G. Hewlett, Oscar E. Anderson, and Francis Duncan. Atomic shield, 1947/1952, volume 2 of A history of the United States Atomic Energy Commission. University of California Press, Berkeley, CA, USA, 1990. ISBN 0-520-07187-5. xviii + 718 pp. LCCN HD9698.U52 H55 1989 vol. 2; QC792.7.

Hahn:1958:DF

[Hah58] Otto Hahn. The discovery of fission. Scientific American, 198(??):76–84, February 1958. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

crownedanarchist.com/emc2/discovery_of_fission.doc;
http://www.nature.com/scientificamerican/journal/v198/
n2/pdf/scientificamerican0258-76.pdf.

Hahn:1962:EFU

[Hah62] Otto Hahn. Enrico Fermi and uranium fission. International Atomic Energy Agency Bulletin, 4(0):9–11, ???? 1962. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-

2690 (electronic). URL http://www.iaea.org/Publications/
Magazines/Bulletin/Bull040su/04004700911su.pdf.

Hahn:1975:EGM

[Hah75a] Otto Hahn. Erinnerungen 1901–1945. (German) [Memories 1901– 1945]. In Erlebnisse und Erkenntnisse. (German) [Experiences and insights] [Hah75b], pages 15–73. ISBN 3-430-13732-2. LCCN QD22.H2 A26. DM36.00. With an introduction by Karl-Erik Zimen. Edited by Dietrich Hahn.

Hahn:1975:EEG

[Hah75b] Otto Hahn. Erlebnisse und Erkenntnisse. (German) [Experiences and insights]. Econ-Verlag, Düsseldorf, West Germany, 1975. ISBN 3-430-13732-2. 320 pp. LCCN QD22.H2 A26. DM36.00. With an introduction by Karl-Erik Zimen. Edited by Dietrich Hahn.

Hall:1928:FSP

[Hal28] Edwin H. Hall. The Fermi statistical postulate: Examination of the evidence in its favor. Proceedings of the National Academy of Sciences of the United States of America, 14(5):366–370, May 15, 1928. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490

(electronic). URL http://www.jstor.org/stable/85209.

Halse:1969:FSN

[Hal69] M. R. Halse. The Fermi surfaces of the noble metals. Philosophical Transactions of the Royal Society of London. Series A, Mathematical and Physical Sciences, 265(1167):507–532, December 31,

1969. ISSN 0080-4614. URL http://www.jstor.org/stable/
73758.

Hamilton:2002:LMP

[Ham02] Janet Hamilton. Lise Meitner: pioneer of nuclear fission. Great minds of science. Enslow Publishers, Berkeley Heights, NJ, USA, 2002. ISBN 0-7660-1756-7. 128 pp. LCCN QC774 .M4 H35 2002.

Harrison:1961:FS

[Har61] Walter A. Harrison. The Fermi surface. Science, 134(3483): 915–920, September 29, 1961. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

jstor.org/stable/1708204;   http://www.sciencemag.org/
content/134/3483/915.full.pdf.

Hart:1975:EAE

[Har75] Michael H. Hart. An explanation of the absence of extraterrestrials on Earth. Quarterly Journal of the Royal Astronomical Society, 16(6):128–135, ???? 1975. CODEN QJRAAK. ISSN

0035-8738. URL http://adsabs.harvard.edu/abs/1975QJRAS.
.16..128H.

Hargittai:2002:RSN

[Har02] István Hargittai. The road to Stockholm: Nobel Prizes, science, and scientists. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-850912-X. xvii + 342 + 24 pp. LCCN Q141 .H267 2002. URL http://www.loc.gov/

catdir/enhancements/fy0614/2002283888-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0724/2002283888-b.
html;   http://www.loc.gov/catdir/toc/fy033/2002283888.
html.

Haynes:2016:BRH

[Hay16] John Earl Haynes. Book review: Half-Life: The Divided Life of Bruno Pontecorvo, Physicist or Spy by Frank Close. Journal of Cold War Studies, 18(3):233–236, Summer 2016. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/632304;   http://muse.jhu.
edu/article/632304/pdf.

Hewlett:1989:HUS

[HDHA89] Richard G. Hewlett, Francis Duncan, Jack M. Holl, and Oscar E. Anderson, editors. A history of the United States Atomic Energy Commission. University of California Press, Berkeley, CA, USA, 1989. ISBN 0-520-07187-5 (v. 2), 0-520-06018-0 (v. 3). ???? pp. LCCN HD9698.U52 H55 1989.

Heisenberg:1955:PMV

[Hei55] Werner Heisenberg. The production of mesons in very high energy collisions. Il Nuovo Cimento (10), 2(Supplement 1):96–103, January 1955. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-

6121 (electronic). URL http://link.springer.com/article/

10.1007/BF02746079. Reproduced from the magnetophone-tape after the lecture of the author.

Heisenberg:1983:PMV

[Hei83] Werner Heisenberg. The production of mesons in very high energy collisions. In Anonymous, editor, Celebrazioni del Trentennale della Scuola Internazionale di Fisica ≪Enrico Fermi≫, pages 96– 103. Editrice Compositori e Società Italiana di Fisica, Bologna, Italia, 1983. ISBN ???? LCCN ???? Posthumous republication of lecture given in 1954 at the II Course of the International School of Physics at Villa Monastero, Varenna, Italy. Also available in [Hei55].

Heilbron:2003:ENP

[Hei03] John L. Heilbron. Experimental nuclear physics in the thirties and forties. In Anonymous [Ano03], pages 341–360. ISBN 88-

8286-032-9. LCCN ???? URL http:old.enea.it/produzione_
scientifica/pdf.../V2003_AttiFermi.pdf.

Hermann:1979:NPR

[Her79] Armin Hermann. The new physics: the route into the atomic age: in memory of Albert Einstein, Max von Laue, Otto Hahn, Lise Meitner. Inter Nationes, Bonn-Bad Godesberg, West Germany, 1979. 175 pp. LCCN QC773 .H4813.

Hargittai:2004:CSI

[HH04a] Magdolna Hargittai and István Hargittai, editors. Candid science IV: conversations with famous physicists. Imperial College Press, London, UK, 2004. ISBN 1-86094-414-0, 1-86094-416-7 (paperback). xvi + 711 pp. LCCN QC15 .H295 2004. URL http://

www.worldscibooks.com/physics/p304.html.

Hargittai:2004:EPW

[HH04b] Magdolna Hargittai and István Hargittai. Eugene P. Wigner. In Candid science IV: conversations with famous physicists [HH04a], pages 1–19. ISBN 1-86094-414-0, 1-86094-416-7 (paperback).

LCCN QC15 .H295 2004. URL http://www.worldscibooks.
com/physics/p304.html.

Howes:1999:TDS

[HHW99] Ruth (Ruth Hege) Howes, Caroline L. Herzenberg, and Ellen C. Weaver. Their day in the sun: women of the Manhattan Project. Labor and social change. Temple University Press, Philadelphia, PA, USA, 1999. ISBN 1-56639-719-7 (hardcover), 1-59213-192-1 (paperback), 0-585-38881-4 (e-book). viii + 264 pp. LCCN QC773.3.U5 H68 1999.

Hille:1969:TFE

[Hil69] Einar Hille. On the Thomas–Fermi equation. Proceedings of the National Academy of Sciences of the United States of America, 62(1):7–10, January 15, 1969. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.jstor.

org/stable/58804.

Hooley:2007:HFI

[HM07] Christopher A. Hooley and Andrew P. Mackenzie. Heavy Fermi ions in the original Fermi liquid. Science, 317(5843):1332– 1333, September 7, 2007. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/317/5843/1332.full.pdf.

Haqq-Misra:2009:SSF

[HMB09] Jacob D. Haqq-Misra and Seth D. Baum. The sustainability solution to the Fermi Paradox. Journal of the British Interplanetary Society, 62(??):47–51, 2009. CODEN JBISAW. ISSN 0007-084X.

URL https://arxiv.org/abs/0906.0568.

Hoch:1983:KCE

[Hoc83] P. K. Hoch. A key concept from the electron theory of metals: history of the Fermi surface 1933–60. Contemporary Physics, 24 (1):3–23, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hoffman:1990:CEW

[Hof90] Tony Hoffman. Chatting with extraterrestrials — what are the odds? Chance, 3(2):20–31, Spring 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Holton:1974:SGS

[Hol74] Gerald Holton. Striking gold in science: Fermi’s group and the recapture of Italy’s place in physics. Minerva [Washington, DC], 12(2):159–198, April 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://www.

springerlink.com/content/k7641k32648127h5/.

Holton:1978:SIC

[Hol78] Gerald James Holton. The scientific imagination: case studies. Cambridge University Press, Cambridge, UK, 1978. ISBN 0-521-21700-8, 0-521-29237-9 (paperback). xvi + 382 pp. LCCN Q175 .H775.

Holden:1990:FA

[Hol90a] Constance Holden. Fermi Award. Science, 250(4980):503, October 26, 1990. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/

2878433. The US Department of Energy’s Enrico Fermi Award for 1990 is given to George A. Cowan and Robley D. Evans.

Holmes:1990:BV

[Hol90b] Frederic L. Holmes. Bush, Vannevar. In Holmes [Hol90d], pages 134–138. ISBN 0-684-19177-6. LCCN Q141 .D5 1981.

Holmes:1990:FE

[Hol90c] Frederic L. Holmes. Fermi, Enrico. In Holmes [Hol90d], page ?? ISBN 0-684-19177-6. LCCN Q141 .D5 1981.

Holmes:1990:DSB

[Hol90d] Frederic Lawrence Holmes, editor. Dictionary of scientific biography: Leason Heberling Adams–Fritz H. Laves, volume 17 (supplement 2). Charles Scribner’s Sons, New York, NY, USA, 1990. ISBN 0-684-19177-6. xii + 532 pp. LCCN Q141 .D5 1981.

Holton:1998:SIN

[Hol98] Gerald James Holton. The scientific imagination: with a new introduction. Harvard University Press, Cambridge, MA, USA, 1998. ISBN 0-674-79488-5. xlii + 382 pp. LCCN Q175 .H775 1998.

Holton:2002:MTT

[Hol02] Gerald Holton. The miracle of the two tables. Enrico Fermi, a piece of paraffin and the way towards nuclear fission. Times literary supplement, ??(??):12–13, January 11, 2002.

Holton:2003:BED

[Hol03] Gerald Holton. The birth and early days of the Fermi group in Rome. In Anonymous [Ano03], pages 53–70. ISBN 88-8286-

032-9. LCCN ???? URL http:old.enea.it/produzione_
scientifica/pdf.../V2003_AttiFermi.pdf.

Holton:2007:WEF

[Hol07] Gerald Holton. The world of Enrico Fermi. Film, 2007. URL

http://adsabs.harvard.edu/abs/1970PhTea...8..200H;   http:
//www.aapt.org/aboutaapt/New-Enrico-Fermi-DVD.cfm.

Hook:2003:GBI

[Hoo03] Ernest B. Hook. Gender bias and Ida Noddack. Science, 301 (5636):1045, August 2003. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). See response [Ore03b].

Hopper:1990:SWI

[Hop90] Teri Hopper. She was ignored. Master’s thesis, Stanford University, Stanford, CA, USA, 1990.

Hahn:1939:NEA

[HS39a] Otto Hahn and Fritz Strassmann. Nachweis der Entstehung aktiver Bariumisotope aus Uran und Thorium durch Neutronenbestrahlung; Nachweis weiterer aktiver Bruchstücke bei der Uranspaltung. (German) [Evidence of formation of active barium isotopes of uranium and thorium by neutron irradiation: further evidence of active fragments from the fission of uranium]. Naturwissenschaften, 27(6):89–95, February 1939. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://

www.springerlink.com/content/w406757166152l83/.

Hahn:1939:NVB

[HS39b] Otto Hahn and Fritz Strassmann. Über den Nachweis und das Verhalten der bei der Bestrahlung des Urans mittels Neutronen entstehenden Erdalkalimetalle. (German) [Concerning the existence of alkaline earth metals resulting from the neutron irradiation of uranium]. Naturwissenschaften, 27(1):11–15, January 1939. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). A facsimile is also available in [Bey49, pages 87– 91] and in [Gra64]. Abridged English translation in [GA71, pages 44–47].

Hahn:1989:PFA

[HS89] Otto Hahn and Fritz Strassmann. Proof of the formation of active isotopes of barium from uranium and thorium irradiated with neutrons; proof of the existence of more active fragments produced by uranium fission. Journal of Chemical Education, 66(5): 362–363, May 1989. CODEN JCEDA8. ISSN 0021-9584 (print),

1938-1328 (electronic). URL http://pubs.acs.org/doi/abs/

10.1021/ed066p362. English translation of [HS39b]. Special issue commemorating fify years of nuclear fission. See also [Gra64].

Haber-Schaim:2011:RVS

[HS11] Uri Haber-Schaim. Recollections of the 1954 Varenna School. Il Nuovo Saggiatore, 27(1–2):58–60, January/April 2011. ISSN 0393-4578 (print), 1827-6148 (electronic). URL http://

prometeo.sif.it/papers/online/sag/027/01-02/pdf/09-il-
nostro-mondo2.pdf.

Hayen:2019:BSG

[HS19] L. Hayen and N. Severijns. Beta spectrum generator: High precision allowed β spectrum shapes. Computer Physics Communications, 240(??):152–164, July 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519300645.

Hawkins:1983:MDH

[HTS83] David Hawkins, Edith C. Truslow, and Ralph Carlisle Smith, editors. Manhattan District history, Project Y, the Los Alamos story, volume 2 of History of modern physics, 1800–1950. Tomash Publishers, Los Angeles, CA, USA, 1983. ISBN 0-938228-08-0. xxvi + 506 + 9 pp. LCCN QC773.3.U5 M25 1983.

Huber:2001:PSP

[Hub01] Greg Huber. Postage stamp poses a Fermi problem. Science, 294(5540):53, October 5, 2001. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor. org/stable/3084756. See [Blu01] for earlier errors on the stamp.

Hubisz:2013:MBR

[Hub13] John L. Hubisz. MicroReviews by the book review editor: Elementary Particles: Enrico Fermi. The Physics Teacher, 51(2): 126, February 2013. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

Hughes:2003:NPC

[Hug03] Jeff Hughes. Nuclear physics at the Cavendish Laboratory in the thirties. In Anonymous [Ano03], pages 105–118. ISBN 88-

8286-032-9. LCCN ???? URL http:old.enea.it/produzione_
scientifica/pdf.../V2003_AttiFermi.pdf.

Hart:1982:EWT

[HZ82] Michael H. Hart and Ben Zuckerman, editors. Extraterrestrials — where are they? Pergamon Press, New York, NY, USA, 1982. ISBN 0-08-026342-9, 0-08-026341-0 (paperback). LCCN QB54 .E95 1982.

IAEA:1997:IAE

[IAE97] IAEA. International Atomic Energy Agency: personal reflections. IAEA, Vienna, Austria, 1997. ISBN 92-0-102497-5. xii + 311 pp. LCCN C770 .I49543 1997. URL http://www-pub.iaea.org/

MTCD/publications/PDF/Pub1033_web.pdf.

Ishak:2019:BRSb

[Ish19] B. Ishak. Book review: Solving Fermi’s paradox, by D. H. Forgan, Cambridge, Cambridge University Press, 2019, 426 pp., £120 (hardback), ISBN: 978-1-107-16365-2. Contemporary Physics, 60 (3):269–271, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jackson:1962:CE

[Jac62] John David Jackson. Classical Electrodynamics. Wiley, New York, NY, USA, 1962. xvii + 641 pp. LCCN QC670 .J2.

Jain:2009:BFL

[JAH09] J. K. Jain, P. W. Anderson, and Bertrand I. Halperin. Beyond the Fermi liquid paradigm: Hidden Fermi liquids. Proceedings of the National Academy of Sciences of the United States of America, 106(23):9131–9134, June 9, 2009. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.jstor.

org/stable/40483179.

James:2004:EF

[Jam04a] I. M. (Ioan Mackenzie) James. Enrico Fermi (1901–1954). In Remarkable physicists: from Galileo to Yukawa [Jam04b], chapter 9.5, pages 335–342. ISBN 0-521-81687-4, 0-521-01706-8 (paperback). LCCN QC15 .J36 2004. URL http://www.loc.

gov/catdir/description/cam032/2003055423.html;   http:
//www.loc.gov/catdir/enhancements/fy0732/2003055423-b.
html;   http://www.loc.gov/catdir/samples/cam041/2003055423.
html;   http://www.loc.gov/catdir/toc/cam032/2003055423.
html.

James:2004:RPG

[Jam04b] I. M. (Ioan Mackenzie) James. Remarkable physicists: from Galileo to Yukawa. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-81687-4, 0-521-01706-8 (paperback). xv + 389 pp. LCCN QC15 .J36 2004. URL http://www.loc.

gov/catdir/description/cam032/2003055423.html;   http:
//www.loc.gov/catdir/enhancements/fy0732/2003055423-b.
html;   http://www.loc.gov/catdir/samples/cam041/2003055423.
html;   http://www.loc.gov/catdir/toc/cam032/2003055423.
html.

Jamil:2011:NPN

[Jam11] M. Jamil. Neutron Physics for Nuclear Reactors: Unpublished Writings by Enrico Fermi, edited by S. Esposito and O. Pisanti, Scope: monograph. Level: nuclear physicists, nuclear engineers and researchers. Contemporary Physics, 52(6):618–619, 2011. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jackman:1983:MFH

[JB83] Jarrell C. Jackman and Carla M. Borden, editors. The Muses flee Hitler: cultural transfer and adaptation, 1930–1945. Smithsonian Institution Press, Washington, DC, USA, 1983. ISBN 0-87474-554-3, 0-87474-555-1 (paperback). 347 pp. LCCN DD68 .M85 1983.

Jensen:2000:CCN

[Jen00] Carsten Jensen. Controversy and Consensus: Nuclear Beta Decay 1911–1934, volume 24 of Science networks historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2000. ISBN 3-0348-9569-0 (paperback), 3-7643-5313-9 (hardcover), 3-0348-8444-3 (e-book). xv + 217 pp. LCCN

QC793.5.B425 J46 2000. URL http://www.springerlink.com/

content/978-3-0348-8444-0. Carsten Jensen died of cancer a few months after presenting his doctoral dissertation in 1990 at the University of Copenhagen. Finn Aaserud, Helge Kragh, Erik Rüdinger, and Roger H. Stuewer produced this book as a slightly edited version of that work, supplying additional figures, but leaving the prose largely untouched.

Jacob:2002:LEF

[JM02] Maurice Jacob and Luciano Maiani. L’eredità di Enrico Fermi nella fisica delle particelle. (Italian) [The legacy of Enrico Fermi in particle physics]. In Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901–2001] [Ber02a], pages 242– 273. ISBN 88-7438-000-3. LCCN QC774.F4. 10.33 EUR. URL

http://prometeo.sif.it:8080/libri/fermi/maiani.pdf. En-

glish translation in [BB04b].

Jones:1976:CG

[Jon76] Eric M. Jones. Colonization of the galaxy. Icarus: International Journal of Solar System Studies, 28(3):421–422, July 1976. CO-DEN ICRSA5. ISSN 0019-1035 (print), 1090-2643 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0019103576901561.

Jones:1985:WE

[Jon85a] Eric M. Jones. Where is everybody? Physics Today, 38(8): 11, 13, August 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://scitation.aip.

org/content/aip/magazine/physicstoday/article/38/8/10.
1063/1.2814654.

Jones:1985:WEA

[Jon85b] Eric M. Jones. “Where is everybody?” an account of Fermi’s question. Report LA-10311-MS / DE85 011898, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, March 1985. 20 pp. URL

http://www.fas.org/sgp/othergov/doe/lanl/la-10311-ms.
pdf;   http://www.osti.gov/accomplishments/documents/
fullText/ACC0055.pdf.

Jones:2010:THR

[Jon10] Derry W. Jones. Titanic Hungarian refugee physicists. Contemporary Physics, 51(3):267–271, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kaempffert:1935:YSP

[Kae35] Waldemar Kaempffert. The year in science: Progress is achieved in many fields: Advances in chemistry, physics, biology and astronomy have made 1935 notable — the need is emphasized for a science of sciences to correlate discoveries. New York Times, ??(??):XX4, December 29, 1935. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL https://search.

proquest.com/hnpnewyorktimes/docview/101272291/.

Kaempffert:1939:WSW

[Kae39] Waldemar Kaempffert. This week in science: When uranium splits: Doubtful source of power; cancer and X-rays; neutron possibilities. New York Times, ??(??):D9, March 5, 1939. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/102937542.

Kaempffert:1945:SSB

[Kae45] Waldemar Kaempffert. Story of scientists’ ‘battle’ for atom bomb secret revealed in Smyth Report. New York Times, ??(??):8, August 16, 1945. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/107074385/.

Kaempffert:1948:RRB

[Kae48] Waldemar Kaempffert. The revolution that radium began: Fifty years after the Curies’ great discovery, nuclear physics is still a realm unbounded. New York Times, ??(??):SM13, SM25, SM27, December 26, 1948. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/108348269/.

Kaiser:2004:BRP

[Kai04] David Kaiser. Book review: “A Physicist with a Capital F”: Fermi Remembered, by James Cronin. American Scientist, 92 (6):569–570, ???? 2004. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://www.jstor.org/

stable/27858491. See [Cro04].

Kubasiak:2005:FDS

[KaKKZL05] Anna Kubasiak, Jaros aw K. Korbicz, Jakub Zakrzewski, and Maciej Lewenstein. Fermi–Dirac statistics and the number theory. Europhysics Letters, 72(??):4, 506–?? 2005. CODEN EULEEJ. ISSN 0295-5075 (print), 1286-4854 (electronic). URL http://

iopscience.iop.org/0295-5075/72/4/506/.

Kapur:1940:DMO

[Kap40] P. L. Kapur. Does the mesotron obey Bose–Einstein or Fermi– Dirac statistics? Nature, 145(3663):69, January 13, 1940. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v145/
n3663/pdf/145069a0.pdf.

Kathren:1975:LP

[Kat75] Ronald L. Kathren. Letters: Also present. Bulletin of the Atomic Scientists, 31(1):3, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). The author notes that Carl C. Gamertsfelder was omitted from the list of participants at the first test of the Chicago pile (CP-1) on 2 December 1942 given in [And74b, And74a].

Katz:2021:FT

[Kat21] J. I. Katz. Fermi at Trinity. Nuclear Technology, 207(S1):S326– S334, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Kuiper:1989:ECc

[KB89] Thomas B. H. (Thomas Bernardus Henricus) Kuiper and Glen David Brin, editors. Extraterrestrial civilization. American Association of Physics Teachers, College Park, MD, USA, 1989. ISBN 0-917853-38-5. 121 pp. LCCN QB54 .E945; QB54 .E97 1989.

Kean:2010:DSO

[Kea10] Sam Kean. The disappearing spoon: and other true tales of madness, love, and the history of the world from the periodic table of the elements. Little, Brown and Co., Boston, MA, USA, 2010. ISBN 0-316-05164-0. vi + 391 pp. LCCN QD466 .K37 2010.

Kenworthy:1954:DHB

[Ken54] E. W. Kenworthy. The drama of the hydrogen bomb — and Dr. Oppenheimer’s key role: Security case focuses attention on disputes that preceded first successful test of H-bomb at Pacific Proving Ground. New York Times, ??(??):E5, April 18, 1954. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL https://search.proquest.com/hnpnewyorktimes/
docview/113104238/.

Kevles:1977:PHS

[Kev77] Daniel J. Kevles. The Physicists: the History of a Scientific Community in Modern America. Alfred A. Knopf, New York, NY, USA, 1977. ISBN 0-394-46631-4. xi + 489 pp. LCCN QC9.U5 K48 1977.

Kevles:1987:PHS

[Kev87] Daniel J. Kevles. The physicists: the history of a scientific community in modern America. Harvard University Press, Cambridge, MA, USA, 1987. ISBN 0-674-66655-0 (paperback). xv + 489 pp. LCCN QC9.U5 K48 1987. URL http://catalog.

hathitrust.org/api/volumes/oclc/15860449.html.

Kevles:1995:PHS

[Kev95] Daniel J. Kevles. The physicists: the history of a scientific community in modern America. Harvard University Press, Cambridge, MA, USA, 1995. ISBN 0-674-66655-0 (paperback). xlv + 489 pp. LCCN QC9.U5 K48 1995.

Kruger:1955:VMP

[KGSO55] H. Krüger, H. Gartmann, H. Schardin, and W. Oppelt. Verleihung der Max-Planck-Medaille an Enrico Fermi in Heidelberg/Der 5. Internationale Astronautische Kongreß in Innsbruck/2. Internationales Symposium über Kurzzeitphotographie und Hochfrequenzkinematographie in Paris/FA Regelungsmathematik über ,,Anwendung der Laplace-Transformation” in Essen. (German) [Awarding the Max Planck Medal to Enrico Fermi in Heidelberg /The 5th International Astronautical Congress in Innsbruck /2nd International Symposium on short-time photography and highfrequency cinematography in Paris /FA mathematical control application on “Use of Laplace transform” in Essen]. Physikalische Blätter, 11(1):35–39, 1955. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Klintenberg:2012:FSD

[KHB12] M. Klintenberg, J. T. Haraldsen, and A. V. Balatsky. 3D Fermi surface database. CINT: The Center for Integrated Technologies, Los Alamos, NM, USA, June 11, 2012. URL http://www.lanl.

gov/cint/viz3.html.

Kuhn:1967:SHQ

[KHFA67] Thomas S. Kuhn, John L. Heilbron, Paul Forman, and Lini Allen. Sources for history of quantum physics: an inventory and report, volume 68 of Memoirs of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1967. ix + 176 pp. LCCN QC174.1 .S66.

URL http://www.amphilsoc.org/guides/ahqp/;   http://
www.amphilsoc.org/guides/ahqp/s-t.htm#schrodinger.

Kiernan:2013:GAC

[Kie13] Denise Kiernan. The girls of Atomic City: the untold story of the women who helped win World War II. Simon and Schuster, New York, NY, USA, 2013. ISBN 1-4516-1752-6 (hardcover), 1-4516-1754-2 (e-book). xvii + 373 + 16 pp. LCCN F444.O3 K54 2013.

Klein:1971:EFP

[Kle71] Martin J. Klein. Enrico Fermi, Physicist by Emilio Segrè (review). Technology and Culture, 12(3):507–508, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic).  URL http://www.jstor.org/stable/3103011;
https://muse.jhu.edu/pub/1/article/893607/pdf.

Kone:1950:EF

[Kon50] Eugene H. Kone. Enrico Fermi. American Scientist, 38(3):442– 445, ???? 1950. CODEN AMSCAC. ISSN 0003-0996 (print),

1545-2786 (electronic). URL http://www.jstor.org/stable/
27826326.

Konopinski:1955:FTB

[Kon55] E. J. Konopinski. Fermi’s theory of beta-decay. Reviews of Modern Physics, 27(3):254–257, July 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.27.254;   http:/
/rmp.aps.org/abstract/RMP/v27/i3/p254_1.

Kopff:1923:FDR

[Kop23] Augusto Kopff. I fondamenti della relatività Einsteiniana: valore e interpretazione della teoria. (Italian) [The foundations of Einsteinian relativity: value and interpretation of the theory]. Ulrico Hoepli, Milano, Italy, 1923. xxx + 455 pp. LCCN ???? Italian translation of German original Grundzüge der Einsteinschen Relativitätstheorie (1921 and 1923). Edited by R. Contu and T. Bembo.

Kothari:1954:FTT

[Kot54] D. S. Kothari. Fermi’s thermodynamic theory of the production of pions. Nature, 173(4404):590, March 27, 1954. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v173/
n4404/pdf/173590a0.pdf.

Kowarski:1953:HAN

[Kow53] Lew Kowarski. Hitting the atomic nucleus. UNESCO Courier, 6(12):3–4, December 1953. ISSN 0041-5278. URL http://

unesdoc.unesco.org/images/0007/000708/070862eo.pdf.

Kowald:2015:WTN

[Kow15] Axel Kowald. Why is there no von Neumann probe on Ceres? Error catastrophe can explain the Fermi–Hart paradox. Journal of the British Interplanetary Society, 68(??):383–338, ???? 2015. CODEN JBISAW. ISSN 0007-084X. URL https://arxiv.org/

abs/1605.02169.

Krafft:1978:LMH

[Kra78] Fritz Krafft. Lise Meitner: Her life and times — on the centenary of the great scientist’s birth. Angewandte Chemie, International Edition in English, 17(11):826–842, ???? 1978. CODEN ACIEAY. ISSN 0570-0833.

Krafft:1988:LMB

[Kra88] Fritz Krafft. Lise Meitner: eine Biographie, volume 448 of HMI-B, 0175-8349. Hans-Meitner-Institut, Berlin, Germany, 1988. 24 pp. LCCN MLCS 88/06254 (Q).

Kragh:1992:RCW

[Kra92] Helge Kragh. Relativistic collisions: The work of Christian Møller in the early 1930s. Archive for History of Exact Sciences, 43(4):299–328, December 1992. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0003-
9519&volume=43&issue=4&spage=299.

Kragh:1999:QGH

[Kra99] Helge Kragh. Quantum generations: a history of physics in the twentieth century. Princeton University Press, Princeton, NJ, USA, 1999. ISBN 0-691-01206-7 (hardcover), 0-691-09552-3 (paperback). xiv + 494 pp. LCCN QC7 .K7 2002. URL http://

press.princeton.edu/titles/6683.html.

Kramer:2015:NPW

[Kra15] David Kramer. New park will honor US atomic heritage. Physics Today, 68(3):24–25, March 2015. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Kragh:2018:TSE

[Kra18] Helge Kragh. From Transuranic to Superheavy Elements: a Story of Dispute and Creation. SpringerBriefs in history of science and technology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2018. ISBN 3-319-75812-8, 3-319-75813-6 (electronic). viii + 106 pp. LCCN QD172.T7 K73

2018. URL http://link.springer.com/10.1007/978-3-319-
75813-8.

Kreifeldt:1971:FNC

[Kre71] J. C. Kreifeldt. A formulation for the number of communicative civilizations in the Galaxy. Icarus: International Journal of Solar System Studies, 14(3):419–430, June 1971. CO-DEN ICRSA5. ISSN 0019-1035 (print), 1090-2643 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
001910357190011X.

Kothari:1942:TRF

[KS42] D. S. Kothari and B. N. Singh. Thermodynamics of a relativistic Fermi–Dirac gas. Proceedings of the Royal Society of London. Series A, Mathematical and Physical Sciences, 180(983):414–423, July 3, 1942. CODEN PRLAAZ. ISSN 0080-4630. URL http:/

/www.jstor.org/stable/97729.

Kaneyuki:1999:NOA

[KS99] Kenji Kaneyuki and Kate Scholberg. Neutrino oscillations: Always elusive, Fermi’s ’little neutral one’ turns out to be a quickchange artist as well, offering answers and new questions for physics and cosmology. American Scientist, 87(3):222–231, ???? 1999. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786

(electronic). URL http://www.jstor.org/stable/27857842.

Kennedy:1946:P

[KSSW46] J. W. Kennedy, Glenn T. Seaborg, Emilio Segrè, and Arthur C. Wahl. Properties of 94(239). Physical Review, 70(7–8):555–556, October 1946. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/

v70/i7-8/p555_1. Received May 29, 1941, but withheld from publication until the end of World War II.

Kubbinga:2009:TEF

[Kub09] Henk Kubbinga. A tribute to Enrico Fermi. Europhysics News, 40(6):27–29, November/December 2009. CODEN EU-PNAS. ISSN 0531-7479 (print), 1432-1092 (electronic). URL

http://www.europhysicsnews.org/articles/epn/abs/2009/
06/epn20096p27/epn20096p27.html.

Kurian:2002:NSB

[Kur02] George Thomas Kurian. The Nobel scientists: a biographical encyclopedia. Prometheus Books, Amherst, NY, USA, 2002. ISBN 1-57392-927-1. 420 pp. LCCN Q141 .K78 2002.

LAnnunziata:2007:RIH

[L’A07] Michael F. L’Annunziata. Radioactivity: introduction and history. Elsevier, Amsterdam, The Netherlands, 2007. ISBN 0-444-52715-X (hardcover). xxi + 609 pp. LCCN QC795

.L264 2007. URL http://www.loc.gov/catdir/toc/fy0805/

2008925731.html. Foreword by Werner Burkart (Deputy Director General, International Atomic Energy Agency, Vienna).

Lafferty:1971:NPE

[Laf71] P. E. Lafferty. Names in physics: Enrico Fermi 1901–54. Physics Education, 6(4):216–217, 1971. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic). URL http://stacks.iop.

org/0031-9120/6/i=4/a=305.

Lamont:1965:DT

[Lam65] Lansing Lamont. Day of Trinity. Atheneum, New York, NY, USA, 1965. xi + 333 pp. LCCN QC773.A1 L3.

Lamont:1985:DT

[Lam85] Lansing Lamont. Day of Trinity. Atheneum, New York, NY, USA, 1985. ISBN 0-689-70686-3 (paperback). 363 + 10 pp. LCCN QC773.A1 L3 1985.

Lanouette:1992:ISP

[Lan92] William Lanouette. Ideas by Szilard, physics by Fermi. Bulletin of the Atomic Scientists, 48(10):16–23, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special edition for the 50th Anniversary of the Chain Reaction (2 December 1942). An article about Leo Szilard’s role as co-designer of the world’s first nuclear reactor, and his invention of atomic secrecy.

Lanouette:1994:AS

[Lan94] William Lanouette. Atomic spies. Debate with the authors of Special Tasks by Soviet spymaster Pavel Sudoplatov [SSSS95], which falsely alleged that Szilard, Niels Bohr, J. Robert Oppenheimer and Enrico Fermi were Soviet agents within the Manhattan Project. MacNeil/Lehrer NewsHour, April 26, 1994.

Lanouette:2001:FSE

[Lan01] William Lanouette. Fermi, Szilard und der erste Atomreaktor. (German) [Fermi, Szilard and the first atomic reactor]. Spektrum der Wissenschaft (German translation of Scientific American), ?? (??):78–83, January 1, 2001. CODEN SPEKDI. ISSN 0170-2971.

Landis:2012:LBF

[Lan12] Geoffrey A. Landis. Last byte: Future tense: Fermi’s Paradox and the end of the Universe. Communications of the ACM, 55 (10):112, October 2012. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lanouette:2017:LMM

[Lan17] William Lanouette. The little match’s momentous legacy: The Pope of Physics: Enrico Fermi and the Birth of the Atomic Age by Gino Segrè and Bettina Hoerlin. New York: Henry Holt and Company, 2016, 351 pp. Issues in Science and Technology, 33(3):93–94, Spring 2017. CODEN ???? ISSN 0748-5492 (print), 1938-1557

(electronic). URL http://issues.org/33-3/book-review-the-
little-matchs-momentous-legacy/.

Lancaster:2018:BRL

[Lan18] Tom Lancaster. Book review: The last man who knew everything: the life and times of Enrico Fermi, father of the nuclear age, by D. N. Schwartz. Contemporary Physics, 59(2):219–220, 2018. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Larkoski:2019:EPP

[Lar19] Andrew J. Larkoski. Elementary Particle Physics: an Intuitive Introduction. Cambridge University Press, Cambridge, UK, 2019. ISBN 1-108-49698-9 (hardcover). 540 (est.) pp. LCCN QC793.2 .L37 2019.

Laurence:1946:BAA

[Lau46a] William L. Laurence. The birth of the atomic age (December 2, 1942). The New York Times Magazine, VI(??):11–??, December 1, 1946. ISSN 0362-1308.

Laurence:1946:DBA

[Lau46b] William L. Laurence. Dec. 2, 1942 — the birth of the Atomic Age: Story of the great experiment which first released the energy that runs the universe. New York Times, ??(??):SM6, SM60, December 1, 1946. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/107694608/.

Laurence:1951:DHR

[Lau51] William L. Laurence. Day of Hiroshima recalls atom race: Two leaders in atomic work at Columbia University. New York Times, ??(??):3, August 6, 1951. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/112188459/fulltextPDF.

Laucht:2013:BRP

[Lau13a] Christoph Laucht. Book review: The Pontecorvo Affair: A Cold War Defection and Nuclear Physics, Simone Turchetti, Chicago, IL, The University of Chicago Press, 2012. 292 pp. ISBN 0-226-81664-8 (hbk) (£31.50). Contemporary British History, 27(4): 539–541, December 2013. ISSN 1361-9462 (print), 1743-7997 (electronic).

Laurence:2013:FMS

[Lau13b] William L. Laurence. Fermi measures speed of neutron / Bohr and Einstein at odds. In Dean [Dea13], page ?? ISBN 1-4027-9320-0 (hardcover). LCCN QC7 .D43 2013. Foreword by Neil deGrasse Tyson.

Lewis:2017:FUL

[LBS17] Geraint F. Lewis, Luke A. Barnes, and Brian Schmidt. A fortunate universe: life in a finely-tuned cosmos. Cambridge University Press, Cambridge, UK, 2017. ISBN 1-107-15661-0 (hardcover), 1-316-71648-1 (e-book), 1-316-71634-1, 1-316-71606-6, 1-316-66141-5. xvii + 373 pp. LCCN Q172.5.C45 L4845 2017.

Levi-Civita:2002:SIA

[LC02] Tullio Levi-Civita. Sugli invarianti adiabatici. (Italian) [on adiabatic invariants]. In Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901–2001] [Ber02a], pages 85– 113. ISBN 88-7438-000-3. LCCN QC774.F4. 10.33 EUR. URL

http://prometeo.sif.it:8080/libri/fermi/06.pdf. English

translation in [BB04b].

Leverett:1942:DHC

[LCM^{+}42] M. C. Leverett, C. M. Cooper, T. V. Moore, E. P. Wigner, E. S. Steinbach, E. Fermi, J. A. Wheeler, S. K. Allison, and Leo Szilard. Discussion of helium cooled power plant. Report CS-267, US Atomic Energy Commission, Washington, DC, USA, September 16, 1942.

Lederman:2003:BPM

[Led03] Leon Lederman. The beginnings of pion and muon physics. In Anonymous [Ano03], pages 361–364. ISBN 88-8286-032-9. LCCN

???? URL http:old.enea.it/produzione_scientifica/pdf.
../V2003_AttiFermi.pdf.

Lederman:2004:FCC

[Led04] Leon Lederman. Fermi Centennial comments. In Orear [Ore04], chapter 32, pages 159–160. ISBN ???? LCCN ????

URL http://dspace.library.cornell.edu/handle/1813/62;
 http://hdl.handle.net/1813/74.

Lee:2001:RMM

[Lee01] Tsung Dao Lee. Ricordi del mio maestro. Il Nuovo Saggiatore, 17(5–6):45–47, September/December 2, 2001. ISSN 0393-4578 (print), 1827-6148 (electronic). URL http://prometeo.sif.it:

8080/papers/online/sag/2001/05-06/pdf/05.pdf. Special is-

sue in honor of the centennial of the birth of Enrico Fermi.

Leffler:1995:TDD

[Lef95] Melvyn P. Leffler. Truman’s decision to drop the atomic bomb. IHJ Bulletin: A Quarterly Publication of the International House of Japan, 15(3):1–7, Summer 1995. ISSN 0285-2608.

Lemarchand:1998:TIL

[Lem98] Guillermo A. Lemarchand. Is there intelligent life out there? Scientific American Presents, ??(??):96–104, Winter 1998. URL

http://www.iar.unlp.edu.ar/SETI/GAL-Scient-Ame.pdf.

Lewis:2021:TNC

[Lew21] Nicholas Lewis. Trinity by the numbers: The computing effort that made Trinity possible. Nuclear Technology, 207(S1): S176–S189, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Leverett:1962:NR

[LF62] Miles C. Leverett and Enrico Fermi. Nuclear reactor. Canadian Patent 637992., March 13, 1962. URL http://brevets-

patents.ic.gc.ca/opic-cipo/cpd/eng/patent/637992/summary.
html.

Lundby:1952:MSN

[LFA^{+}52] Arné Lundby, Enrico Fermi, Herbert L. Anderson, Darragh E. Nagle, and G. Yodh. M8. scattering of negative pions by hydrogen [abstract only]. Physical Review, 86(4):603, May 1952. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.86.579.2. Min-

utes of the 1952 Annual Meeting Held at New York City, January 31, February 1–2, 1952.

Libby:1979:UP

[Lib79] Leona Marshall Libby. The Uranium People. Crane Russak, New York, NY, USA, 1979. ISBN 0-8448-1300-1 (Crane Russak), 0-684-16242-3 (Scribners). x + 341 + 16 pp. LCCN QC773.3.U5 L52.

Liboff:1985:GPF

[Lib85] Richard L. Liboff. Geometrical properties of the Fermi energy. Foundations of Physics, 15(3):339–352, March 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00737322.

Lieb:1981:TFR

[Lie81] Elliott H. Lieb. Thomas–Fermi and related theories of atoms and molecules. Reviews of Modern Physics, 53(4):603–641, October 1981. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.53.603;   http://rmp.aps.org/abstract/

RMP/v53/i4/p603_1. See erratum [Lie82].

Lieb:1982:ETF

[Lie82] Elliott H. Lieb. Erratum: Thomas–Fermi and related theories of atoms and molecules. Reviews of Modern Physics, 54(1):311, January 1982. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.54.311;   http://
rmp.aps.org/abstract/RMP/v54/i1/p311_1. See [Lie81].

Lundqvist:1983:TIE

[LM83] Stig Lundqvist and Norman H. (Norman Henry) March, editors. Theory of the Inhomogeneous Electron Gas. Physics of Solids and Liquids. Plenum Press, New York, NY, USA; London, UK, 1983. ISBN 0-306-41207-1, 1-4899-0417-4, 1-4899-0415-8. xiii + 395 pp. LCCN QC175.16.E6 T46 1983. URL http://www.

springerlink.com/content/978-1-4899-0415-7.

Leone:2006:BRE

[LMR06] Matteo Leone, Angelo Mastroianni, and Nadia Robotti. Bruno Rossi and the entrance of the Geiger–Müller counter in the Italian physics: 1929–1934. Physis: Rivista Internazionale di Storia della Scienza. Nuova Serie, 43(??):445–472, 2006. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Lombardi:1993:CRN

[Lom93] Giovanni Lombardi. Città raccontate: i nuovi ragazzi di via Panisperna. (Italian) [The story of a city: the new Via Panisperna boys]. International Maxpress, Roma, Italy, 1993. ISBN ???? 69 pp. LCCN ????

Lombardi:2001:ENF

[Lom01] C. Lombardi. Energia nucleare da fissione. (Italian) [Nuclear energy from fission]. Il Nuovo Saggiatore, 17(5–6):66–69, September/December 2001. ISSN 0393-4578 (print), 1827-6148 (elec-

tronic). URL http://prometeo.sif.it:8080/papers/online/

sag/2001/05-06/pdf/05.pdf. Special issue in honor of the centennial of the birth of Enrico Fermi.

Loudon:1983:QTL

[Lou83] Rodney Loudon. The quantum theory of light. Oxford science publications. Clarendon Press, Oxford, UK, second edition, 1983. ISBN 0-19-851155-8 (paperback), 0-19-851152-3 (hardcover). xiv + 393 pp. LCCN QC446.2 .L68 1983.

Leone:2000:FAD

[LRS00] M. Leone, N. Robotti, and C. A. Segnini. Fermi Archives at the Domus Galilaeana in Pisa. Physis: Rivista Internazionale di Storia della Scienza. Nuova Serie, 37(??):501–533, 2000. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic). URL

http://www.olschki.it/riviste/physis.htm.

Lieb:1977:TFT

[LS77] Elliott H. Lieb and Barry Simon. The Thomas–Fermi theory of atoms, molecules and solids. Advances in Mathematics, 23(1):22– 116, January 1977. CODEN ADMTA4. ISSN 0001-8708 (print),

1090-2082 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0001870877901086.

Lanouette:1992:GSB

[LS92] William Lanouette and Bela A. Silard. Genius in the shadows: a biography of Leo Szilard: the man behind the bomb. C. Scribner’s Sons, New York, NY, USA, 1992. ISBN 0-684-19011-7. xix + 587 + 16 pp. LCCN QC16.S95 L36 1992. US$35.00 (US$44.50 Can.).

URL http://adsabs.harvard.edu/abs/1994gsbl.book.....L.

Lieb:2005:OER

[LT05] Elliott H. Lieb and Walter Thirring, editors. One-Electron Relativistic Molecules with Coulomb Interaction. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 2005. ISBN 3-540-22212-X (hardcover), 3-540-27056-6. xv + 932 pp. LCCN QC173.4.T48 L54 2005. URL

http://link.springer.com/10.1007/b138553.

Luttinger:1966:BRF

[Lut66] J. M. Luttinger. Book review: Fermi’s Molecole e Cristalli. Science, 152(3730):1735–1736, June 24, 1966. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.jstor.org/stable/1718368;   http://www.sciencemag.
org/content/152/3730/1735.3.full.pdf.

Lewis:1971:APT

[LW71] Richard S. Lewis and Jane Wilson, editors. Alamogordo plus twenty-five years; the impact of atomic energy on science, technology, and world politics. Viking Press, New York, NY, USA, 1971. ISBN 0-670-11151-1. vi + 281 pp. LCCN QC792 .A43 1971. With Eugene Rabinowitch.

McMillan:1940:RE

[MA40] Edwin McMillan and Philip Hauge Abelson. Radioactive element 93. Physical Review, 57(12):1185–1186, June 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v57/i12/p1185_2. This

paper, on the discovery of neptunium, was the last paper on transuranic elements and fission in the Physical Review until the after the end of World War II in 1945; see [Wal04].

MacPherson:1986:TBF

[Mac86] Malcolm MacPherson. Time bomb: Fermi, Heisenberg, and the race for the atomic bomb. Dutton, New York, NY, USA, 1986. ISBN 0-525-24409-3. xiii + 316 + 8 pp. LCCN QC773 .M24 1986. US$18.95.

MacPherson:1987:TBF

[Mac87] Malcolm MacPherson. Time bomb: Fermi, Heisenberg, and the race for the atomic bomb. Berkley Books, New York, NY, USA, 1987. ISBN 0-425-10423-0. xi + 284 + 8 pp. LCCN QC773 .M24 1987.

Maddox:1987:RHF

[Mad87] John Maddox. Relics of Hutchins and Fermi linger on. Nature, 326 (6114):650, April 22, 1987. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v326/n6114/pdf/326650a0.pdf.

Mafai:1992:LFS

[Maf92] Miriam Mafai. Il lungo freddo: storia di Bruno Pontecorvo, lo scienziato che scelse l’URSS. (Italian) [The long cold: story of Bruno Pontecorvo, the scientist who chose the USSR]. Le scie. A. Mondadori, Milano, Italia, 1992. ISBN 88-04-33922-5. 318 pp. LCCN QC774.P66 M34 1992.

Magnuson:1961:BRF

[Mag61] G. D. Magnuson. Book review: The Fermi Surface by W. A. Harrison; M. B. Webb. American Scientist, 49(3):332A–332A, September 1961. CODEN AMSCAC. ISSN 0003-0996 (print),

1545-2786 (electronic). URL http://www.jstor.org/stable/
27827893.

Magueijo:2009:BDE

[Mag09] João Magueijo. A brilliant darkness: the extraordinary life and disappearance of Ettore Majorana, the troubled genius of the nuclear age. Basic Books, New York, NY, USA, 2009. ISBN 0-465-00903-4. xxi + 280 pp. LCCN QC774.M34 M35 2009.

Maiocchi:1991:NSF

[Mai91] Roberto Maiocchi. Non solo Fermi: i fondamenti della meccanica quantica quantistica nella cultura italiana tra le due guerre. (Italian) [Not only Fermi: the fundamentals of quantum mechanics in Italian culture between the two Wars], volume 2 of Giornale critico della filosofia italiana / Quaderni. Le Lettere, Firenze, Italia, 1991. ISBN 88-7166-050-1. 261 pp. LCCN ????

Maiani:2003:PHE

[Mai03] Luciano Maiani. Perspectives in high energy particle physics. In Anonymous [Ano03], pages 365–388. ISBN 88-8286-032-9. LCCN

???? URL http:old.enea.it/produzione_scientifica/pdf.
../V2003_AttiFermi.pdf.

Majorana:1928:TQN

[Maj28] Ettore Majorana. La teoria quantistica dei nuclei radioattivi. (Italian) [The quantum theory of radioactive nuclei]. Master’s thesis., Università di Roma “La Sapienza”, Roma, Italia, July 6, 1928.

Majorana:1937:TSD

[Maj37] Ettore Majorana. Teoria simmetrica dell elettrone e del positrone. (Italian) [Symmetrical theory of the electron and the positron]. Il Nuovo Cimento (8), 14(4):171–184, April 1937. CODEN NU-CIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL

http://en.wikipedia.org/wiki/Majorana_mass;   http://
nl.wikipedia.org/wiki/Majorana-deeltje. In this paper, Ma-

jorana predicted the existence of a new type of particle, now called a Majorana fermion, which is its own antiparticle, and whose existence may have finally been confirmed by experiment seventy years later [FK08, MZF^{+}12, Ser12]. See also comments in [ACM10]. Esposito [Esp08a] reports about this paper “With amazing farsightedness Majorana suggested that the neutrino, which had just been postulated by Wolfgang Pauli and Fermi to explain puzzling features of radioactive beta decay, could be such a particle. This would make the neutrino unique among the elementary particles and, moreover, enable it to have mass. Today many experiments are still devoted to detect these peculiar properties, which include the phenomenon of neutrino oscillations: we have not yet succeeded to find a definite answer to Majorana’s proposal.”.

Maltese:2003:EFB

[Mal03a] Giulio Maltese. Enrico Fermi and the birth of high-energy physics after World War II. In Anonymous [Ano03], pages 221–258. ISBN 88-8286-032-9. LCCN ???? URL http:old.enea.it/

produzione_scientifica/pdf.../V2003_AttiFermi.pdf.

Maltese:2003:EFA

[Mal03b] Giulio Maltese. Enrico Fermi in America: una biografia scientifica: 1938–1954. (German) [Enrico Fermi in America: a scientific biography 1938–1954]. Zanichelli, Bologna, Italy, 2003. ISBN 88-08-07727-6 (paperback). xxvi + 510 + 20 pp. LCCN ????

Maltese:2005:BRF

[Mal05] Giulio Maltese. Book reviews: Fermi’s legacy: Fermi Remembered, edited by James W. Cronin and Enrico Fermi: His Work and Legacy, edited by Carlo Bernardini and Luisa Bonolis. Nature, 433(7025):460–461, February 2, 2005. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Maltese:2008:BRS

[Mal08] Giulio Maltese. Book review: Simone Turchetti, Il caso Pontecorvo. Fisica nucleare, politica e servizi di sicurezza nella guerra fredda. Sironi: Milano, 2007. 285 pp., ISBN 978-88-518-0081-9. Nuncius, 23(1):183–184, ???? 2008. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00508.

Maltese:2010:PIE

[Mal10] Giulio Maltese. Il Papa e l’Inquisitore: Enrico Fermi, Ettore Majorana, via Panisperna. (Italian) [The Pope and the Inquisitor: Enrico Fermi, Ettore Majorana, via Panisperna]. Zanichelli, Bologna, 2010. ISBN 88-08-16814-X. 398 pp. LCCN QC774.F4.

Manley:1957:BRL

[Man57] John H. Manley. Book review: Laura Fermi, Atoms for the World: US Participation in the Conference on the Peaceful Uses of Atomic Energy. Physics Today, 10(7):28, July 1957. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://www.physicstoday.org/resource/1/
phtoad/v10/i7/p28_s1.

Martin:1959:GCD

[Mar59] J. L. Martin. Generalized classical dynamics, and the ’classical analogue’ of a Fermi oscillator. Proceedings of the Royal Society of London. Series A, Mathematical and Physical Sciences, 251 (1267):536–542, June 23, 1959. CODEN PRLAAZ. ISSN 0080-

4630. URL http://www.jstor.org/stable/100729.

March:1975:SCF

[Mar75] Norman H. (Norman Henry) March. Self-consistent fields in atoms: Hartree and Thomas–Fermi atoms. Commonwealth and international library. Selected readings in physics. Pergamon, New York, NY, USA, 1975. ISBN 0-08-017819-7, 0-08-017820-0 (fleXicover). ix + 233 pp. LCCN QC173 .M3633 1975. See pages 205–213 for an English translation of [Fer28b].

March:1983:OTF

[Mar83] N. H. March. Origins — the Thomas–Fermi theory. In Lundqvist and March [LM83], chapter 1, page ?? ISBN 0-306-41207-1, 1-4899-0417-4, 1-4899-0415-8. LCCN QC175.16.E6

T46 1983. URL http://link.springer.com/chapter/10.1007/
978-1-4899-0415-7_1/.

Marx:1996:MMG

[Mar96] George Marx. The myth of the Martians and the golden age of Hungarian science. Science & Education (Springer), 5(3):225– 234, July 1996. CODEN SCEDE9. ISSN 0926-7220 (print),

1573-1901 (electronic). URL http://adsabs.harvard.edu/abs/
1996Sc%26Ed...5..225M.

Maric:2000:ETF

[Mar00] Vojislav Marić. Equations of Thomas–Fermi type. Lecture Notes in Mathematics, 1726:71–104, 2000. CODEN LNMAA2. ISBN 3-540-67160-9 (print), 3-540-46520-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103956/.

Michaudon:2000:FMW

[MB00] André F. Michaudon and Ileana G. Buican. A factor of millions: Why we made plutonium. Los Alamos Science, ??(26):4–9, November 2000. CODEN LASCDI. ISSN 0273-7116. URL http:

//www.fas.org/sgp/othergov/doe/lanl/pubs/00818004.pdf.

McMillan:1951:TEE

[McM51] Edwin M. McMillan. The transuranium elements: Early history. Report UCRL-1619, Radiation Laboratory, University of California, Berkeley, Berkeley, CA, USA, Decem-

ber 1951. URL http://www.osti.gov/accomplishments/
documents/fullText/ACC0361.pdf. Nobel Lecture given at

Stockholm on December 12, 1951.

McMillan:1994:SFF

[McM94] Priscilla Johnson McMillan. The Sudoplatov file: Flimsy memories. Bulletin of the Atomic Scientists, 50(4):30–33, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.highbeam.com/doc/1G1-15587910.
html.

McDayter:1967:GBB

[MD67] Walt McDayter and Norman Drew. The giants: The bomb builders. Denver Post, ??(??):??, February 3, 1967. URL http:

//library.ucsd.edu/dc/object/bb0103915g. This is a reason-

ably accurate 83-frame comic strip on the history of the building of the atomic bomb, with Leo Szilard as the central figure of the story.

Mehra:1973:PCN

[Meh73] Jagdish Mehra, editor. The physicist’s conception of nature: Symposium on the Development of the Physicist’s Conception of Nature in the 20th century. Held at the International Centre for Theoretical Physics, Miramare, Trieste, Italy, 18–25 September 1972. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1973. ISBN 90-277-0345-0, 90-277-2536-5. LCCN QC173.96 .S95 1972. URL http://www.springer.

com/us/book/9789027703453.

Mehra:1975:SCP

[Meh75] Jagdish Mehra, editor. The Solvay conferences on physics: aspects of the development of physics since 1911. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1975. ISBN 90-277-0635-2. xxxii + 415 pp. LCCN QC1.S792 M43.

Meitner:1939:NPF

[Mei39] Lise Meitner. New products of the fission of the thorium nucleus. Nature, 143(3624):637, April 15, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v143/n3624/pdf/143637a0.pdf.

Meitner:1962:RWR

[Mei62] Lise Meitner. Right and wrong roads to the discovery of nuclear energy. International Atomic Energy Agency Bulletin, 4(0):4–6, ???? 1962. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-

2690 (electronic). URL http://www.iaea.org/Publications/
Magazines/Bulletin/Bull040su/04004790608su.pdf.

Meitner:1964:LMLa

[Mei64a] Lise Meitner. Lise Meitner looks back. Advancement of Science, 21(??):39–46, ???? 1964. CODEN ADSCAH. ISSN 0001-866X.

Meitner:1964:LMLb

[Mei64b] Lise Meitner. Lise Meitner looks back. International Atomic Energy Agency Bulletin, 6(1):4–12, ???? 1964. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-2690 (electronic). URL http:/

/www.iaea.org/Publications/Magazines/Bulletin/Bull061/
06101400412.pdf.

Mertens:2001:BB

[Mer01] Richard Mertens. Beyond the bomb. The University of Chicago Magazine, 94(2):??, December 2001. ISSN 0041-9508. URL http:

//magazine.uchicago.edu/0112/features/beyond.html.

Metropolis:1954:EFO

[Met54] N. Metropolis. Enrico Fermi offprints (1922–1954). Technical report, University of Chicago, Chicago, IL, USA, 1954. From the papers of Herbert Anderson.

Metropolis:1955:EF

[Met55] Nicholas Metropolis. Enrico Fermi. Physics Today, 8(11):10–12, November 1955. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://www.physicstoday.org/
resource/1/phtoad/v8/i11/p10_s1.

Metz:1974:RTS

[Met74] William D. Metz. Research trends: Others suffer for the good of the Fermi Lab. Science, 185(4152):686, August 23, 1974. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/185/
4152/686.full.pdf.

Meyer:1937:FTL

[MF37] Stefan Meyer and Enrico Fermi. Further tributes to the late Lord Rutherford. Energia elettrica. Tribute to Lord Rutherford. Nature, 140(3555):1052, December 18, 1937. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v140/

n3555/index.html. Not in Nature online database.

Meitner:1939:DUN

[MF39a] Lise Meitner and Otto Robert Frisch. Disintegration of uranium by neutrons: a new type of nuclear reaction. Nature, 143(3615): 239–240, February 11, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v143/n3615/pdf/143239a0.pdf. This paper,

and [MF39c], both submitted 16 January 1939 (see [Mei62]), provided the first published explanation of nuclear disintegration, called ‘nuclear fission’ by Frisch, that was first observed experimentally by Hahn and Strassmann in December 1939 [HS39b]. Frisch’s paper [Fri39] describes the first experimental confirmation. It was these results that Niels Bohr intended to hold confidential until their journal publication during his January 1939 trip to the USA, but his traveling companion Léon Rosenfeld [Ros72] wasn’t informed of that intent, and the news escaped and spread quickly.

Meitner:1939:PFUa

[MF39b] Lise Meitner and Otto Robert Frisch. On the products of the fission of uranium and thorium under neutron bombardment. Math.- fys. Meddr, 17(5):1–13, 1939.

Meitner:1939:PFUb

[MF39c] Lise Meitner and Otto Robert Frisch. Products of the fission of the uranium nucleus. Nature, 143(3620):471–472, March 18, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v143/
n3620/pdf/143471a0.pdf. See note in [MF39a].

Marconi:196x:VPN

[MFS^{+}6x] Guglielmo Marconi, Enrico Fermi, Emilio Segrè, Daniele Bovet, Giulio Natta, Grazia Deledda, Luigi Pirandello, Salvatore Quadimodo, and Angelo Monteverdi, editors. Le Voci dei Premi Nobel italiani. (Italian). [The voices of Italian Nobel Prize winners]. Discoteca di Stato DSM 251, ????, 196x. Sound recording (33 1/3 rpm).

Millet:2005:OPR

[Mil05] Lydia Millet. Oh pure and radiant heart. Soft Skull Press, Brooklyn, NY, USA, 2005. ISBN 1-932360-85-9. 489 pp. LCCN

PS3563.I42175 O37 2005. URL http://www.loc.gov/catdir/
toc/ecip056/2005001028.htm.

Milotti:2007:EFV

[Mil07] Edoardo Milotti. Enrico Fermi’s view of identical particles. arxiv.org, ??(??):9, May 9, 2007. URL http://arxiv.org/abs/ 0705.1363. See [Fer24b, Fer26g] for the original Italian and German versions.

Miniati:2002:BRS

[Min02] Mara Miniati. Book review: Salvo D’Agostino, Arcangelo Rossi (a cura di), Enrico Fermi e l’Enciclopedia Italiana, Roma, Istituto della Enciclopedia Italiana, 2001, 191 pp., ill. Nuncius, 17(2):723– 724, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911

(electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539102x00315.

Miranda:1934:TMP

[Mir34] Carlo Miranda. Teoremi e metodi per l’integrazione numerica della equazione differenziale di Fermi. (Italian) [Theorems and methods for the numerical integration of the differential equation of the Fermi]. R. Accad. d’Italia. Memoire, 5(??):285–322, ????

Metropolis:1987:NDP

[MKR87] N. (Nicholas) Metropolis, Donald M. Kerr, and Gian-Carlo Rota, editors. New Directions in Physics: The Los Alamos 40th Anniversary Volume. Academic Press, New York, NY, USA, 1987. ISBN 0-12-492155-8. xii + 292 pp. LCCN QC44 .N49

1987. URL http://www.osti.gov/energycitations/product.
biblio.jsp?osti_id=6120718&query_id=0.

Magnusson:1948:FIE

[ML48] L. B. Magnusson and T. J. LaChapelle. The first isolation of element 93 in pure compounds and a determination of the halflife of_{93}Np^{237}. Journal of the American Chemical Society, 70(11): 3534–3538, November 1948. CODEN JACSAT. ISSN 0002-7863 (print), 1520-5126 (electronic), 1943-2984. URL http://pubs.

acs.org/doi/abs/10.1021/ja01191a002.

Mladjenovic:1998:DYN

[Mla98] Milorad Mladjenović. The Defining Years in Nuclear Physics, 1932–1960s. IOP Publishing, Bristol, UK, 1998. ISBN 0-7503-0472-3 (hardcover). xx + 441 pp. LCCN QC773 .M54 1998.

Mongredien:1966:AOS

[Mon66] André Mongredien. Analyses d’ouvrages: The Story of Atomic Theory and Atomic Energy (Formerly titled: The Atom Story) par J. G. Feinberg; Histoire de l’atome, ^{≪}Les grandes découvertes scientifiques ^{≫}(traduction française de la première version de l’ouvrage ci-dessus) par J. G. Feinberg; Stephen Spriel; L’histoire de l’énergie atomique par Laura Fermi; Nicole Rey; La découverte de l’atome, (Petite Bibliothèque Payot, n^{o} 26) par Alfred Romer; J. Métadier; The Discovery of Radioactivity and Transmutation (Classics of Science, vol. 2) par Alfred Romer; Becquerel; Rutherford; Crookes; P. Soddy; M. Curie; Laborde; Ramsay. Revue d’Histoire des Sciences et de Leurs Applications, 19(1):77–81, January 1966. CODEN RHSAAM. ISSN 0048-7996 (print),

1969-6582 (electronic). URL http://www.jstor.org/stable/
23904840.

Moore:1992:ENCd

[Moo92a] Mike Moore. Editor’s note: Consequences. Bulletin of the Atomic Scientists, 48(10):2, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1992:ISF

[Moo92b] Mike Moore. The incident at Stagg Field. Bulletin of the Atomic Scientists, 48(10):11–15, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). 50th anniversary of the first controlled nuclear chain reaction at the University of Chicago, 2 December 1942. See comment [Phi93].

Moore:1995:BRG

[Moo95] Mike Moore. Book review: The gadgets made real: Picturing the Bomb: Photographs from the Secret World of the Manhattan Project, by Rachel Fermi and Esther Samra, introduction by Richard Rhodes. Bulletin of the Atomic Scientists, 51(5):65–66, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.highbeam.

com/doc/1G1-17459604.html. See [FS95].

Morton:1957:DUA

[Mor57] Louis Morton. The decision to use the atomic bomb. Foreign Affairs, 35(2):334–353, January 1957. CODEN FRNAA3. ISSN

0015-7120. URL http://www.jstor.org/stable/20031230.

Morrison:1958:GRA

[Mor58] Philip Morrison. On gamma-ray astronomy. Il Nuovo Cimento (10), 7(6):858–865, March 16, 1958. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://link.

springer.com/article/10.1007/BF02745590.

Morrison:1963:EFC

[Mor63a] Philip Morrison. Enrico Fermi: Collected Papers (Note e Memorie) Vol. I. American Journal of Physics, 31(9):740, September 1963. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://ajp.aapt.org/resource/
1/ajpias/v31/i9/p740_s1;   http://link.aip.org/link/?

AJP/31/740/1. See [Fer62a].

Morrison:1963:FQ

[Mor63b] Philip Morrison. Fermi questions. American Journal of Physics, 31(8):626–627, August 1963. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.aapt.

org/resource/1/ajpias/v31/i8/p626_s2.

Moshinsky:1950:IPP

[Mos50] Marcos Moshinsky. Interference phenomena for particles obeying Bose or Fermi statistics. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 94(1):53–58, February 27, 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://www.jstor.

org/stable/3143251.

Mott:1981:ERE

[Mot81] Sir Nevill Mott. Essay review: Electrons at the Fermi surface. A Festschrift to honour Professor David Shoenberg, edited by M. Springford. Contemporary Physics, 22(2):249–250, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

March:1956:RBW

[MP56] Norman H. March and J. S. Plaskett. The relation between the Wentzel–Kramers–Brillouin and the Thomas–Fermi approximations. Proceedings of the Royal Society of London. Series A, Mathematical and Physical Sciences, 235(1202):419–431, May 8, 1956. CODEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor.

org/stable/99980.

Mott:1977:WHD

[MP77] Sir Nevill Mott and Sir Rudolf Peierls. Werner Heisenberg, 5 December 1901–1 February 1976. Biographical Memoirs of Fellows of the Royal Society, 23:212–251, November 1977. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL http://www.jstor.org/stable/769614;
 https://royalsocietypublishing.org/doi/epdf/10.1098/
rsbm.1977.0009.

Melchiorri:1986:GC

[MR86] F. Melchiorri and Remo Ruffini, editors. Gamow cosmology: Proceedings of the International School of Physics “Enrico-Fermi”, Course 86, held at Varenna, Italy, 13–23 July 1982, volume Course 86 of Proceedings of the International School of Physics “Enrico Fermi”. North-Holland Publishing Co., Amsterdam, The Netherlands, 1986. ISBN 0-444-87004-0. LCCN QB980 .I58 1982.

URL http://adsabs.harvard.edu/abs/1986gaco.conf.....M.

Modugno:2002:CDF

[MRR^{+}02] Giovanni Modugno, Giacomo Roati, Francesco Riboli, Francesca Ferlaino, Robert J. Brecha, and Massimo Inguscio. Collapse of a degenerate Fermi gas. Science, 297(5590):2240–2243, September 27, 2002. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/3832374.

McDougall:1938:CFD

[MS38] J. McDougall and Edmund C. Stoner. The computation of Fermi– Dirac functions. Philosophical Transactions of the Royal Society of London. Series A, Mathematical and Physical Sciences, 237 (773):67–104, February 7, 1938. ISSN 0080-4614. URL http://

www.jstor.org/stable/91333.

Myers:1994:NTF

[MŚ94] W. D. Myers and W. J. Światecki. The nuclear Thomas– Fermi model. Report LBL-36004 / UC-413, Nuclear Science Division Lawrence Berkeley Laboratory University of California, Berkeley, California 94720, USA, August 1994. 44

pp. URL http://www.osti.gov/accomplishments/documents/
fullText/ACC0060.pdf.

Messore:2001:RSG

[MS01] D. Messore and F. Sebastiani. Le ricerche spettroscopiche del gruppo Fermi. (Italian) [The spectroscopic research of the Fermi group]. Giornale di Fisica, 42(2):67–95, April/June 2001. CO-DEN GFSIAD. ISSN 0017-0283 (print), 1827-6156 (electronic).

URL http://prometeo.sif.it/papers/?pid=gdf0385.

Mourik:2012:SMF

[MZF^{+}12] V. Mourik, K. Zuo, S. M. Frolov, S. R. Plissard, E. P. A. M. Bakkers, and L. P. Kouwenhoven. Signatures of Majorana fermions in hybrid superconductor–semiconductor nanowire devices. Science, 336(6084):1003–1007, May 25, 2012. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://arxiv.org/abs/1204.2792;   http://
www.sciencemag.org/content/336/6084/1003.

Nagle:1952:SNP

[NAF^{+}52] Darragh E. Nagle, Herbert L. Anderson, Enrico Fermi, E. A. Long, and R. L. Martin. Scattering of negative pions by hydrogen [abstract only]. Physical Review, 86(4):603, May 1952. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.86.579.2. Min-

utes of the 1952 Annual Meeting Held at New York City, January 31, February 1–2, 1952.

Nambu:1985:QFE

[Nam85] Yoichiro Nambu. Quarks: Frontiers in Elementary Particle Physics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1985. ISBN 9971-966-65-4, 9971-966-66-2 (paperback). xi + 228 pp. LCCN QC793.2 .N36 1985.

Nier:1940:NFS

[NBDG40] Alfred O. Nier, E. T. Booth, J. R. Dunning, and A. V. Grosse. Nuclear fission of separated uranium isotopes. Physical Review, 57(6):546, March 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.org/ doi/10.1103/PhysRev.57.546. This paper reports experiments that confirm the Bohr–Wheeler prediction [BW39b] that U-235 is much more fissile than U-238. That information was critical for the production of both a nuclear reactor, and a nuclear bomb.

Neal:2014:PEC

[Nea14] Mark Neal. Preparing for extraterrestrial contact. Risk Management, 16(2):63–87, May 2014. CODEN ???? ISSN 1460-3799 (print), 1743-4637 (electronic). URL https://www.jstor.org/

stable/43695437.

Nier:1989:SRM

[Nie89] Alfred O. Nier. Some reminiscences of mass spectrometry and the Manhattan Project. Journal of Chemical Education, 66(5):385– 388, 1989. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

Niss:2018:WPP

[Nis18] Martin Niss. What is physics problem-solving competency? The views of Arnold Sommerfeld and Enrico Fermi. Science & Education (Springer), 27(3–4):357–369, May 2018. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic). URL https://

link.springer.com/article/10.1007/s11191-018-9973-z.

Noddack:1934:EGE

[Nod34] Ida Noddack. Über das Element 93. (German) [On element 93 [neptunium]]. Zeitschrift für Angewandte Chemie, 47(37):653– 655, September 15, 1934. ISSN 0932-2132. URL http://en.

wikipedia.org/wiki/Neptunium;   http://onlinelibrary.
wiley.com/doi/10.1002/ange.19340473707/abstract;   http:

//www.chemteam.info/Chem-History/Noddack-1934.html. Ac-

cording to Frisch [FW67, page 48], “Ida Noddack, a German chemist, quite rightly pointed out that they might be lighter elements [after bombardment of uranium by neutrons]; but her comments (published in a journal not much read by chemists and hardly at all by physicists) were regarded as mere pedantry. She did not indicate how such light elements could be formed; her paper had probably no effect whatever on later work.”. English translation in [GA71, pages 16–20].

Oakes:2018:RDV

[Oak18] Andrew Oakes. Review: Douglas A. Vakoch and Matthew F. Dowd. The Drake Equation: Estimating the Prevalence of Extraterrestrial Life through the Ages. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1):186–188, ???? 2018. CODEN ???? ISSN 1913-0465. URL https:/

/spontaneousgenerations.library.utoronto.ca/index.php/
SpontaneousGenerations/article/view/29357.

Oppenheimer:1949:AAC

[OBC^{+}49] J. Robert Oppenheimer, Oliver E. Buckley, James B. Conant, Lee A. DuBridge, Enrico Fermi, I. I. Rabi, Hartley Rowe, Glenn T. Seaborg, and Cyril S. Smith. AEC Advisory Committee statement on fellowships. Bulletin of the Atomic Scientists, 5(8–9):210, 254, August/September 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oliver:1975:PGC

[Oli75] B. M. Oliver. Proximity of galactic civilizations. Icarus: International Journal of Solar System Studies, 25(2):360–367, June 1975. CODEN ICRSA5. ISSN 0019-1035 (print), 1090-2643 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0019103575900317.

Olsen:1963:TGS

[Ols63] Arthur J. Olsen. Trackdown of the German scientist: Nazism and defeat scattered the leaders of a once-great scientific establishment. Herewith a review of where they went and what some of them did when they got there. New York Times, ??(??): 214, September 22, 1963. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/116463805/.

ONeill:1974:CS

[O’N74] Gerard K. O’Neill. The colonization of space. Physics Today, 27(9):32–40, September 1974. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v27/i9/p32_s1.

Orear:1958:NSP

[Ore58] Jay Orear. Notes on statistics for physicists, revised. Report UCRL-8417, University of California Radiation Laboratory, Berkeley, CA, USA, August 13, 1958. 36 pp. URL http:

//cds.cern.ch/record/104881/files/SCAN-9709037.pdf;
http://hdl.handle.net/2027/mdp.39015077593799. From the

preface: “The primary source for the basic material and approach presented here was Enrico Fermi.” See also revised and typographically-improved version [Ore82].

Orear:1982:NSP

[Ore82] Jay Orear. Notes on statistics for physicists, revised. Report CLNS 82/511, Laboratory for Nuclear Studies, Cornell University, Ithaca, NY 14853, USA, July 28, 1982. iv + 45 pp. URL

http://pages.physics.cornell.edu/p510/w/images/p510/6/
62/Notes_on_Statistics_for_Physicists.pdf. From the pref-

ace: “The primary source for the basic material and approach presented here was Enrico Fermi.” Revision of original version [Ore58].

Orear:2001:EFM

[Ore01a] Jay Orear. Enrico Fermi, the man. Il Nuovo Saggiatore, 17(5–6): 30–38, September/December 2, 2001. ISSN 0393-4578 (print),

1827-6148 (electronic). URL http://prometeo.sif.it:8080/
papers/online/sag/2001/05-06/pdf/05.pdf. Special issue in

honor of the centennial of the birth of Enrico Fermi.

Orear:2001:NSP

[Ore01b] Jay Orear. Notes on statistics for physicists, revised. Report CLNS 82/511, Laboratory for Nuclear Studies, Cornell University, Ithaca, NY 14853, USA, September 24, 2001. 32 pp. URL

http://pages.physics.cornell.edu/p510/w/images/p510/6/
62/Notes_on_Statistics_for_Physicists.pdf. From the pref-

ace: “The primary source for the basic material and approach presented here was Enrico Fermi.” Revision of earlier versions [Ore58, Ore82].

Orear:2003:EFM

[Ore03a] Jay Orear. Enrico Fermi, the man. Excerpts from some documents. In Anonymous [Ano03], pages 317–340. ISBN 88-

8286-032-9. LCCN ???? URL http:old.enea.it/produzione_
scientifica/pdf.../V2003_AttiFermi.pdf.

Oreskes:2003:GBI

[Ore03b] Naomi Oreskes. Gender bias and Ida Noddack: Response. Science, 301(5636):1045–1046, August 2003. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). See [Hoo03].

Orear:2004:EFM

[Ore04] Jay Orear, editor. Enrico Fermi — The Master Scientist. The Internet-First University Press, Ithaca, NY, USA, 2004.

ISBN ????  iv + 165 pp. LCCN ????  URL http:

//dspace.library.cornell.edu/handle/1813/62;   http://
hdl.handle.net/1813/74.

Orlando:1998:PJP

[Orl98] Lucida Orlando. Physics in the 1930s: Jewish physicists’ contributions to the realization of the ‘new tasks’ of physics in Italy. Historical Studies in the Physical and Biological Sciences, 29(1): 141–182, ???? 1998. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Onorato:2015:RTF

[OVPL15] Miguel Onorato, Lara Vozella, Davide Proment, and Yuri V. Lvov. Route to thermalization in the α-fermi–pasta–ulam system. Proceedings of the National Academy of Sciences of the United States of America, 2015. CODEN PNASA6. ISSN 0027-8424 (print),

1091-6490 (electronic). URL http://www.pnas.org/content/
early/2015/03/23/1404397112.abstract.

Pontecorvo:1942:RAI

[PAF^{+}42] Bruno Pontecorvo, Edoardo Amaldi, Enrico Fermi, Emilio Segrè, and Franco Rasetti. Radio-active isotope production. Canadian Patent 407558., September 22, 1942.

Pais:1986:IBM

[Pai86] Abraham Pais. Inward bound: of matter and forces in the physical world. Clarendon Press, Oxford, UK, 1986. ISBN 0-19-851971-0, 0-19-851997-4 (paperback). xiv + 666 pp. LCCN QC7 .P27 1986. US$24.95.

Paldy:1969:BFR

[Pal69] Lester G. Paldy. Book and film reviews: Nuclear physics by a master: The Collected Papers of Enrico Fermi, Volume I, Italy 1921–1938, Volume II, United States 1939–1954, Emilio Segrè. The Physics Teacher, 7(8):470, November 1969. CO-DEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

URL http://tpt.aapt.org/resource/1/phteah/v7/i8/p470_

s1. See [Fer62a].

Papagiannis:1978:WAA

[Pap78] M. D. Papagiannis. Are we all alone, or could they be in the asteroid belt? Quarterly Journal of the Royal Astronomical Society, 19 (??):277, September 1978. CODEN QJRAAK. ISSN 0035-8738.

URL http://adsabs.harvard.edu/abs/1978QJRAS..19..277P.

Papagiannis:1985:RPF

[Pap85] Michael D. Papagiannis. Recent progress and future plans on the search for extraterrestrial intelligence. Nature, 318(6042):135–140, November 14, 1985. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v318/n6042/pdf/318135a0.pdf.

Parisi:2002:SFI

[Par02] Giorgio Parisi. La statistica di Fermi. (Italian) [Fermi’s statistics]. In Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901–2001] [Ber02a], pages 68–75. ISBN 88-7438-000-3. LCCN QC774.F4. 10.33 EUR. URL http://

prometeo.sif.it:8080/libri/fermi/04.pdf. English transla-

tion in [BB04b].

Pauli:1925:ZAE

[Pau25] Wolfgang Pauli. Über den Zusammenhang des Abschlusses der Elektronengruppen im Atom mit der Komplexstruktur der Spektren. (German) [On the relation of the completion of groups of electrons in the atom with the complex structure of spectra]. Zeitschrift für Physik, 31(1):765–783, February 1925. CODEN ZEPYAA. ISSN 0044-3328. This is the paper in which Pauli introduced the famous Exclusion Principle, for which he received the Nobel Prize in Physics 1945 [Pau45].

Pauli:1930:LTC

[Pau30] Wolfgang Pauli. Letter to Tübingen conference participants. Web document., December 4, 1930. English translation in [Bro78, page 27].

Pauli:1934:DPM

[Pau34] Wolfgang Pauli. Discussion of paper by M. Heisenberg on “La Structure du Noyau”. In Cockcroft et al. [CCJ^{+}34], pages 324– 325. LCCN ???? Publiés par la commission administrative de l’institut.

Pauli:1945:NPP

[Pau45] Wolfgang Pauli. The Nobel Prize in Physics 1945. Nobelprize.org, 1945. The Nobel Prize in Physics 1945 was awarded to Wolfgang Pauli “for the discovery of the Exclusion Principle, also called the Pauli Principle”.

Pais:2006:JRO

[PC06] Abraham Pais and Robert P. Crease. J. Robert Oppenheimer: a life. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-516673-6. xxii + 353 + 16 pp. LCCN QC16.O62 P35 2006. URL http://www.loc.gov/

catdir/enhancements/fy0636/2005002173-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0723/2005002173-b.
html;   http://www.loc.gov/catdir/toc/ecip056/2005002173.
html.

Pegram:1939:EFN

[Peg39] George B. Pegram. Enrico Fermi — Nobel Prize man in physics for 1938. The Scientific Monthly, 49(2):182–184, August 1939. CO-DEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic).

URL http://www.jstor.org/stable/17044.

Peierls:1986:FFW

[Pei86] Rudolf Peierls. Fact or fission [wild speculation]. London Review of Books, 8(18):4, October 23, 1986. ISSN 0260-9592. URL https:

//www.lrb.co.uk/v08/n18/letters#letter4.

Peierls:1993:P

[Pei93] Rudolf Peierls. Physics 1920–45. Atti Convegni Lincei, 104(??): 23–33, ???? 1993. Proceedings of the Symposium “Enrico Fermi” (50th anniversary of the first reactor) held 10 December 1992 in Rome, Italy.

Penzias:1979:OE

[Pen79] Arno A. Penzias. The origin of the elements. Reviews of Modern Physics, 51(3):425–431, July 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.51.425;   http:/
/rmp.aps.org/abstract/RMP/v51/i3/p425_1.

Penny:2013:SED

[Pen13] Alan John Penny. The SETI episode in the 1967 discovery of pulsars. European Physical Journal H, 38(4):535–547, September 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/
10.1140/epjh/e2012-30052-6.

Peoples:2004:FT

[Peo04] John Peoples. Fermi and technology. In Orear [Ore04], chapter 25, pages 125–128. ISBN ???? LCCN ???? URL http:

//dspace.library.cornell.edu/handle/1813/62;   http://
hdl.handle.net/1813/74.

Persico:1955:SEF

[Per55] Enrico Persico. Souvenir de Enrico Fermi. (French) [Memory of Enrico Fermi]. Scientia (Milan), 90(522):316–324, 1955. CODEN SCIMAI. ISSN 0036-8687 (print), 1825-4373 (electronic). URL

http://amshistorica.unibo.it/7.

Persico:2002:CEF

[Per02] Enrico Persico. Commemorazione di Enrico Fermi. (Italian) [commemoration of Enrico Fermi ]. In Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901– 2001] [Ber02a], pages 37–45. ISBN 88-7438-000-3. LCCN

QC774.F4. 10.33 EUR. URL http://prometeo.sif.it:8080/

libri/fermi/18.pdf. English translation in [BB04b].

Pestre:2003:NLA

[Pes03] Dominique Pestre. New large accelerators in the world in the forties and early fifties. In Anonymous [Ano03], pages 219–220. ISBN 88-8286-032-9. LCCN ???? URL http:old.enea.it/

produzione_scientifica/pdf.../V2003_AttiFermi.pdf.

Polvani:1922:RIR

[PFM^{+}22] G. Polvani, Enrico Fermi, I. Maghieru, M. Pierucci, A. Pontremoli, et al. Rivista. (Italian) [Reviews]. Il Nuovo Cimento (6), 23(1):i–xl, ???? 1922. CODEN NUCIAD. ISSN 0029-6341 (print),

1827-6121 (electronic). URL http://www.springerlink.com/
content/g8r59x5431702m9g/.

Parr:1986:TFT

[PG86] Robert G. Parr and Swapan K. Ghosh. Thomas–Fermi theory for atomic systems. Proceedings of the National Academy of Sciences of the United States of America, 83(11):3577–3579, June 1, 1986. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490

(electronic). URL http://www.jstor.org/stable/27728.

Phillips:1993:LMD

[Phi93] Alan F. Phillips. Letter: Moral distinctions. Bulletin of the Atomic Scientists, 49(2):45, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Moo92b].

Park:2025:MNG

[PHK25] Patrick J. Park, Sebastian Herzele, and Timothy W. Koeth. Myths of nuclear graphite in World War II, with original translations. European Physical Journal H, 50(1):??, December 2025. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-025-00098-7.

Pietronero:2001:CEF

[Pie01] Luciano Pietronero. Il centro Enrico Fermi di via Panisperna: Museo storico e centro studi e ricerche. (Italian) [The Enrico Fermi Center on via Panisperna: Historical museum and study and research center]. Il Nuovo Saggiatore, 17(5–6):97–98, September/December 2001. ISSN 0393-4578 (print), 1827-6148 (elec-

tronic). URL http://prometeo.sif.it:8080/papers/online/

sag/2001/05-06/pdf/05.pdf. Special issue in honor of the centennial of the birth of Enrico Fermi.

Pincherle:1974:FS

[Pin74] L. Pincherle. Fermi surfaces. Nature, 247(5435):78, January 4, 1974. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v247/n5435/pdf/247078a0.pdf.

Pinault:2003:CCA

[Pin03] Michel Pinault. Cooperation and competition among nuclear physics laboratories during the thirties: the role of Frédéric Joliot. In Anonymous [Ano03], pages 119–132. ISBN 88-8286-

032-9. LCCN ???? URL http:old.enea.it/produzione_
scientifica/pdf.../V2003_AttiFermi.pdf.

Pontecorvo:1955:EFK

[Pon55] Bruno M. Pontecorvo. Enrico Fermi (1901–1954) (k godovschine cmerti). (Russian) [Enrico Fermi (1901–1954) (the anniversary of his death)]. Uspekhi Fizicheskikh Nauk, 57(11):349–359, November 1955. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652

(electronic). URL http://ufn.ru/ru/articles/1955/11/a/.

Pontekorvo:1972:FFM

[Pon72] B. (Bruno) Pontekorvo. Fermi e la fisica moderna. (Italian) [Fermi and modern physics]. Editori riuniti, Roma, Italiy, 1972. 86 pp. LCCN QC16.F46 P633.

Pontecorvo:1992:AEI

[Pon92] B. Pontecorvo. Alcuni episodi istruttivi nella vita e nel lavoro di Enrico Fermi. (Italian) [Some instructive episodes in the life and work of Enrico Fermi]. In Galbiata et al. [GER^{+}92], pages 33–?? ISBN 88-7794-051-4. LCCN ????

Pontecorvo:1993:EFR

[Pon93] B. (Bruno) Pontecorvo. Enrico Fermi: ricordi di allievi e amici. (Italian) [Enrico Fermi: memories of students and friends], volume 4 of Collezione Biblioteca contemporanea. Edizioni Studio Tesi, Pordenone, Italy, 1993. ISBN 88-7692-420-5. xii + 210 pp. LCCN QC16.F46 P6316 1993. Translated from the Russian original by Giovanna Marinella Tabet and Marina Battazzi.

Pontekorvo:2004:FFM

[Pon04] B. (Bruno) Pontekorvo. Fermi e la fisica moderna. (Italian) [Fermi and modern physics], volume 4 of Universale di base. La città del sole, Napoli, Italy, 2004. ISBN 88-8292-260-X. 80 pp. LCCN QC16.F46 P635 2004. Introduction by Andrea Martocchia.

Pontecorvo:2013:AEI

[Pon13] Bruno Pontecorvo. Alcuni episodi istruttivi nella vita e nel lavoro di Enrico Fermi. (Italian) [Some instructive episodes in the life and work of Enrico Fermi]. In Bilenkij et al. [B^{+}13], pages 523–526. ISBN 88-7438-080-1. LCCN ???? URL http://www.sif.it/

libri/pontecorvo_2nd/contents.

Pontecorvo:1972:EFV

[PP72] B. M. Pontecorvo and V. N. Pokrovskij. Enriko Fermi v vospominanijakh uchenikov i druzej. (Russian) [Enrico Fermi remembered by his students and friends]. Nauka, Moscow, Russia, 1972. ISBN ???? 323 pp. LCCN ????

Pang:1994:PPF

[PQQ94] Yang Pang, Jianwei Qiu, and Zhaoming Qiu, editors. Particle physics at the Fermi scale, volume 10 of China Center of Advanced Science and Technology (World Laboratory) Symposium/Workshop proceedings, 0894-2536. Gordon and Breach Science Publishers, Switzerland, 1994. ISBN 2-88449-108-2. LCCN QC793 .C39 1993.

Prasar:2004:EFF

[Pra04] Vigyan Prasar. Enrico Fermi: Father of nuclear physics. Government of India Web site, 2004. URL http://www.vigyanprasar.

gov.in/scientists/EFermi.htm.

Price:1995:RDC

[Pri95] Matt Price. Roots of dissent: The Chicago Met Lab and the origins of the Franck Report. Isis, 86(2):222–244, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236323.

Petrucci:2007:LDD

[PS07] C. Petrucci and F. Sebastiani. L’opera di divulgazione della fisica quantistica svolta in Italia da Enrico Fermi negli anni Venti. (Italian) [The work of dissemination of physics quantum conducted in Italy by Enrico Fermi in the Twenties]. Quaderni di Storia della fisica, 14:49–70, 2007. CODEN ???? ISSN 1594-9974 (print),

1827-6164 (electronic). URL http://en.sif.it/journals/qsf/
econtents.

Pasta:1965:SNP

[PUF65] John R. Pasta, Stanisław M. Ulam, and Enrico Fermi. Studies on nonlinear problems. In Collected works of Enrico Fermi, volume 2, page 978. University of Chicago Press, Chicago, IL, USA, 1965.

Purcell:1963:BKS

[Pur63] John Francis Purcell. The best-kept secret; the story of the atomic bomb. Vanguard Press, New York, NY, USA, 1963. 188 pp. LCCN QC773.A1 P8.

Porter:2009:FPU

[PZHC09] Mason A. Porter, Norman J. Zabusky, Bambi Hu, and David K. Campbell. Fermi, Pasta, Ulam and the birth of experimental mathematics: a numerical experiment that Enrico Fermi, John Pasta, and Stanisław Ulam reported 54 years ago continues to inspire discovery. American Scientist, 97(3):214–221, May/ June 2009. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-

2786 (electronic). URL http://www.americanscientist.org/
issues/feature/2009/3/fermi-pasta-ulam-and-the-birth-
of-experimental-mathematics. The work referred to in the

title is [FPU55].

Rabi:1963:BRR

[Rab63] I. I. Rabi. Book review: A real physicist: The Collected Papers of Enrico Fermi. vol. 1. Italy, 1921–1938 by Enrico Fermi and Emilo Segrè. Science, 139(3556):746, February 22, 1963. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1710208;
http://www.sciencemag.org/content/139/3556/746.1.full.

pdf. See [Fer62a].

Rabi:1966:BRE

[Rab66] I. I. Rabi. Book review: Enrico Fermi’s papers, 1939 to 1954: The Collected Papers of Enrico Fermi. vol. 2. Science, 152(3724): 950, May 13, 1966. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
1718593;   http://www.sciencemag.org/content/152/3724/
950.1.full.pdf. See [Fer65b].

Rasetti:1955:EF

[Ras55] Franco Rasetti. Enrico Fermi. Science, 121(3144):449–451, April 1, 1955. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
1681653;   http://www.sciencemag.org/content/121/3144/
449.full.pdf.

Rasetti:1968:EFF

[Ras68] Franco Rasetti, editor. Enrico Fermi e la fisica italiana. Accademia nazionale dei Lincei. Celebrazioni lincee. Accademia nazionale dei Lincei, Roma, Italia, 1968. 18 pp. LCCN QC16.F4 E5.

Rasetti:2002:EFF

[Ras02] Franco Rasetti. Enrico Fermi e la fisica italiana. (Italian) [Enrico Fermi and Italian physics]. In Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901– 2001] [Ber02a], pages 46–56. ISBN 88-7438-000-3. LCCN

QC774.F4. 10.33 EUR. URL http://prometeo.sif.it:8080/

libri/fermi/19.pdf. English translation in [BB04b].

Reines:1953:DFN

[RC53] Frederick Reines and Clyde L. Cowan. Detection of the free neutrino. Physical Review, 92(3):830–831, November 1953. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.92.830. See con-

firmation [CRH^{+}56].

Russell:2012:F

[RC12] Jesse Russell and Ronald Cohn. FERMIAC. Bookvika Publishing, Moscow, Russia, 2012. ISBN ???? ???? pp.

Rees:2000:JSN

[Ree00] Martin J. Rees. Just six numbers: the deep forces that shape the universe. Basic Books, New York, NY, USA, 2000. ISBN 0-465-03673-2. xi + 195 pp. LCCN QB 981 R434 2000.

Reed:2017:RAP

[Ree17] B. Cameron Reed. Revisiting The Los Alamos Primer. Physics Today, 70(9):42–49, September 2017. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Regis:1985:ESA

[Reg87] Edward Regis, Jr., editor. Extraterrestrials: Science and Alien Intelligence. Cambridge University Press, Cambridge, UK, 1985– 1987. ISBN 0-521-26227-5 (hardcover), 0-521-34852-8 (paperback). x + 278 pp. LCCN QB54 .E947 1987.

Regge:2003:FGR

[Reg03] Tullio Regge. Fermi and General Relativity. In Anonymous [Ano03], pages 303–304. ISBN 88-8286-032-9. LCCN

???? URL http:old.enea.it/produzione_scientifica/pdf.
../V2003_AttiFermi.pdf.

Reilly:1969:BRI

[Rei69] Kevin Reilly. Book review: Illustrious Immigrants: The Intellectual Migration From Europe 1930–41, by Laura Fermi. Journal of Social History, 3(1):81, ???? 1969. ISSN 0022-4529 (print),

1527-1897 (electronic). URL http://www.jstor.org/stable/
3786649.

Reines:1972:CFOa

[Rei72] Frederick Reines, editor. Cosmology, fusion and other matters: George Gamow memorial volume. Colorado Associated University Press, Boulder, CO, USA, 1972. ISBN 0-87081-025-1. xiv + 320 pp. LCCN QC780 .C65.

Reines:1996:NPP

[Rei96] F. Reines. The neutrino: from poltergeist to particle. Reviews of Modern Physics, 68(2):317–327, April 1996. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756.  URL http://link.aps.org/doi/10.
1103/RevModPhys.68.317;   http://rmp.aps.org/abstract/
RMP/v68/i2/p317_1;   http://www.nobelprize.org/nobel_
prizes/physics/laureates/1995/.

Rasetti:1926:SER

[RF26] Franco Rasetti and Enrico Fermi. Sopra l’elettrone rotante. (Italian) [On the spinning electron]. Il Nuovo Cimento (8), 3(1–2): 226–235, ???? 1926. CODEN NUCIAD. ISSN 0029-6341 (print),

1827-6121 (electronic). URL http://www.springerlink.com/
content/7h082w4562804802/.

Rhodes:1950:FDF

[Rho50] P. Rhodes. Fermi–Dirac functions of integral order. Proceedings of the Royal Society of London. Series A, Mathematical and Physical Sciences, 204(1078):396–405, December 22, 1950. CO-DEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor.org/

stable/98693.

Rhodes:1986:MAB

[Rho86] Richard Rhodes. The Making of the Atomic Bomb. Simon and Schuster, New York, NY, USA, 1986. ISBN 0-671-44133-7 (paperback), 0-671-65719-4 (hardcover). 886 + 42 pp. LCCN QC773 .R46 1986.

Rhodes:1995:DSM

[Rho95] Richard Rhodes. Dark Sun: The Making of the Hydrogen Bomb. Sloan technology series. Simon and Schuster, New York, NY, USA, 1995. ISBN 0-684-80400-X. 731 pp. LCCN UG1282.A8 R46 1995.

Rhodes:1999:APE

[Rho99] Richard Rhodes. Atomic physicist: Enrico Fermi. Time, 153 (12):154–??, March 29, 1999. CODEN TYMEA9. ISSN 0040-781x (print), 2169-1665 (electronic). URL http://www.time.

com/time/magazine/article/0,9171,990618,00.html.

Rhodes:2018:EHH

[Rho18a] Richard Rhodes. Energy: a human history. Simon and Schuster, New York, NY, USA, 2018. ISBN 1-5011-0535-3 (hardcover), 1-5011-0537-X (e-book). xiv + 464 pp. LCCN TJ163.2 .R56 2018; TJ163.2 .R498 2018.

Rhodes:2018:FPF

[Rho18b] Richard Rhodes. Full power in fifty-seven. In Energy: a human history [Rho18a], chapter 17, pages 272–292. ISBN 1-5011-0535-3 (hardcover), 1-5011-0537-X (e-book). LCCN TJ163.2 .R56 2018; TJ163.2 .R498 2018.

Ricci:2001:FV

[Ric01] Renato Angelo Ricci. Fermi in Varenna. Il Nuovo Saggiatore, 17(5–6):40–45, September/December 2, 2001. ISSN 0393-4578 (print), 1827-6148 (electronic). URL http://prometeo.sif.it:

8080/papers/online/sag/2001/05-06/pdf/05.pdf. Special is-

sue in honor of the centennial of the birth of Enrico Fermi.

Ricci:2002:ULF

[Ric02] Renato Angelo Ricci. Le ultime lezioni di Fermi. (Italian) [Fermi’s last lectures]. In Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901–2001] [Ber02a], pages 290– 317. ISBN 88-7438-000-3. LCCN QC774.F4. 10.33 EUR. URL

http://prometeo.sif.it:8080/libri/fermi/16.pdf. English

translation in [BB04b].

Rider:1984:AOE

[Rid84] Robin E. Rider. Alarm and opportunity: Emigration of mathematicians and physicists to Britain and the United States, 1933– 1945. Historical Studies in the Physical Sciences, 15(1):107–176, ???? 1984. CODEN HSPSAS. ISSN 0073-2672. URL http://

www.jstor.org/stable/27757544.

Rife:2006:LMD

[Rif06] Patricia Rife. Lise Meitner and the dawn of the nuclear age. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2006. ISBN 0-8176-4559-4 (softcover), 0-8176-3732-X (hardcover). ???? pp. LCCN ???? URL http:

//www.loc.gov/catdir/enhancements/fy0812/2006935949-d.
html;   http://www.loc.gov/catdir/enhancements/fy0812/

2006935949-t.html. Foreword by John Archibald Wheeler.

Rigden:1984:EHW

[Rig84] John S. Rigden. Editorial — how will Fermi be remembered? American Journal of Physics, 52(10):875, October 1984. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://ajp.aapt.org/resource/1/ajpias/v52/i10/p875_s1.

Rigden:2000:IIR

[Rig00] John S. Rigden. Isidor Isaac Rabi: 29 July 1898–11 January 1988. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 144(1):113–118, March 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Rijnierse:1966:HDE

[Rij66] P. J. Rijnierse. High density expansion for the Thomas–Fermi– Dirac function. Proceedings of the Royal Society of London. Series A, Mathematical and Physical Sciences, 292(1429):288–297, May 17, 1966. CODEN PRLAAZ. ISSN 0080-4630. URL http://

www.jstor.org/stable/2415718.

Riley:1970:WEF

[Ril70] William Riley. The world of Enrico Fermi. 47 minute film, from Harvard Project Physics., 1970.

Ryutova-Kemoklidze:1995:QGH

[RK95] Margarita Ryutova-Kemoklidze. The quantum generation: highlights and tragedies of the golden age of physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-53298-6 (Berlin), 0-387-53298-6 (New York). xix + 327 pp. LCCN QC7 .R9613 1995. Translated from Russian original by J. Hine.

Ringhofer:2004:SAE

[RNKS04] C. Ringhofer, M. Nedjalkov, H. Kosina, and S. Selberherr. Semiclassical approximation of electron–phonon scattering beyond Fermi’s Golden Rule. SIAM Journal on Applied Mathematics, 64(6):1933–1953, August/September 2004. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://

www.jstor.org/stable/4096104.

Robson:1950:RDN

[Rob50] J. M. Robson. Radioactive decay of the neutron. Physical Review, 78(3):311–312, May 1950. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRev.78.311.

Robson:1983:ESB

[Rob83] J. M. Robson. Experimental studies of the beta decay of the neutron: a historical review. Contemporary Physics, 24(2):129–141, ???? 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00107518308210669.

Robison:1995:RMX

[Rob95] Roger F. Robison. The race for megavoltage X-rays versus Telegamma. Acta Oncologica (Stockholm), 34(8):1055–1074, 1995. CODEN ACTOEL. ISSN 0284-186X (print), 1651-226X

(electronic). URL https://www.tandfonline.com/doi/abs/10.
3109/02841869509127233.

Robinson:2008:DJS

[Rob08] A. W. Robinson. Don’t just stand there — teach Fermi problems! Physics Education, 43(1):83–87, 2008. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic). URL http://

iopscience.iop.org/0031-9120/43/1/009.

Robinson:2012:SED

[Rob12] Andrew Robinson, editor. The scientists: an epic of discovery. Thames and Hudson, New York, NY, USA, 2012. ISBN 0-500-25191-6. 304 pp. LCCN Q141 .S3712 2012.

Rodgers:2019:TAS

[Rod19] Glen E. Rodgers. Travelling with the Atom: a Scientific Guide to Europe and Beyond. Royal Society of Chemistry, Cambridge, UK, 2019. ISBN 1-78801-528-2 (paperback), 1-78801-702-1 (e-book). xxxii + 551 pp. LCCN QC171.2 .R63 2020.

Rogers:2010:MIS

[Rog10] Kara Rogers, editor. The 100 Most Influential Scientists of All Time. The Britannica guide to the world’s most influential people. Britannica Educational Publishers, in association with Rosen Educational Services, New York, NY, USA, 2010. ISBN 1-61530-002-3 (library binding). 360 pp. LCCN Q162 .A15 2010.

Rogers:2013:NDY

[Rog13] J. D. Rogers. The neutron’s discovery — 80 years on. Physics Procedia, 43:1–9, 2013. CODEN PPHRCK. ISSN 1875-3892. URL

http://adsabs.harvard.edu/abs/2013PhPro..43....1R.

Romer:1991:EMP

[Rom91] Robert H. Romer. Editorial: Memorable papers from the American Journal of Physics, 1933–1990. American Journal of Physics, 59(3):201–207, March 1991. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Roque:1992:MSN

[Roq92] Xavier Roqué. Møller scattering: a neglected application of early quantum electrodynamics. Archive for History of Exact Sciences, 44(3):197–264, September 1992. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=44&issue=3&spage=197.

Rosenfeld:1972:NR

[Ros72] Léon Rosenfeld. Nuclear reminiscences. In Reines [Rei72], pages 289–299. ISBN 0-87081-025-1. LCCN QC780 .C65. URL http:/

/adsabs.harvard.edu/abs/1972cht..conf..289R.

Rosenfeld:2004:CSF

[Ros04] Art Rosenfeld. Comments of some former grad students. In Orear [Ore04], chapter 27, pages 139–140. ISBN ???? LCCN ????

URL http://dspace.library.cornell.edu/handle/1813/62;
 http://hdl.handle.net/1813/74.

Rubbia:2003:FCW

[Rub03] Carlo Rubbia. Fermi’s contribution to the world energy supply. In Anonymous [Ano03], pages 33–42. ISBN 88-8286-032-9. LCCN

???? URL http:old.enea.it/produzione_scientifica/pdf.
../V2003_AttiFermi.pdf.

Rutherford:1920:BLN

[Rut20] Sir Ernest Rutherford. Bakerian Lecture: Nuclear constitution of atoms. Proc. Roy. Soc. A, 97(686):374–400, July

1, 1920. URL http://web.lemoyne.edu/~giunta/ruth1920.
html;   http://www.jstor.org/stable/93888.

Schaerf:1979:PFP

[SA79] Carlo Schaerf and Edoardo Amaldi, editors. Perspectives of fundamental physics: proceedings of the conference held at the University of Rome 7–9 September 1978, dedicated to Edoardo Amaldi on the occasion of his retirement from his teaching duties at the University of Rome, volume 1 of Studies in high energy physics. Harwood Academic Publishers, Chur, Switzerland; New York, NY, USA, 1979. ISBN 3-7186-0007-2. LCCN QC770 .C748 1979.

Sandberg:2016:DWE

[SAĆ16] Anders Sandberg, Stuart Armstrong, and Milan M. Ćirkoviç. That is not dead which eternal lie: The aestivation hypothesis for resolving Fermi’s Paradox. Journal of the British Interplanetary Society, 69:406–415, 2016. CODEN JBISAW. ISSN 0007-

084X. URL https://arxiv.org/abs/1705.03394;   https://
www.jbis.org.uk/paper/2016.69.406. See comment [BHR19].

Sagan:1973:CEI

[Sag73] Carl Sagan, editor. Communication with Extraterrestrial Intelligence (CETI). MIT Press, Cambridge, MA, USA, 1973. ISBN 0-262-19106-7. LCCN QB54 .S68 1971a.

Sagan:1983:NWC

[Sag83] Carl Sagan. Nuclear war and climatic catastrophe: Some policy implications. Foreign Affairs, 62(2):257–292, Winter 1983. CODEN FRNAA3. ISSN 0003-0554. URL http://www.jstor.

org/stable/20041818.

Sagan:2004:IN

[Sag04] Carl Sagan. The Italian Navigator. In Orear [Ore04], chapter 17, pages 71–74. ISBN ???? LCCN ???? URL http:

//dspace.library.cornell.edu/handle/1813/62;   http://
hdl.handle.net/1813/74.

Salvini:2001:EFM

[Sal01] Giorgio Salvini. Enrico Fermi il maestro sperimentale e teorico del secolo ora trascorso. Alcuni personali ricordi. (Italian) [Enrico Fermi, the experimental and theoretical master of the last century. Some personal memories]. Il Nuovo Saggiatore, 17(5–6): 20–23, September/December 2, 2001. ISSN 0393-4578 (print),

1827-6148 (electronic). URL http://prometeo.sif.it:8080/
papers/online/sag/2001/05-06/pdf/05.pdf. Special issue in

honor of the centennial of the birth of Enrico Fermi.

Salvetti:2002:NDN

[Sal02a] Carlo Salvetti. Nascita dell’energia nucleare: La pila di Fermi. (Italian) [birth of nuclear energy: Fermi’s pile]. In Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901–2001] [Ber02a], pages 178–204. ISBN 88-7438-000-3. LCCN QC774.F4. 10.33 EUR. URL http://prometeo.sif.it: 8080/libri/fermi/11.pdf. English translation in [BB04b].

Salvini:2002:EFS

[Sal02b] Giorgio Salvini. Enrico Fermi. La sua vita, ed un commento alla sua opera. (Italian) [Enrico Fermi. his life, and a comment on his work]. In Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901–2001] [Ber02a], pages 1– 22. ISBN 88-7438-000-3. LCCN QC774.F4. 10.33 EUR. URL

http://prometeo.sif.it:8080/libri/fermi/02.pdf. English

translation in [BB04b].

Salvini:2003:CR

[Sal03a] Giorgio Salvini. Concluding remarks. In Anonymous [Ano03], pages 395–398. ISBN 88-8286-032-9. LCCN ???? URL http:old.

enea.it/produzione_scientifica/pdf.../V2003_AttiFermi.
pdf.

Salvini:2003:EFG

[Sal03b] Giorgio Salvini. Enrico Fermi: a guiding light in an anguished century. In Anonymous [Ano03], pages 13–32. ISBN 88-8286-

032-9. LCCN ???? URL http:old.enea.it/produzione_
scientifica/pdf.../V2003_AttiFermi.pdf.

Santilli:1981:ILE

[San81] Ruggero Maria Santilli. An intriguing legacy of Einstein, Fermi, Jordan, and others: The possible invalidation of quark conjectures. Foundations of Physics, 11(5–6):383–472, June 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00727075.

Sebastiani:2003:FTQ

[SC03] Fabio Sebastiani and Francesco Cordella. Fermi toward quantum statistics (1923–1925). In Anonymous [Ano03], pages 71– 96. ISBN 88-8286-032-9. LCCN ???? URL http:old.enea.it/

produzione_scientifica/pdf.../V2003_AttiFermi.pdf.

Schwinger:1958:SPQ

[Sch58] Julian Schwinger, editor. Selected Papers on Quantum Electrodynamics. Dover books on engineering and engineering physics. Dover, New York, NY, USA, 1958. ISBN 0-486-60444-6. xvii + 424 pp. LCCN QC680 .S35.

Schorn:1981:EBD

[Sch81] R. Schorn. Extraterrestrial beings don’t exist. Sky and Telescope, ??(1):??, January 1981. CODEN SKTEA3. ISSN 0037-6604.

Schmidt:1989:SDD

[Sch89] Helmut Schmidt. A simple derivation of distribution functions for Bose and Fermi statistics. American Journal of Physics, 57(12): 1150–1151, December 1989. CODEN AJPIAS. ISSN 1943-2909.

Scheffer:1994:MIC

[Sch94] L. K. Scheffer. Machine intelligence, the cost of interstellar travel and Fermi’s Paradox. Quarterly Journal of the Royal Astronomical Society, 35(??):157, June 1994. CODEN QJRAAK. ISSN

0035-8738. URL http://adsabs.harvard.edu/abs/1994QJRAS.
.35..157S.

Schofield:1999:NFL

[Sch99a] A. J. Schofield. Non-Fermi liquids. Contemporary Physics, 40 (2):95–115, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Schucking:1999:JPP

[Sch99b] Engelbert L. Schucking. Jordan, Pauli, politics, Brecht, and a variable gravitational constant. Physics Today, 52(10):26–31, October 1999. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://adsabs.harvard.edu/abs/
1999PhT....52j..26S;   http://link.aip.org/link/phtoad/
v52/i10/p26/s1;   http://www.physicstoday.org/resource/
1/phtoad/v52/i10/p26_s1.

Schwarzschild:2000:BHS

[Sch00] Bertram Schwarzschild. Bohr–Heisenberg Symposium marks Broadway opening of Copenhagen. Physics Today, 53(5):51–52, May 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v53/i5/
p51/s1.

Schweber:2002:EFQ

[Sch02] Silvan S. Schweber. Enrico Fermi and quantum electrodynamics, 1929–32. Physics Today, 55(6):31–36, June 2002. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://www.physicstoday.org/resource/1/
phtoad/v55/i6/p31_s1.

Schweber:2003:FQE

[Sch03a] Sam Schweber. Fermi and quantum electrodynamics (QED). In Anonymous [Ano03], pages 185–216. ISBN 88-8286-032-9. LCCN

???? URL http:old.enea.it/produzione_scientifica/pdf.
../V2003_AttiFermi.pdf.

Schwinger:2003:SPQ

[Sch03b] Julian Schwinger. Selected Papers on Quantum Electrodynamics. Dover, New York, NY, USA, 2003. ISBN 0-486-60444-6. xvii + 424 pp. LCCN QC680 .S35. Reprint of [Sch58] with ISBN.

Schwarcz:2015:RCH

[Sch15] Joe Schwarcz. The right chemistry: How the Manhattan Project and ‘the bomb’ came to be. The Montreal Gazette, ??(??):??, May

10, 2015. ISSN 0384-1294. URL http://montrealgazette.com/
technology/science/the-right-chemistry-how-the-manhattan-
project-and-the-bomb-came-to-be.

Schwartz:2017:LMW

[Sch17] David N. Schwartz. The Last Man Who Knew Everything: the Life and Times of Enrico Fermi, Father of the Nuclear Age. Basic Books, New York, NY, USA, 2017. ISBN 0-465-07292-5 (hardcover), 0-465-09312-4 (e-book). xxiii + 453 pp. LCCN QC16.F46 S39 2017.

Sciuti:2001:EFD

[Sci01] Sebastiano Sciuti. Enrico Fermi divulgatore e persuasore scientifico. (Italian) [Enrico Fermi, popularizer and scientific persuader]. Il Nuovo Saggiatore, 17(5–6):23–26, September/ December 2, 2001. ISSN 0393-4578 (print), 1827-6148 (elec-

tronic). URL http://prometeo.sif.it:8080/papers/online/

sag/2001/05-06/pdf/05.pdf. Special issue in honor of the centennial of the birth of Enrico Fermi.

Sagan:1975:SEI

[SD75] Carl Sagan and Frank Drake. The search for extraterrestrial intelligence. Scientific American, 232(5):80–89, ???? 1975. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/
v232/n5/pdf/scientificamerican0575-80.pdf.

Sandberg:2018:DFP

[SDO18] Anders Sandberg, Eric Drexler, and Toby Ord. Dissolving the Fermi Paradox. arXiv.org, ??(??):??, January 6, 2018. URL

https://arxiv.org/abs/1806.02404.

Sebastiani:1997:SDC

[Seb97] F. Sebastiani. Storia e didattica: il caso della fisica quantistica. (Italian) [History and education: the case of quantum physics].

Giornale di Fisica, 38(??):145–150, ???? 1997. CODEN GFSIAD. ISSN 0017-0283 (print), 1827-6156 (electronic).

Segre:1955:FNP

[Seg55] Emilio Segrè. Fermi and neutron physics. Reviews of Modern Physics, 27(3):257–263, July 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.27.257;   http:/
/rmp.aps.org/abstract/RMP/v27/i3/p257_1.

Segre:1962:BI

[Seg62] Emilio Segrè. Biographical introduction. In Collected papers (Note e memorie). Vol. I: Italy, 1921–1938 [Fer62a], pages xvii–xlv.

Segre:1970:EFPa

[Seg70a] Emilio Segrè. Enrico Fermi: Physicist. Bulletin of the Atomic Scientists, 26(9):32, 37–39, November 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpts from [Seg70b].

Segre:1970:EFPb

[Seg70b] Emilio Segrè. Enrico Fermi: physicist. University of Chicago Press, Chicago, IL, USA, 1970. ISBN 0-226-74472-8. x + 276 pp. LCCN QC16.F46 S4.

Segre:1971:EFF

[Seg71] Emilio Segrè. Enrico Fermi: fisico: una biographia scientifica. (Italian) [Enrico Fermi: physicist: a scientific biography]. Zanichelli, Bologna, Italia, 1971. vi + 284 pp. LCCN QC16.F46 S416.

Segre:1976:PSN

[Seg76] Emilio Segrè. Personaggi e scoperte nella fisica contemporanea. (Italian) [Personalities and discoveries in contemporary physics]. Biblioteca della EST. Edizioni scientifiche e tecniche Mondadori, Milano, Italia, 1976. 297 pp. LCCN QC7 .S44. Ciclo di lezioni tenute dal nov. 1972 fino al marzo 1973.

Segre:1979:IPA

[Seg79] Emilio Segrè. Italian physics in Amaldi’s time. In Schaerf and Amaldi [SA79], pages 348–377. ISBN 3-7186-0007-2. LCCN QC770 .C748 1979.

Segre:1980:XRQ

[Seg80] Emilio Segrè. From X-rays to quarks: modern physicists and their discoveries. W. H. Freeman, New York, NY, USA, 1980. ISBN 0-7167-1146-X, 0-7167-1147-8 (paperback). ix + 337 pp. LCCN QC7 .S4413. English translation of [Seg76].

Segre:1983:PSN

[Seg83] Emilio Segrè. Personaggi e scoperte nella fisica contemporanea: dai raggi X ai quark. Biblioteca della EST, 0303-2752. Edizioni scientifiche e tecniche Mondadori (IS), Milano, Italia, second edition, 1983. 297 pp. LCCN ????

Segre:1984:PML

[Seg84] Emilio Segrè. Les physiciens modernes et leurs découvertes: des rayons X aux quarks. (French) [Modern physicists and their discoveries: from X-rays to quarks]. Temps des sciences. Fayard, Paris, France, 1984. 456 pp. LCCN QC7 .S4414 1984. French translation of [Seg76].

Segre:1985:HPR

[Seg85] Emilio Segrè. Historical perspective: Refugee scientists and nuclear energy. Annals of the New York Academy of Sciences, 452(1):xv–xix, 1985. CODEN ANYAA9. ISBN 0-89766-298-9, 0-89766-299-7 (paperback). ISSN 0077-8923 (print), 1749-6632 (electronic). Sixth International Conference on Collective Phenomena: reports from the Moscow Refusnik Seminar / edited by Inga Fischer-Hjalmars and Joel L. Lebowitz. Contributions from the Moscow Refusnik Seminar and from two International Conferences on Collective Phenomena, one held in Stockholm, Sweden, 1–2 December 1983, and the other in Tel Aviv, Israel, 31 May–1 June 1984.

Segre:1986:MKR

[Seg86a] Emilio Segre. Mi-karne Ranotgen ove-ad kvarkim: fisikaim modernim ove-tagliyotehe. Keter, Yerushalayim, Israel, 1986. 331 pp. LCCN ???? Hebrew translation of [Seg76].

Segre:1986:WLM

[Seg86b] Emilio Segrè. Wu li ming ren he wu li fa xian. Zhi shi chu ban she, Shanghai, People’s Republic of China, 1986. iii + 366 pp. LCCN QC7 .S4412. Mandarin Chinese translation by Zuwei Liu of [Seg76].

Segre:1987:RXA

[Seg87a] Emilio Segre. Dos raios X aos quarks: fisicos modernos e suas descobertas. (Portuguese) [From X-rays to quarks: modern physicists and their discoveries], volume 24 of Pensamento cientifico. Ed. UnB, Brasilia, Brazil, 1987. ISBN 85-230-0078-X. 345 pp. LCCN ????

Segre:1987:EFF

[Seg87b] Emilio Segrè. Enrico Fermi: fisico: una biographia scientifica. (Italian) [Enrico Fermi: physicist: a scientific biography]. Le ellissi. Zanichelli, Bologna, Italia, second edition, 1987. ISBN 88-08-02238-2. 316 pp. LCCN ????

Segre:1988:FSR

[Seg88] Emilio Segrè. The Fermi school in Rome. European Journal of Physics, 9(2):83–87, April 1988. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic). URL http://iopscience.

iop.org/0143-0807/9/2/001.

Segre:1993:MAM

[Seg93] Emilio Segrè. A mind always in motion: the autobiography of Emilio Segrè. University of California Press, Berkeley, CA, USA, 1993. ISBN 0-520-07627-3. xii + 332 pp. LCCN QC16.S35 A3

1993. URL http://ark.cdlib.org/ark:/13030/ft700007rb;
 http://www.loc.gov/catdir/description/ucal041/92010722.
html.

Segre:1996:FE

[Seg96] Emilio Segré. Fermi Enrico. In Dizionario Biografico degli Italiani, volume XLVI [Feducci–Ferrerio], page ?? Istituto dell’Enciclopedia Italiana, Roma, Italy, 1996. ISBN ???? LCCN ????

Segev:2003:FGR

[Seg03] Bilha Segev. Fermi’s golden rule in the Wigner representation. Journal of Optics B: Quantum and Semiclassical Optics, 5(3):S219–S220, 2003. CODEN JOBOFD. ISSN 1464-4266 (print), 1741-3575 (electronic). URL http://iopscience.iop.

org/1464-4266/5/3/373.

Segre:2007:XRQ

[Seg07] Emilio Segrè. From X-rays to quarks: modern physicists and their discoveries. Dover classics of science and mathematics. Dover, New York, NY, USA, 2007. ISBN 0-486-45783-4. ix + 339 pp.

LCCN QC7 .S4413 2007. URL http://www.loc.gov/catdir/
enhancements/fy0702/2006102450-d.html.

Segre:2009:EFF

[Seg09] Emilio Segrè. Enrico Fermi: fisico: una biographia scientifica. (Italian) [Enrico Fermi: physicist: a scientific biography]. Le ellissi. Zanichelli, Bologna, Italia, second (reprinted) edition, 2009. ISBN 88-08-26174-3. 316 + 20 pp. LCCN ????

Segre:2015:BRF

[Seg15] Gino Segrè. Book review: Frank Close: Half-Life: The Divided Life of Bruno Pontecorvo, Physicist or Spy. New York: Basic Books, 2015, xix + 366 pages. $29.99 (cloth). Physics in Perspective (PIP), 17(1):70–73, March 2015. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-015-0155-y.

Seitz:1955:FSA

[Sei55] Frederick Seitz. Fermi statistics and its applications. Reviews of Modern Physics, 27(3):249–253, July 1955. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
27.249;   http://rmp.aps.org/abstract/RMP/v27/i3/p249_
2.

Seidel:1990:BRB

[Sei90] Robert Seidel. Book reviews: Books on the bomb. Atomic Bomb Scientists: Memoirs, 1939-1945 by Joseph J. Ermenc. The End of the World That Was: Six Lives in the Atomic Age by Peter Goldman. Manhattan: The Army and the Atomic Bomb by Vincent C. Jones. Day of the Bomb: Countdown to Hiroshima by Dan Kurzman. The General and the Bomb: A Biography of General Leslie R. Groves, Director of the Manhattan Project by William Lawren. Time Bomb: Fermi, Heisenberg, and the Race for the Atomic Bomb by Malcolm C. MacPherson. The Making of the Atomic Age by Alwyn McKay. The Road to Trinity: A Personal Account of How America’s Nuclear Policies Were Made by K. D. Nichols. The Making of the Atomic Bomb by Richard Rhodes. Stallion Gate by Martin Cruz Smith. The Atomic Scientists: A Biographical History by Henry A. Boorse. Lloyd Motz, and Jefferson Hane Weaver. Forging the Atomic Shield: Excerpts from the Office Diary of Gordon E. Dean by Gordon E. Dean and Roger M. Anders. The Nuclear Oracles: A Political History of the General Advisory Committee of the Atomic Energy Commission, 1947-1977 by Richard T. Sylves. Better a Shield Than a Sword by Edward Teller. Klaus Fuchs, Atom Spy by Robert Chadwell Williams. Justice Downwind: America’s Atomic Testing Program in the 1950s by Howard Ball. The Atomic Papers: A Citizen’s Guide to Selected Books and Articles on the Bomb, the Arms Race, Nuclear Power, the Peace Movement, and Related Issues by Grant Burns. Physics, Technology and the Nuclear Arms Race by D. W. Hafemeister and D. Schroeer. Under the Cloud: The Decades of Nuclear Testing by Richard L. Miller. Bombs in the Backyard: Atomic Testing and American Politics by A. Costandina Titus. Nuclear Fear: A History of Images by Spencer R. Weart. Isis, 81(3):519–537, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233429.

Seidel:2003:EFH

[Sei03] Robert Seidel. Enrico Fermi, high-energy physics and high speed computing. In Anonymous [Ano03], pages 259–268. ISBN 88-

8286-032-9. LCCN ???? URL http:old.enea.it/produzione_
scientifica/pdf.../V2003_AttiFermi.pdf.

Seife:2004:ECC

[Sei04] Charles Seife. Energy curve confirms paired-up Fermi condensate. Science, 305(5683):459–460, July 23, 2004. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.jstor.org/stable/3837434.

Selove:1953:SCE

[Sel53] Walter Selove. Summary of communication by Enrico Fermi on meson physics. Proceedings of the American Academy of Arts and Sciences, 82(7):350–351, December 1953. CODEN PAAAAV. ISSN 0065-6836. URL http://www.jstor.org/

stable/20023737.

Service:2012:PDN

[Ser12] Robert F. Service. Physicists discover new type of particle—sort of. Science Now, ??(??):??, April 12, 2012. URL http://news.

sciencemag.org/sciencenow/2012/04/majorana-fermions-found.
html.

Szilard:1936:CNC

[SF44] Leo Szilard and Enrico Fermi. [correspondence on nuclear chain reactions], 1936–1944. URL http://library.ucsd.edu/

dc/object/bb6929916v.

Segre:2016:PPE

[SH16] Gino Segrè and Bettina Hoerlin. The Pope of Physics: Enrico Fermi and the birth of the atomic age. Henry Holt and Company, New York, NY, USA, 2016. ISBN 1-62779-005-5, 1-62779-006-3 (e-book). xi + 351 pp. LCCN QC16.F46 S44 2016.

Shankland:1967:BRC

[Sha67] R. S. Shankland. Book review: The Collected Papers of Enrico Fermi. vol. II: US 1939–1954. Technology and Culture, 8(1):119– 120, January 1967. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL http://www.jstor.org/stable/
3101549;   https://muse.jhu.edu/pub/1/article/894513/

pdf. See [Fer65b].

Shirokov:1981:FFP

[Shi81] M. Shirokov. Fermi–Ferretti problem and signal velocity. Foundations of Physics, 11(1–2):21–36, February 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00715193.

Shils:1991:RUC

[Shi91] Edward Shils, editor. Remembering the University of Chicago: teachers, scientists, and scholars. University of Chicago Press, Chicago, IL, USA, 1991. ISBN 0-226-75335-2. xxi + 593 pp. LCCN LD920 .R46 1991. URL http://www.loc.

gov/catdir/description/uchi051/91016741.html;   http:/
/www.loc.gov/catdir/enhancements/fy0608/91016741-t.html.

Schwalbe:2018:FLO

[SHL^{+}18] Sebastian Schwalbe, Torsten Hahn, Simon Liebing, Kai Trepte, and Jens Kortus. Fermi–Löwdin orbital self-interaction corrected density functional theory: Ionization potentials and enthalpies of formation. Journal of Computational Chemistry, 39(29):2463– 2471, November 5, 2018. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Simons:2019:TMM

[SHM19] S. Stoney Simons and Jacob D. Haqq-Misra. A trip to the moon might constrain the Fermi Paradox. Futures, 106:24–32, February 2019. CODEN FUTUBD. ISSN 0016-3287 (print), 1873-6378 (electronic).

Shostak:2009:CAH

[Sho09] G. Seth Shostak. Confessions of an alien hunter: a scientist’s search for extraterrestrial intelligence. National Geographic, Washington, DC, USA, 2009. ISBN 1-4262-0392-6. x + 309 pp. LCCN QB54 .S549 2009. URL http://www.loc.gov/

catdir/enhancements/fy0910/2008046731-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0910/2008046731-d.
html;   http://www.loc.gov/catdir/toc/fy0906/2008046731.
html.

Sime:1989:LMD

[Sim89] Ruth Lewin Sime. Lise Meitner and the discovery of fission. Journal of Chemical Education, 66(5):373–375, May 1989. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic). URL

http://pubs.acs.org/doi/abs/10.1021/ed066p373.

Sime:1996:LML

[Sim96] Ruth Lewin Sime. Lise Meitner: a life in physics, volume 13 of California studies in the history of science. University of California Press, Berkeley, CA, USA, 1996. ISBN 0-520-08906-5, 0-520-20860-9. xiii + 526 pp. LCCN QC774.M4 S56 1996.

URL http://www.loc.gov/catdir/bios/ucal051/95035246.
html;   http://www.loc.gov/catdir/description/ucal041/
95035246.html.

Sime:2000:STE

[Sim00] Ruth Lewin Sime. The search for transuranium elements and the discovery of nuclear fission. Physics in Perspective (PIP), 2(1): 48–62, March 2000. CODEN PHPEF2. ISSN 1422-6944 (print),

1422-6960 (electronic). URL http://www.springerlink.com/
content/nr2t13tndn6t9t72/.

Sime:2003:FFM

[Sim03] Ruth Lewin Sime. From Fermi to fission: Meitner, Hahn and Strassmann in Berlin. In Anonymous [Ano03], pages 133–144. ISBN 88-8286-032-9. LCCN ???? URL http:old.enea.it/

produzione_scientifica/pdf.../V2003_AttiFermi.pdf.

Sime:2010:IHN

[Sim10] Ruth Lewin Sime. An inconvenient history: the nuclearfission display in the Deutsches Museum. Physics in Perspective (PIP), 12(2):190–218, June 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://www.

springerlink.com/content/w716842562715257/.

Sime:2012:PFO

[Sim12] Ruth Lewin Sime. The politics of forgetting: Otto Hahn and the German Nuclear-Fission Project in World War II. Physics in Perspective (PIP), 14(1):59–94, March 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-011-0065-6;
http://www.springerlink.com/content/k12202vg92147h68/.

Six:1988:PNB

[Six88] Jules Six. Pourquoi ni Bothe ni les Joliot–Curie n’ont découvert le neutron. (French) [Why neither [Walther] Bothe nor [Pierre] Joliot–Curie discovered the neutron]. Revue d’Histoire des Sciences, 41(1):3–24, January 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.

org/stable/23632789.

Stoddart:1967:ETF

[SM67] J. C. Stoddart and N. H. March. Exact Thomas–Fermi method in perturbation theory. Proceedings of the Royal Society of London. Series A, Mathematical and Physical Sciences, 299(1457): 279–286, June 14, 1967. CODEN PRLAAZ. ISSN 0080-4630.

URL http://www.jstor.org/stable/2415728.

Schilling:2009:CFA

[SM09] Govert Schilling and Alan M. MacRobert. The chance of finding aliens. Sky and Telescope, ??(??):??, June 3, 2009. CODEN

SKTEA3. ISSN 0037-6604. URL https://skyandtelescope.
org/astronomy-news/the-chance-of-finding-aliens/.

Smith:1968:BRT

[Smi68] Alice Kimball Smith. Book review: The transplantation of European intellectuals: Illustrious Immigrants: The Intellectual Migration From Europe 1930–41, Laura Fermi. Science, 160(3828):636–638, May 10, 1968. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

jstor.org/stable/1724702;   http://www.sciencemag.org/
content/160/3828/636.full.pdf.

Seaborg:1946:REDa

[SMKW46] Glenn T. Seaborg, Edwin M. Mcmillan, J. W. Kennedy, and Arthur C. Wahl. Radioactive element 94 from deuterons on uranium. Physical Review, 69(7–8):366–367, April 1946. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://physics.aps.org/story/v14/st17;
http://prola.aps.org/abstract/PR/v69/i7-8/p366_2. This

first paper on the discovery of plutonium was submitted January 28, 1942, but withheld from publication until after the end of World War II. See also [SWK46].

Smyth:1945:AEMb

[Smy45a] Henry DeWolf Smyth. Atomic energy for military purposes. Reviews of Modern Physics, 17(4):351–471, October 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
17.351;   http://rmp.aps.org/abstract/RMP/v17/i4/p351_
1.

Smyth:1945:AEMa

[Smy45b] Henry DeWolf Smyth. Atomic energy for military purposes; the official report on the development of the atomic bomb under the auspices of the United States Government, 1940–1945. Princeton University Press, Princeton, NJ, USA, 1945. ix + 1 + 264 pp. LCCN QC173 .S4735 1945a.

Sagan:1983:SAE

[SN83] Carl Sagan and William Newman. The solipsist approach to extraterrestrial intelligence. Quarterly Journal of the Royal Astronomical Society, 24(113):??, ???? 1983. CODEN QJRAAK. ISSN 0035-8738.

Solomon:1968:BRA

[Sol68] Eric Solomon. Book reviews: The American Writer and the Great Depression, by Harvey Swados; Years of Protest, by Jack Salzman; Hard-Hitting Songs for Hard-Hit People, by Alan Lomax; Just Around the Corner, by Robert Bendiner; All the Things We Were, by Louise Tanner; The Invisible Scar, by Caroline Bird; Illustrious Immigrants, by Laura Fermi; As We Saw the Thirties, by Rita Simons; The Thirties, by Morton J. Frisch; Martin Diamond; Proletarian Writers of the Thirties, by David Madden; Tough Guy Writers of the Thirties, by David Madden; The Thirties: Fiction, Poetry, Drama, by Warren French; The Last Great Cause, by Stanley Weintraub; Writers in Arms, by Frederick Benson; Writers and Partisans, by James Gilbert. American Quarterly, 20(4):810–819, 1968. ISSN 0003-0678 (print), 1080-6490

(electronic). URL http://www.jstor.org/stable/2711410.

Solovej:2003:EFS

[Sol03] Jan Philip Solovej. The evolution of Fermi’s statistical theory of atoms. In Anonymous [Ano03], pages 97–104. ISBN 88-

8286-032-9. LCCN ???? URL http:old.enea.it/produzione_
scientifica/pdf.../V2003_AttiFermi.pdf.

Sommerfeld:1927:EMV

[Som27] Arnold Sommerfeld. Elektronentheorie der Metalle und des Volta-Effektes nach der Fermi’schen Statistik. (German) [Electron theory of metals and the Volta effect according to Fermi’s statistics]. In ????, editor, Atti Congr. Intern. dei Fisici Como–Pavia–Roma, Sept. 1927, volume 2, pages 449–473. Zanichelli, Bologna, Italy, 1927.

Sommerfeld:1928:EMGb

[Som28a] Arnold Sommerfeld. Zur Elektronentheorie der Metalle auf Grund der Fermischen Statistik. I. Teil: Allgemeines, Strömungs- und Austrittsvorgänge. (German) [On the electron theory of metals on the basis of Fermi statistics. Part I: General, flow and discharge operations]. Zeitschrift für Physik, 47(1–2):1–32, January 1928. CODEN ZEPYAA. ISSN 0044-3328. URL http://link.

springer.com/article/10.1007/BF01391052.

Sommerfeld:1928:EMGc

[Som28b] Arnold Sommerfeld. Zur Elektronentheorie der Metalle auf Grund der Fermischen Statistik. II. Teil: Thermo-elektrische, galvanomagnetische und thermo-magnetische Vorgänge. (German) [On the electron theory of metals on the basis of Fermi statistics. Part II: Thermo-electric, electro-magnetic and thermo-magnetic processes]. Zeitschrift für Physik, 47(1–2):43–60, January 1928. CO-DEN ZEPYAA. ISSN 0044-3328. URL http://link.springer.

com/article/10.1007/BF01391055.

Sommerfeld:1932:AID

[Som32a] Arnold Sommerfeld. Asymptotische Integration der Differentialgleichung des Thomas-Fermischen Atoms. (German) [Asymptotic integration of the differential equation of the Thomas–Fermi atom]. Zeitschrift für Physik, 78(5–6):283–308, May 1932. CO-DEN ZEPYAA. ISSN 0044-3328. URL http://link.springer.

com/article/10.1007/BF01342197.

Sommerfeld:1932:IAD

[Som32b] Arnold Sommerfeld. Integrazione asintotica dell’equazione differenziale di Thomas–Fermi. (Italian) [Asymptotic integration of the Thomas–Fermi differential equation]. Rendiconti dell’Accademia Nazionale dei Lincei, 15(??):293–308, ???? 1932. CODEN AANLAW. ISSN 0001-4435.

Sommerfeld:1933:HIA

[Som33] Arnold Sommerfeld. Über die höheren Ionisierungsspannungen der Atome im Thomas–Fermischen Modell. (German) [On the higher ionization voltages of the atoms in the Thomas–Fermi model]. Zeitschrift für Physik, 80(7–8):415–422, February 6, 1933. CODEN ZEPYAA. ISSN 0044-3328. URL http://link.

springer.com/article/10.1007/BF02057304.

Sopka:1976:QPA

[Sop76] Katherine Russell Sopka. Quantum physics in America, 1920– 1935. Ph.D. in Education thesis, Graduate School of Arts and Sciences, and the Graduate School of Education, Harvard University, Cambridge, MA, USA, May 1976. various pp. URL

http://hollis.harvard.edu/?itemid=%7Clibrary/m/aleph%
7C000979886;   http://search.proquest.com/docview/302811033.

Reprinted in [Sop80], and revised in [Sop88].

Sopka:1980:QPA

[Sop80] Katherine Russell Sopka. Quantum physics in America, 1920– 1935. Three centuries of science in America. Arno Press, New York, NY, USA, 1980. ISBN 0-405-12585-2. xvi + 543 pp. LCCN QC173.98 .S66 1980.

Sopka:1988:QPA

[Sop88] Katherine Russell Sopka. Quantum physics in America: the years through 1935, volume 10 of The history of modern physics, 1800-1950. Tomash, Los Angeles, CA, USA, 1988. ISBN 0-88318-553-9. xxvi + 376 pp. LCCN QC173.98 .S67 1988. Foreword by Abraham Pais. Introduction by Gerald Holton.

Seaborg:1948:SEN

[SP48] Glenn T. Seaborg and Morris L. Perlman. Search for elements 94 and 93 in nature. presence of 94^{239} in pitchblende1. Journal of the American Chemical Society, 70(4):1571–1573, April 1948. CODEN JACSAT. ISSN 0002-7863 (print), 1520-5126

(electronic), 1943-2984. URL http://pubs.acs.org/doi/abs/
10.1021/ja01184a083.

Spruch:1991:PNT

[Spr91] Larry Spruch. Pedagogic notes on Thomas–Fermi theory (and on some improvements): atoms, stars, and the stability of bulk matter. Reviews of Modern Physics, 63(1):151–209, January 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.63.151;   http://rmp.aps.org/abstract/
RMP/v63/i1/p151_1.

Serber:1992:APF

[SR92] R. (Robert) Serber and Richard Rhodes. The Los Alamos primer: the first lectures on how to build an atomic bomb. University of California Press, Berkeley, CA, USA, 1992. ISBN 0-520-07576-5. xxxiii + 98 + 8 pp. LCCN QC773.A1 S47 1992. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http://
www.loc.gov/catdir/bios/ucal051/91014068.html;   http:
//www.loc.gov/catdir/description/ucal041/91014068.html.

Shklovskii:1977:ILU

[SS77] I. S. Shklovskiĭ and Carl Sagan. Intelligent life in the universe. Picador. Pan Books, London, UK, 1977. ISBN 0-330-25125-2 (paperback). xiv + 509 pp. LCCN QB54 .S523 1977.

Sassi:1999:FSE

[SS99] M. C. Sassi and F. Sebastiani. La formazione scientifica di Enrico Fermi. (Italian) [The scientific formation of Enrico Fermi]. Giornale di Fisica, 40(??):89–113, ???? 1999. CODEN GFSIAD. ISSN 0017-0283 (print), 1827-6156 (electronic).

Sidje:2011:RAF

[SS11] Roger B. Sidje and Yousef Saad. Rational approximation to the Fermi–Dirac function with applications in density functional theory. Numerical Algorithms, 56(3):455–479, March 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=1017-1398&volume=56&issue=3&spage=455.

Sudoplatov:1995:STM

[SSSS95] Pavel Sudoplatov, Anatolii Pavlovich Sudoplatov, Jerrold L. Schecter, and Leona Schecter. Special tasks: the memoirs of an unwanted witness, a Soviet spymaster. Little, Brown and Co., Boston, MA, USA, updated edition, 1995. ISBN 0-316-82115-2. xxxi + 527 pp. LCCN JN6529.I6 S83 1995.

Stewart:1959:EFA

[Ste59] George W. Stewart. Enrico Fermi and the Atomic Age (“… the first man in all the world to achieve nuclear chain reaction”). Pamphlet., 1959.

Stedman:1971:FGR

[Ste71] G. E. Stedman. Fermi’s Golden Rule — an exercise in quantum field theory. American Journal of Physics, 39(2):205–214, February 1971. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://ajp.aapt.org/resource/1/ajpias/
v39/i2/p205_s1.

Stephenson:1977:FLI

[Ste77] D. G. Stephenson. Factors limiting the interaction between twentieth century man and interstellar cultures. Journal of the British Interplanetary Society, 30(??):105–108, ???? 1977. CODEN JBI-SAW. ISSN 0007-084X.

Stewart:1993:BRF

[Ste93] A. B. Stewart. Book review: The Fermi Solution, by Hans Christian von Baeyer. The Antioch Review, 51(4):649, Autumn 1993.

ISSN 0003-5769. URL http://www.jstor.org/stable/4612847.

Steinberger:2001:EFM

[Ste01] Jack Steinberger. Enrico Fermi, my master and teacher. Il Nuovo Saggiatore, 17(5–6):38–40, September/December 2, 2001. ISSN 0393-4578 (print), 1827-6148 (electronic). URL http://

prometeo.sif.it:8080/papers/online/sag/2001/05-06/pdf/

05.pdf. Special issue in honor of the centennial of the birth of Enrico Fermi.

Stimson:1947:DUB

[Sti47] H. L. Stimson. The decision to use the bomb. Harper’s Magazine, 194(2):97–107, February 1947. CODEN HAMAA3. ISSN

1045-7143. URL http://www.harpers.org/archive/1947/02/

0032863. Reprinted in [Sti76]. See also [Lef95] for a contrary view based on a half-century of historical scholarship, and access to previously-secret US documents.

Stimson:1976:DUB

[Sti76] H. L. Stimson. The decision to use the bomb. In Baker [Bak76], pages 14–28. ISBN 0-03-089873-0. LCCN ???? Reprinted from [Sti47]. See also [Lef95] for a contrary view based on a half-century of historical scholarship, and access to previously-secret US documents.

Strauss:1963:F

[Str63a] Lewis L. Strauss. Fermi. In Men and decisions [Str63b], pages 236–241. LCCN E741 .S78.

Strauss:1963:MDa

[Str63b] Lewis L. Strauss. Men and decisions. Doubleday, Garden City, NY, USA, 1963. viii + 468 + 16 pp. LCCN E741 .S78.

Strauss:1963:MDb

[Str63c] Lewis L. Strauss. Men and decisions. Macmillan Publishing Company, New York, NY, USA, 1963. 468 pp. LCCN ????

Strogatz:1998:BRT

[Str98] Steven H. Strogatz. Book review: Thomas P. Weissert, The genesis of simulation in dynamics: Pursuing the Fermi–Pasta–Ulam problem Lattice-gas cellular automata: Simple models of complex hydrodynamics. Computers in Physics, 12(6):574–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208

(electronic). URL https://aip.scitation.org/doi/10.1063/
1.168754.

Stuewer:1979:NPR

[Stu79] Roger H. Stuewer, editor. Nuclear physics in retrospect: proceedings of a symposium on the 1930s. University of Minnesota Press, Minneapolis, MN, USA, 1979. ISBN 0-8166-0869-5. LCCN QC773 .S95 1977.

Stuewer:1984:NPN

[Stu84] Roger H. Stuewer. Nuclear physicists in a new world. The Émigrés of the 1930s in America. Berichte zur Wissenschaftsgeschichte, 7(1):23–40, ???? 1984. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/bewi.19840070104/abstract.

Stuewer:1985:BNF

[Stu85] Roger H. Stuewer. Bringing the news of fission to America. Physics Today, 38(10):48–56, October 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

www.physicstoday.org/resource/1/phtoad/v38/i10/p48_s1.

Stuewer:1986:ND

[Stu86] Roger H. Stuewer. The naming of the deuteron. American Journal of Physics, 54(3):206–218, March 1986. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://

ajp.aapt.org/resource/1/ajpias/v54/i3/p206_s1.

Stuewer:1993:MEN

[Stu93] Roger H. Stuewer. Mass-energy and the neutron in the Early Thirties. Science in Context, 6(1):195–238, Spring 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Stuewer:2006:BRC

[Stu06] Roger H. Stuewer. Book review: Carlo Bernardini and Luisa Bonolis, ed., Enrico Fermi: His Work and Legacy. Bologna: Società Italiana di Fisica and Berlin, Heidelberg, New York: Springer Verlag, 2004, xii + 411 pages. $49.95 (cloth). Physics in Perspective (PIP), 8(1):104–105, March 2006. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Stuewer:2018:AIN

[Stu18] Roger H. Stuewer. The Age of Innocence: Nuclear Physics Between the First and Second World Wars. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2018. ISBN 0-19-186658-X, 0-19-882787-3 (hardback), 0-19-256290-8 (e-book). xv + 484 pp. LCCN QC773 .S78 2018.

Sumner:1965:EF

[Sum65] W. L. Sumner. Enrico Fermi. Nature, 207(5000):900, August 28, 1965. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v207/n5000/pdf/207900a0.pdf.

Sutton:1992:SN

[Sut92] Christine Sutton. Spaceship Neutrino. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-36404-3 (hardcover), 0-521-36703-4 (paperback). xiv + 244 pp. LCCN QC793.5.N42 S88 1992. URL http://www.loc.gov/

catdir/description/cam021/92004215.html;   http://www.
loc.gov/catdir/samples/cam034/92004215.html;   http://
www.loc.gov/catdir/toc/cam021/92004215.html.

Seaborg:1948:CPE

[SW48] Glenn T. Seaborg and Arthur C. Wahl. The chemical properties of elements 94 and 93. Journal of the American Chemical Society, 70 (3):1128–1134, March 1948. CODEN JACSAT. ISSN 0002-7863 (print), 1520-5126 (electronic), 1943-2984. URL http://pubs.

acs.org/doi/abs/10.1021/ja01183a076.

Sagan:1966:IDD

[SW66] Carl Sagan and Russell G. Walker. The infrared detectability of Dyson civilizations. Astrophysical Journal, 144(??):1216–1218, ???? 1966. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357

(electronic). URL http://adsabs.harvard.edu/abs/1966ApJ.

..144.1216S. See [Dys60].

Salam:1972:AQT

[SW72] Abdus Salam and Eugene Paul Wigner, editors. Aspects of quantum theory. Cambridge University Press, Cambridge, UK, 1972. ISBN 0-521-08600-0. xvi + 268 pp. LCCN QC174.1

.A85 1972. URL http://hooke.lib.cam.ac.uk/cgi-bin/
bib_seek.cgi?cat=ul&bib=1733506;   http://www.loc.gov/
catdir/enhancements/fy1001/72075298-d.html;   http://
www.loc.gov/catdir/enhancements/fy1001/72075298-t.html.

Siedentop:1987:LEC

[SW87] Heinz Siedentop and Rudi Weikard. On the leading energy correction for the statistical model of the atom: interacting case. Communications in Mathematical Physics, 112(3):471–490, September 1987. CODEN CMPHAY. ISSN 0010-3616 (print), 1432-

0916 (electronic). URL http://link.springer.com/article/
10.1007/BF01218487.

Swing:1945:RSB

[Swi45] Raymond Swing. Raymond Swing’s broadcast[: History of atomic bomb]. Script for radio broadcast on station WMAL on ABC network., September 7, 1945. URL http://library.ucsd.edu/

dc/object/bb0718253r.

Seaborg:1946:REDb

[SWK46] Glenn T. Seaborg, Arthur C. Wahl, and J. W. Kennedy. Radioactive element 94 from deuterons on uranium. Physical Review, 69(7-8):367, April 1946. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v69/i7-8/p367_1. Received 7 March 1941,

but withheld from publication until the end of World War II. See [SMKW46].

Szilard:1939:IEF

[SZ39] Leo Szilard and Walter H. Zinn. Instantaneous emission of fast neutrons in the interaction of slow neutrons with uranium. Physical Review, 55(8):799–800, April 15, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http:/

/adsabs.harvard.edu/abs/1939PhRv...55..799S;   http://
prola.aps.org/abstract/PR/v55/i8/p799_1.

Stephen:1962:CAI

[SZ62] M. J. Stephen and K. Zalewski. On the classical approximation involved in the Thomas–Fermi theory. Proceedings of the Royal Society of London. Series A, Mathematical and Physical Sciences, 270(1342):435–442, November 27, 1962. CODEN PRLAAZ. ISSN

0080-4630. URL http://www.jstor.org/stable/2414544.

Szasz:1992:BSM

[Sza92] Ferenc Morton Szasz. British scientists and the Manhattan Project: the Los Alamos years. St. Martin’s Press, New York, NY, USA, 1992. ISBN 0-312-06167-6. xx + 167 pp. LCCN QC773.3.U5

S97 1991. URL http://www.loc.gov/catdir/enhancements/
fy0809/91019904-d.html.

Szilard:1971:CLC

[Szi71] Gertrud Weiss Szilard. Communications: Lunar craters. Bulletin of the Atomic Scientists, 27(4):4, April 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Talas:2017:BRG

[Tal17] Sofia Talas. Book review: Giuseppe Bruzzaniti. Enrico Fermi: The Obedient Genius. Isis, 108(2):473–475, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tamm:1934:EFB

[Tam34] Ig. Tamm. Exchange forces between neutrons and protons, and Fermi’s theory. Nature, 133(3374):981, June 30, 1934. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v133/
n3374/pdf/133981a0.pdf.

Tarter:2003:BRO

[Tar03] Jill Tarter. Book review: Ongoing debate over cosmic neighbors: If the Universe Is Teeming with Aliens … Where Is Everybody?: Fifty Solutions to the Fermi Paradox and the Problem of Extraterrestrial Life by Stephen Webb. Science, 299(5603):46–47, January 3, 2003. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/3833213. See

[Web02].

Taub:1963:JNCb

[Tau63] A. H. Taub, editor. John von Neumann: Collected Works. Volume VI: Theory of Games, Astrophysics, Hydrodynamics and Meteorology. Pergamon, New York, NY, USA, 1963. x + 538 pp. LCCN ????

Teller:1957:NNWb

[Tel57] Edward Teller. The nature of nuclear warfare. Bulletin of the Atomic Scientists, 13(5):162–165, May 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from Air Force Magazine, January 1957. See comment [Fer57c].

Teller:1987:LL

[Tel87] E. Teller. The lunar laboratory. In Metropolis et al. [MKR87], pages 77–85. ISBN 0-12-492155-8. LCCN QC44 .N49

1987. URL http://www.osti.gov/energycitations/product.
biblio.jsp?osti_id=6120718&query_id=0.

Telegdi:1991:EF

[Tel91] Valentine L. Telegdi. Enrico Fermi. In Shils [Shi91], chapter 9, pages 110–129. ISBN 0-226-75335-2. LCCN LD920

.R46 1991. URL http://www.loc.gov/catdir/description/
uchi051/91016741.html;   http://www.loc.gov/catdir/enhancements/
fy0608/91016741-t.html.

Telegdi:2001:REF

[Tel01] Valentine Telegdi. Ricordi di Enrico Fermi. (Italian) [Memories of Enrico Fermi]. Il Nuovo Saggiatore, 17(5–6):27–30, September/December 2, 2001. ISSN 0393-4578 (print), 1827-6148 (elec-

tronic). URL http://prometeo.sif.it:8080/papers/online/

sag/2001/05-06/pdf/05.pdf. Special issue in honor of the centennial of the birth of Enrico Fermi.

Telegdi:2002:EFA

[Tel02] Valentine L. Telegdi. Enrico Fermi in America. Physics Today, 55(6):38–43, June 2002. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v55/i6/p38_s1.

Telegdi:2004:FC

[Tel04] Valentine Telegdi. Fermi at Chicago. In Orear [Ore04], chapter 20, pages 89–100. ISBN ???? LCCN ???? URL http:

//dspace.library.cornell.edu/handle/1813/62;   http://
hdl.handle.net/1813/74.

Thomas:1927:CAF

[Tho27] Llewellyn H. Thomas. The calculation of atomic fields. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 23(5):542–548, January 1927. CODEN PCPSA4. ISSN 0008-1981. This is the Thomas side of the Thomas–Fermi (or Fermi–Thomas) model. See [Fer27e] for the Fermi side.

Tipler:1980:EBD

[Tip80] Frank J. Tipler. Extraterrestrial beings do not exist. Quarterly Journal of the Royal Astronomical Society, 21(??):267–281, August 1980. CODEN QJRAAK. ISSN 0035-8738.

Tipler:1981:BHE

[Tip81a] Frank J. Tipler. A brief history of the extraterrestrial intelligence concept. Quarterly Journal of the Royal Astronomical Society, 22 (??):133–??, June 1981. CODEN QJRAAK. ISSN 0035-8738.

Tipler:1981:EIBa

[Tip81b] Frank J. Tipler. Extraterrestrial intelligent beings do not exist. Quarterly Journal of the Royal Astronomical Society, 22(??):267– 281, September 1981. CODEN QJRAAK. ISSN 0035-8738.

Tipler:1981:EIBb

[Tip81c] Frank J. Tipler. Extraterrestrial intelligent beings do not exist. Physics Today, 34(4):9, 70–71, April 1981. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Treder:1976:GUF

[Tre76] Hans-Jürgen Treder. Gravitation and universal Fermi coupling in general relativity. Foundations of Physics, 6(5):527–538, October 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/
10.1007/BF00715106.

Trepte:2019:AAG

[TSH^{+}19] Kai Trepte, Sebastian Schwalbe, Torsten Hahn, Jens Kortus, Der-You Kao, Yoh Yamamoto, Tunna Baruah, Rajendra R. Zope, Kushantha P. K. Withanage, Juan E. Peralta, and Koblar A. Jackson. Analytic atomic gradients in the Fermi–Löwdin orbital self-interaction correction. Journal of Computational Chemistry, 40(6):820–825, March 5, 2019. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Turchetti:2006:SNS

[Tur06a] Simone Turchetti. “For slow neutrons, slow pay”: Enrico Fermi’s patent and the U.S. Atomic Energy Program, 1938–1953. Isis, 97 (1):1–27, March 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
10.1086/501097;   http://www.jstor.org/stable/pdfplus/
10.1086/501097.pdf.

Turchetti:2006:IBN

[Tur06b] Simone Turchetti. The invisible businessman: Nuclear physics, patenting practices, and trading activities in the 1930s. Historical Studies in the Physical and Biological Sciences, 37(1):153–172, September 2006. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Turchetti:2008:BRV

[Tur08] Simone Turchetti. Book review: Valeria Del Gamba, Il Ragazzo di Via Panisperna. l’avventurosa vita del fisico Franco Rasetti. Torino: Bollati Boringhieri, 2007. 167 pp., ISBN 978-88-339-1746-7. Nuncius, 23(1):174–175, ???? 2008. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00427.

Turchetti:2012:PAC

[Tur12] Simone Turchetti. The Pontecorvo affair: a cold war defection and nuclear physics. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-81664-8 (hardcover). 292 pp. LCCN QC774.P66 T8713 2012.

Ulam:1955:HF

[Ula55] Stanisław M. Ulam. Homage to Fermi. Santa Fe New Mexican, ??(??):??, January 6, 1955.

Valdevit:2006:OFS

[Val06] Giampaolo Valdevit. Oppenheimer fra scienza e potere. una storia americana. (Italian) [Oppenheimer between science and power. An American story]. Studi Storici, 47(1):115–142, January/March 2006. ISSN 0039-3037 (print), 2036-458X (electronic). URL http:

//www.jstor.org/stable/20567340.

VanDeMark:2003:PKN

[Van03] Brian VanDeMark. Pandora’s keepers: nine men and the atomic bomb. Little, Brown and Co., Boston, MA, USA, 2003. ISBN 0-316-73833-6. xii + 399 + 16 pp. LCCN QC774.A2 V36 2003. URL

http://www.loc.gov/catdir/toc/fy045/2002043646.html.

vonBaeyer:1993:FSE

[vB93] Hans Christian von Baeyer. The Fermi solution: essays on science. Random House, New York, NY, USA, 1993. ISBN 0-679-40031-1. 172 pp. LCCN Q158.5 .V66 1993.

US $19.00; CAN $25.00. URL http://www.loc.gov/catdir/
enhancements/fy0904/92056853-d.html.

Vakoch:2015:DEE

[VD15] Douglas A. Vakoch and Matthew F. Dowd, editors. The Drake Equation: Estimating the Prevalence of Extraterrestrial Life Through the Ages, volume 8 of Cambridge astrobiology. Cambridge University Press, Cambridge, UK, 2015. ISBN 1-107-07365-0 (hardcover), 1-107-42326-0, 1-139-68359-4 (e-book). xxii + 319

pp. LCCN QB54 .D73 2015. URL http://assets.cambridge.
org/97811070/73654/cover/9781107073654.jpg.

VergaraCaffarelli:2001:EFA

[Ver01a] Roberto Vergara Caffarelli. Enrico Fermi al Liceo Umberto I di Roma e all’Università di Pisa. (Italian) [Enrico Fermi at the Umberto I School in Rome and at the University of Pisa]. Il Nuovo Saggiatore, 17(5–6):8–15, September/December 2, 2001. ISSN 0393-4578 (print), 1827-6148 (electronic). URL http://

prometeo.sif.it:8080/papers/online/sag/2001/05-06/pdf/

05.pdf. Special issue in honor of the centennial of the birth of Enrico Fermi.

VergaraCaffarelli:2001:EFI

[Ver01b] Roberto Vergara Caffarelli, editor. Enrico Fermi: immagini e documenti. (Italian) [Enrico Fermi: images and documents]. Limonaia, Pisa, Italia, 2001. ISBN 88-8492-092-2. 111 pp. LCCN QC16.F46 E56 2001.

VergaraCaffarelli:2002:EFI

[Ver02] Roberto Vergara Caffarelli. Enrico Fermi, Immagini e Documenti inediti. (Italian) [Enrico Fermi, Unpublished Pictures and Documents]. La Limonaia and Edizioni Plus, Università di Pisa, Pisa, Italy, 2002. ISBN ???? ???? pp. LCCN ???? Catalogue of the Exhibition at Limonaia di Palazzo Ruschi, Pisa, October 18–28, 2001.

vonHippel:2001:WFS

[vH01] Frank N. von Hippel. Where Fermi stood. Bulletin of the Atomic Scientists, 57(5):26–29, September 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/57/5/26.full.

vonHippel:2012:NPT

[vHEGM12] Frank von Hippel, Rodney Ewing, Richard Garwin, and Allison Macfarlane. Nuclear proliferation: Time to bury plutonium. Nature, 485(7397):167–168, May 10, 2012. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v485/

n7397/full/485167a.html. See [MB00] for the history of plutonium production.

vonNeumann:1996:PJN

[vN96] John von Neumann. Papers of John von Neumann, 1912–1996 (bulk 1935–1957). US Library of Congress archival manuscript material (collection)., 1996. 11,660 items. 34 containers plus 1 vault container. 13.4 linear feet. Manuscript number MSS44180. Correspondence, memoranda, journals, speeches, article and book drafts, notes, charts, graphs, patent, biographical material, family papers, printed materials, newspaper clippings, photographs, and other materials pertaining primarily to von Neumann’s career as professor of mathematics at the Institute for Advanced Study including his directorship of the Electronic Computer Project; adviser and commissioner on the U.S. Atomic Energy Commission; scientific consultant to government and private concerns, including the Los Alamos Scientific Laboratory, Los Alamos, New Mexico, and the U.S. Army Ballistic Research Laboratory, Aberdeen, Maryland; and author of works on ballistic research, computers, continuous geometries, logic, operator theory, quantum mechanics, and the theory of games. Includes evaluations of his work written after his death by colleagues including Herman Heine Goldstine, Paul R. Halmos, and Abraham H. Taub. Of special interest are an Albert Einstein letter and report on theoretical physics (1937). Also includes a small amount of material pertaining to Eva and Peter Aldor. Correspondents include Eva Aldor, Frank Aydelotte, Hans Albrecht Bethe, Garrett Birkhoff, S. Chandrasekhar, George Bernard Dantzig, P. A. M. Dirac, Carl Eckart, Enrico Fermi, Abraham Flexner, George Gamow, Kurt Gödel, Herman Heine Goldstine, Werner Heisenberg, L. van Hove, Cuthbert Corwin Hurd, Pascual Jordan, R. H. Kent, George B. Kistiakowsky, Oskar Morgenstern, J. Robert Oppenheimer, Rudolf Ortvay, Wolfgang Pauli, Marshall H. Stone, Lewis L. Strauss, Abraham Haskel Taub, Edward Teller, Stanislaw M. Ulam, Oswald Veblen, Klara Dan Von Neumann, Warren Weaver, Hermann Weyl, Norbert Wiener, and Eugene Paul Wigner. Gift, Marina Von Neumann Whitman, 1974–1975. Gift, Nicholas A. Vonneuman, 1993.

Volovik:2008:EPF

[Vol08] Grigory Volovik. Emergent physics: Fermi-point scenario. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 366(1877):2935–2951, August 28, 2008. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print),

1471-2962 (electronic). URL http://www.jstor.org/stable/
25197303.

Walsh:1964:FAR

[Wal64a] John Walsh. Fermi Award [for 1964]: Rickover honored; selection signals some changes. Science, 146(3648):1149–1150, November 27, 1964. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1713922;
 http://www.sciencemag.org/content/146/3648/1149.full.
pdf.

Walter:1964:CFP

[Wal64b] Donald O. Walter. Congress and the Fermi Prize. Science, 144 (3620):796, May 15, 1964. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/ stable/1712708. See also [G.64].

Walsh:1974:FNAb

[Wal74a] John Walsh. Fermi National Accelerator Lab: Making the users more at home. Science, 185(4154):841–844, September 6, 1974. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1738994;
 http://www.sciencemag.org/content/185/4154/841.full.
pdf.

Walsh:1974:FNAa

[Wal74b] John Walsh. Fermi National Accelerator Lab: Progress on a grand design. Science, 185(4153):766–768, 770, August 30, 1974. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1738477;
 http://www.sciencemag.org/content/185/4153/766.full.
pdf.

Wald:2004:FLP

[Wal04] Chelsea Wald. Focus: Landmarks: The Physical Review’s explosive secret. Physical Review Focus, 14(??):17–??, ???? 2004. CODEN PRFHAQ. ISSN 1539-0748. URL http://physics.

aps.org/story/print/v14/st17.

Warner:1961:FLD

[War61] R. M. Warner, Jr. Fermi level demonstration. American Journal of Physics, 29(8):529–531, August 1961. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://

ajp.aapt.org/resource/1/ajpias/v29/i8/p529_s1.

Wattenberg:1974:AOT

[Wat74] Albert Wattenberg. ‘All in our time’: The building of the first chain reaction pile. Bulletin of the Atomic Scientists, 30(6):51–57, June 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wattenberg:1982:DEP

[Wat82] Albert Wattenberg. December 2, 1942: the event and the people. Bulletin of the Atomic Scientists, 38(10):22–32, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wattenberg:1988:FSU

[Wat88] Albert Wattenberg. The Fermi school in the United States. European Journal of Physics, 9(2):88–93, April 1988. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic). URL

http://iopscience.iop.org/0143-0807/9/2/002.

Wattenberg:1992:LE

[Wat92] Albert Wattenberg. A lovely experiment. Bulletin of the Atomic Scientists, 48(10):41–43, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wattenberg:1993:BNA

[Wat93] Albert Wattenberg. The birth of the nuclear age. Physics Today, 46(1):44–51, January 1993. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://physicstoday.

scitation.org/doi/pdf/10.1063/1.881378.

Wattenberg:2004:FNA

[Wat04] Al Wattenberg. Fermi and the nuclear age. In Orear [Ore04], chapter 19, pages 83–88. ISBN ???? LCCN ????

URL http://dspace.library.cornell.edu/handle/1813/62;
 http://hdl.handle.net/1813/74.

Weisskopf:1964:FAD

[WB64] V. F. Weisskopf and H. A. Bethe. Fermi Award defended. New York Times, page 28, March 23, 1964. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Williams:1984:AAD

[WC84] Robert Chadwell Williams and Philip L. (Philip Louis) Cantelon, editors. The American atom: a documentary history of nuclear policies from the discovery of fission to the present, 1939–1984. University of Pennsylvania Press, Philadelphia, 1984. ISBN 0-585-17297-8 (e-book), 0-8122-7920-4, 0-8122-1169-3 (paperback). xv + 333 pp. LCCN U264 .A44 1984.

Wilczek:2006:FRM

[WD06] Frank Wilczek and Betsy Devine. Fantastic realities: 49 mind journeys and a trip to Stockholm. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2006. ISBN 981-256-649-X (hardcover), 981-256-655-4 (paperback), 981-277-430-0 (ebook). ix + 522 pp. LCCN QC75 .W55 2006.

Weart:1976:SS

[Wea76] Spencer R. Weart. Scientists with a secret. Physics Today, 29(2):23–30, February 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v29/i2/p23_s1.

Webb:2002:IUT

[Web02] Stephen Webb. If the universe is teeming with aliens… where is everybody?: Fifty solutions to the Fermi paradox and the problem of extraterrestrial life. Copernicus Books in association with Praxis Pub., New York, NY, USA, 2002. ISBN 0-387-95501-1. xi + 288 pp. LCCN QB54 .W384 2002. URL http:

//www.loc.gov/catdir/enhancements/fy0817/2002073910-d.
html;   http://www.loc.gov/catdir/enhancements/fy0817/
2002073910-t.html.

Webb:2015:IUT

[Web15] Stephen Webb. If the Universe Is Teeming with Aliens … Where is Everybody?: Seventy-Five Solutions to the Fermi Paradox and the Problem of Extraterrestrial Life. Science and Fiction. Springer International Publishing, Cham, Switzerland, second edition, 2015. ISBN 3-319-13235-0 (hardcover), 3-319-13236-9 (e-book). ISSN 2197-1188 (print), 2197-1196 (electronic). xv + 434 + 70 pp. LCCN QH327-328. URL http://link.springer.

com/10.1007/978-3-319-13236-5.

Weintroub:1963:FCP

[Wei63] S. Weintroub. The Fermi Chicago Pile. Nature, 197(4866):439, February 2, 1963. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v197/n4866/pdf/197439a0.pdf.

Weiner:1970:PGD

[Wei70] Charles Weiner. Physics in the Great Depression. Physics Today, 23(10):31–38, October 1970. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

www.physicstoday.org/resource/1/phtoad/v23/i10/p31_s1.

Reprinted in [WP85, pages 115–121].

Weinberg:1971:BRE

[Wei71] Alvin M. Weinberg. Book review: Enrico Fermi, Physicist by Emilio Segrè. American Scientist, 59(3):382, May 1971. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL

http://www.jstor.org/stable/27829684.

Weiner:1977:HTC

[Wei77] Charles Weiner, editor. History of twentieth century physics: Storia della fisica del XX secolo, Proceedings of the International School of Physics “Enrico Fermi” = Rendiconti della Scuola internazionale di fisica “Enrico Fermi”, course 57, July 31–August 12, 1972. Academic Press, New York, NY, USA, 1977. ISBN 0-12-368857-4. LCCN QC7 .V37 1977.

Weiner:1994:LFD

[Wei94] Nella Fermi Weiner. Letter: Fermi: Didn’t even tell his wife. Bulletin of the Atomic Scientists, 50(4):3, July/August 1994. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weissert:1997:GSD

[Wei97] Thomas P. Weissert. The Genesis of Simulation in Dynamics: Pursuing the Fermi–Pasta–Ulam Problem. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-98236-1, 0-387-98237-X. xiii + 176 pp. LCCN QC133 .W45 1997.

Wesson:1990:CEI

[Wes90] Paul S. Wesson. Cosmology, extraterrestrial intelligence, and a resolution of the Fermi–Hart paradox. Quarterly Journal of the Royal Astronomical Society, 31(??):161–170, June 1990. CODEN

QJRAAK. ISSN 0035-8738. URL http://adsabs.harvard.edu/
abs/1990QJRAS..31..161W.

Westfall:2016:BRP

[Wes16] Catherine Westfall. Book review: The Pope of Physics: Enrico Fermi and the Birth of the Atomic Age, by Gino Segrè and Bettina Hoerlin, Henry Holt and Co, 2016. $30.00 (368pp.). ISBN 978-1-62779-005-5. Physics Today, 69(12):57–58, December 2016. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Weiner:1972:EHN

[WH72] Charles Weiner and Elspeth Hart, editors. Exploring the history of nuclear physics: proceedings of the American Institute of Physics–American Academy of Arts and Sciences conferences on the history of nuclear physics, 1967 and 1969, volume 7 of AIP conference proceedings. American Institute of Physics, Woodbury, NY, USA, 1972. ISSN 0094-243X, 1551-7616, 1935-0465. LCCN QC173 .E88 1972.

Wheeler:1962:FTN

[Whe62] John Archibald Wheeler. Fission then and now. International Atomic Energy Agency Bulletin, 4(0):33–36, ???? 1962. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-

2690 (electronic). URL http://www.iaea.org/Publications/
Magazines/Bulletin/Bull040su/04005093336su.pdf.

Wheatley:1993:FMV

[Whe93] Joseph Wheatley. Fermi meets van Hove for a new HTC theory. Physics World, 6(1):23–24, January 1993. CO-DEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (elec-

tronic). URL http://physicsworldarchive.iop.org/full/
pwa-pdf/6/1/phwv6i1a20.pdf.

Wheeler:2009:MF

[Whe09] John A. Wheeler. Mechanism of fission. Physics Today, 62(4): 35–38, April 2009. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Wigner:1943:RCW

[Wig43] Eugene P. Wigner. Radioactivity of the cooling water. Report CP-499, United States Department of Energy, Washington, DC, USA, March 1, 1943. 10 pp. URL http://www.osti.gov/

accomplishments/documents/fullText/ACC0143.pdf.

Wigner:1955:EF

[Wig55] Eugene P. Wigner. Enrico Fermi, 1901–1954. In ????, editor, Yearbook of the American Philosophical Society, pages 435–439. American Philosophical Society, Philadelphia, PA, USA, 1955. ISSN 0065-9762. LCCN ???? URL http://www.aps-pub.com/.

Wigner:1962:FAA

[Wig62] Eugene Wigner. Fermi Award: AEC honors Teller for contributions to nuclear science. Science, 138(3545):1087–1088, December 7, 1962. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
1709475;   http://www.sciencemag.org/content/138/3545/
1087.full.pdf.

Wigner:1967:SRS

[Wig67] Eugene Paul Wigner. Symmetries and reflections; scientific essays of Eugene P. Wigner. Indiana University Press, Bloomington, IN, USA, 1967. viii + 280 pp. LCCN Q171 .W65.

Wigner:1996:EF

[Wig96a] E. P. Wigner. Enrico Fermi (1901–1954). In Wigner [Wig96b], pages 115–119. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:CWE

[Wig96b] Eugene Paul Wigner, editor. The collected works of Eugene Paul Wigner. Part A. The scientific papers. Vol. II. Nuclear physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-56972-3. x + 574 pp. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wilson:1968:FTB

[Wil68] Fred L. Wilson. Fermi’s theory of beta decay. American Journal of Physics, 36(12):1150–1160, December 1968. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http:/

/microboone-docdb.fnal.gov/cgi-bin/RetrieveFile?docid=
953;filename=FermiBetaDecay1934.pdf;version=1.

Wilson:1969:HNP

[Wil69] Mitchell Wilson. How Nobel prizewinners get that way. Atlantic Monthly, ??(??):69–74, December 1969. ISSN 1072-7825 (print), 2151-9463 (electronic). URL http://www.theatlantic.

com/author/mitchell-wilson/.

Wilson:1970:CP

[Wil70] Robert Rathbun Wilson. The conscience of a physicist. Bulletin of the Atomic Scientists, 26(6):30–34, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1971:BRS

[Wil71] Jane Wilson. Book reviews: Several lives and more: My World Line, by George Gamow, The Viking Press, New York, 1970, 178 pages; Enrico Fermi: Physicist, by Emilio Segrè, The University of Chicago Press, Chicago, 1970, 276 pages; My Several Lives, Memoirs of a Social Inventor, by James B. Conant, Harper & Row, New York, 1970, 701 pages. Bulletin of the Atomic Scientists, 27(2):47–48, February 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1975:AOT

[Wil75] Jane Wilson, editor. All in our time: the reminiscences of twelve nuclear pioneers. Bulletin of the Atomic Scientists, Chicago, IL, USA, 1975. 236 pp. LCCN QC773.A1 A44. Reprinted by the Educational Foundations for Nuclear Science, Chicago, IL, USA (1975).

Wilson:2001:SEI

[Wil01] T. L. Wilson. The search for extraterrestrial intelligence. Nature, 409(6823):1110–1114, February 22, 2001. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://www.nature.com/nature/journal/v409/n6823/full/
4091110a0.html.

Wilson:2004:FF

[Wil04a] Jane Wilson. The Fermi family. In Orear [Ore04], chapter 23, pages 117–120. ISBN ???? LCCN ???? URL http:

//dspace.library.cornell.edu/handle/1813/62;   http://
hdl.handle.net/1813/74.

Wilson:2004:FP

[Wil04b] Robert Wilson. Fermi and politics. In Orear [Ore04], chapter 22-2, pages 113–116. ISBN ???? LCCN ???? URL http:

//dspace.library.cornell.edu/handle/1813/62;   http://
hdl.handle.net/1813/74.

Wilson:2004:WF

[Wil04c] Robert Wilson. Working with Fermi. In Orear [Ore04], chapter 22-1, pages 107–112. ISBN ???? LCCN ????

URL http://dspace.library.cornell.edu/handle/1813/62;
 http://hdl.handle.net/1813/74.

Winter:1986:BRF

[Win86] M. J. Winter. Book review: The Fermi–Pico–Bagels Logo Game: Developing Thinking Skills Using Words and Lists. Commodore 64 diskette and book, by Larry Wiley. Arithmetic Teacher, 33 (7):55, March 1986. CODEN ARITBF. ISSN 0004-136X. URL

http://www.jstor.org/stable/41192867.

Wollan:1980:ORF

[Wol80] E. O. Wollan. The other record of the first nuclear reactor start-up. American Journal of Physics, 48(11):979–, November 1980. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://scitation.aip.org/content/aapt/
journal/ajp/48/11/10.1119/1.12210.

Wolfenstein:2004:WC

[Wol04] Lincoln Wolfenstein. Wolfenstein comment. In Orear [Ore04], chapter 29, pages 143–144. ISBN ???? LCCN ???? URL http:

//dspace.library.cornell.edu/handle/1813/62;   http://
hdl.handle.net/1813/74.

Wolk:2009:MHB

[Wol09] Herman S. Wolk. Making the H-bomb. Air Force magazine, 95 (3):66–69, March 2009. CODEN AFORCO. ISSN 0730-6784 (print), 1943-4782 (electronic). URL http://www.airforce-

magazine.com/MagazineArchive/Pages/2009/March%202009/
0309H-Bomb.aspx.

Weart:1985:HP

[WP85] Spencer R. Weart and Melba Phillips, editors. History of physics, volume 2 of Readings from Physics Today. American Institute of Physics, Woodbury, NY, USA, 1985. ISBN 0-88318-468-0 (paperback). 375 pp. LCCN QC7 .H694 1985.

Werner:1963:AFI

[WR63] H. Werner and G. Raymann. An approximation to the Fermi integral F_{1/2}(x). Mathematics of Computation, 17(82):193–194, April 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842

(electronic). URL http://www.jstor.org/stable/2003641.

Wright:1964:CSP

[Wri64] Stephen Wright, editor. Classical scientific papers: physics: facsimlie reproductions of famous scientific papers. Mills & Boon, Amsterdam, The Netherlands, 1964. 393 pp. LCCN ????

Wright:1965:CSP

[Wri65] Stephen Wright, editor. Classical scientific papers: physics: facsimlie reproductions of famous scientific papers. American Elsevier Publishing Company, New York, NY, USA, 1965. xix + 393 pp. LCCN ????

Wahl:1948:NP

[WS48] Arthur C. Wahl and Glenn T. Seaborg. Nuclear properties of 93^{237}. Physical Review, 73(9):940–941, May 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.73.940.

Weismann:2009:SFS

[WWL^{+}09] Alexander Weismann, Martin Wenderoth, Samir Lounis, Peter Zahn, Norbert Quaas, Rainer G. Ulbrich, Peter H. Dederichs, and Stefan Blügel. Seeing the Fermi surface in real space by nanoscale electron focusing. Science, 323(5918):1190–1193, February 27, 2009. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.sciencemag.org/content/
323/5918/1190.full.pdf.

Yang:2003:EF

[Yan03] Chen Ning Yang. Enrico Fermi. In Anonymous [Ano03], pages 389–394. ISBN 88-8286-032-9. LCCN ???? URL http:old.enea.

it/produzione_scientifica/pdf.../V2003_AttiFermi.pdf.

Yang:2004:EF

[Yan04] C. N. Yang. Enrico Fermi. In Orear [Ore04], chapter 31,

pages 155–158. ISBN ????  LCCN ????  URL http:

//dspace.library.cornell.edu/handle/1813/62;   http://
hdl.handle.net/1813/74.

Yang:2013:EF

[Yan13a] Chen Ning Yang. Enrico Fermi. In Selected papers II, with commentaries [Yan13c], pages 243–246. ISBN 981-4449-00-8 (hardcover), 981-4449-01-6 (paperback), 981-4449-02-4 (e-book).

LCCN QC21.3. URL http://www.worldscientific.com/doi/
abs/10.1142/9789814449021_0031.

Yang:2013:FDT

[Yan13b] Chen Ning Yang. Fermi’s β-decay theory. In Selected papers II, with commentaries [Yan13c], pages 326–332. ISBN 981-4449-00-8 (hardcover), 981-4449-01-6 (paperback), 981-4449-02-4 (e-book).

LCCN QC21.3. URL http://www.worldscientific.com/doi/
abs/10.1142/9789814449021_0045.

Yang:2013:SPI

[Yan13c] Chen Ning Yang, editor. Selected papers II, with commentaries. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2013. ISBN 981-4449-00-8 (hardcover), 981-4449-01-6 (paperback), 981-4449-02-4 (e-book). x +

34 pp. LCCN QC21.3. URL http://www.worldscientific.com/
worldscibooks/10.1142/8640.

York:1987:MWT

[Yor87] Herbert F. (Herbert Frank) York. Making weapons, talking peace: a physicist’s odyssey from Hiroshima to Geneva. Alfred P. Sloan Foundation series. Basic Books, New York, NY, USA, 1987. ISBN 0-465-04338-0. xiv + 359 + 8 pp. LCCN JX1974.7 .Y575 1987. US$22.95.

Zaghloul:2011:IFP

[Zag11] M. R. Zaghloul. Inconsistency in Fermi’s probability of the quantum states. European Physical Journal H, 36(3):401–406, November 2011. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/
10.1140/epjh/e2011-20018-8. See erratum [Zag13].

Zaghloul:2013:EIF

[Zag13] Mofreh R. Zaghloul. Erratum to: Inconsistency in Fermi’s probability of the quantum states. European Physical Journal H, 38(2): 279, March 2013. CODEN EPJHAD. ISSN 2102-6459 (print),

2102-6467 (electronic). URL http://link.springer.com/
content/pdf/10.1140/epjh/e2013-40005-1.pdf. See [Zag11].

Zuckerman:1995:EWT

[ZH95] Ben Zuckerman and Michael H. Hart, editors. Extraterrestrials — where are they? Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-44335-0 (hardback), 0-521-44803-4 (paperback). LCCN QB54 .E95 1995. URL http://www.loc.gov/

catdir/description/cam026/94043739.html;   http://www.
loc.gov/catdir/toc/cam021/94043739.html.

Ziman:1962:EMSa

[Zim62a] J. M. Ziman. Electrons in metals: A short guide to the Fermi surface. Contemporary Physics, 3(4):241–256, 1962. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ziman:1962:EMSb

[Zim62b] J. M. Ziman. Electrons in metals: A short guide to the Fermi surface. Contemporary Physics, 3(5):321–333, 1962. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ziman:1962:EMSc

[Zim62c] J. M. Ziman. Electrons in metals: A short guide to the Fermi surface. Contemporary Physics, 3(6):401–414, 1962. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ziman:1962:EMSd

[Zim62d] J. M. Ziman. Electrons in metals: A short guide to the Fermi surface. Contemporary Physics, 4(1):1–14, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ziman:1962:EMSe

[Zim62e] J. M. Ziman. Electrons in metals: A short guide to the Fermi surface. Contemporary Physics, 4(2):81–99, 1962. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zinn:1955:FAE

[Zin55] Walter H. Zinn. Fermi and atomic energy. Reviews of Modern Physics, 27(3):263–268, July 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.27.263;   http:/
/rmp.aps.org/abstract/RMP/v27/i3/p263_1.

Zuckerman:1970:EPA

[Zuc70] Harriet Zuckerman. European physicist in America: Book review: Laura Fermi, Illustrious Immigrants: The Intellectual Migration From Europe 1930–41 and Donald Fleming and Bernard Bailyn (eds.) The Intellectual Migration: Europe and America, 1930–1960. Physics Today, 23(8):56–57, August 1970. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://www.physicstoday.org/resource/1/
phtoad/v23/i8/p56_s1.

Zhu:2012:AAT

[ZZWK12] Shengfeng Zhu, Hancan Zhu, Qingbiao Wu, and Yasir Khan. An adaptive algorithm for the Thomas–Fermi equation. Numerical Algorithms, 59(3):359–372, March 2012. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
1017-1398&volume=59&issue=3&spage=359.